A Complete Bibliography of Publications in Technology and Culture: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 September 2023 Version 1.00

Title word cross-reference

1 [Whi63]. 11 [Woo60a]. 111 [And65b]. 1603 [Goo69a]. 1640 [Fus67b]. 1692 [Lan68]. 1700 [Bed64b, Esp67, Rob69, Wil60b]. 1720 [Col61]. 1750 [Bar65b, Wat63]. 175th [Gil66]. 1772 [Sch63b]. 1786 [Bur69a]. 1789 [Hea61]. 1791 [Bra66]. 18 [Hub64]. 1800 [Jaf65a, Wol66]. 1817 [Bur65c]. 1825 [Rya67, War68]. 1830 [Wil60a]. 1837 [Bur68]. 1840 [Bar61a]. 1850 [For62]. 1860 [Cal68d, Koh66, Mil61, Ras61]. 1864 [Bat67]. 1865 [Fer65b]. 1867 [Joh69a]. 1868 [Con66b]. 1876 [McH65, Wil61a]. 1877 [Whi68b]. 1880 [Lev61a, Lon67, Sto69b]. 1885 [Bar66d]. 1890 [Bed61b, Meh61]. 1898 [Bra65a]. 18th [Hol65c, La 64, Ogd66]. 1900 [Bak65, Hin66, Joh65a, K.61, Koh68, Ste68a]. 1903 [Sca68a]. 1909 [Bry68b]. 1911 [Con64a]. 1912 [Gai67]. 1914 [Bar62, San68, Sco61]. 1915 [Gue61]. 1918 [Bry64a, Hug68, Ros65b]. 1920 [Gar67, Har65b]. 1921 [Ols69, Wie64]. 1928 [Bla64]. 1932 [Bla63]. 1933 [Kni67, Zim64]. 1939 [Woo61a]. 1940 [Mul69, Mur69]. 1941 [Emm62c]. 1942 [Wil69b]. 1947 [Gha68b]. 1950 [Fle65, Hid62, Ihd69]. 1954 [Sha67a]. 1959 [Ano60a, Bar63b, hC66, Dau60a]. 1960 [Ano61d, Bar63c, Bru68, De 67a, Fer64b, Fer64d, Goo65b, Sny62]. 1961 [Ano62j, Hei68, Car69, Duv64, Jew64]. 1962 [Ano63g, De 67a, McH66b, Ros65a, Sto64a, Sto65a, Wik64]. 1963 [Ano64e, Sto64a, Gru67, Hof65]. 1964 [Ano65d, Vat68]. 1965 [Ano66g, Duv66, Rei67]. 1966 [Ano67k, Cha68, Hig69a, Pur68a]. 1967 [Ano68g, Ben68]. 1968 [Ano69b, Dor68a, Liv69]. 1980 [Ger66b]. 19th [Jen69, Sha63a, Ros64b]. 19th-Century [Ros64b].

2 [Dor63]. 2000 [Tho65b]. 20th [Str64, Jen69].

41 [Rod61]. 47 [Hew66].

59 [Har62a].

A. [Ogd66]. A.D [Con64a, Nee69]. Aaron [Bee66, Ben66, Rav66, Smi69d]. Abbe [Kin68]. Abbot [Hin62]. Abbott [Smi61c, Ano69a, Smi65c]. Abram [Ham69]. Abrams [Lub65a]. Abriss [Ret61]. Academics [Gul64]. Académie [Cha68]. Accelerators [Shr68]. Accidental [Tof66]. Account [Smi67b]. Accounts [Che61b]. Achieving [Dav62]. Acquisition [Hol65a, Sto63]. Acreage [Ger66b]. Act [Emm63a, Hal62, Whi62]. Acta [Jac69]. Action [Gai67, Gue61]. Activities [All62, Mon67, Sto65a]. Activity [Bar66a]. Acute [Chi63]. Adams [Mon67, Kra66]. Adapting [Bar63a, Ang67]. Adelman [Ayr67]. Adjustment [Jaf67]. Adler [Con64d, Koh69]. Administering [Lew65]. Administration [Dea69, Smi66b, Sto64a, Hof65, Sto64a, Sto65a]. Administrative [Gru67]. administrators [Sto64a]. Admiral [Wil68, Wil68]. Adney [Pru65]. Adrián [Gli66]. Adulteration [Bar67b]. Advance [Car65]. Advances [Ras68a]. Adventures [Smi67b, Hin62, Spi60]. AEA [Hei68]. Aerial [Bed67a]. Aeronautical [Hig69a, Rae60a, Sto64a]. Aeronautics [Emm63a, Emm68a, Ful63, Hig69a, Hof65, Smi66b, Sto64a, Sto65a, Per68, Sny62]. Aeronauts [Sca68a]. Aeroplane [Hun63a, Raz64]. Aeroplanes [Hig65b, Hig66a, Hub64]. Aerospace [Swe67]. Aesthetics [Con66a]. Affair [Tag69]. Affairs [Kle65, All62]. Africa [DeG65a, IU65, Les60]. African [DeG65a]. Africans [DeG65a]. After [Col66, Mes66, Sha67b]. AG [Bry66a]. Agassi [Wis66, Wis67]. Age [Aus63b, Bry66b, Dib64, Fin62a, Gha68a, Gra61, Hei62, Hut66, Jaf65b, Mic66, Ras61, Rie65, Sin69b, Smi64d, Smi66a, Spi68b, Ste68c, Nie62, Ste68b, Mon67]. Agenda [Wei66b]. Ages [Bow66, Whi60, Fin64a, Fus67a, Hal60b]. Agnes [dC61a]. Agraria [Dow64]. Agrarian [Fus66a, Fus67b, Fus67a, Jac64]. Agriculturae [Jac69]. Agricultural [Dal69, DeG65a, Fus66c, Koh67, Ras64b, Süs65, Wik60, Fus68]. Agriculture [Dal64, KH61, Koh66, Koh67, Ras61, Ras64a, Ras67, Ras68a, Spe62b, Bri61, Par66, Min65a]. Agrindus [Ras64a]. Ahead [Tho65b]. Ahmed [dC63b].

Aid [Jun68, Bur62b]. Ainsi [Bry68a]. Air [CB64, Hig65b, Hig66b, Hig66e, Nix65, San61a, Rae69]. Aircraft [Hig66b, Hol69, McC63, Rae62, Top66a]. Airliner [Nil61]. Airlines [Hig65a, Nix65]. Airplane [Sca61, Sca65b]. Airpower [Emm67]. Airship [Top65a]. Airships [Top66a, Top65b]. Aitchison [Smi61a]. Aitken [Cha66, Gue61]. Aix [De 67a]. Aix-en-Provence [De 67a]. Aken [Bak66]. Akron [Top66a]. al [Dow64, Len60]. al. [Arm62, Arn69, Cal68b, Con63a, Con64c, Dav68, Dun68, Fin62c, Goo61, Hal68, Hew67, Hig65b, Hof64, Ihd67, Joh62, Kea68, Lew65, McC66a, McK64b, Pau65, Pau67a, Rae61a, Ras68b, Rav66, Sev65, Sko61, Smi69a, Uye68b, Whi68e, Wol67]. Alan [Bed65a, Ben68, Jac68, Kou66, Man65, Nea64]. Alaska [Hin66]. Albany [Joh69a]. Albert [Bar62, Con61b, Con62c, Far63b, Had61, Had64a, Pet65]. Albertus [Jol68]. Albi [Boy66]. Albrecht [Har65a]. Alchemy [Nee69, Sin60b]. Alexander [Bla64, Bra66, Con69b, Con69a, De 68c, Hug67b, Jaf68, Smi66a]. Alexandria [dC65]. Alfred [Ros65a, Tho65a, Wil64b]. Alice [Hew66, Rie67]. Alienation [Gue65]. Alison [Emm63a]. Alistair [Smi64c]. Allan [Koh64, Wik64, Sil69]. Allchin [Gha69a, Gha69a]. Allegiance [Che61a]. Allen [Hig66a, Kle62, Pau67b, Sca65a]. Allerley [Ede64]. Allied [Fly68]. Allocation [Sko67]. Alloys [Fai62]. all’Unità [Lor67a]. Alois [She63]. Amber [Rei63, Rei63]. America [Aus65, Bed64b, Con61b, Con62c, Con63b, Fer65b, Gra60, Hea63, Hew66, Kur61b, McH61, Men66, Mil61, New66b, Nie67c, Pau65, Pau67a, Pur68b, Ras66, Rya67, Wik62, Pru65, Sal64, Sca68b, Wei63, Wil61c, Woo65, Nie62]. Américain [Bar69b]. American [Bak65, Bar69b, Bir62, Bir66, Bra65a, Bra65b, Bri61, Bur62b, Cal68d, Cha69a, Dav64, De 68d, Dob62, Fin62a, Gar65, Hew67, Hid60, Hol69, Jaf65a, K.61, Kil62, Kre67, Mar69b, Pau67b, PS68, Rub67, Sny62, Svo67, Tho61a, Wat61b, Wei62, Wel65, Wis63, Bil69, Dal64, Gil60a, Hip69, Jew69, Pur62, Rae60b, Ras68a, Sin69a, Spe62b, Wel63, Wel66, BB62, Con68b, Had64a, Hec67, Her62, Hug63, Mah61, Rot69, Sin69b, Sto69b]. Americanizers [Ols69]. Ames [Gil63]. Ami [Bar66a]. Amin [Rod61]. Amitai [Swe68]. Amman [Smi65b]. Ammoniac [Mul65b]. among [Lei69]. Analyses [Ros64b]. Analysis [Che61b, Sto63, Had63, Smi65a]. Anatomy [Rei63]. Anciennes [Smi67a]. Ancient [Con69b, Dru63, Fai62, For64b, Hac68, Had63, Mei66a, Rav65, Smi65a, Wer67, Pac68, For65a, Had64b, Mei68]. Anderson [Kle63a, Car61c]. Andrew [Con64a, Mid63, Hid68]. Andrews [Bri62b]. Andrus [Rod67]. Anfängen [Ros69]. Angus [Pat62]. Anniversary [Gil66]. Annotated [DeG69b]. Announcements [Ano61a, Ano61b, Ano61e, Ano62a, Ano62b, Ano62d, Ano63a, Ano63b, Ano63f, Ano64d, Ano65e, Ano65f, Ano66c, Ano66d, Ano66e, Ano67a, Ano67d, Ano67e, Ano67f, Ano68d, Ano68e, Ano68f, Ano69f, Ano69g]. Annual [Ano63h, Ano64f]. Anoshchenko [Bra69]. Antebellum [Hun63b]. Antecedents [Dor68b, Jew67]. Anthology [Hol69]. Anthony [Hoo65].

Anthropology [Ben68]. Anti [Lok68]. Anti-Utopians [Lok68]. Anticipation [Ram69, Wil65c]. Antiqua [Fer64b]. antiquity [Smi64c, For64b, Kel69c]. Anton [dSP61]. Apollonios [dC61b]. Appalachias [Hun67]. Appier [Wil65b]. Appliances [Dor68d]. Application [Bar61a]. Applied [Bun66b, Fei61, Kur61a]. Appraisal [Fin61]. Appreciation [Wat61a]. Approach [Shu67, Con67a, Rae69]. April [Dor68a, Ano67h, Ano68i, Ano69j, Hig68c, Lew67, Rae68]. Aqueduct [Fer68c]. Arabic [Lei63]. Aratri [Dow64]. Arc [Eis66]. Archaeology [For64a, Gra61, Hud65b, Hud65a, Lea66, Lea69, McC69, Buc68a, Gre65]. Archeological [Sam65]. Archéologique [For64b]. archéologiques [For65b]. Archeology [Vog67]. Arches [She69]. Architect [Con67a]. Architects [Con65a]. Architectural [Con68a]. Architecture [Bur65a, Con61b, Con62c, Con64a, Con66c, Con67b, Con69a, Esp68, Mei66b, Pac68, Con64a, Con64b, Con65a, Kan65, Bur62a, Con69b]. Architektur [San68]. Archives [Sch63a, Svo67, Boy64]. Areas [DeG69b]. Aregger [Con68c]. Arendt [Car60]. Argument [How69]. Aridity [Sal65]. Armement [Smi67a]. Armes [Smi67a]. Armor [Bla62, Hol61a]. Armorican [Bac69]. Armory [Bla62]. Armour [Gra61, Gra65, Hoo65, Wil60b]. Arms [Gra61, Hol61b]. Armstrong [Mil62]. Army [Hig66b, Sko61, Spe62a, Esp65, Hug68]. Armytage [Bar63c, Kil64, Lay66]. Arnold [Bri68, Gra60, Hug61b, Nie65]. Aron [Ces69a, Le 65, Woo68]. Arsenal [Gue61, Hig64]. Art [Aus68, BB62, Con66d, Con69c, Esp68, Gul69, Hal66, Mah61, Ran63, Smi67a, Smi69c, Viv68, SS61, Bur64b]. Arthur [Deu65, Deu66, Esp68, Ros61]. Artificial [Nel66]. Artillery [Jor62]. Artisans [Hea61]. Arts [Cha68, For60, Gul69, Sil69]. Arturo [Gli69a]. Aryan [Puh64]. Asa [Hut63, Hut66]. Ashby [Gul64]. Ashton [Rei65a]. Asian [Wer64a]. Aspecten [Alt65]. Aspects [Bar61b, Bed68a, Car61a, Car62a]. Aspin [Hip67]. Assaying [Dib65]. Assembling [Vog65]. Assertation [Smi68c]. Assistance [Bar67a]. Assyria [Dav61]. Astronautical [Hof64, Sto64a]. Astronautics [Emm68a, Per68, Sny62, Hof65, Smi66b]. Astronomical [Goo61]. Atlas [Per63]. Atmosphere [Dal69]. Atmospheric [Bis69, Fer68a]. Atom [Nie65, Tho61b, Hie63]. Atomic [And62b, Fly68, Kle63a, Hug61b]. Atoms [Nie67a]. Atomuhr [dSP61]. Atramentarius [Moe68]. Attempt [Fei61, Gil60a]. attempts [Sin60b]. Atthreya [Dea69]. Attic [Bul66, Bul62]. Attorney [Kun60]. Attributes [Deu66]. Atwood [Cha69a]. Aubrey [Fer64a]. Aufgaben [For61a]. August [Wei62]. Augustin [Pay64]. Augustin-Normand [Pay64]. Australian [Joh65a, McH66b]. Australians [Pil66]. Authoritarian [Mum64]. Autobiographical [Smi67b]. Autobiography [Hig63, Hig66d, Hig66e, Tho60, Hig66c]. Autogiros [De 68a]. Automata [Bed64d, dSP64]. Automation [Jaf61, Mic64, Ros63, Kol62, Dib64, Kol62, Lay68, Mic66, Rei63, Chr62, Kol64]. Automobile [Rae61b, Rae66, Tho65a, Kyr64]. Automotive [Rae60a, Rae61a]. Autos [Rae61a]. Availability [DeG65b]. Aviation [Bil69, Bra69, Emm69, Pur68a, Rae61c, Smi68c, Süs68a, Emm62a, Tay67]. Aviator [Whi61]. aviatsii [Bra69]. Axe [dC61a]. Axel [Arn69]. Axles [Boy60, Hal61b]. Ayres [Mor63b]. Azizali [Rod67]. Aztecs [Out65].

B [Alt65, Bar65b, Bar66c, Bri63, Cha69b, Dau64, Far63b, Fer67, Fer69a, Fin68, Hin66, Joh62, Jor62, K.61, Kou61, Kre67, Lan68, Min65b, Pau67b, PSD67, Rae69, Rol68, Sko61, Smi68a, Ste63]. B.C [Con64a]. B.C. [Gha69a]. Babbage [Cal68a]. Babel [Kni67]. Babylon [And62a, Bur64a]. Background [Emm67, Hei62, Hig64, Sch63b]. Badawy [Con69b, Con69a]. Baghdad [Kra66]. Bagrit [Mic66]. Baier [New66a]. Bain [Ros67]. Baker [Ber65a]. Balázs [Gab66]. Baldwin [Whi68b, Sca68b]. Ballinger [Tag69]. Ballinger-Pinchot [Tag69]. Ballooning [Sca68a, Sca68b]. Baltimore [Gla62]. Banani [Rod61]. Banat [Fre69a]. Baranson [DeG69b, Mor69a]. Barber [Dun63]. Bark [Pru65]. Barna [Ble63]. Barnard [Fai62]. Barnett [DeG65b]. Barnouw [Kni67]. Barock [Dib69c]. Barometer [Hal65a]. Barrel [Wil64a]. Barrie [Spi69]. Barry [Shr63]. Barton [Bil68, Fer69a, Rol68, Smi68a]. Basil [Top65b]. Basile [Flo69a]. Basin [Tow64]. Bass [Goi69]. Bassermann [Bed65a]. Bassermann-Jordan [Bed65a]. Bates [Sin66, Wie64]. Bath [Buc68c]. Bathe [Cha64]. Battle [Sny66]. Baugeometrie [She69]. Bäumler [Far64b]. Be [Bur69b]. Beals [Ben68, Nea64]. Beam [Dor68b, Jew67, Jew69, Rol68, Con62b]. Beardsley [Bak65]. Bedini [Gar65, Hoo64]. Beer [Kat63, Ihd64]. Beers [Sch63a]. before [Emm67, Gha69a, Meh61]. Beginning [Ihd64]. Beginnings [Bur64b, Bur65a, Fer67, Kin66, Mei60, Spi64b, Sto69a]. Behavior [Lew65, She65c]. behind [Kra66]. Being [Sin60a]. Beings [Wal69]. Beitrag [May67a, May69c]. Bell [Svo67]. Bellifortis [Whi69]. Beloved [Hug61c]. Below [Bar63c]. Belt [Koh64, Roa63]. Ben [Bar66a]. Ben-Ami [Bar66a]. beneath [Cal66]. Benfey [Jen67]. Benito [Con67b]. Benjamin [Kin66, Rie67, Bur69c]. Bennett [Lon65, Pie64]. Bennis [Tei67]. Benton [Bry64b, May69a]. Bereiding [Alt65]. Berengo [Lor67a]. Berenice [De 68d]. Berg [Han67]. Bergson [Ham69]. Berkner [Rie65]. Bern [Hug62b]. Bernal [Bar67c]. Bernard [Ben68, Car61a, Dun63, Wat61a]. Bernardo [Emm62a]. Berndt [Pil66, Pil66]. Bernstein [Ham68]. Berry [Con64c]. Bert [Bar61b, Kan66]. Bertrand [Kel65]. Besson [Bat66c, Dav66]. bet [Alt65]. between [Aga66, Con65b, Mos67]. Bewegung [Ros69]. Beyond [Rae69, Süs68a, Gha67]. Bibliographia [Fer64b]. Bibliography [Ano64a, Bra65b, Fer62a, Fer62b, Fer62c, Fer63b, Fer64e, Fer64f, Fer65c, Goo65c, Goo66b, Goo67b, Goo68c, Goo68b, Goo69b, Hig61a, Hig61b, Dau64, DeG69b, Fer65a, Fer63a, Goo65b]. Biblioteca [Ret67b]. Biennial [Ben68]. Big [Sca61]. Biggest [Gra60]. Billion [Sny66]. Biographical [Hig61a, Hig61b]. Biography [Sca61, Hug60]. Biological [Car62a, Nie69]. Biringuccio [Cah67]. Birkenfeld [May67a]. Birmingham [Hol65c, Tha69]. Birth [Gha69a, Hut63, Vir68]. Births [Mec66]. bis [Bry64a, Woo61a]. Bituminous [Spo63]. Black [Dun69, Hat62, Hou61].

Blackmore [Hoo66b]. Blair [Wil60b]. Blaisdell [And65b]. Blaser [Con66d]. Blätter [Har63, Har64, May68]. Blauner [Gue65]. Blazers [Fin63a]. Blockmaking [Woo66b]. Blowers [Ret65]. Blue [Che61a]. Blunted [Hig65c]. Boaga [Con67b]. Boats [Pru65]. Bobbin [Qui65]. Boeckmann [vW67]. Bogue [Koh64]. Boguslaw [Rie68b]. Bohemia [Ush64]. Boiler [Bur68]. Boilers [Bur66b]. Bold [Hig68a]. Bolino [Wei62]. Bollettino [Kel69a]. Bolognese [Dow64]. Bombay [Gha68b]. Boni [Con67b]. Bonner [Koh66]. Book [Bed65a, Dun69, Ede64, Ihd69, New66a, Jol68]. Booker [Chi65]. Bookmaking [Lei63]. Books [Dau64, Esp66a, Fer64c]. Boorstin [Wis63]. Boris [Con63b]. Borrowings [Wil63]. Boserup [Fus66c]. Bosnia [Ros65b]. Bosnia-Hercegovina [Ros65b]. Boston [Bur65c, Joh69a, Gla62]. Bottles [Had64a]. Boucher [Wat65]. Boughton [Raz64]. Bouguer [Sch62a]. Boulton [Sch65a]. Bowley [Str67]. Bowman [Pea65]. Boyd [Con64a, Edw68]. Bradbury [Bed68a, Mid68]. Brady [Ros63]. Bragdon [Hou63]. Brahmins [Mar69a]. Branch [Zvo61b]. Brand [Gai67]. Brandywine [Wil63]. Brass [Sca68b]. Braun [Emm68b, Sto68]. Bredo [Bur61a]. Breimeyer [Par66]. Brewery [Wil61a]. Brewing [Wil60a]. Brewington [Bed64c]. Brian [Che61b]. Bridge [Boy66, Kou66, McN64]. Bridgebuilder [Kou61]. Bridget [Gha69a]. Brief [Pur65b, Rae62]. Brigadier [Hig64]. Briggs [Hut63, Hut66]. Brinton [Hig66b]. Britain [Bar65b, Car62b, Fle65, Nix65, Ste68c, Ste68a, Ush60, Bar67b, Bat62, Car67, Gai67, PS67, Nix65]. Britinsh [Raz64]. British [Aus63a, Bar61a, Bar64, Bat67, Ble63, Che61b, For64a, Gai67, Gue68, Hig68b, Hig69a, Hug63, Nil61, Str67, Top65a, Hug62c, Kra68, Bar63b, Bar66c, Lam61, Rob69]. Britkin [Ogd66, Kel69b]. Broad [Kou61]. Broadcasting [Hut63, Hut66, Kni67]. Brody [McH61]. Broehl [Woo61b]. Bronze [Fai62, Fai62]. Brooke [Wil65a]. Brooklyn [Kou66]. Brophy [Sko61]. Brough [Sco66]. Brown [Hid68, Con61b, Con62c, Kle62, Man65, Mul69]. Bruce [Bur68, Tho61a, Tof67]. Bruchey [Bir66]. Brumbaugh [Mei68]. Bruno [Kle67]. Brunschwig [Ger66a]. Brush [Eis66]. Buckingham [Kol62]. Buckinghamshire [Fus66b]. Buckminster [Bur65b, Bur66a]. Builder [Ogd66, Con64b]. Builders [Con66b, Fer69b, Kel69b, Mey60]. Building [BB62, Con60c, Con60b, Con62a, Con64a, Con64d, Con68b, Con68d, Con68c, Mah61, Rot69, Str67, Con66a, Hig66a]. Buildings [Bau64, Rot60]. Builds [Car69]. Bulfinch [Bur65c]. Bulleid [Bur65d]. Burchard [Con61b, Con62c, McK64a]. Bureau [Mur69, Pur68c, Mul62]. Burgess [Smi60]. Burnap [Bed61a]. Burstall [Fer64a]. Bursting [Bur66b]. Burton [Pen66]. Bush [Con61b, Con62c]. Bush-Brown [Con61b, Con62c]. Business [Bil69, Fer65b, Goo66a, Goo68a, Kol62, Gra60, Rin64]. Butlin [Joh65a]. Buying [Hig66b].

C [Aus63a, Bar67a, Bat67, Bea66, Bra66, Cal68c, Dav62, Edw67, Fus66b, Gul69, Hea63, Hip67, Koh66, Lay69, Le 64, Lon67, Nie67c, Nix64, Pur65b, Ras66, Ras67, Ric68, Rob65, Sca68a, Sil69, Spe62a, Spo63, Tho61a, Wei62, Wik61, Woo66a, Zim64, Zuc68]. Cables [Dib69b]. Cableships [Dib69b]. Caboose [Whi68c]. caduta [Lor67a]. Cahier [Whi63]. Calder [Hie63]. Caley [Had63, Smi65a]. Calhoun [Wat61b]. California [Sto64b]. Calvert [Pur68b]. Calvin [Pru62]. Cambridge [Col66, Rei67, Fus67a]. Came [Tyl62, Sin60a]. Cameron [Dob62, Sco61]. Camp [For64b, Fin62a]. Campbell [Bra65b, Car62a]. Can [Kat63]. Canada [Goo65b, Arm67, Gra62, Car69]. Canadian [O’D69, Bro65a, Con63e]. Canal [Gla62, Mey60, Skr69, McN64, Gla62]. Canals [Gli68]. Candela [Con64b]. Candler [Col61]. Cannon [Jor62]. Canoeing [Cha69a]. Canoes [Pru65]. Capital [Arm62, Min65b]. Capitalism [PS64]. Car [Gue68, Tow63]. Carbon [Smi64e]. Card [Mur69]. Career [Fin63b]. Carl [Bat66b, Bou67, BB62, Kan65, Mah61, Min65a, Rot69, Kin68]. Carle [Sal65]. Carleton [Lay69]. Carlo [Dow64, Esp67]. Carlos [Gli69a, Gli69b]. Carlton [Lay69]. Carol [Bak66]. Caroline [Har69]. Carr [Aus63a, dC61a]. Carrier [Pau65, Pau67a, Wil68]. Carriers [Top66a]. Carroll [Buc68b]. Carstens [Whi68a]. Case [Bar66a, Car61b, Gab66, Gue62a, Har69, Joh62, Mul65a, Mul65b, Ras68a, Whi61]. Casey [Sca65b]. Cassier [Pur65a]. Cast [Bat62, Fer68c, Nee64]. Cast-Iron [Fer68c]. Casting [Fai62]. Castles [Jor62]. Castro [Dun69]. Catalog [Whi68b]. Catalogs [K.61]. Catalogue [Kil62, Woo69b, Bla62]. Catapult [Hac68]. Catapults [Hac68]. Cathedrals [She61a]. Catherine [Pil66]. Causation [Bir66]. Causes [Bou61, Bur68, Whi62]. Cavour [Lor67b]. Cayley [Hun63a]. CCLE [Tra64]. Cecil [Bed67b]. Celso [Bar65a]. Census [Mur69]. Cent [Hei68]. Centenary [Hig69a]. Centennial [Ano64g, Bur67, Jen67]. Center [Koh67, Out62, Sto65a]. Central [DeG65a, Ble61]. Centre [Buc68c]. Centuries [Ste68c]. Century [Alt69a, Aus65, Bar67b, Car67, Dau64, Fer67, Gra65, Had61, Hig69a, Hol65c, Hug63, Jen69, Kur61a, La 64, Men66, Mul68, Sca68b, Spe63, Str64, Ush64, Whi61, Bis69, Ede66, Pea66b, Ros64b, Sha63a, Whi69, BB62, Hid60, Koh64, Kra64, Mah61, Mul68, Ogd66, Tof66, Yab67]. Cepede [Sev65]. Ceramics [Had60]. Chadwick [Gai67]. Challenge [Bar69b, Bri68, Gul61b, Smi67d]. Challenges [CB64, Emm64]. Chandler [DeG65b, Tho65a, Wil64b]. Change [And68a, Bar61a, Bur61c, De 64, DeG69a, Fus66c, Hig68c, Jaf63, Lon67, Mes65, Mum61, PSD64, Ram69, Smi64c, Wei66a, Wel63, Col66, Har69, Kol64, Rie68b, Rod65, Str67, Str63, dM67, Bri68]. Changing [Hal62, McG65, Sel69, Tei67]. Chanute [Fer65a]. Chaos [Con63b]. Chapelle [Pru65]. Chaplin [Coh69]. Chapman [Hip67]. Chapters [Wil64b]. Character [Jar66, Big64]. charlatans [Sin60b]. Charles [Bed64b, Bry66b, Bur62b, Cal68a, De 68a, Eva65, Fer68a, Fin64a, Fle65, For60, Hec67, Hin62, Kin66, Lew64, Lew65, Lub65a, McC66b, Out65, PS68, Rae66, Ras67, Tei68, Boy64, Kra60]. Charleston [Hea61]. Checkland [Bar66d]. Chemical [Lei61b, LD60, Nie69, Sko61]. Chemicals [Duv66]. Chemie [Far64b]. Chemistry [Bak65, Duv64, Ihd69, Lei61b, Lei63, Bee66, Ben66, Far61, Far63a, Far63b, Far64a, Far65, Far68a]. chemists [Sin60b].

Chevreul [Far63b]. Ch’i [tZS65]. Ch’i-chün [tZS65]. chia [hC66]. Chibnall [Fus66b]. Chicago [Kan65, Ano60a]. Childers [Pet65]. China [Bil68, Nee69, tZS64, Cha61, Did62, Fai62, Goo61, Goo65a]. China-Clay [Bil68]. Chinese [hC66, Yab67, Nee64, Sam65, tZS65, Con64a]. Chivalry [Bac69, Gra61]. Christensen [Gra62]. Christenson [Spo63]. Christian [Bro62, Had61, Had64a, Cla62, Bry68b]. Christopher [Con63b, Pur69a]. Chronicle [Sca68b]. Chronology [Per68, Smi66b, Sny62, Sto64a]. chün [tZS65]. chung [hC66]. Cincinnati [Con68d, Con66b]. Cipolla [Esp67]. circa [Wil60b]. Cirurgia [Ger66a]. Cities [Gha68c, War68, Gut64]. City [Bur64a, Goi69, Liv69, McK64a]. Ciudad [Gli69a]. Civil [Con60c, Mer65, Pat62, Rei62a, Sch63a, Vog65, Wat61b, Wei66a, Fin64a]. Civilian [Sch64b]. Civilization [Gha69a, Ras68b, tZS64, Buc68b, For65a, Gha69b, Mei62b, Mor63a, Spi64b]. Civilizations [She68, Gha67]. Clack [Gue68]. Clagett [Hal60b]. Clairmonte [Kam63]. Clapham [Bar65b]. Clapp [Rei65b]. Clara [Dor68a]. Clarence [Mor63b]. Clark [Agn67, Fin62c, Rae65, Wik61, Wil68, Bur61d, Car68, Erm64, Sco66, Spi68b, Ste68a, d’A68]. Clarke [Nie69, Top65b]. Class [Smi67e, Shu67]. Classical [Bur66d, Gil62, Kel69c, Lev61b, Fus67c, Lev61b]. Classrooms [Agn67]. Claude [Wil60b]. Claus [Koh69]. Clawson [Tan65]. Clay [Bil68]. Clayburn [Wol66]. Clayton [Bur66d]. Clepsydra [Bed62]. Cleveland [Tas66]. Clifford [McK64b]. Climate [Ful63]. Clock [Bed61b]. Clockmaker [Bed61a]. Clockmakers [Bed64a]. Clocks [Bed65a, Gli69c, Goo61, Bed65a]. Clockwork [Goo61]. Clough [Phi63]. Clue [Bur61c, Mum61]. Clutton [Bed67b]. Clyde [McC66a]. Coal [Spo63]. Coast [Bra65b, Bur69c, Sca65b]. Coast-to-Coast [Sca65b]. Cockcroft [Lew66]. Codding [Kni65]. Codes [Ben63]. Codex [Ret67a]. Codices [Ret67b]. Cogs [Fer64c]. Cohen [McC67]. Cold [And68b, Jaf61]. Cole [Bar63b]. Coler [Col65]. Colin [Con64b, d’A68]. Collar [Che61a]. Collected [Sha67a]. Collection [Bed64c, Fer64c, New66b]. Collections [Svo67]. Collector [Bed65a]. College [Lay69, Ras67, Smi69b]. Colles [Pur69a]. Collins [Gli69a]. Colonial [Rot69]. Colony [Lan68]. Colt [Hol61b]. Colvin [She65a]. Coman [Goo66a]. Combustion [Sch63b, Har67]. Comes [Dal64, Smi69c]. Comet [Nil61]. Coming [Aus63b, Bri62a, Smi64d, Wik61]. Comment [Bur69b, Con65b, Dru69, Dup69, Flo69b, Mes69a, Ram69, Sch65b, Wal69]. Commentary [Bar66c, Cal62a, Hug61a, Hut62, Hux62, Kra62, Mei62a, Ong62, Pie62, Rae63, Sto62a, Tha62, Wer68a]. Comments [Dru61a, Jar66, Mes69a, Wis66]. Commerce [Bil69]. Commission [Mon67, Kle63a]. Committee [Ano68b]. Common [Ran63]. Communication [Gol63, Lei69, Dib66]. Communications [FMG61, Sto69a]. Communist [Did62]. Community [PR62]. comp [Bee66, Dau64, Mar63, Pur65a, Svo67]. Companies [Woo61b]. Company [Che61b, Dau60a, Pau67b, McH63, Fer67, Hol61b]. Compartmented [Bed62]. Compass [She65b]. Comprehensive [Bur66a, Had63].

Computable [Man65]. Computer [Smi67d, Chr61]. Computers [Tei68]. Concept [Kel68, Zvo61a]. Concerning [Pur68d, Zvo61a]. Concise [Bar65b, Lev61b]. Concrete [Con67b, Con68d]. Condit [BB62, Kan65, Mah61, Rot69]. Condition [Car60]. Conditions [Fus66c, Gai67, Smi64c]. Conduct [Bil69]. Conestoga [Pau65, Pau67a]. Conference [De 67a, Emm69, Sch64b, Sha67b, Tra64, Uye64]. Conflict [Pur68b]. Confronting [Mor67]. Confusion [Aga66]. Congress [Per68]. Connecticut [Tra65]. Conover [Pie66]. Conquest [Dow65, Tav69]. Consensual [Smi68c]. Consequences [De 64, Whi62, Tav69]. Conservation [McC69, Tag69, Zim64, Pen66]. Conservatoire [Cha68]. Considerations [Bar63a, Emm69]. Construction [Hun63b, She61a]. Consumption [Ber65b]. Contact [Fin68]. Contemporary [Hou63, San60]. Context [Whi61]. Contexts [Whi62]. Continents [WW66]. Continuities [Whi62, Arn65]. Continuity [Wer68a]. Contracting [Nie67a]. Contribution [Pac68, Wel66]. Contributions [Fer62a, Fer62b, Fer62c, Fer63b, Fer64e, Fer64f, Fer65c, Rae61a, Woo60a, Woo69a]. Control [Bar67b, Bur62b, Con63b, Mil64, Pur68a]. Controlled [Dal69]. Controlled-Atmosphere [Dal69]. Controversial [Smi69c]. Convergent [Mul68]. Conversion [Tho66]. Cooke [Bra66]. Cooling [Eic64]. Cooper [Con62b]. Cooperation [Ano68b, Sko67, Lon69]. Coordinator [Kni65]. Copland [McG65]. Copper [Wer65]. Corn [Koh64]. Cornelius [Rie67]. Corning [Had68, Lew69a]. Cornish [Bil68, Fer69a, Rol68]. Cornwall [Fer69b, Smi68a]. Corporate [Hea63]. Corroboration [Wis66]. Cosmonautics [Bra69]. Costa [Far63b]. Costly [Dib69a]. Cotton [Bar61a, Fed60, Gha68b]. Coulston [For60]. Council [Rin65, Rin66, Rin65]. Counseling [Gue67]. Countdown [Sca66]. Countermeasures [Fly68]. Countries [Bar63a, De 68b, Gol62, Bar69a, De 68c]. Country [McC66b, Wel67]. County [Hud65a]. Court [Bar65b, Bar66c]. Cowan [Con68a]. Cox [Dav68]. Craft [Wel63, Hou63]. Crafts [PS66, She68, Smi69a]. Craftsman [Hig66e]. Craftsmen [Kel69b, Wel67]. Craig [Spi60]. Crank [HS69]. Creamer [Arm62]. Created [And68b]. Creation [Kel68]. Creative [Tay67]. Creativity [Col65, Maw65, Car61c]. Credit [Min65b]. Crisis [Jaf61]. Critical [Ede64, Had63]. Crombie [Smi64c]. Crosfield [Duv66]. Cross [Hub64]. Crossbow [Las66]. Crosson [Rie68a]. Crossroads [Gul61a]. Crouzet [And64]. Crowleys [Wer64b]. Crucial [Sch63b]. Cubbedge [Mic64]. Cueco [Gli66]. Culbertson [Rei64]. Culp [New66b]. Cultivation [Wik61, Car61c]. Cultural [Arn65, Bur62b, Ces65, Con61b, Con62c, Lei67, PSD64, Son63, Flo69a]. Culture [And68a, Big64, Dav68, Dru61b, Les60, Spi69, Wer68a, Hou63, Ben62a, Bro62, Mar69b, Ano60b, Ano61c, Ano62c, Ano63d, Ano64b, Ano65b, Ano66a, Ano67b, Ano68a, Ano69d]. Culturelle [Smi67a]. Cultures [Le 64, Pur68b, Rob65]. Cummins [Mor69a]. Cup [Fra64]. Current [Ano64a, Fra64, Goo65c, Goo66b, Goo67b, Goo68c, Goo69b]. Curt [Con64a]. Curtiss [Joh69b]. Custom [dC64a, Ben63]. Cutting [Har60, Bat66c].

Cybernetics [Rie68a]. Cycle [Bry67b]. Cycles [Sco66]. Cyclopaedia [Fer68c]. Cylindrical [Bed62]. Cypress [Hid68]. Cyril [Bur60a, Bur67, Bur69a, Meh61, Wat65]. Czechoslovakia [Ble61, Ros65a].

D [Bar67c, Bra69, Bri61, Con63d, Fer69a, Fus68, Gha68a, Ham68, Hip67, Hug68, Hun62, Lew64, Lew65, McH65, McN64, Mur69, Nil61, Rae62, Rol68, Sil69, Smi68a, Ste68c, Svo67, Swa67b, Tho65a, War68, Wil64b, Wil67]. Dairy [Koh68]. dal [Dow64, Len60]. Dale [Gue62b]. Dales [Phi63]. dalla [Lor67a]. Dallas [Ano69b]. Damascene [Wul65]. Dams [Swa67a]. Dan [Hid62]. Daniel [And65a, And65b, Arm62, Bed61a, De 68c, Flo69a, Wat61b, Wis63, Fra64]. Daniels [Bed67b, Net65, Sin69b]. Danilevskii [Alt69a]. Dankert [McC66a]. Darkhouse [dC65]. Darrah [New66b]. Darrett [Lan68]. Dart [Spi60]. Daumas [For63a]. Davenport [Flo69a]. Davey [Con62a]. David [Bla63, Coh69, Con63e, Con69c, Dav62, Fer65b, Fly68, Gha68b, Gue60, Hig65c, Hug61c, Kou61, Las67, McH61, Nie62, Ras67, Swa67b, Wei66a, Whi67, Wil69a, Wat61a, Wil65a]. Davidson [Bul62, Smi66a]. Davies [Hig65a]. Dawn [Mei62b]. Dawson [Goo65a]. Day [Bur65d, Rav65, Hew66]. Days [Fer61]. Dead [For65a, PR62]. Dean [Smi69d]. Deane [Bar63b]. Debate [Bla64]. Decade [Bur65b, Bur66a]. December [Ano60a, Ano61d, Ano62j, Ano63g, Ano64e, Ano65d, Ano66g, Ano67k, Ano68g, Ano69b, Bra66, Hei68, Sto65a]. Decision [Gue62a, Smi69c]. Decline [Cha64, Gil60a, Wik64]. Decorating [Hou63]. Decoration [Con66c]. Decorative [Bat62]. Defense [San63, Spe64, Hig68b]. Défi [Bar69b]. Definitions [Fei61]. Degan [Pie64]. Degna [Sha63b]. della [Bon69, Kel69a, Len60, Lor67a]. dell’Istituto [Kel69a]. Demand [Ger66b]. Democratic [Dau62, Jac64, Mum64, Rin66]. Demontage [Ros68]. Dempster [Nil61]. Denis [For60]. Dennis [Spi60]. Department [Aus68, Spe62a, Mul65a]. Derek [And62a, Con66c, Nil61]. Derry [Pay63]. Description [Spe63]. Descriptive [Dau64]. Design [Ayr67, Bur65b, Bur66a, Hig65b, McC63, Pac68, Rie68b, Wel66, Con64a, Con68c, McK64b]. Designers [Pie66]. d’Espagne [Smi63]. Destiny [Le 65]. Detection [Bar67b]. Determinants [Bir62]. Deterrence [Emm67]. d’Ethnologie [Spi68a]. Detroit [Deu65]. Deutsche [Whi68e]. deutschen [Bry64a, For66, Hug69, Ros69]. Deutschland [Fry61]. Deutz [Bry66a]. Developed [Bar69a, Bou66]. Developing [Bar63a, De 68b, De 68c]. Development [Ben66, Bur62b, Bur61d, Cha61, hC66, Col61, Dal69, DeG65a, De 65, DeG69a, Emm62b, Emm62c, Emm63a, Emm69, Erm64, Flo69a, Fry61, Gha65, Hig65b, Hug67b, Joh65a, Koh67, Lew65, Lon67, Lor67b, Meh61, Mur69, Nie67c, Pea66b, Pen63, Pur62, Pur68c, Rae61a, Ras66, Rod65, Rod67, Sco61, Sha63a, She68, Shr68, Wei62, Wil69a, Wol66, Zvo62, Koh67, Sko67, And68a, Ayr67, Dea69, Dru68, Eva65, Fro66, Had61, Hat62, McC66b, Min65a, Rod62b, Tan65, Ush61, Wol67, Woo68, Bar65a]. Developments [Ful63, Bak67]. Dexter [Ano69c]. Diagnosis [Chi63]. Dialectics [Ces69a].

Dibner [Hug62b]. Dickey [Wil69b]. Dickinson [Rol65]. Dickson [Gue67]. Dictionaries [Fer63a]. Dictionary [Bed65a, Dav64]. Dictionnaire [For60, For64b, For65b]. Diderot [For60, For60]. Diego [Mul69]. dient [Süs68b]. Diesel [Bry66b, Mor69a, Rae61a, Whi68a, Whi68a]. Differences [Jaf68, Ros67]. Diffusion [La 64, Puh64, Rog69, WW66]. Dilemma [Bri62b]. Dimensional [Mor67]. Dimensions [Liv69]. Diplomacy [Hat62]. Direct [Net65]. Directe [Alt65]. Directions [Dau62]. Directory [Gra62]. Discontinuity [Maz67]. Discoveries [Wis67]. Discovery [Goo67a, Hug62b, Smi64e, Smi64c]. Discussion [Pre62a, Pre62b, Pre62c, Pre62d, Pre62e, Ret65]. Disillusion [Ces69a]. Disintegration [Kam63]. Divine [Hig65c]. Division [Per68]. Diyala [Kra66]. Dizard [Wei66b]. Do [Hei67]. Dobson [Bur66d]. Dobzhansky [Gre63]. Doctors [Gai67]. Doctrine [Jor61]. Document [Bur65b, Bur66a]. Documentary [Men66]. Documents [Ste67, Ste68b, Bar66c, Whi63]. Dodds [Bur66d]. Dodge [Bak67]. Doggle [Swe68]. Doig [Uye68a]. Dollar [Sny66]. Dome [Wie64]. Domestic [Swe68]. Don [Lon69, Smi69c]. Donald [All62, Con64c, Gol65, Mul62, Pie66, Pur68a, Smi62, Smi67e, Zim64]. Donnell [Mec66]. Donovan [Kol66]. Door [Wul66a]. Dora [Cha69b]. Double [Eis66, dC61a]. Double-Arc [Eis66]. Douglas [Lyn67, McG65, Smi67d]. Dowd [Lyn67]. Down [Hud65a]. Dr. [Gil60b, Wis67]. Drachmann [For64b]. Drainage [Dib69a]. Draper [Sha65]. Drawing [Qui65, Chi65]. Dream [Fer61, Wis63]. Dreamer [Ste63]. Drehbank [Woo61a]. Drew [Cla66]. Drill [HS69]. Drive [Bou67]. driven [Kel69b]. Droze [Swa67a]. Drucker [Dru61a, Zvo61a]. Drug [Bar67b]. Dual [Che61a]. Dubos [Ang67]. Duck [Piz68]. Due [Arm67, Wil61c]. Duncan [Gul61b]. Dunlop [Kol64]. Dunsheath [Sha63a]. DuPont [Ros66]. During [Gar67, Hid60]. Dutch [Hou61]. Dwiggins [Smi69c]. Dyeing [Ede64, Ede66]. Dyke [Swe65]. Dynamics [Spe69, Vat68, Rae66].

Earl [Wat68]. Earle [Con61a, Had63, Smi65a]. Earliest [Bar65b, Ede64]. Early [Bed67c, Esp67, Fer64c, Fus66a, Gor68, Hal63, Hew64, Jor62, Kil63, Kin66, Lei63, Lev61a, O’D69, Pia61, Ret67c, Rot60, Spe62b, Sto65a, Wel65, Wer64b, dC64b, Bri62a, Dav61, Bur68, Dib66, Gar65, Hoo65, Hub64, Sch66]. East [DeG65a, Pea65, Ros65a, Con69b, Gha61]. Eastern [She68]. Eat [Bor65b]. Ebenezer [Pur69b]. Eberhard [Wer68b]. Eckhardt [Bed61b]. Ecology [Ihd67]. Econometric [Rub67]. Economia [Dow64]. Economic [Aus65, Ayr67, Bak67, Bar63a, Bar61b, Bar63b, Bar65b, Bar66c, Bir66, Ces69b, Che61b, hC66, Col61, Col66, De 64, DeG65a, De 67a, De 67b, DeG69a, Eva65, Fus67a, Ger66b, Goo68a, Hat62, Hea63, Hei68, Hol65a, Jaf65a, Joh65a, Kea68, Koh67, Lub65b, Man65, McC66b, Mil62, Min65a, Mor61, Par66, Pea66a, Rod62b, Rub67, Sco61, Sko67, Sol68, Sto63, Tho61a, Tra65, Wil69a, WW66, Bar69a, Fle65, Kam63, Koh67, Spo63, Wei66a]. Economics [DeG65b, Fus66c, Ham68, d’A68, Ber67, Fro66, Kan66]. Economische [Alt65]. Economist [Sch60]. Economy [Bak67, Bru68, hC66, Jac64, Sch64b, Tho61a, Vat68, Wil61a, Bar64, Ble61, Ces67, Lon65, McG65, Meu64, Pea65, PS64, Sel69, Wei62]. ed [And64, And65a, Arn69, Ayr67, Bak66, Bar66b, Bed68a, Ben68, Bra69, Bra66, Bri61, Bro62, Buc68b, Bur67, Bur69a, Cal67, Cal69b, Car61a, Car62a, Car61c, Ces65, Ces67, Cha66, Che61b, Chr61, Cla66, Col65, Col66, Con63a, Con63c, Con63e, Con65a, De 68c, Dea69, Dun63, Emm64, Flo69a, For60, For63a, Fro66, Fus67b, Fus67a, Gli69a, Goo65a, Goo65b, Goo68a, Gra62, Gue60, Gul61b, Gut64, Had69, Hal68, Han67, Har67, Hew67, Hof64, Hol69, Hug67a, Ihd67, Ihd69, Jac69, Kan66, Kel69a, Kol62, Kol64, Lay69, Le 65, Lew66, Man65, May69b, May68, McC66a, McC66b, McC67, McH66a, McK64a, McK64b, Mei62b, Men66, Meu64, Min65a, Min65b, Nee69, Nie62, Pen66, Pet65, Pie64, Rae69, Ras68b, Rav66, Rie67, Rie68a, Roc66, Rol67, Sal65, San61a, San64, Sch66]. ed [Sel69, She65a, She65d, Shr68, Smi64c, Smi69d, Son63, Spe69, Spi64a, Sto65b, Swa67b, Tan65, Tei68, Tof67, Tow64, War68, Wei66a, Whi68a, Whi68e, Zim65, dM67]. Edelstein [Kel69c]. Edison [Hug62a, Hug60]. edited [And65b]. editor [Fus67b, Ano63c, Ano65a]. Editorial [Hal60a]. Edmund [Fer69b]. Edson [Süs68a]. Eduard [Duv64, Jen69]. Eduardo [Con60a]. Education [De 68d, Did62, Gul61a, How69, McC66b, Cal69a, Eva65]. Edward [Bak65, Boe63, Kle62, Nie67b, Pie64, Ras66, dC63a, dM67]. Edwards [Rae66]. Edwin [Gai67, Goo66a, Ihd64, Pru65, Smi68c]. Effect [Jaf63]. Efficiency [Ret67c, Ros66, Har65b]. Effort [Nie67c]. Egypt [Con69b, Dav61, San61b]. Egyptian [Con69a, HS69]. Ehrenfried [Lev61b]. Eicher [Min65a]. Eight [Sca61]. Eighteenth [Kra64, Ush64, Bis69]. Eilmer [Whi61]. Eisenbud [Kle64]. Eisenmenger [Vat68]. Eisner [Lay68]. EJC [Kle65]. Elbers [Gul61b]. Electric [Arm67, Gor68, Wil61c]. Electrical [Dib64, Dib66, Hig60, Hig61a, Hig61b, Hug62c, Sha63a, Wri68]. Electricity [Spe62b]. Electromagnetism [Hug62b]. Electron [Kra68]. Electron-Tube [Kra68]. Electrophysicists [Hig61a, Hig61b]. Elektrische [Dib69c]. Elektrodynamik [Süs68b]. Eleventh [Whi61]. Eli [De 68b, McC67, Fed60, Woo60b]. Elihu [Hug61c]. Elite [Lay67]. Elizabeth [Ber65a]. Elizabethan [Esp66a]. Ellul [Fal65]. Elman [Smi64b]. Elting [Rie67]. Embellished [Sca68b]. Emergence [Gha68b]. Emerges [Wat63]. Emerging [Gol62]. Emil [Bea66]. Emme [San61a, Sca67, Sny62, Sto65b]. Empire [Bra65a, Con69a, Mei66b]. Empires [Esp67]. Employment [Sha67b]. Encyclopedia [For60]. Encyclopedie [For60]. End [Bis69, Flo69b]. Energies [Car62a]. Energy [And62b, Hug61b, Kle63a, Shr68, Tho61b, Net65, Tho61a, Tho66]. Enfield [Gil63]. Engine [Cle61, Fer69b, Ker61, Pur69a, Pur69b, Rae61a, Rae62, Sch65a, Wil69b, Bur61d, Erm64, Nix64, Rol65, Rol68, Vog69]. Engineer [Pur68b, Ste63, Wat61b, Wat65]. Engineering [Cal68b, Chi65, Con60c, Con63a, Emm69, Fei61, Fer64d, Fin61, Fin62b, Fin64a, Hal61a, Hig60, Hof64, How69, Kle65, KH61, McK64b, Pie66, Pur65a, Rae61c, Ret63, Sch66, Sto69b, Wal69, Flo69a, Sch64b, Vog65, Fer64a, Fin65a, For61c, Hal68, Har67, Kil64, McK65, Mer65, Pet65, Sha63a, Sut66b, Fin63b, Gul69, Rei67]. Engineers [Flo69b, Hig61a, Hig61b, Kle65, Lew65, Rol67, Wri68, For64b, Bur65e]. Engines [Dor68d, Rae61a]. England [Bar63c, Bar66d, Fus67b, Goo69a, Vat68, Wil60a, Bri62a, Buc68c, Hol65c, Jor62, Wik61]. Engler [Dau62]. English [Bur61d, Erm64, Koh68, Esp65, Boy69]. Enos [Bur63]. Enrico [Sha67a]. Enterprise [Lon67, Tho65a, Car61b, Cha66, Wil64b, Duv66]. Entrepreneurship [Lon66, Ush64]. Entstehung [Ros68]. Entwicklung [Fry61, Hoo62, Rot65, Woo61a, May69d]. Environmental [Ces67, Con63a, Kle64]. Epilogue [Kim64]. Epoch [Hig68b]. Equilibrium [Sch65b]. Equipment [McK64b, Pie66]. Era [Har65b, Lev61b, Fus67c, Con63d, McH61, Wei66a]. Erasmus [Bur62b]. Ercker [Dib65]. Erection [She61a]. Erfinderleben [Ste62]. Erfindung [Bry67a]. Eric [Gul64, Haw66, Wel65]. Erich [Bri65, Tow65]. Ericsson [Fer61]. Erie [Gla62, McN64]. Erik [Ayr67, Kni67]. Erlich [Bla64]. Ernest [Gue62b, McK65]. Ernst [Bar67b, Bed65a, Far64b, Fry61, Hoo62, Kin68]. Ernst-Otto [Fry61]. erste [Kle63b]. Erwin [And62b, Gra62]. Escol [Sch66]. Esmond [Con62b]. espagnol [Smi63]. Essay [Ben63, Bir66, Gul64]. Essays [Rub67, Woo68, Lay69, Ber67, Col65, Fer69a]. Establishments [DPS66, Lew66]. Estates [Bur61a]. Ester [Fus66c]. Esthetics [Bur62a]. Eternal [Bur64b, Bur65a]. Ethiopia [De 68b]. Ethnography [Jac64]. Ethnologia [Spi68a]. Ethnology [Smi64b]. Études [Smi67a]. Etzioni [Swe68]. Eugen [Rae61a, Sca66, Sut66b]. Eugene [Hat62, Hig66e, Kre67, Pur65a, San61a, Sca67, Sch66, Sny62, Sto65b, Far63b]. Eurasian [Sam65]. Europaea [Spi68a]. Europe [Bru68, Col66, Fus67a, Nie67c, Sco61, Fox69, Hun67, Wai68, Jaf67, Wat63]. European [Cal68d, Esp67, Puh64, Wer64a, Wil63, Wil60b]. Européenne [Spi68a]. européens [Smi63]. Evaporation [Bis69]. Eve [Ful63]. Events [Sto64a]. Everywhere [Bur64a]. Evidences [Out62]. Evolution [And64, Bed65b, Gre63, Mid68, Arn65]. Evolving [Gre63, Rie67]. Ewart [Gra61]. Excellence [Lea68]. Excitement [Rol64]. Executives [Flo69a]. Exhibit [Lea68]. Exhibition [End68]. Exhibitions [Fer65d]. Exhibits [Woo69b]. Expanding [Ste68a]. Expansion [Bra65a, Esp67, Pau65, Pau67a]. Experience [DeG69a, Mor69a]. Experiment [Mes66]. Experiments [Bar63c, Bis69, Sch63b]. Exploitation [Buc62]. Exploration [Bry64b, Dry61, Sny62, Emm62b, Hin66]. Explorations [Cha66]. Explosions [Bur68]. Extensions [Wer66b].

F [Agn67, Arm67, Bar61b, Bee66, Boy64, Bra65b, Bur69c, Cha69a, Che61b, Cla66, De 64, Esp66b, Fer64a, For64a, Hec67, Hig64, Jac68, Kan66, Le 64, Lyn67, Par66, Rod67, Ros65b, Smi66a, Smi67d, Smi67b, Ste68c, Sut66a, Whi68c, Wil61c]. F. [Ras64b]. Faber [Con64b]. Fable [Tra65]. fabricating [Wil69a]. Fabrics [Ede66]. Fabulous [Sca65a]. Face [Jaf61]. Faces [Nie62]. Facsimile [Ede64]. Fact [Kou66]. Factors [McK65, Pie64]. Factory [Gue62a, Gue65, Gue68]. Failed [Hig65b]. Fairchild [Nie62]. Fakhry [dC63b]. Fall [Ano60f, Ano61k, Ano62i, Ano63l, Ano64k, Ano65j, Ano66k]. Family [Bed64a]. Far [Hig68a, Smi64a]. Farben [Hug66]. Farber [Ras68b, Duv64, Jen69]. Farewell [Whi67]. Farmer [Bri62b, Koh68, Ras61]. Farming [Fus67c, Koh64, Koh67]. Farms [Lan68]. Farrand [Fra64]. Farrington [Net65]. fascination [Sin60b]. Fate [Bur64a]. Father [Dob62, Sha63b]. Faulkner [Ber65a]. February [Bra66]. Federal [Hec67, Pur68a, Sch63a, Bur66b, Zim64]. Feldhaus [Har61, May67b]. Feldman [Hug68]. Fenced [Wil67]. Fens [Dib69a]. Ferdinand [Ben63]. Ferguson [Pur65a, Sch66]. Fermi [Sha67a]. Ferriere [Spe63]. Ferries [Tow63]. Fertilizer [Lor67b]. Fever [Hig63]. Fidelle [SS61]. Fiefs [Fus66b]. Field [Sko61]. Fields [Fus66b]. Fifteen [For61b]. Fifteenth [Spe63, Whi69]. fighting [Dor68d]. Figure [Had64a]. Filarete [Eic64, Ret65, Smi64f, Spe63, Spe64]. Filler [Dav64]. filosofi [Bun66a]. Finance [Che61b, Rod67]. Financing [Arm62]. Finbar [Wat68]. Finberg [Fus67b]. Finch [Fit68, For61c]. find [Sin60b]. Finder [Mum65]. Findings [Mon67]. Finest [Gue60]. Finishes [Hou63]. Fire [Dor68d, Fox69, Lei61a, Dor68d]. Fire-fighting [Dor68d]. Firearm [Bat67]. Firearms [Esp65]. Firms [Ble63]. First [Con68d, Dru66, End68, Etz66, Hun62, Jen69, Jew69, Pil66, Roa63, Rot69, Sca65b, Sch63b, Whi69, De 67a]. Firth [Min65b]. Fisch [Jaf65b]. Fischer [McH63]. Fisher [Sko61, Cal68d, De 68d]. Fitch [Bur62a, Con68b]. Fitchen [She61a]. Five [Hip69, Lea69, Rae61a]. Flaherty [Car61a]. Flesh [Mar69b]. Flew [Sca65a]. Flieger [Kle63b]. Flight [Sca65b, Sca66, Sto64a, Sto65a, Sut66a, Sut66b, Car61a, Hun62, Sca67, Ste63]. Flights [Hig68a, Sut66a]. Flinn [Wer64b]. Florence [Bon69, Ros64a]. Flores [Gli69a]. Florman [Gul69]. Flying [Hig66d, Top66a]. Focus [Cal62b]. Fogel [Car61b, Rub67]. Folklore [Hou61, McN64]. Font [Gli66]. Fontaine [Rei62a]. Food [Ras60, Ras66, Sev65, Bor65a, Gra60, Ras68b]. Forbes [Hal68, Lev61a, Fer64b, Lev61a, Tav69]. Force [Gha68b, Hig65b, Ces65, Coh69]. Forces [Con68b, Hig66b, Hol61a]. Ford [Rae61b, Tho65a, Wik62, Wik64]. Foreign [Joh65b, Kre67, Nie65]. Forest [Mil62]. Foreword [Ben62b]. Form [Con64a]. Formation [Arm62]. Forrest [Wei64]. Forsyth [Fer65b]. Forti [Len60]. Fortification [dlC63]. Fortifications [Jor62]. Fortified [Gha68c]. Fotografie [New66a]. Foundation [Pur62, Sch64b, DeG64]. Foundations [Fin65a]. Founder [Sil69]. Four [Bry67b, Mor67, Str67]. Four-Dimensional [Mor67]. Four-Stroke [Bry67b]. Fourastié [Bou61]. Fourth [Maz67, Sto64a]. France [Fle65, Mil64, Sco61]. Frances [Wil67]. Francesco [Ret63]. Francis [Hec67, Rei62b]. Francisco [Ano66g]. Francois [And64]. Frank [Kol66, Rod62b, Wik64]. Franklin [Bur68, Kin66, Sin69a, Tof67]. Franz [Har61, May67b, Ret61]. französische [May69c]. Fraunberger [Dib69c]. Fred [Ste65]. Frederic [Had64b, Hol65a, Sto63]. Frederick [Emm68b, Eva65, Her62, Kam63, McC66b, Rie68a, Wil65c]. Free [Sch64a, Sch64a]. Freedom [Gue65, Par66, Jac68]. Freeman [Nie67c].

Freight [Pau65, Pau67a]. French [Col61]. Freudenberger [Ush64]. Friederich [For63b]. Friedrich [Bry64a, Oli60]. Fritz [Dib69c, Far68b, Gue67, Hug67a, Gor69]. Frontier [Sto68, McH65]. Frontiers [Hew64, Smi64a]. Fruit [Dal69]. Fryling [Har67]. Fuchs [Jaf68]. Fuel [Emm62c, Tok68]. Fuller [Bur65b, Bur66a]. Fulton [Vir68]. Function [Bar67c, Flo69a]. Functional [Rot60]. Functions [Gha68a]. Fundamental [Sko67]. Funds [Hug67b]. Furniture [Hec67]. Furtado [Bar65a]. Fussell [Koh68]. Future [Dor68a, Dry61, Gut64, Lok68, Rei65b, Woo60c, Mic64, Sca66, Tof67]. Fysh [Hig66d, Hig69b, Hig66d].

G [Bak66, Bar69a, Bar63c, Bar66d, Bed68a, Bou67, Cha66, Fer66a, For64b, Gue61, Hig64, Hig65a, Hol69, Hug66, Hug67b, Jaf65b, Joh65a, Kel69a, Kil64, Kle63a, Koh64, Koh68, Lay66, Lok68, Mid63, Ros61, Rya67, San60, Sca61, Sil69, Smi67b, Ste68a, Swa67b, Tei67, Wat65, Wil68, Woo61b]. G.M [Tho65a]. Gablehouse [De 68a]. Gadgets [Mei68]. Galaxies [Ste65]. Garden [Woo65]. Gardiner [Hea63]. Garfield [Gue68]. Garretts [Bur66c]. Garrick [Con64d]. Gartmann [Cle61]. Gas [Dau64, Sca68b]. Gasindustrie [For66]. Gasoline [Hug66]. Gates [Ras61]. Gauge [Whi68b]. Gauthier [Mor69b]. Gauthier-Larouche [Mor69b]. Gear [Har60]. Gear-Cutting [Har60]. Gebändigte [Bry67a]. geistige [Ros69]. Gen [San60]. Gendai [Yam68]. General [Ful63, Fus67b, Mes69b, Fer64d]. Generale [Dau60b, For63a]. Generation [Gol65]. Genèse [Pay64]. Genesis [Had61, Fre69b]. Genevieve [Goo65b]. Genius [Cal68a]. Geoffrey [Hig63]. Geography [IU65, May69a, Smi68b, Son63]. Geometric [Bul62]. Geometry [Har60]. Georg [Far61]. George [Bed61b, Bed67b, Ber65b, Ber67, Chr61, Emm67, Gil64, Gli69a, Hou61, Hun63a, Jaf61, Kni65, Les60, Mil62, Pau65, Pau67a, Rei63, Sch66, Sin69b, Sko61, Tow63, Vog69, Wil61c, Wol67, McH63]. Georges [Mor69b]. Georgetown [Skr69]. Georgia [Koh66]. Geotechnics [May69a]. Geraint [Wel67]. Gerald [Gul61b, Hug68, Jaf65b]. Gerd [Ols69]. Gerhard [Kle63b]. Gerlach [Süs68b]. German [Dor63, Jac64]. Germany [Fly68, Fry61, Hug68, Ros65a]. Gershenson [And65a, And65b]. Geschichte [Bat66b, Bry67a, For63b, Gol66a, Har65a, New66a, Ret61, Bry64a, For66, Har61]. geschichtliche [Hoo62]. geschichtlichen [May67b]. Gesellschaftliche [May69d]. Gesenken [Hoo62]. Giant [Bur65d, Tho65a]. Gibson [Out65]. Giedion [Bur64b, Bur65a]. Gilbert [Dor68d, Woo66b, Woo69b]. Gilchrist [Wei66a, War68]. Giles [Mul69]. Gilfillan [Gil60b, Zuc68]. Gille [Kel65]. Gillian [Rin68]. Gillispie [For60]. Gin [Fed60]. Ginestier [Boe62]. Ginzberg [De 68b]. Giorgio [Con67b, Ret63]. Girl [Qui65]. Gizitsu [Yam68]. Gizitsu-ron [Yam68]. Gladius. [Smi67a]. Glass [Had60, Had61, Had64a, Had67, Had68, Had69, Per69, Sch62b, Had68, Had64b, Lew69a]. Glasses [Had63]. Glaus [Con68c]. Glazier [Smi65b]. Glenn [Har67, Hun62, Joh69b]. God [Bro65b, Wik61]. Goddard [Sto65a, Emm62c, Hol65b]. Golay [Rod62b]. Gold [Dib65, Smi66a, Sin60b]. Goldbeck [Bry66a, Bry67a, Bry68b, Ros69, Ste62]. Golden [Hut66, Ste68b]. Golem [Bro65b]. Gollin [Rin68]. Goodman [Ihd67, Fer66b, Rob69]. Goodwin [Swe68]. Gopalpur [Nea64]. Goran [Far68b]. Gordon [Bur66d, Hug62a, Ihd67, Jaf67]. Gothic [She61a]. Gould [Lay67]. Goumaz [Smi63]. Governing [Wat68]. Government [Gai67, Hig65b, Hig68c, Lon67, Nie67b, Pur68c, Pur68d, Rae68, Sko67, Sko67, Car67, Rin65, Sko67]. Governments [And62b]. Grabar [Con66c]. Gradation [Sch62a]. Graf [Bat66b]. Graffiti [Boy69]. Graham [Mon67]. Grahame [Spi68b]. Grainger [Bur66d]. Grancsay [Bla62]. Granick [Wil69a]. Granulating [Smi64f]. Grazia [Kur63]. Great [Bat62, Car68, Gai67, Goo61, Gue62b, Piz68, Tow63, Tow65, Wat65, Tow64]. Greater [Smi68b]. Greek [For64b, Hoo65, Lei61a, Mei68, Hac68, Smi66a]. Greeks [Lei69]. Greeley [Hin62]. Green [Hud65a]. Greenberg [And65b, And65a]. Greenleaf [Rae61b]. Greville [Cha64]. Griffith [Emm63a]. Grimwood [Sto64a]. Grinding [Bed67c]. Groniowski [Mos67]. Ground [Ran63]. Groups [And62b]. Growing [Ces67, Hud65b]. Growth [Bar63b, Bir66, Ble61, Ble63, DeG65b, De 67b, Eva65, Fle65, Fus66c, Hei68, Jaf65a, Kea68, Man65, Rod65, Rub67, Tho65b, Vat68, Ver64, War68, WW66, Bar61b, Ces69b, Goi69, Man65, Spe69]. Grubbe [Bea66]. Gründe [Bry69]. Grundlagen [Gol66a]. Grunwald [Gha61]. Guatemala [PS66]. Guatemalan [PS64]. Guerlac [Sch63b]. Guest [Jaf63]. Guidance [Kin66]. Guide [Bar66a, Gul69, K.61, McK64b, Pie66, Fou64, Whi68a, Sch63a, Svo67]. Guidelines [Bar67a]. Guinea [De 64]. Guinness [Wil61a]. Gunner [dC61b]. Gunpowder [Lei61a]. Guns [Esp67, Hoo66b]. Gustav [Bry66a, Bry67a, Bry68b, Ros69, Ste62]. Gustavson [Bou67]. Guthrie [Mil62]. Guy [And64]. Gwembe [Spi69]. Gyorgy [McH66a].

H [And68a, Bak65, Bar63c, Bar65b, Bar66c, Bat67, Bea66, Bed61b, Bed65a, Boy67, Bur61d, Car68, Car61c, Col66, Con60b, Con63c, Con63e, Con66b, Con67a, Dea69, Dob62, Emm62c, Emm67, Erm64, Fer67, Flo69a, Fus66a, Fus67b, Gha68a, Hig65b, Hig66a, Hol65b, Hug63, Joh69b, Jor62, Kil64, Lam61, Lay66, Lok68, PSD67, Pil66, Put67, Rei63, Rod65, Rol65, Rot65, Sco66, She65a, Sin69b, Smi68c, Sto69b, Tho61a, Tho65b, Tof67, Wat61b, Whi68a, dM67]. Habakkuk [Col66, Hug63]. Haber [Gor69, Har65b, Sel69, Far68b]. Habit [Bri68]. Hadfield [Fer68a]. Hägerstrand [Rog69]. Haigerty [Mar63]. Haigh [Dib69b]. Haim [Ras64a]. Hal [Ras64b]. Hall [Swa67b, Vog65, Fer68b]. Hallam [Fus66a]. Halle [Kle63b]. Halperin [Ras64a]. Hamilton [Bra66]. Hampe [Ras66]. Hance [IU65]. Hand [Sha65, Wel66]. Hand-Tool [Wel66]. Handlin [McK64a]. Handwerkerbilder [Whi68e]. Hannah [Car60]. Hans [Con68c, She68, Tho65b, Wer68b]. Hanzelet [Wil65b]. Happened [Wis63]. Harald [Kin68]. Harbison [Eva65, McC66b]. Hard [Ste67]. Hargraves [Alt67]. Hargreaves [Hip67, Alt67]. Harlan [Tow65]. Harland [Fin63a]. Harold [Agn67, Bra66, Bur65c, Car61c, DeG65b, Dib64, Joh65b, May69b, Mul68, Nie67a, Par66, Sny66, Swe68, Whi68a]. Harrington [Tof66]. Harris [Wer65]. Harrison [Mil61]. Harrod [Bar64]. Harry [Bar69a, Ces65, Con65a, Spe62a]. Hart [PSD67, PSD67, Ste63]. Harteveld [Alt65]. Harvard [Mes65, Mes66, Mes69b]. Harvester [Ras68a]. Harvey [Dav68]. Harveys [Fer69b]. Hatcher [Tow65]. Hausbuch [Whi68e]. Havilland [Hig63]. Hawaii [Koh69]. Hawkes [Spi64b]. Hayes [Cha69c]. Hayle [Fer69b]. Haynes [Pea65]. Health [Deu65, Gai67]. Heating [Fer67]. Heavenly [Goo61]. Heavens [Bar63c]. Hebron [Hid68]. Hedges [Pea66a]. Heinz [Cle61, Hin64]. Heissluftmaschine [Bry69]. Helen [McH66b]. Helfgott [Joh62]. helice [Pay64]. Helicopters [De 68a]. Heller [Con64b]. Helmets [Gra65]. Henderson [Che61b]. Hendrichs [Ret61]. Hendricks [Hug62a]. Henry [Car67, Ces67, Con68a, Edw67, Esp66a, Gol66b, Hal65b, Hec67, Ihd69, Jaf63, Lay68, Rae61b, Rei62a, Sch63a, Sch63b, Swa67a, Tra65, Wei66b, Wil67, Fra64, Wik62]. Herausforderung [May69d]. Herbert [De 68b, Hig68a, Smi66a, Rae65]. Hercegovina [Ros65b]. Here [Smi69c]. Heritage [O’D69]. Herman [Swe67, Ush64]. Heroic [Fin62a]. Heron [Rae62]. Heterogeneous [Sch65b]. Hewlett [Kle63a]. Hiebert [And62b]. Higgins [Bla62, Ras64b]. High [Fry61, Hol65b, Shr68, Swa67a]. High-Energy [Shr68]. Higham [Nix65, PS67, Top65a]. Highlands [Mos67]. Highrise [Con68c]. Highway [Wol67, Kat63]. Highways [Kou61]. Hilken [Rei67]. Hill [Con66c, Wik64]. Hillegas [Lok68]. Hills [Dib69a]. Hilton [Rie67, Wil61c, Tow63]. Hindle [Wil65a]. Hiroshima [Emm67, Kle62]. Hirsch [Dun63]. Hirschmeier [Lon66]. Histoire [Whi63, Cha68, For63a]. Histones [Cal68b]. Historia [Hug67a]. Historian [McK64a]. Historic [Ric68, Sca65b]. Historical [Bee65, BT69, Bur61c, Car62b, Con68b, Con68a, Dup69, Emm69, Fin61, Har60, Hat66, Kle67, KH61, Smi64c, Smi66b, Sto65a, Tha69, Woo69b, Bed68a]. Historically [dSP65]. Histories [Meu66]. Historiography [Dre67, dSP65]. History [And62b, And68b, Ano64a, Ano64g, Ano68b, Aus63a, Bar65b, Bar66c, Bed64d, Bee66, Ben62a, Ber67, Big64, Bil68, Bri61, Buc68c, Bur63, Bur67, Bur69a, Cal69a, Car68, Chi65, Col66, Con62a, Con68d, Con69a, Dau60a, De 67a, Dib69a, Dre67, Dun68, Emm63b, Far63a, Far64a, Far65, Fer62a, Fer62b, Fer62c, Fer63b, Fer64e, Fer64f, Fer64d, Fer64a, Fer65c, Fer66b, Fer68b, Fin62b, Fus66a, Fus67b, Fus67a, Gab66, Goo65b, Goo65c, Goo66b, Goo67b, Goo68c, Goo68b, Goo69b, Gru67, Gul69, Hal60a, Hal65a, Hei67, Hid62, Hig65a, Hin64, Hol61a, Hol65c, Hug61a, Hut63, Hut66, Jor61, Kil63, Kil64, Kin66, Kle63a, Kni67, Koh66, Kra66, Kra60, Lev61a, Lev61b, Meh61, Mer65, Mul62, Mul65a, Mul65b, Mum61, New66a, New66b, Nun63, Oli60, Pay63, Pur64, Rae60b, Rae61c, Rei60]. History [Rol65, Sca67, Sca68a, Sha63a, She65a, Shu67, Sin60a, Smi61c, Smi61a, Smi61b, Smi62, Smi68a, Smi69b, Sto65b, Sto69b, Tho61a, Tib67, Top65b, Tow65, Whi68c, Wil64a, Wil64b, Woo60a, Woo60c, Woo66a, Woo69a, Zap62, Zvo60, Zvo61b, Bri65, Cla66, Con60b, Con61b, Con62c, Fer69a, Goo68a, Hig64, Lay66, Les60, Lev61a, Men66, Pea66a, Pur65b, Rub67, Bra69, Bra65b, Emm68b, For62, Gha65, Har60, Hol69, Lei61a, Mei66a, Rae62, Ush61]. Hitler [Hug66]. Hochdruckdampfes [Fry61]. Hodge [Sal65]. Hodges [Bea66]. Hoffmann [Smi66a]. Hofmann [Lev61b]. Hogg [Hig64]. Hollander [Ros66]. Holley [Hig66b]. Holmyard [Sin60b]. Home [O’D69]. Homer [Qui65]. Hoopes [Bed61a]. Hoos [Chr62]. Hoover [Rae65]. Hope [Hew66]. Hopewell [Ric68]. Horizons [Hig68a]. Horizontal [Ret65, Ret67c]. Hoselitz [Bar61b, Kan66]. Hospitals [Mec66]. Hour [Gue60]. Hours [McC66a]. House [Fly68, dC61a]. Housework [Rav65]. Howard [Con67a, Gar67, Hoo66b, Pru65, Tow64, Zim65]. Hower [Lew64]. Hoyle [Ste65]. hsing [Yab67]. Hubler [Sca61]. Hudson [Hig66d, Hig69b, Gre65]. Hugh [Cha66, Emm64, Gue61, Jun68]. Hughes [McH66b, Rol67]. Huguenots [Col61]. Human [All62, Bur65b, Car60, Dru68, Eva65, Flo69a, Gre63, Le 65, Lew64, Liv69, Rie67, Ste67, Ste68b, Tra64, Wal69, Wer66a, McK64b, McK65, Meu64, Pie64, Pie66, Spi64a, dM67]. Humanism [Har62b]. Humboldt [Bry66a]. Hundred [Fer64c]. Hung [Nee69]. Hunger [Dun69]. Hunt [Sha65]. Hunters [Spi68b]. Hurford [Wil64a]. Husa [Smi69a]. Husbandmen [Lan68]. Hy [Hun62].

I-Beam [Dor68b, Jew67]. I. [Hug66]. Ian [Pen66]. Ickx [Bry68a]. ICOHTEC [Ano68b]. Ida [Chr62]. Idea [Hir64, Kan66, Kel69c, Kam63]. Ideal [Woo65]. Ideals [De 68d]. Ideas [Cal68b, Hug61a, Meh61, Tra65, Tay67]. Ideology [Woo68]. Ihde [Bee66, Ben66]. Ihr [Kin68]. ihrer [May69c]. II [Bur65a, Col66, Fer64b, Hut66, Mei66a, Sha67a, Sko61, Spe62a, Ano61c, Fer62b, Hig61b, Pat62, She65b, Cal67]. III [And65b, Bur66a, Lew65, Ano62c, Emm68b, Fer62c, Lew64, Wil69b, Cal69b]. Ijzer [Alt65]. Ikonologie [May69c]. Illinois [Koh64, Cal68b]. Illustrated [Con60b, Mer65, Sca68b]. im [Dib69c, Fre69a, Ros68]. Image [Wis63]. Images [Bou66, Swe68]. Imitation [Gla62]. Immigrants [Ols69]. Impact [And62b, CB64, Kol62, Rav66, Rie65, Roc66, San61a, Tei68, Wol67, Bru68]. Imperial [Nix65, Pac68, Smi69b]. Implements [Fus68]. Implications [All62, Hei68, Mes69b, Swe68]. Improvement [Wei66b]. Inaccurate [Top66b]. Inc. [Bro65b]. Incandescent [Gab66]. Incentives [Hol65a]. Incline [Skr69]. Including [Sto64a]. Income [hC66, Sha67b]. Increased [Ros66]. Incredible [Hun62]. Independent [dSP65]. Index [Ano60b, Ano61c, Ano62c, Ano63d, Ano64b, Ano65b, Ano66a, Ano67b, Ano68a, Ano69d, Fer64d, Pur65a]. India [Gha68b, Gha69a, Mor69a, Gor68, Mor67, Gha68c, Gha65]. Indian [Gha69a, Nea64, PS64, PS66]. Individual [Par66]. Indo [Puh64]. Indo-Aryan [Puh64]. Indo-European [Puh64]. Indonesian [Fra63]. Indus [Gha67, Gha69b]. Industrial [And68a, Bar61a, Bar66d, Ble63, Buc68a, Coh69, Col66, Deu65, Fin62c, Gab66, Gha68b, Har69, Her62, Hig65b, Hip69, Hud65b, Hud65a, Ihd67, Kra64, Lea66, Lew64, Lon67, McC69, Mon67, Rei65a, Ros67, Rot60, Sal64, Smi68b, Ste67, Vog67, Wai67, Wat63, Wil64b, Woo68, Lea69, Bur61a, De 68d, Deu66, Gre65, Gue68, Maw65, Ros65a, Ver64]. Industrialism [Fin62c]. Industrialization [Bla64, Cal68d, Mul65b, Rod62a, Ush60, Bur61a, Gha61, Ols69, Ros65b]. Industrie [Hug69]. Industries [Ras64a]. Industriestaates [Ros69]. Industry [Arm67, Bar61a, Bir62, For60, Ham68, Hig68c, Hol69, Hol65c, Hug62c, Hug68, Kra68, La 64, Lor67b, McH66b, Mil62, Mil64, Mos67, Pur62, Rae68, Str67, Wil60a, Wis61, Wol66, Aus63a, Bil68, Che61a, Fin64b, Gru68, Gue65, Nie67b, Rae61a, Rae66, Ras66, Ros63, Swe67, Tho65a, Wer64b]. Inez [Sha65, Wei63]. Influence [She68]. Influences [Wer64a]. Informal [Fou64]. Information [Goo66a]. Ingalls [Con68d]. Ingenieur [For61a]. Ingénieurs [Kel65]. Ingenuity [Cal66]. Initiative [Bur65b]. Injury [Chi63]. Inland [McC65]. Innocence [Flo69b]. Innovation [Esp67, Fin62b, Gla62, Hid60, Sch65a, Sko67, Smi69d, Str64, Whi61, dC64a, Bur63, Rog69]. Inquest [Wei66b]. Inquiry [Aus65, Kur61b]. Institute [Bur68, Sto64b, Sin69a, Cal69a, Smi67e]. Institution [Vog67, Wri68, Woo69a]. Institutional [Bou67]. Institutions [De 68d]. Instruments [Bed64b, Bed64c, Gar65, Ger66a, Kil62]. Insull [Wei64]. Integration [Ras64a]. intellectual [Smi64c]. Intellectuals [PS67]. Intelligence [Nel66]. Interaction [Ham68, Smi61b]. Interchangeable [Woo60b]. Intercity [Arm67]. Interim [Vog67]. Interlingual [Fer63a]. International [De 67a, Fer65d, Kni65, Swe68, Tan65, Ano68b, Jun68, Rod65, Ros67]. Internationale [Spi68a]. Interplanetary [Sut66a]. Interpretation [Bra65a, Che61b]. Interrelatedness [WW66]. Interstate [Dau60a]. Interurban [Con63d, Wil61c]. Introduction [Emm63b, Sto62b, Woo69b, Car62b]. Introductory [Mei66b]. Intruders [Nie67b]. Invasion [Nie67b]. Inventing [Gil60a]. Invention [Han67, Kin66, Lei67, Rei60, Sch65a, Whi62, Whi68f, Smi64c, Fin62a, Ces69b, Zuc68]. Inventions [Hug63, Smi61c]. Inventive [Gil62]. Inventor [Hun63a, Cal68a, Maw65]. Inventors [Bur65e]. Inventory [Bur65b]. Investigation [Bur68, Deu66]. Investment [Ros64a, Wol67, Ble63, Joh65a]. Investor [Joh69a]. Iowa [Koh64, Koh67]. Iran [Rod61]. Iraq [Rod62a]. Irascible [Cal68a]. Ireland [McC65, McC69]. Irish [Wil61a]. Irma [Ayr67]. Iron [Cha61, Con62b, Fer68c, McH65, McH66b, Nee64, Pia61, SS61, Smi64f, Spe63, Wer64b, Zap62, McH63, Aus68, Aus65, Rei65a]. Ironwork [Bat62]. Irrigated [Ger66b]. Irrigation [Gli68, Gli69c, d’A68]. Irving [Hig66b, Fly68, Las67]. Isaac [Gol66b]. Ishino [Lon65]. Islamic [Con66c]. Isles [For64a]. Issue [Kun60, Sch60, Wil61b]. Issues [CB64]. Istituto [Bon69]. Istokov [Wyn60]. istorii [Bra69]. Italian [Spe63]. Italiano [Kel69a]. Italy [dlC63]. IV [Bur66a, Fus67b, tZS64, Ano63d, Fer63b]. Ivor [Ste63]. Iwao [Lon65]. IX [Bra66]. Iz [Bra69]. Izobretenie [Han67].

J [And62a, Aus63a, Bar69b, Bar61a, Bar67c, Bee66, Ben66, Ben68, Ber67, Bri68, Buc68a, Bur62b, Cha66, Chi65, Col66, Con68a, DeG65b, Esp66a, Far63a, Far64a, Far65, Fer64b, Fin63b, Fin64a, Goo68b, Goo68a, Gue61, Gue67, Hal68, Hal65b, Hug63, Hun63a, Ihd64, Jor62, Lei61a, Lev61a, Mar63, McK65, Mer65, Mor61, Nee69, Rei67, Rei62b, Rie68a, Rod67, Rot60, Sin60b, Sko61, Smi68b, Smi69c, Sto63, Sto69a, Svo67, Tav69, Tow64, Uye68b, Wei66b, Wel67, Wer65, Wie64, Wil68, dC63a, vW67]. J.-J [Bar69b]. Jack [Car62b, DeG69b, Mor69a]. Jackson [Sin69b]. Jacob [Bra66, Ces69b, Koh69]. Jacquelyn [Bak66]. Jacques [Bat66c, Bry68a, Fal65, Dav66]. Jacquetta [Spi64b]. Jahre [Woo61a]. Jahrhundert [Far64b, Fre69a, Hug67a]. Jahrhunderts [May69c, Whi68e]. James [Boy67, Bur62a, Bur65c, Con68b, Emm62a, For61c, Gar67, Gru68, Hew67, Hol61b, Koh66, Las66, McH65, Pat62, Pau67b, Pet65, Pie64, Rei64, Sha63a, Smi67e, Sto64a, Tag69, Wol66, Fit68, Gar67, Hip67]. Jameson [Uye68a]. Jamieson [Cha69a]. Jamnitzer [Bed68b]. Jan [Ble61, Mos67]. Janes [Wil64a]. January [Ano67g, Ano68h, Ano69i]. Japan [Lon67, Gra65, Loc69, Lon66]. Japanese [Lon65]. Jarrett [Ces67]. Jay [Lay67]. Jean [Bou61, Bul62, Jaf68, Pea65, Wil65b]. Jeanne [Gai67]. Jenkins [Wel67]. Jennett [Wyn60]. Jenny [Hip67]. Jerry [Whi68a]. Jet [Nil61, Sto64b, Sut66a]. Jeweler [Smi67c]. Jewelry [Smi66a]. Jiri [Gue62a]. Joachim [Gha61]. Joan [Fus67b]. Jobs [Deu66, Dru63]. Jochmann [Fry61]. Jocko [Wil68]. Joe [Ros67]. Joedicke [Con65a]. Johannes [For66, Lon66]. John [Arm67, Bar65b, Bla62, Bra65b, Bur65b, Bur66a, Con60b, Con61b, Con62c, Con63a, Con66b, Dau60a, Dow65, Fer61, Hid68, Kat63, Kol64, Lew66, Lon65, McK64a, Mil62, Mil64, Mor63a, Ric68, Rol64, She61a, Wil61a, Wil61c, Wat65]. Johns [Rei62a]. Johnson [Nie62, Smi67e, Bar69a]. Johnston [Pat62]. Joint [Car68]. Jones [Fou64]. Jongh [PS66]. Joravsky [Bla63]. Jordan [Bed65a, Fre69a]. José [Gli66]. Josef [May68]. Joseph [Bul66, Cha61, Duv66, Flo69a, Goo61, Gul61a, Lev61b, Maw65, Pie64, Rei62a, Rin64, tZS64]. Josephson [Hug60]. Josiah [Sch62b]. Jost [Smi65b]. Josué [Dun69]. Journal [Hig69a, Gil66, Had67, Had68, Had69, Tha69]. Journey [Mec66]. Jousse [SS61]. Jr [Dun68, Hew67, Ras66, Shr63, Svo67, Str63]. Jr. [Buc68b, Con67a, Goi69, Goo66a, Gut64, Hig66b, Hig68a, Hun64, Kle63a, Kni65, Lay68, Mon67, Sto69b, Tag69, Tho65a, Wil64b, Wol66, Woo61b]. Juanelo [Ret65, Ret67a, Wul66b]. Julius [Gla62]. July [Bra66, Ano67i, Ano68j, Ano69k]. Juno [vB63]. Jupiter [vB63]. Jürgen [Con65a].

Kōzan [Gra65]. Kahler [Bri65]. K’ai [Yab67]. K’ai-wu [Yab67]. Kaiserliche [Fre69a]. Kaplan [Bar66b, Kur61b, San64]. Kapp [McC67]. Karinya [And68a]. Karl [And68a, Rot65, Woo61a]. Kármán [Süs68a, Mal64]. Karoly [Viv68]. Kash [Lon69]. Kates [Pen66]. Kathleen [Rod62a]. Katona [Ber65b]. Kauffman [Hal65b]. Kazimier [Mos67]. Keenleyside [Jun68]. Keiichi [Yam68]. Keith [San61b]. Kekulé [Jen67]. Keller [Fer66a]. Kellermann [Bat66a]. Kellogg [Ras67]. Kenneth [Gre65, Lew69a, Nel66, Rie68a, Sch63a]. Kentucky [Hal65b, Pea65]. Kepes [McH66a]. Kephart [Pru62]. Kerr [Fin62c]. Kettering [Boy64, Edw68]. Keyboard [Bow66]. Keystones [She69]. Kieffer [Bee66]. Kimball [Hew66]. Kindleberger [Fle65]. King [Koh69, She65a, Sha63a, Wer65]. Kingdom [Con69a, Hut63, Hut66]. Kingery [Cal68b]. Kingsford [Bur65e]. Kip [Fit68, For61c]. Kipp [Goo68a]. Kircher [Jaf61]. Kirker [Bur65c, Bur65c]. Kisch [Kle67]. Kitson [Ste68a]. Kittler [Hun62]. Klaus [May69c, May69d]. Klaw [Mar69a]. Kleine [San68, Har65a]. Kleingewerbemaschinen [Bry69]. Klemm [For63b, Oli60]. Klinckowstroem [Bat66b]. Klöckner [Bry66a]. Knabbe [Esp66b]. Knapke [Whi68c]. Knapp [Ras67]. Knaurs [Bat66b]. Knibb [Bed64a]. Knigopechatanija [Wyn60]. Knossos [dC61a]. Know [Rae60b]. Know-How [Rae60b]. Knowles [Hal65a, Sch62a]. Ko [Nee69]. Kolaja [Gue62a]. Kompfner [Kin66]. Konrad [Con64a]. Korman [Ols69]. Kornhauser [Deu65, Fin64b]. Korol [Hug67b]. Korson [Hou61]. Körting [For66]. kosmonautiki [Bra69]. Kouwenhoven [Kat63]. Kraft [Bry67a, Bry66a]. Kramer [Big64]. Kramish [Hug61b, Nie65]. Kranzberg [Buc68b]. Kuhn [Gru68]. Kulturgeschichte [Bat66a, She63]. Kung [hC66, Yab67]. Kung-chia [hC66]. Kunkel [dC63a]. Kurt [Bry69, Gha61]. Kurze [For63b]. Kuznets [De 67b]. Kyeser [Whi69]. Kyle [Con60b].

L [Cal68c, Chr61, Dau60a, De 68c, Dib69a, Fer66b, Fin62a, For64b, Gai67, Gru67, Hew67, Hol61b, Hoo66b, Ihd65, Joh65b, Jun68, Mei66b, Nix64, Put67, Rie65, Rob69, Sca68a, Son63, Spo63, Wat63, Woo66a, vW67]. Labor [Coh69, Gha68b, Hug68, Jaf67, Mec66, Sel69]. Laboratories [Her62]. Laboratory [Sko61, Sto64b, Gul61a]. Labour [Hug63]. Labour-Saving [Hug63]. Labouring [Gai67]. LaFeber [Bra65a]. LaForce [Wol66]. Lag [Gil62, Hug62c]. Lagging [Pea65]. L’agricoltura [Lor67a]. Lakes [Tow63, Tow64, Tow65]. Lammot [Wil65c]. Lamont [Hew66]. Lamp [Eis66, Gab66, O’D69]. Lancashire [Har69]. Land [DeG65a, Joh69b, Gut64, Kra66, She65d, Zim65]. Landman [Nie62]. Landmarks [Hew64]. Landsberg [Tho65b]. Landscape [Ihd65]. Langley [Rod62a]. Lanifricariae [Moe66]. Lanifricarius [Moe66]. Lansdell [Tho61b]. Lansing [Hew66]. Lapidary [Bed65c]. Larouche [Mor69b]. Lass [Pur64]. Last [Bur65d]. Late [Bow66]. Later [PSD67]. Lathe [Bed65c, Bed67c, For62, Bat66c]. Latin [PS68]. Laurence [Goo68a, Hun63a]. laurentienne [Mor69b]. l’automobile [Bry68a]. Lavoisier [Sch63b]. Law [dC64b]. Lawrence [Bur63, K.61, Min65a, Deu66, Lea66]. Laws [Zvo62]. Lazarus [Dib65]. Lead [Phi63]. Leaders [Emm62b, Pet65]. Leadership [How69, Jaf63]. Lear [Cla66]. Learning [Zvo61b]. Leather [PR62]. Leavis [Le 64]. Leben [Kin68]. Lebergott [Jaf65a]. Lee [Süs68a, Bed64a]. Legacy [Goo65a, Kle62]. Legend [Whi68c, Woo60b]. Legislation [Pur68c]. Lehman [Hol65b]. Leicester [Ihd69]. Leiston [Bur66c]. Leisure [Kur63, Kur61b]. Leland [Swe68, Edw67, Edw68, Edw67]. Lens [Bed67c]. Lens-Grinding [Bed67c]. Lenzen [Bur69c]. Leo [Did62, For65a, Mar63, Sko61, Woo65].

Leon [Mic66, Mur69]. Leonard [Han61, She65c, Spi64b, Wie64]. Leonardo [Ste63, Ano63e, Ano64c, Ano65c, Ano66b, Ano67c, Ano68c, Ano69e, Ret65, Ret67b]. Leslie [Smi61a]. Less [Bar69a]. Lessons [De 68b, Dru66]. Lester [Pur69b]. Let [Bur69b]. Level [She61b]. Levelers [Gli68]. Levels [Gli68]. Levey [Lei63]. Levine [Goo69a]. L’Évolution [Mor69b]. Levy [Roc66]. Lewis [Dru68, Bur65d, Wei66a]. L’Histoire [Dau60b]. Liberal [Gul69]. Liberalism [Kam63]. Liberty [Hea61, Wil69b]. Library [DeG64, Per68, Ras64b, McH63, Wri68, Rei65b]. Liddell [PSD67]. Lido [Spe62a]. Life [Bea66, Cal68a, Fus67a, Hol65b, Kou61, Sha65, Smi69a, Wat65, Bri63, Uye64]. Lifeline [Ras66]. Light [O’D69, Raz64, Sch62a]. Lighting [O’D69]. Lightning [Sha65]. Lilienthal [Kle63b]. Lillian [Roc66]. Lilly [PS66]. Limited [Duv66]. Lincolnshire [Fus66a]. Lindt [Rin68]. Line [Dau60a]. Lineal [Gli69a]. Linguistic [Puh64]. Liniger [Smi63]. Liniger-Goumaz [Smi63]. Link [Rob66, Spi60]. Linn [Con63c]. L’inventeur [Gol66b]. Lionel [McN64]. Lipetz [Bar66a]. Liquid [Emm62c]. Listen [Pet65]. Lister [Bat62]. Literature [Gul69, Mar69b, dlC63]. Liu [hC66]. Lives [Smi67e, Rol67]. Living [Hun67, Ihd65, Hie63]. Livingston [Shr68]. Llewellyn [Fer64c]. Lloyd [Bed65a]. Location [Ros64a]. Lockemann [Far61]. Lockheeds [Sca65a]. Locks [Wul66a]. Loco [Whi68a]. Locomotive [Con66b, Whi68d]. Locomotives [Sto69b, Whi68b, Con63c]. Logan [Con63a]. London [Smi68b, Woo69b, Smi69b, Wri68]. Long [The62, Nie67b]. Long-Term [The62]. Longbow [Esp65]. Look [Tho65b, Wer67, Rob65]. Loos [Dau60a]. L’orange [Smi63]. Lore [Whi68c]. Loris [O’D69]. Louis [Dav64, McC67, Sca65b]. Love [Pet65]. Low [McN64]. Lowdon [Gut64]. Loyd [Dun68]. Lübke [dSP61]. Lucy [PS65]. Ludwig [Kel69c]. Luigi [Con66a, Mei66a]. Lumbering [Hid68]. Lunar [Hol65c, Tha69]. Lyall [Bur66d]. Lynch [Wil61a]. Lynn [Str63, BB66].

M [Bil68, Ble61, Buc68a, Bul62, Cal67, Cal69b, Col66, De 68d, Edw67, Esp67, Fer68b, Fus67a, Gol65, Hig65b, Hol61a, Hol65a, Ihd69, Lay67, Lew64, Lew69a, Mec66, Mer65, Nel66, Pil66, Ras68b, Rie68a, Rod62a, Rot60, San61a, Sca67, She65a, Shr68, Sny62, Sto63, Sto64a, Sto65b, Wer64b]. M.D. [Shr63]. Maboth [Cal68a]. macchine [Bun66a]. MacDonald [Mei66b, Ste68c]. Machinarum [Alt69a]. Machine [Alt69a, Bed68b, Dru68, Har60, Kel69a, Mar69b, Ogd66, Ush61, Woo61b, Woo66a, Woo65, Gil63, Boe62, Woo69b]. Machinery [Bur60b, Fer68b, Gil63, Hip69, Rog60, Wik60, Kel69b, Woo66b]. Machines [Hei67, Rie67, Ros64b, Fer66a, Har60, Mei68, Rei62b, Dib69a, Kol62]. MacKaye [Bry64b, May69a]. Mackel [Ede65]. Macon [Top66a]. Made [Aus68, Con63b, Joh62]. Madrid [Ret67b]. Mae [Rob66]. Magazine [Pur65a]. Magnificent [Hig66a]. Magnus [Bor65a, Kur61a, Le 63, Jol68]. Mail [Bur60a]. Mails [Kni65]. Main [Agn67]. Mainland [hC66]. Mair [PS65]. Mairi [Hou63]. maison [Mor69b]. Maizels [Ver64]. Maj [San60].

Make [Hei67]. Makers [Bed64b, Rob69, Bed64c, Gar65]. Making [Bur60a, Dib64, Smi60, Wis67, Gue62a, Wil65b]. Male [Moe69]. Malmesbury [Whi61]. Man [Bru68, Bur64a, Bur69b, Che61a, Cla62, Cle61, Dor68a, Hol65b, Hol61b, Kol62, Mum60, Mum65, Mum66, Roc66, Ste63, Wer66b, Fin62c, Sal65, Tho66, Ang67, Bor65b, Bur62b, Con63b, Lub65a]. Man-Made [Con63b]. Management [Fre69b, Gue61, Har65b, Jaf65b, Pen66, Boy67, Tei68, Chr61]. Managerial [Gru68, She65c]. Managers [Flo69a, Lew64]. Manchester [Fin63a]. Manflight [Gar67]. Mankind [Hir64, Mei66a, Pru62, Kan66, Pru62, Gre63]. Manley [Cha69a]. Mann [Wil64a]. Manned [Sto64a]. Manpower [Bir62, Did62, Eva65, Gru68, Rin65, Rin66, Wer66a, De 68b, Jaf65a, McC66b, McC67, Rin66]. Mansfield [Sny66]. Mantoux [Kra64]. Manufacture [Gra65]. Manufactures [Dob62]. Manufacturing [Arm62, Bar62, DeG69a, For60, Hid60, Out62, Sch62b, Str64, Joh62, Mor69a]. Manuscript [Svo67]. Maramec [McH65]. marchés [Smi63]. Marconi [Sha63b, Sha63b]. Marcson [Bir62]. Marcus [Ste62]. Mare [Las67]. Margaret [Jaf67, Spi64a]. Margery [Kel68]. Maria [Har61, May67b, Gli66]. Marian [Str67]. Marine [Mul69]. Mariner [Sto64b]. Marino [Lor67a]. Mario [Con64b]. Marion [Bed64c, Tan65]. Mark [Lok68, Meu64, Tra65]. Market [Jaf67]. Markets [Sch64a]. Marland [Dib66]. Marshall [Hal60b, Wei63, Wer66b]. Marston [Bur62a, Con68b]. Martin [Lei63, Smi60, Smi68b]. Martini [Ret63]. Marvin [Cal68d, Ham68, Son63]. Marx [Woo65]. Marxism [Bla63]. Marxist [Jor61]. Mary [Pea65, Rie67]. Mason [Hig68a, She61b, She65b]. Mass [Ber65b]. Massachusetts [Lea66]. Master [dC61b, Edw67]. Mata [Gli69a]. Material [Dow65, Spi69]. Materials [Goo68a, Rot69, Con62a]. Matériel [Hig66b]. Mathematical [Rei62b]. Mathematics [Lev61b, Lev61b]. Mathias [Wil60a]. Mathurin [SS61]. Matkel [Ede64, Ede65]. Mattfeld [Bak66]. Matthew [Hug60, Rei62a]. Matthias [Ros68]. Mature [Ble61]. Mauel [Bry69]. Maurice [Ben62a, For63a, May69c]. Maury [Rei62a]. Maus [Hin64]. Max [Car62c, Kur61b, Smi63]. May [Liv69, Sto64a]. Mayo [Spe62a]. Mazlish [Wer68a]. McC [Kra66]. McCallum [Wil67, Wil67]. McClelland [Dav62]. McClintock [Wei63, Wei63]. McCormick [McK65]. McDonald [Smi62, Wei64]. McGovern [Las66]. McGuire [Rin64]. McHale [Bur65b, Bur66a]. Mcluhan [Wer66b]. McNamara [Smi69c]. Mead [Arn65]. Meaning [Bri65, Sol68]. Means [Gar67, Gar67, Hea63]. Measure [Gil60a, Sol68]. Measures [Fou64]. Mechanical [Fer64a, For64b, Pur68b, Ras68a, Smi61c]. Mechanicals [Cal68c]. Mechanics [Alt69b, Hal60b]. Mechanism [Bow66, dSP64]. Mechanistic [dSP64]. Mechanization [Ros64b]. Mechanized [Hol61a]. Mecklenburg [Bau64]. Medal [Ano63e, Ano64c, Ano65c, Ano66b, Ano67c, Ano68c, Ano69e]. Mediaeval [Smi69a, Lei63]. Medical [Gai67, Car62a]. Medicine [Goo65b, Nee69, Rob66, Mil61]. Medieval [Boy60, Boy69, Cla66, Esp68, Gli69c, Goo61, Hal61b, She61a, She61b, She65b, She69, Gli68, Str63]. Medioevo [Len60]. Mediterranean [Gil62]. Meeting [Ano60a, Ano61d, Ano62j, Ano63h, Ano63g, Ano64f, Ano64e, Ano65d, Ano66g, Ano67k, Ano68g, Ano69b]. Meiji [Lon66]. Melsen [Mid63]. Melvin [Buc68b]. Members [Smi67e]. Memoirs [PSD67]. Men [And68b, Hig66a, Hig68a, Ros64b, Sch64a, Ste65, Cal66, Cal68b, Emm62b, Rie67, Wer64b]. Mendelschen [Whi68e]. Mental [Deu65]. Merchants [Fer69b]. Mercury [Sto64a, Sto64a, Dun68, Rob66]. Merdinger [Fin64a]. Merle [Ras66]. Merril [Kle64]. Merrimack [Lea67]. Merton [Sto63]. Mesopotamia [For65a]. Metal [Wil69a, Hou63]. Metal-fabricating [Wil69a]. Metallic [Wel66]. Metallography [Meh61]. Metallurgical [Zap62]. Metallurgy [Ano64g, For64a, Nee64, Pan60, Smi61b, Wer64a, Bur67, For64a]. Metals [Meh61, Smi61a, Smi65a]. Meter [Fra64]. Methodology [Ger66b]. Methods [Bur60a, Hid60, Smi60, Str64]. Métiers [For60, Cha68]. Metraux [And64]. Metropolitan [Con63a, Joh62, Uye68a]. Mexican [Ham68]. Mexico [DeG69a]. Meyer [Uye68b]. Meyers [McC66b]. Miasmas [Shr63]. Michael [Swa67b, Tof66, All62, Gol65]. Michal [Ble61]. Michel [Far63b]. Michele [Sev65]. Michelet [Gol66b]. Microscope [Mid68]. Microscopy [Bed68a]. Mid [Kur61a, McC63, Sca68b]. Mid-1930 [McC63]. Mid-Century [Kur61a]. Mid-nineteenth [Sca68b]. Middle [Bow66, Fus67a, Gha61, Hal60b, Whi60]. Middleton [Con63d, Hal65a, Sch62a]. Midland [Car68]. Miedzywojennym [Mos67]. Mies [Con66d]. Miessner [Kin66]. Mikesell [Son63]. Miles [Hig66c, Har69]. Militaire [Smi67a]. Military [DPS66, Esp66a, PS67, Smi68c, Smi69c]. Mill [Ush64, Hid62]. Miller [Goo65b]. Milling [Ush61]. Mills [Dib69a, Gha68b, Hid62, Hun67, Phi63, Rob66, Wul66b]. Milo [Cal67, Cal69b]. Milton [Hol65b]. Milwaukee [Ols69]. Minds [Rei64]. Minerals [Jol68]. minería [Gli69b]. Mining [Arm62, Fer69a, Ham68, Hou61, Smi68a, Smi64a]. Ministry [Hig65b]. Missiles [Hun64]. Missing [Spi60]. Mission [Sto64b]. Missouri [Pur64]. Mitchell [Hol61b]. Mobility [Wol65]. Model [Man65]. Models [Kil62]. Modern [Ben66, Ben62a, Ces69a, Con64a, Dru63, Emm61, Emm62a, Emm67, Fre69b, IU65, Rav65, Rie67, WW62, Ush60, De 67b]. Modernization [Koh67, Rod61, Spe69]. Mohammed [Rod67]. Mohan [Bro62]. Molecules [Shr63]. Molloy [Hig68a]. Monopoly [Rae61b]. Montague [DeG65a]. Monte [Pur68b]. Montgomery [Hec67]. Monthly [Pur65a]. Montreal [Ano65d]. Monuments [Wai67]. Moon [Swe68]. Moon-Doggle [Swe68]. Mooney [Smi67e]. Moore [Hid68]. Morandi [Con67b]. Moravians [Rin68]. Morgan [Hin66, McK64b, Con63e, Gue60]. Morison [Rie67]. Morris [Far68b, Gha68b, Gha68b]. Morrow [Bry66b]. Morse [Smi69d, DeG65b]. Mortimer [Gha67, Gha69b, Goo69a]. Morton [San64]. Moseley [Cal68a]. Motion [Hug62a]. Motor [Gor68, Sco66]. Motorenerfinder [Bry68b]. Motorization [Mos67]. motors [Bry67a]. Motoryzacyjna [Mos67]. Movement [Hew66]. Mr. [Smi60]. Mrs [Edw67]. Mule [Lei67]. Mullen [Boy67]. Multhauf [Far68a, Kil62]. Mumford [Dru68, Wer68a]. Munden [Sch63a]. mundi [Hug67a]. Mundo [Gli69b]. Munson [Sca68b]. Murdock [Les60]. Murphy [Svo67, Wat68]. Murray [Rei62b]. Musaeorum [Jac69]. Musée [Cha68]. Museo [Bon69]. Museum [Bed64c, Bou66, Fer68b, Fin65b, Had68, Lew69a, Mul65a, Süs65, Wel65, Woo60a, Woo69a, Woo69b, Lea67, Fer68b, Hec67]. Museums [Bed65b, Bro65a, Fer65d, Hip69, Per69, Gra62]. Music [Gul69]. Musson [Duv66]. Must [Bor65b]. Muzea [Süs65]. My [Sha63b]. Myers [Eva65, Tei68]. Myron [Col65, Spe69]. mystics [Sin60b]. Myth [Dru68, WW66, Hug62a, Le 63].

N [All62, And62b, Bra69, Dea69, Deu66, Esp66b, Had69, Hig68b, Joh65a, Lew69a, Rei67, Tok68]. Nacional [Ret67b]. Naeve [Cal67, Cal69b]. Nagler [May68]. Nail [Smi60]. Nail-Making [Smi60]. Name [Put67]. Names [Gha68a]. naquit [Bry68a]. Narrow [Whi68b]. Narrow-Gauge [Whi68b]. Nartov [Ogd66, Alt69a]. NASA [Gru67]. Nash [Tra65]. Nasser [San61b]. Nathan [Men66]. Nation [Kol66]. National [Cha68, hC66, Emm63a, Fer68b, Hof65, Kle65, Loc69, Ric68, Rin65, Rin66, Rod62b, San61a, Smi66b, Sto64a, Sto65a, Pur68c]. Nationalized [Wis61]. nationalsozialistischen [May67a]. Nations [PS65]. Natural [And65a, And65b, Bla63, DeG65b, Tan65, Tho65b]. Naturalis [Fer64b]. Nature [Con69c, Lea68, Mum66, Spi64a, Tav69, Wat68]. Naturvölker [May67b]. Naval [Ful63, McC63, Pea66b]. Navigating [Bed64c]. Navigation [Vir68]. Navigations [McC65]. Navigator [Bra65b]. Navy [Top66a]. Near [Con69b]. Necessities [Dib69a]. Nedoluha [She63]. Need [Dup69, Wis66]. Needham [Goo61, Cha61, tZS64]. Needs [Bur65b, Mic64, Wer66a]. Nef [Dow65, Mor63a]. Neglected [Bur61c, Mum61]. Nei [Nee69]. Neil [Bar61a]. Nel [Dow64]. Nels [Hun64]. Nelson [Cha69c, Kea68]. Nervi [Con66a]. Nest [Las67]. Netschert [Tho61a]. Neuburg [Had64b]. Neuen [For61a]. Neueren [Ret61]. Neville [Hat66]. Nevins [Wik64]. Newcastle [Bur66d]. Newcomen [Fer64d, Har62a, Jew64, Nix64]. Next [Gol65]. Nieburg [Joh65b, Put67]. Nightmare [Lok68]. Nikola [Sha65]. Nineteenth [Aus65, Bar67b, Car67, Hid60, Hug63, Koh64, Mah61, Men66, Mul68, Ste68c, Sca68b, Pea66b, Had61]. Nineteenth-Century [Bar67b, Car67, Pea66b]. Nitske [Bry66b]. No [Mul69, Whi63]. Noah [Big64]. Nobel [Duv64]. Noble [Ush64, Bul66]. Noel [Fai62]. Nonstop [Sca65b]. Nonunion [McH61]. Norbert [Bro65b]. Nordatlantik [San68]. Nordatlantik-Passagierdampfer [San68]. Norman [Bar66b, Con62a, Sal64, Tho61b]. Normand [Pay64]. Norris [McH65, Whi68d]. North [Nie67c, Pru65, Hip69, McC65]. Northern [Car68, McC69]. Norton [Bak67]. Note [Ano61f, Ano63c, Ano66f, Dor68b, She69]. Notes [Ano60c, Ano61e, Ano61g, Ano62d, Ano62e, Ano63f, Ano64d, Ano65a, Ano65e, Ano65f, Ano66c, Ano66d, Ano66e, Ano67d, Ano67e, Ano67f, Ano68d, Ano68e, Ano68f, Ano69f, Ano69g, Bed64c, Dru61b, Fer60, Wul65, Wul66b]. Nothing [Hig65b]. Novels [Bri62a]. Nuclear [Gru68, Hew64, Hig68b, Nie62, Joh65b]. Nuevo [Gli69b]. Number [Ano60f, Ano60d, Ano60e, Ano61k, Ano61i, Ano61j, Ano61h, Ano62i, Ano62g, Ano62h, Ano62f, Ano63l, Ano63j, Ano63k, Ano63i, Ano64k, Ano64i, Ano64j, Ano64h, Ano65j, Ano65h, Ano65i, Ano65g, Ano66k, Ano66i, Ano66j, Ano66h, Ano67h, Ano67g, Ano67i, Ano67j, Ano68i, Ano68h, Ano68j, Ano68k, Ano69j, Ano69i, Ano69k, Ano69l, Tas66]. Nürnberg [Whi68e].

O [Gha61, Gra62, Hig64, Hug61c, Jen67, Swe67]. OAH [Hig68c, Lew67, Rae68]. Oakeshott [Gra61]. Objects [Sam65]. Occident [Smi67a]. Occupation [Rav65]. Ocean [Dun68]. Octave [Fer65a]. October [Ano67j, Ano68k, Ano69l, Emm69]. Odishaw [Emm64]. OECD [Loc69]. Oersted [Hug62b]. Office [Gil66, Rei60, Chr62]. Offices [Con61a]. Officinae [Moe66]. Ogorkiewicz [Hol61a]. Oil [Dau60a, Dau62, Dau64, And68a, Dau60a]. Okresie [Mos67]. Old [Bed65a, Hid68, Lan68]. Oleg [Con66c]. Oliver [Tay67]. One [Fer64c, Kou65]. O’Neil [Jor62]. Operational [Hig65b]. Oppenheim [For65a]. Optical [Bed67c, Sch62a]. Oral [Nun63]. Orange [Cal66]. oranger [Smi63]. Orbital [Sto64a]. Order [EW62]. Ordnance [Spe62a]. Ordway [Emm68b]. Organic [Far63b]. Organization [Ble61, Chr61, Hug67b, Lew66, Par66, Sko67, Hig65b, Gue67, Dea69, Jaf63, Jaf65b, Ros63]. Organizational [Ano60c, Ano61f, Ano61g, Ano62e, Ano65f, Bir62]. Organizations [Lew64, Lew65, Tei67]. Organizers [Gue62b]. Organizing [Sko61]. Orient [Smi67a]. Origin [Bac69, Bow66, Bry67b, Fra63, HS69, Hig64, Pru62, Sch63b, Sha64]. Original [Gil63]. Origins [Far68a, Fox69, Lon66, MW63, dSP64, Sto65a, Whi68d, Wie64]. Orlans [May69b, Nie67a]. Orleans [Did62]. Orth [Lew64, Lew65]. Osborne [PS66]. Oscar [Kle63a, McK64a]. Ostia [Pac68]. Other [Dor68d]. Otto [Con68c, Fry61, Bry66c, Bry67a, Kle63b]. Otto-motors [Bry67a]. Ottoson [Zim65]. Our [Bri68]. Outer [Dry61]. Outline [Con68a, Kle67, Rae62]. Outlook [Mil62]. Output [Sha67b]. Outstanding [Ogd66]. Ouverture [SS61]. Overcast [Las66]. Overlapping [Meu66]. Overseas [Nix65]. Owen [Wol65]. Ownership [Wis61]. Oxidation [Jen69]. Oxygen [Jen69].

P [Buc68a, Bur65e, Chi65, Chr61, Con63e, Far68a, Fer65b, Fle65, Fus67b, Gue60, Kil62, Kol62, Le 64, Mer65, Rob65, Ros64a, Sko61, Wei66b]. Pacific [Car61b]. Paddle [Cha64]. Paddle-Wheel [Cha64]. Pains [Hud65b]. Painted [Bul66]. Pakistan [Gha69a, Mor67, Rod67]. Palace [dC61a]. Palenque [Out62]. Pamor [Fra63, Wul65]. Pannell [Buc68a, Mer65]. Paolo [Bun66a]. Paper [Jar66, Wis66]. Papers [Bra66, Sha67a, Spi64a, Woo60a]. Pappenfort [Mec66]. Parachute [Whi68f]. Parchment [PR62]. Pareti [Mei66a]. Park [Spi64a]. Parke [Rol67]. Parkhill [Smi67d]. Parmelee [Ben62a]. Parris [Car67]. Parrying [Lew69b]. Parson [Hun64]. Part [Fer62a, Jen69, tZS64, Fer62b, Fer62c, Fer63b, Fer64e, Fer64f, Fer65c, Hig61a, Hig61b]. Participation [Gue62a]. Parties [Wie64]. Partington [Far63a, Far64a, Far65, Lei61a]. Parts [Col66, Woo60b]. Passagierdampfer [San68]. Past [Hun67]. Pastoral [Woo65]. Patent [Fed60, Gil66, Kun60, Rei60, Zuc68, dC64b, Rae61b]. Patenting [Gil60a]. Patents [Bed61b]. Paternalism [Lon65]. Pathfinder [Süs68a]. Patricia [Rei62a, Smi66a]. Patrick [Wil61a]. Patterns [Che61a]. Paul [Bea66, Boe62, Bri63, Bro62, Car62a, Cha69a, DeG69a, Deu66, Gul61b, Had69, Hid60, Jaf61, Kra64, Lew69a, Pay64, Ras61, Rei63, Smi64a]. Pauline [Gra60]. Payne [Mey60]. Payson [Ano69a, Smi61c, Smi65c]. Peabody [Bed64c]. Peaceful [All62, Nie65]. Peacemaker [Pea66b]. Pearl [Fer65a]. Peasant [Jac64, Min65b]. Peck [Sto63]. Peirce [Bur69c]. Pendule [May69c]. Pendulum [Pur69b]. Penick [Hew67, Tag69]. Penn [Tra64]. Pennsylvania [Hal65b, Hou61]. Penny [PS64]. Penrose [Bed61a]. People [Con65a, Mic64, Pea65, Kol62, Jaf64]. Peoples [Les60, Spi69]. Percival [Gra60]. Percy [Sha63a]. Perfect [Tof67]. Performance [Swe67]. Peril [Hew66]. Period [Bri62a, Gar67, Hec67]. Periods [Goi69]. Perlman [Meu64]. Perrot [Lew69a, Lew69b, Had69]. Persecution [Col61]. Persia [She68, Wul66a]. Personality [Boy67]. Personnel [Hug67b]. Perspective [Bed68b, Dup69, KH61, Dav68]. Perspectives [Cla66]. Peruvian [Coh69]. Pesticides [Ihd65]. Peter [Aus65, Fin68, Les60, Pur65b, Ros65b, San68, Wil60a, Con62b]. Peterson [Ros61]. Petroleum [Lev61a, Wie64, Bur63]. Pharaoh [dC63a]. Pharmacology [Lei63]. Phase [Bur65b, Bur66a]. Phases [Esp67]. Philadelphia [Gla62, Goi69, Ano63g, Ano64e]. Philip [Fin65a, Tho66, Wil69b]. Philippines [Rod62b]. Phillip [Son63]. Philosopher [Sin60b, And65a, And65b]. Philosophia [Fer64b]. Philosophical [Kil62, Svo67]. Philosophies [Aga66]. Philosophy [Bry64b, Gul69, Nel66, dSP64, Con60a, Rie68a]. Photographic [New66a]. Phyllis [Bar63b]. Physical [tZS64]. Physics [tZS64, tZS64]. Pictorial [For60, Tow65]. Picture [Hug62a]. Piedmont [Lor67b]. P’ien [Nee69]. Pier [Con66a]. Pierce [Sto69a]. Pierre [Sch62a]. Piggott [Mei62b]. Pike [Ste67, Ste68b]. Pilot [Bra65b]. Pilots [Hig68a, Sca65a]. Pinchot [Tag69]. Pincus [Tow64]. Pinkepank [Whi68a]. Pioneer [Bry66b, Far63b, Hig68a, Kel69b, Pen63, Süs68a, Smi67e]. Pioneering [Hew64]. Pioneers [Rae61a, Wyn60]. Pipe [Dau60a]. Pirotechnia [Cah67]. Piston [Fox69, Rae62]. Pit [Eic64]. Pittsburgh [Lub65b]. Pius [Mar63]. Pivoted [Boy60, Hal61b]. Place [Con68d, Har62b]. Plagiarists [Ret63]. Plains [Kra66]. Plane [Rob69, Wel66]. Planning [Aga67, Ble61, Con64a, Jac60, Bry64b, Con63a, Ham69]. Plant [Ros64a]. Plants [Ros66]. Plate [Pur62]. Plates [For60]. Platinum [Smi62]. Platt [Rol64]. Plenty [Bur62a]. Plough [Puh64]. Plow [Wik61]. Plowden [Whi67]. Plowshare [San63, Nie64]. Plumb [She61b]. Plymouth [Lan68].

Podhalu [Mos67]. Poet [Boe62]. Point [Con64a]. Pointed [She69]. Poirier [For61b]. Pokrovsky [San60]. Poland [Mos67, Ros65a, Pia61]. Polaris [Mil63]. Policies [Bar69a, Ble63, McC67, Rin65, Rin66, Sko67]. Policy [Loc69, May69b, Rod62b, Zim64, Zim65, Emm63a, Joh65b, Kea68, Kre67, Nie65, Per68, Smi66b, Top65a, Wil69a]. Polish [Gue62a, Mos67]. Polishing [Bat66c]. Political [Smi68c, Tra65, Lon67]. Politics [Dau62, Ham68, Hew67, Lon69, Smi68c, Tag69, Uye68a, Owe65, San61a, San64]. Pollack [Sal64]. Polsce [Mos67]. Polvani [Kel69a]. Polytechnic [Cal69a, Smi67e]. Pompeii [Moe66, Moe68, Moe69]. Poni [Dow64]. Pont [Hec67, Wil65c]. Popov [Kin66]. Population [Col66, Fus66c, Gai67, Sev65]. Populist [Sal64]. Portfolio [Cal67, Cal69b, Kou65]. Portrait [For65a]. Portsmouth [Woo66b]. Possession [Kil62]. Possibilities [Mor61]. Possible [Out62]. Post [Ros64a]. Post-War [Ros64a]. Postal [Kni65]. Postan [Hig65b, Col66, Fus67a]. Postscript [Ret65, Wul66b]. Postwar [Mil64]. Potentials [Pea65]. Pottery [Bul66]. pour [Whi63]. Powder [Wil65b]. Powell [Fer64c]. Power [Bur66b, Car65, Dau62, Fer68b, San61a, Swe65, Bry66b]. Pp [And65b]. Práce [Süs65]. Practical [Bra65b]. Practice [Smi61b, Wil69a, Esp66a]. Practices [Wel63]. Prairie [Koh64]. Prairies [Koh64]. Precision [Edw67, Woo61b]. Predictable [Dru69]. Preface [Pur68c]. Prefatory [Ano66f]. Prehistory [For64a, Gra61, Nix64, Spi64b]. Preliminary [Ger66b, Goo67a, Pia61]. Premature [Car61b]. Preparation [PR62]. Present [Bur64b, Rav65, Bur65a, Hew67, Rot69, Smi64c]. Present-Day [Rav65]. Presidents [Smi67b]. Press [Fer65b]. Pressed [Had64a]. Pressure [Fry61, Fus66c]. prestige [Cha68]. Price [And62a]. Pride [Swe65]. Prieto [Gli69b]. Princeton [Pea66b]. Principal [Hal68]. Printed [Ede64]. Printing [Goo67a, Wyn60, Wyn60]. Priorities [Nee64]. Pritchard [Boy69, Hun63a]. Privacy [Nie67b, Jac68]. Private [Dau62, Goi69]. Prize [Ano69c, Ano69a]. Prizewinners [Duv64]. Probable [Gha68a]. Problem [Gar67, Mor67, Uye68b]. problème [Smi63]. Problems [DeG65a, Jar66, Kle65, Lew64, Mor69a, Pea65, The62, Zim65, dM67, Sut66a]. Procede [Alt65]. Procedures [Pea66b]. Process [Bur63, Hol65a, Sto63, Rog69]. Processes [And68a]. Procurement [Hig66b, Smi68c, Spe62a]. Production [Tho66, Gab66, Hoo66a]. Productivity [Boy67, Ham69, Sha67b, Jaf64, Jaf68]. Profession [Bak65, Gai67, Cal68c]. Professional [Pur68b, Did62]. Professionals [Flo69a]. Professions [Bak66]. Professor [Aga67]. Profile [tZS65]. Profiles [Emm62b, Smi64b]. Profit [Jaf65b]. Profits [Bur63]. Program [Ano63h, Ano64f, Hei68, Hig68c, Lew67, Mes65, Mes66, Mes69b, Rae68, Wer67, Bur66a, Swe65]. Programme [Man65]. Progress [Bur63, Cal68c, Cle61, Kel69c, Lei61b, Ros65a, Sol68, Tra65, Zvo61a, Mor61, Ces69a]. Progressive [Har65b, Tag69]. Progressives [Wie64]. Project [Dun68, Rob66, Sto64a, San63, Sto64a, Nie64]. Projections [Ger66b]. Promise [Gul61b]. Promotion [Mil64]. Propagation [Bou66]. Property [Rei60]. Prophets [Lay68]. Proposal [dC64b]. Proposals [Hal63]. Proposed [All62]. Propulsion [Sto64b, Sut66a]. propulsive [Pay64]. Prospects [The62, Tho61a]. Protestantism [Dav66]. Provence [De 67a]. Provincial [Hol65c]. Przemysl [Mos67]. Psychology [Maw65]. Psychophysiological [Car61a]. Public [Emm63a, Gai67, Kea68, Kle65, Ras67, Rod62b, Wis61, Rin65]. Publications [Pur68d]. Puerto [DeG69a]. Puget [Mul69]. Pulse [Tho60]. Pump [Sha64, dC63a]. Punch [Mur69]. Purcell [Che61a]. Pure [Fei61]. Purely [Fin62b]. Purpose [Lea68]. Pursell [Buc68b]. Purver [Kel68]. Pushkarev [Con63b]. Putney [Sch63a]. Puzzle [Jew69]. Pye [Con69c]. Pyke [Bor65a, Kur61a, Le 63]. Pyramids [dC63b].

Qantas [Hig66d, Hig69b]. Quadrant [Bed69]. Qualey [Lay69]. Quality [Ces67]. Quarterly [Kyr64]. Quellendarstellungen [New66a]. Quester [Emm67].

R [Alt65, And68b, Bed61a, Ben68, Bil68, Bur65b, Bur66a, Bur66c, Cal66, Cal68b, Che61b, De 64, Deu66, Dib69b, Dor68d, Far63a, Far64a, Far65, Fer64b, For64a, Fus67b, Gli69a, Gra61, Had63, Hal68, Har67, Hat62, Hig65a, Hig68b, Hol69, Hou63, Hud65a, Jaf68, Kea68, Lam61, Le 64, Lei61a, Lev61a, Lok68, Mul62, Nea64, Nee69, Pur68a, She65c, Smi64b, Smi65a, Ste68a, Sto69a, Tav69, Uye68b, Wer65, Woo66b, Woo69b]. Race [Swe68]. Races [Pru62]. Radar [Tho60]. Rader [Rol64]. Radiation [Chi63]. Radio [Han67, Kin66]. Radioactivity [Kle64]. Radiotelegraphy [Kin66]. Rail [Jew69]. Rail-Beam [Jew69]. Railroad [Bri62a, Car61b, Gla62, Joh69a, Whi68c]. Railroads [Pat62, Rub67]. Railway [Arm67, Car68, Con63e]. Railways [Car62b, Car67, Fer68a, Wil61c]. Raisonné [For60]. Ralph [Fin63b, Lew64, Nie64, Sin66]. Ransom [Con65a]. Rasmussen [Bri61]. Raspail [Cha69b]. Rate [De 67b]. Ratigan [Kou61]. Rationale [Swe65]. Raymond [Bat62, Ces69a, Gha69a, Goo65a, Le 65, Min65b, Spi60, Woo68]. Rayon [Ros66, Ihd64]. Reactor [Hew64]. Readings [Bee66, Bri61, Pen66, Son63]. Ready [Smi69c]. Real [Bar62]. Realities [Nie69]. Reality [Fer61, Hea63, WW66]. Realm [Hig68b]. Reasonable [Mor63b]. Rebirth [Wik64]. Rebuilding [Boy66]. Rebuilds [Swa67a]. Recessing [Gil63]. Recognition [Nel66]. Record [Jaf65a]. Records [Bed61a, Fed60, Rei60]. Red [Wil64a]. Redevelopment [Spo63]. Redfield [Spi64a, Spi64a]. Redivivus [Whi64]. Redstone [vB63]. Reed [Fer67, Mar69b]. Rees [Bar62, Fer68c]. Reference [DeG65a]. Reflections [Hal60a]. Reform [Dav64]. Reformer [Cha69b]. Refrigeration [And68b]. Refutation [Ano69h]. Region [Uye68a, Lub65b]. Regional [Bry64b]. Rehabilitation [Bon69]. Reichswerke [Ros68]. Reid [Con63e]. Reinforced [Con68d]. Reinforced-Concrete [Con68d]. Reingold [Men66]. Reithmann [Bry68b]. Rejoinder [Ano69h, Dor68c]. Relating [Sch63a].

Relation [Lei63]. Relations [Bee65, Gue68, Mon67]. Relationship [Dru61b]. Relationships [Con65b]. Religion [Nee69]. Religious [And62b]. Reminiscences [Sch66]. Removing [Ede64]. Renaissance [Smi69a, Bed65c, dlC63, Kel65]. René [For61b, Ang67]. Rennie [Wat65]. Rensselaer [Cal69a]. Replacement [Esp65]. Replies [Far69]. Reply [Aga67, Bur60a, Gil60b, Smi60, Wis67]. Report [Goo67a, Pia61, Pre62a, Pre62b, Pre62c, Pre62d, Pre62e, Vog67, Gai67]. Repubblica [Lor67a]. Republic [Jac64]. Research [Bur68, Dea69, Emm62b, Emm62c, Fly68, Fro66, Gab66, Han61, Her62, Hig65b, Hug67b, Jac64, Lew64, Lew65, Lew66, Mes69b, Nie67c, Pea66b, Pur68c, Rei65b, Sch64b, Sha67b, Sko67]. Resistance [PSD64, Str67]. Resisted [Wel63]. Resource [DeG65b, Eva65, Pen66]. Resources [Bur65b, Meu64, Sko67, Tan65, Tho65b, Wer66a, Rin65, Pea65]. Response [And68a, Cal68d, Deu66, Gla62, Jew68, Pie63, Sal64, Spe64, Str67]. Responses [And62b]. Ressentiment [Car62c]. Results [Sto64a]. Ret [Wil68]. Reti [Wul66b]. Retraining [Jaf67]. Rev [Mar63]. Reverence [Haw66]. Revi [Had61, Had64a]. Review [Ben68, Fin61, Wil61b, Agn67, All62, Alt65, Alt69a, And62b, And62a, And64, And65a, And65b, And68b, And68a, Ang67, Arm62, Arm67, Arn65, Arn69, Aus63b, Aus63a, Aus65, Aus68, Ayr67, Bak65, Bak66, Bak67, Bar65a, Bar67a, Bar69b, Bar69a, Bar61b, Bar61a, Bar66a, Bar66b, Bar67b, Bar67c, Bar62, Bar63b, Bar63c, Bar64, Bar65b, Bar66c, Bar66d, Bat62, Bat66b, Bat66a, Bat67, Bea66, Bed61a, Bed61b, Bed64a, Bed64b, Bed64c, Bed65a, Bed67b, Bed68a, Bee66, Ben66, Ben62a, Ben63, Ber65a, Ber65b, Ber67, Big64, Bil68, Bir62, Bir66, Bla63, Bla64, Bla62, Ble61, Ble63, Boe62, Boe63, Bor65a, Bor65b, Bou61, Bou67, Boy67, Boy69, Bra69, BT69, Bra66, Bra65a, Bra65b, Bri65, Bri68, Bri61, Bri62b, Bri63, Bro62, Bro65b, Bru68, Bry64a, Bry64b, Bry66a, Bry66b]. review [Bry67a, Bry68a, Bry68b, Bry69, Buc68a, Buc68b, Bul62, Bul66, Bun66a, Bur62a, Bur64a, Bur64b, Bur65a, Bur61a, Bur62b, Bur65b, Bur66a, Bur63, Bur67, Bur68, Bur69a, Bur65c, Bur61d, Bur65d, Bur65e, Bur66c, Bur66d, Bur69c, BB62, Cah67, Cal66, Cal67, Cal68a, Cal68b, Cal68c, Cal68d, Cal69a, Cal69b, Car69, Car61a, Car61b, Car62a, Car62b, Car67, Car68, Car60, Car61c, Car62c, Ces65, Ces67, Ces69b, Ces69a, Cha66, Cha61, Cha64, Cha69a, Cha68, Cha69b, Cha69c, Che61a, Che61b, Chi63, Chi65, hC66, Chr61, Chr62, Cla66, Cle61, Coh69, Col61, Col65, Col66, Con60b, Con60a, Con61b, Con61a, Con62c, Con62a, Con62b, Con63a, Con63d, Con63b, Con63c, Con63e, Con64a, Con64b, Con64c, Con65a, Con66a, Con66b, Con66c, Con66d, Con67b, Con67a, Con68b, Con68c, Con68a, Con69b]. review [Con69a, Con69c, Dau60a, Dau62, Dau64, Dav68, Dav62, Dav64, De 68a, De 64, DeG65a, DeG65b, De 67a, De 67b, De 68b, De 68c, DeG69a, DeG69b, De 68d, Dea69, Deu65, Deu66, Dib69a, Dib64, Dib66, Dib69b, Dib69c, Did62, Dob62, Dor68d, Dow64, Dow65, Dru68, Dun69, Dun63, Dun68, Duv64, Duv66, Edw67, Emm61, Emm62a, Emm62b, Emm62c, Emm63a, Emm64, Emm67, Emm68b, Erm64, Esp66a, Esp66b, Esp67, Esp68, Eva65, Fai62, Fal65, Far61, Far63a, Far63b, Far64a, Far64b, Far65, Far68a, Far68b, Fer61, Fer63a, Fer64b, Fer64c, Fer64d, Fer64a, Fer65b, Fer65a, Fer66b, Fer66a, Fer67, Fer68a, Fer68b, Fer69a, Fer69b, Fin62a, Fin63b, Fin63a, Fin64a, Fin65a, Fin62c, Fin64b, Fin68, Fle65, Flo69a, Fly68, For60, For61b, For61a, For61c, For62, For63a, For63b, For64b, For64a]. review [For65a, For65b, For66, Fou64, Fre69a, Fro66, Fry61, Fus66c, Fus66b, Fus66a, Fus67b, Fus67a, Fus68, Gab66, Gai67, Gar67, Gar65, Ger66b, Gha61, Gha65, Gha67, Gha68b, Gha68c, Gha68a, Gha69a, Gha69b, Gil64, Gil66, Gla62, Gli66, Gli69a, Gli69b, Goi69, Gol66a, Gol66b, Gol65, Goo61, Goo65a, Goo65b, Goo66a, Goo68b, Goo68a, Goo69a, Gra60, Gra62, Gra61, Gra65, Gre65, Gre63, Gru67, Gru68, Gue60, Gue61, Gue62b, Gue62a, Gue65, Gue67, Gue68, Gul61a, Gul61b, Gul64, Gul69, Gut64, Had61, Had63, Had64b, Had64a, Had67, Had68, Had69, Hal60b, Hal65a, Hal66, Hal68, Hal65b, Ham68, Ham69, Han61, Han67, Har69, Har60, Har61, Har62a, Har63, Har64, Har65a, Har66, Har67, Har65b, Hat62, Haw66, Hea61, Hea63, Hec67, Her62, Hew66]. review [Hew67, Hid60, Hid62, Hid68, Hie63, Hig63, Hig64, Hig65c, Hig65a, Hig65b, Hig66a, Hig66b, Hig66d, Hig66e, Hig66c, Hig68a, Hig68b, Hig69a, Hig69b, Hin62, Hin66, Hin64, Hip67, Hof64, Hof65, Hol61a, Hol65b, Hol65a, Hol69, Hol61b, Hol65c, Hoo62, Hoo64, Hoo65, Hoo66b, Hoo66a, Hou61, Hou63, Hub64, Hud65a, Hug60, Hug61b, Hug61c, Hug62a, Hug62b, Hug63, Hug67a, Hug67b, Hug68, Hug69, Hun62, Hun63a, Hun64, Hut63, Hut66, Ihd64, Ihd65, Ihd67, Ihd69, IU65, Jac68, Jac69, Jaf61, Jaf63, Jaf64, Jaf65a, Jaf65b, Jaf67, Jaf68, Jen67, Jen69, Jew64, Joh62, Joh65a, Joh69a, Joh65b, Joh69b, Jol68, Jor62, Jun68, K.61, Kam63, Kan65, Kan66, Kat63, Kea68, Kel65, Kel68, Kel69b, Kel69c, Kel69a, Kil62, Kil64, Kin66, Kin68, Kle62, Kle63a, Kle64, Kle67, Kle63b]. review [Kni65, Kni67, Koh64, Koh66, Koh67, Koh68, Koh69, Kol62, Kol64, Kol66, Kou61, Kou65, Kou66, Kra66, Kra64, Kre67, Kur61b, Kur61a, Kur63, Kyr64, Lam61, Lan68, Las66, Las67, Lay66, Lay67, Lay68, Lay69, Le 63, Le 64, Le 65, Lei61a, Lei63, Les60, Lev61a, Lev61b, Lew64, Lew65, Lew66, Lew69a, Loc69, Lok68, Lon65, Lon66, Lon67, Lon69, Lor67a, Lub65a, Lub65b, Lyn67, Mah61, Man65, Mar69a, Mar63, Mar69b, Maw65, May69a, May69b, May67a, May67b, May68, May69c, May69d, McC66a, McC66b, McC67, McG65, McH66a, McH61, McH65, McH66b, McK64a, McK64b, McK65, McN64, Mec66, Meh61, Mei62b, Mei66b, Mei66a, Mei68, Men66, Mer65, Meu64, Mey60, Mic64, Mic66, Mid63, Mid68, Mil61, Mil62, Mil64, Min65a, Min65b, Mon67, Mor69a]. review [Mor63a, Mor63b, Mor69b, Mor61, Mos67, Mul68, Mul69, Mul62, Mur69, Nea64, Nee69, Nel66, Net65, New66a, New66b, Nie62, Nie64, Nie65, Nie67a, Nie67b, Nie67c, Nie69, Nil61, Nix64, Nix65, O’D69, Ogd66, Oli60, Ols69, Out65, Owe65, Par66, Pat62, Pau65, Pau67b, Pau67a, Pay63, Pay64, Pea66a, Pea65, Pen66, Per68, Pet65, Phi63, PS64, PS65, PS66, PSD67, PS67, PS68, Pie64, Pie66, Pil66, dSP61, Pru62, Pru65, Pur64, Pur65a, Pur65b, Pur68b, Pur68a, Put67, Rae61a, Rae61b, Rae62, Rae66, Rae69, Ran63, Ras61, Ras64a, Ras66, Ras67, Ras68b, Rav66, Raz64, Rei65a, Rei62a, Rei65b, Rei67, Rei62b, Rei63, Rei64, Ret61, Ric68, Rie65, Rie67, Rie68b, Rie68a, Rin64, Rin65, Rin66, Rin68, Rob69, Rob65, Rob66, Roc66, Rod61, Rod62a, Rod62b]. review [Rod65, Rod67, Rog69, Rol64, Rol65, Rol67, Rol68, Ros65b, Ros65a, Ros67, Ros68, Ros69, Ros63, Ros64a, Ros66, Ros61, Rot60, Rot65, Rot69, Rub67, Rya67, Sal64, Sal65, San60, San61a, San61b, San64, San68, Sca61, Sca65b, Sca65a, Sca66, Sca67, Sca68a, Sca68b, Sch64a, Sch62a, Sch63a, Sch63b, Sch66, Sco66, Sco61, Sel69, Sev65, Sha67a, Sha67b, Sha63a, Sha63b, Sha65, She61a, She63, She65a, She65c, She68, She65d, Shr68, Shr63, Sil69, Sin66, Sin69b, Sin60b, Sko61, Sko67, Smi64a, Smi61c, Smi61a, Smi62, Smi63, Smi64d, Smi64b, Smi64c, Smi65a, Smi66b, Smi66a, Smi67d, Smi67a, Smi67e, Smi67b, Smi68b, Smi68a, Smi68c, Smi69a, Smi69c, Smi69d, Sny62, Sny66, Son63, Spe62a, Spe69, Spi60, Spi64a, Spi64b, Spi68a, Spi68b, Spi69, Spo63, Ste67]. review [Ste68c, Ste68a, Ste68b, Ste63, Ste65, Ste62, Sto63, Sto64a, Sto64b, Sto65a, Sto65b, Sto68, Sto69a, Sto69b, Str67, Str63, tZS64, Süs65, Süs68a, Süs68b, Sut66a, Sut66b, Svo67, Swa67a, Swa67b, Swe65, Swe67, Swe68, Tag69, Tan65, Tav69, Tay67, Tei67, Tei68, Tha69, Tho65a, Tho60, Tho61b, Tho61a, Tho65b, Tho66, Tof66, Tof67, Tok68, Top65a, Top65b, Top66a, Tow63, Tow64, Tow65, Tra65, Ush61, Ush64, Uye64, Uye68a, Uye68b, Vat68, Ver64, Viv68, Vog69, War68, Wat61b, Wat63, Wat65, Wat68, Wei62, Wei63, Wei64, Wei66a, Wei66b, Wel65, Wel67, Wer64b, Wer65, Whi63, Whi67, Whi68a, Whi68e, Whi68b, Whi68c, Wie64, Wik61, Wik64, Wil69a, Wil65a, Wil60a, Wil60b, Wil61a, Wil64a, Wil61c, Wil64b, Wil67, Wil68, Wil69b, Wis63, Wis61, Wol65]. review [Wol67, Wol66, Woo60a, Woo61a, Woo61b, Woo66b, Woo66a, Woo69a, Woo69b, Woo65, Woo68, Wyn60, Yab67, Yam68, Zap62, Zim64, Zim65, Zuc68, d’A68, dC61a, dC63b, dC63a, dM67, vW67]. Reviewer [Far69]. Reviews [Lea68, Wil61b, Loc69]. Revolution [Bar61a, Dru61b, Dru66, Etz66, Gul61b, Hal61a, Hea63, Hun64, Kra64, Ros63, San64, Gul64, Han61, Rei65a, Ste67, Tho61b, Sca65a]. Revolutions [Col66]. Revue [Spi68a]. Reychman [Mos67]. Reynolds [Spi69, Wat63, Wil68]. Rezneck [Cal69a]. Rhodesia [DeG65a]. Riccardo [Con67b]. Richard [Dib69a, Fro66, Hea61, Hol61a, Kea68, Kle63a, Man65, Mil61, Sca61, Sca65a, Top66a]. Richards [Rot60]. Rico [DeG69a]. Riddle [Ric68]. Rider [Goo68b]. Riedel [Ros68]. Rifle [Gil63, Hal65b]. Rifles [Hoo66b]. Rigid [Top65a]. Rinascimento [Len60]. Rings [Cle61]. Rinman [Alt69b]. Rise [Bak65, Bar66d, Cha64, Gil60a, Lay66]. Rising [Hig66d]. Risk [Hid60]. Risks [Str64]. Ritchie [Hie63, Car69]. Rittenhouse [Wil65a]. Rivalität [Bry69]. Rivaz [Gol66b]. River [Fra64, Hid62, Pur64, Swa67a]. Road [Gha65]. Roads [Bat67]. Robert [Bry66b, Con64b, Dau62, Emm62c, Far68a, Fer68b, For62, Gru67, Har60, Hid62, Hol65b, Hol65c, Ihd65, Jaf63, Jaf64, Joh69b, Kil62, Kol62, Kra66, Mei68, Mey60, Mic64, Pen66, Phi63, Rie68b, Ros63, Rub67, Sch64a, Smi69c, Tho60, Ush61, Vat68, Wat63, Vir68]. Robertson [Hou63]. Robin [Nie69, Nix65, PS67, Top65a]. Robots [Rei64]. Robson [Wis61]. Rock [Hou61]. Rocket [Emm62c, Emm63b, Pen63, Roa63, Sto65b, Tas66, Tok63, Emm62c, Sut66b]. Rocket-Belt [Roa63]. Rocketry [BT69, Emm68a, Emm68b]. Rockets [Tok68]. Rodman [Smi64a]. Roethlisberger [Gue67]. Rohe [Con66d]. Rokoko [Dib69c]. Role [Bak67, Bed64d, Boe63, Bou66, Bru68, Dup69, Fin62b, Flo69a]. Rollo [Rya67]. Rolls [Sco66]. Rolt [Cal68c, Nix64, Sca68a, Woo66a]. Romaine [K.61]. Roman [For64b, Fus68, Mei66b, Pac68]. Romans [Lei69]. ron [Yam68]. Ronald [Bed64a, Bur61d, Car68, Erm64, Pil66, Sco66]. Ronall [Gha61]. Rondo [Sco61]. Roots [Bir66]. Ropp [Emm61]. Rosecrance [Hig68b]. Rosenthal [Flo69a]. Rosholt [Gru67]. Rossi [Bun66a]. Rossman [Maw65]. Routes [Nix65]. Roy [Bar64, Hub64, Joh62, Owe65]. Royal [Hig64, Hig65b, Hig69a, Kel68, Sil69]. Royce [Sco66]. Royston [Ste67, Ste68b]. Rubin [Gla62]. Rudd [Ihd65]. Rudolf [Bat66a, Kin66, Bry66b]. Rule [She61b, Top66b, Out65]. Rules [Wis67]. Rural [Bau64, Mos67, Wik62]. rurale [Mor69b]. Rushton [Cha69a]. Russakoff [Chr62]. Russell [Rod67, O’D69]. Russia [Jor61, Vir68, Esp68]. Russian [Alt69a, Wyn60, BT69, Tok68]. Russkogo [Wyn60]. Rüstungspolitik [May67a]. Ruth [Fer61]. Rutkowski [Smi68c]. Rutman [Lan68]. Ruttan [Ger66b].

S [Agn67, And64, Bed68a, Bur64b, Cla66, Con64c, Con65a, Dun68, For62, Gol66a, Han61, Har60, Hip67, Hoo66a, Jaf67, Joh69b, Kel69b, Mei68, Mid68, Min65b, O’D69, Ogd66, Pea66a, Pur65a, Rae62, Rei65a, Rei63, Ros67, Sca65b, Sch66, Sin66, Sko61, Ste68a, Ush61, Wil69b, Zuc68]. Safer [Pur68a]. Saga [Hun62]. Sail [Mul69]. Sails [Esp67]. Sakakibara [Gra65]. Sal [Mul65b]. Salisbury [De 64]. Salsbury [Joh69a]. Salt [Ros61, Ros61]. Salvador [PS66]. Salvadori [Con64b]. Salzgittergebiet [Ros68]. Sam [Goi69, Tho61a]. Samuel [Big64, Cal69a, Gul69, Har65b, Rol67, Ros66, Dob62]. San [Mul69, Ano66g]. Sanders [Sca65a, Nie64]. Sänger [Sca66, Sut66b]. Sanitary [Gai67]. Sankalia [Gha68a]. Santa [Dor68a]. Sargant [Ros64a]. Sass [Bry64a]. Satellite [Hal63, Sto69a, Cle61]. Saving [Hug63, Min65b]. Saxon [Pet65]. Sayles [She65c]. Sayre [Nel66, Rie68a]. Scales [Kle67]. Scarcity [DeG65b]. Scent [Hoo64]. Scharff [Joh69b]. Scheler [Car62c]. Scherer [Hol65a, Sto63]. Schmookler [Ces69b]. Schneider [Bur64a]. Schofield [Hol65c]. Scholar [tZS65]. Scholar-Technologist [tZS65]. Scholarly [Woo60c]. School [Kan65]. Schraube [Bat66a]. Schreiber [Bar69b]. Schroeder [Fin68]. Schuchardin [Gol66a]. Schultz [Con61a]. Schurr [Tho61a]. Schwerin [And68a]. Science [Aga66, Bed65b, Bee65, Bla63, Bor65a, Bun66b, Bur65b, Bur66a, Bur69a, Car65, Con65b, De 65, Dru61b, Esp66a, Fei61, Fin61, Fin65b, Goo68b, Hac68, Hal60b, Hew67, Hol65c, Ker61, KH61, Le 63, Loc69, Mor67, Mul68, Mul65a, Per68, dSP65, Rae61c, Ran63, Ras67, Rav66, Rie65, Sin69b, Sko67, Smi61b, Smi66b, Sny62, Ste63, Tib67, Uye64, Wik62, Woo69b, Zvo61b, Rie67, And64, Bar67c, Con68a, Dun63, Hin62, Hof64, Kur61a, Put67, Rol64, And62a, Bar66b, Ces65, Gul61a, Kre67, May69b, Men66, Mid63, San60, tZS64, Swa67b]. Sciences [Cha68, For60, Sko67, Col65]. Scientific [Bak66, Bak67, Bar66a, Fer63a, Gab66, Gar65, Gue61, Gul61b, Gul64, Hal61a, Har65b, Lei61b, Rie65, Ros63, Süs65, Smi64c, Gru68, Sin66, Smi64c]. Scientist [Bir62, Cha69b, Hew66, Hug61c]. Scientists [And62b, Lew64, Lew65, Rei62a, Sto64a, Fin64b]. Scienza [Bon69]. Scoville [Col61]. Screw [Sin69a]. Scrolls [PR62]. Sea [Joh69b, PR62, Fin68]. Sean [Bur65d, Wyn60]. Seaport [War68]. Search [Hug63, Mor63a]. Sears [Bri63]. Sebastian [Kur63]. Secolo [Dow64]. Second [Rob65, De 67a]. Secrecy [Joh65b]. Secret [Lei69]. Sector [Jaf68]. Security [San61a]. Segó [Gli66]. Seifert [Rae69]. Seine [For61a, Bry69]. Selden [Rae61b]. Select [Goo68b]. Selected [For60, Bee66]. Selections [Rol67]. Self [Boy68]. Self-Starter [Boy68]. Sellers [Sin69a, Sch66]. Semiconductor [Kra68]. Seminar [Lay69]. Seonaid [Hou63]. Septimus [Whi68d]. Sergeevich [Esp66b]. Serrurier [SS61]. Servan [Bar69b]. Servan-Schreiber [Bar69b]. Service [Jaf68, Sko61, Tho66, Smi64b, Ras67]. Services [Sko61, Spe62a]. Sesquicentennial [Cal69a]. Setting [She69]. Settlement [Kra66, Fus66a]. Settlements [Rot69]. Seventeenth [Yab67]. Seymour [Con67a, Ras68b]. Shakers [Ros61]. Shaped [Con68b]. Sharlin [Dib64, Mul68]. Shaw [Con62b]. Sheahan [Mil64]. Shell [Con64b, Con65a]. Shelter [Lub65a]. Sherington [Fus66b]. Sherwood [Hin66]. Shipley [Sil69]. Shipping [Lam61]. Ships [Mul69]. Shipwrights [Fer69b]. Shirley [Emm62b]. Shop [Smi67c, Gil64, Bed61a]. Short [Car68, Dib69a, Hin64, Pay63, Rol65, Woo66a]. SHOT [Hei68, Hig68c, Lew67, Rae68]. SHOT-AEA [Hei68]. SHOT-OAH [Hig68c, Lew67, Rae68]. Shriever [Rae69]. Shryock [Mil61]. Shultz [Chr61]. Shumway [Pau65, Pau67a]. Shute [Hat66]. Sidney [Bar67a, Bar66d, Gha68c, Hid62]. Sieg [Bry69]. Siegel [Ben68, Con64a]. Siegfried [Ste62]. Siemens [Hug69]. Sigfrid [Hug69]. Sigfried [Bur65a]. Significance [Le 64]. Silent [Bry66c, Fer68a, Nie69]. Silk [Han61]. Silver [Rya67]. Silvio [Gar65, Hoo64]. Sim [Net65]. Simmons [Car62b, Con61a]. Simon [Bir62, De 67b, Hal66, Hal68]. Simonson [Hol69]. Since [And62a, Bed64b, Jaf65a]. Sinclair [Bur68]. Singer [Kra60, Rod65]. Sir [Bar65b, Bor65b, Gha67, Gha69b, Hig63, Hig66c, Hig69b, Lew66, McG65, Mic66, Spi64b, Tho60, Hun63a]. Site [Ric68]. Sixteenth [Gra65, Ede66]. Sixteenth-Century [Ede66]. Sixty [Dav61]. Sixty-Ton [Dav61]. Size [Ros64a]. Sketches [Smi67e]. Ski [Dre67]. Skies [Cha69c]. Skills [Fei66, Smi69a]. Skin [Pru65]. Skolimowski [Jar66]. Skolnikoff [Kre67]. Skornia [Wei66b]. Sky [Sca65a, Con61a, Hig63]. Skyscraper [Con68d]. Skyways [Pur68a]. Slack [Swa67a]. Slater [Dob62, Bor65b]. Slaves [Kur61a]. Slide [Top66b]. Sliderule [Hat66]. Slipstream [Hig66e]. Sloan [Agn67]. Sloane [Haw66, Wel65]. Slovakia [Tib67]. Slukhai [BT69]. Smart [Bed64b]. Smelhaus [Jac69]. Smelser [Bar61a]. Smelter [Eic64, Ret65, Smi64f, Spe63]. Smelting [Phi63, Smi68a]. Smiles [Rol67]. Smith [Fer67, Hew66, Bur60a, Bur67, Bur69a, De 68b, Fin63b, Hid62, Lon67, Meh61, Svo67, Top66a, Tra65]. Smithsonian [Woo69a, Vog67, Wer67]. Snodgrass [Hoo65]. Snow [Rob65, Le 64]. Soap [Duv66]. Social [Bar63a, Bar67c, Ben63, Bir66, Con61b, Con62c, Dav64, Hol65c, How69, Jar66, Kil64, Kur61b, Lay66, Liv69, Pea66a, Rie68b, Sch65b, Sko67, Str63, Smi64c, Bar61a]. Socialism [Ham69]. Societies [Sin66, Min65b]. Society [Bar66d, Cal69a, Dup69, Fer64d, Fus66a, Gil62, Har62a, Hig69a, Hol65c, Jew64, Kel68, Mil61, Mor67, Rav65, Ros63, Sil69, Spi64a, Ste68a, Swa67b, Tas66, Tha69, Wat63, Wei66b, Woo68, Flo69a, Hig69a, Mes65, Mes69b, Rie67, Tas66, Zap62, Bar66b, Ber65b, Ces69a, Dav62, Fal65, Ihd67, Kil62, Mor63b, Rie65, Rin64, Rin66, Roc66, Smi69d, Svo67, Gil66]. Sociological [Bar61b]. Sociology [Dun63, Hin64]. Sokol’skii [Tok68]. Sol [PS64]. Soldier [Bat67]. Solid [Tok68]. Solid-Fuel [Tok68]. Solla [And62a]. Solving [Jew69]. Some [Bir62, Ger66b, Hal60a, Lew64, Mes69b, Pur68d, Ros64b]. Sonja [Fre69a]. Sonnenuhr [dSP61]. Sons [Duv66, Hea61]. Sorby [Ano64g, Bur67]. Soria [Gli69a]. Sound [Mul69]. Source [Hat66, New66a, Goo68a, Ihd69]. Sourcebook [Kol62]. Sources [Rei60, Ros66, Wyn60, Bur69a, Goo66a]. South [Fus66a, Nea64]. Southern [DeG65a, Hun67, Wai68]. Southwest [Hid68]. Soviet [Bak67, Bla64, Ham69, Hug61b, Nie67c, Dal64, Jor61, Tok63, Zvo60, Zvo61b, Bla63, Hug67b, Wil69a]. Space [All62, Car61a, Dry61, Emm62b, Emm63a, Emm68b, Gut64, Hof65, Lon69, MW63, Sca67, Smi66b, Sny62, Sto64a, Sto65a, Swe65, Swe68, Uye64, Swe68, Car62a, Emm64, Süs68a, Rob66, Roc66, Sca66, Sto68]. Spanish [Out65, Wol66, La 64]. Spatial [Rog69]. Special [DeG65a]. Species [Gre63]. Speed [Wik61, Fer68a]. Spence [Wik61]. Spencer [De 68c, Mar69a]. Spicer [dM67]. Spiegel [Pie64]. Spielereien [Dib69c]. Spinning [Hip67]. Sporn [Fin65a, Tho66]. Spot [Ede64]. Spotter [Whi68a]. Sprague [Fin62a, For64b]. Spread [De 67b]. Spreckels [Koh69]. Spring [Ano60d, Ano61i, Ano62g, Ano63j, Ano64i, Ano65h, Ano66i]. Springfield [Woo61b]. Square [She65b]. SST [Smi69c]. St [Aus68, Jor62]. Staal [Alt65]. Stacy [Fou64]. Staff [Sto64b, Sto65a, Smi66b]. Stages [Con65b]. Stahlhochbaus [Rot65]. Standard [Aga66, Lea68, Sin69a]. Standards [Pur68c]. Stanley [Bri62b, Bur67, Bur69a, Jaf65a, Meh61, Shr68, Bur60a]. Start [Hou63]. Starter [Boy68]. State [Gai67, Joh69a, Koh67, Liv69, Tra64]. States [Bed61b, Goo65b, Gra62, Kle63a, Kni67, Mon67, Mul62, Mul65a, Pur65b, Rae66, Sha67b, Spe62a, Top66a, Pen63, Sin66, Swa67b, Zap62, Zim65]. Stationary [Vog69]. Statistical [dSP65]. Statues [Dav61]. Steam [Bur68, Cle61, Fer67, Fry61, Ker61, Nix64, Piz68, Pur69a, Pur69b, Rol65, Sch65a, Vir68, Vog69, Wik61, Con63e, Bur65d, Whi67, Con63c, Gue60]. Steam-Engine [Sch65a]. Steamboating [Pur64]. Steel [And68a, Aus63a, Aus65, Bur69a, Cha61, De 64, Mar69b, McH66b, Rei65a, SS61, Smi64e, Wul65, Aus63b, Smi64d]. Steele [Jaf61]. Steelmaking [Zap62]. Steelworkers [McH61]. Steensberg [Arn69]. Stefan [Mos67]. Steinman [Kou61, Wat61a]. Steirman [Hun62]. Stekler [Swe67]. Stephen [Bla62, Joh69a]. Stereo [New66b]. Stereographs [New66b]. Stevin [Hal68, Hal66]. Stewart [Tay67]. Stieb [Bar67b]. Stifler [Fer67]. Stigler [Ber67]. Still [Gul61a]. Stockholm [De 67a]. Stone [De 64, Man65, Spi68b, Sin60b, Bat66c, Gha68a]. Stone-Cutting [Bat66c]. Stones [Per69]. Storage [Dal69]. Storia [Bon69, Kel69a, Len60]. Story [Far61, Far68b, For61c, Gor69, Joh69b, Nil61, Nix65, Raz64, Sha65, Smi69c, Tas66, Sin60b]. Strassman [Hid60]. Strassmann [DeG69a]. Strategies [Eva65]. Strategy [De 68b, Hig68b, Emm67, Wil64b, Wol65]. Stream [Dau60a]. Street [Agn67]. Stroke [Bry67b]. Structural [Con64d, Con68d, Dor68b, Jew67, Jew69]. Structure [Con64a, De 67b, McG65, Meh61, Pac68, Sko66, Swe67, Wil64b, Con64a, Con66d, Ros67, Con64b, Con67a]. Structures [Con60a]. Struggle [Lub65a]. Stuart [Bir66, Mei62b]. Students [Goo68b]. Studies [All62, Bar66a, Bau64, Hig65b, Joh62, Kam63, Lev61a, McC66b, Sam65, Sha67b, Str67, Süs65, Smi64c, Had68, Had69, Che61b, Had67, Lev61a]. Study [Bru68, Buc68c, Dau62, Ede64, Emm63a, Fer64d, Fus66a, Gab66, Gha68b, Gue62a, Ham68, Har69, Har60, Hea61, Jor62, Lub65b, Mec66, Nel66, Pac68, Pea66b, dSP65, Puh64, Ras68a, Rie68b, Ros66, She61a, Smi68c, Spi64a, Tho61a, Top65a, Ush61, Whi61, Wil65c, Deu65, Mei66b]. Stuhlinger [Hof64]. Sturt [Gil64]. Subject [Pur65a]. Submarine [Dib69b]. Subsequent [Hig64]. Success [Aga67]. Successful [Jaf63]. Suction [Sha64]. Sufrin [Bar67a]. Sugar [Koh69, Ros65b]. Sullivan [Con64d]. Sumerians [Big64]. Summary [Sto64a, Jun68]. Summer [Ano60e, Ano61j, Ano62h, Ano63k, Ano64j, Ano65i, Ano66j]. Sung [Yab67]. Superior [Sco66]. Supersonic [Smi69c]. Supplement [Fer64b]. Supply [Spe62a]. Support [Pur68c]. Sure [Bur69b]. Surgical [Ger66a]. Survey [Bro65a, Wai67, BT69, Bur69c, Had63]. Surveying [Bed64b, Gli68]. Survivor [Gil63]. Susan [Fer67]. Susskind [Kin66]. Sutermeister [Jaf64]. Swain [Zim64]. Swanson [Dau64]. Sweden [Fer61]. Swenson [Dun68]. Sword [Hig65c]. Symbol [Kou66]. Symposium [Ano64g, Bur67, Che61b, Gil66, Kle65, Ras68b]. Synthetic [Hug66]. synthetische [May67a]. Syrett [Bra66]. System [Buc62, Con64d, Gil66, Hei62, Rie68b, Sch65b, Zuc68]. Systems [Fus67c, Rae69]. Szechy [Viv68].

T [Bak67, Cal68c, Car69, Goo66a, Ihd67, Kol64, McK64b, Mul69, Nix64, Pay63, Rei67, Rei64, Sca65b, Sca68a, War68, Wat65, Wei66a, Wil67, Woo66a]. T-2 [Sca65b]. Ta-chung [hC66]. Tabulation [Mur69]. Takes [Bur62b, Kun60, Sch60]. Tale [Nil61]. Tanning [Wel63, Pur65b]. Taplin [Aus63a]. Tappan [Pru65]. Tarasov [Hoo66a]. Tariff [Pur62]. Task [Deu66]. Tassel [Swa67b]. Tatnall [Smi67b, Smi67b]. Tatra [Mos67]. Tax [PS64]. Taylor [Joh69b, Wil65c]. Taylorism [Gue61]. Teapot [Wie64]. Technical [End68, Fer63a, Fer65d, Fin62b, For60, Hal62, Lay67, Sam65, Sko61, Sko67, Spe62a, Ush61, Zvo61a, Smi64c, Bar67a]. Technics [Dru68, Mum64, Mum66]. Technik [May67b, May69d, Ret61, Bat66b, For63b, Gol66a, Ros69]. Technika [Mos67]. Technikgeschichte [May68, Har63, Har64, Har66]. Techniques [Bul66, Dau60b, Gha68a, Jen69, Rot69, Whi63, For64b, For63a, Buc68a, For65b]. Technischen [Har61, May69d, She63]. Technocrats [Lay66, Lay68]. Technoculture [Bur61b]. Technological [Bak67, Cal69a, Col66, De 64, Dru61b, Dru66, EW62, Esp67, Etz66, Fal65, Ful63, Har62b, Hei68, Hid60, Hig68c, Jar66, Kol64, La 64, McC69, Mum61, PSD64, Puh64, Rei60, Whi61, Whi69, dM67, DeG69a, Smi69d]. Technologie [Har65a]. Technologies [Bar63a, Car65]. Technologist [Gul69, tZS65]. Technology [Aga66, Ano64a, Ano68b, Bed64d, Bee65, Bil69, Bro65a, Buc62, Buc68c, Bun66b, Bur69b, Bur61c, Cal62b, Cha61, CB64, Cla62, Con64c, Con65b, Con66a, Dal64, Dal69, De 65, DPS66, De 68c, Dru61b, Dru63, Dru69, Dup69, Emm63b, Fei61, Fei66, Fer62a, Fer62b, Fer62c, Fer63b, Fer64e, Fer64f, Fer65c, Fer68b, For64b, Gol62, Goo65c, Goo66b, Goo67b, Goo68c, Goo68b, Goo69b, Hac68, Hal60a, Hew64, Hir64, Hug61a, Hug63, Jor61, Kra60, Kre67, Le 65, Lei61b, LD60, Lew67, Liv69, Mei60, Mes65, Mes69b, Meu66, Mor67, Mul65a, Nun63, Per68, Pia61, dSP65, Pur62, Pur68d, Rae60a, Rae60b, Rae68, Ras60, Rav65, Rie67, San60, Sch62b, Sch64b, Sha63a, She68, Shu67, Sin60a, Sko66, Smi66b, Smi69b, Sny62, Str63, tZS64, Tav69, Tib67, Tok63]. Technology [Tra64, WW62, Wer67, Whi60, Wik62, Wil63, Woo60a, Woo60c, Woo65, WW66, Yab67, Zvo60, Zvo61b, Zvo62, Dor68a, Fer64d, Mul65a, Ras64b, Sto64b, Bor65a, Fin63a, Ham68, Jaf65b, Mar63, Mec66, Mid63, Oli60, Pay63, Pie64, Rav66, Sto65b, Woo69a, Ano60b, Ano61c, Ano62c, Ano63d, Ano64b, Ano65b, Ano66a, Ano67b, Ano68a, Ano69d, Ber65a, Bro62, Buc68b, Dav68, DeG69b, Fer64c, Gul64, Hoo66a, Kea68, Ran63]. Tecnica [Len60, Kel69a]. Telegraph [Dib64]. Telegraphen [vW67]. Telegraphic [Fra64]. Telegraphie [vW67]. Telemetry [MW63]. Telescope [Bed67a]. Television [Wei66b, Wei66b]. Teller [Kle62]. Temin [Aus65]. Tempir [Jac69]. Ten [Bur66a, Whi68d]. Ten-Wheel [Whi68d]. Terence [Raz64]. Term [The62]. Tesla [Sha65]. Testing [Hig65b, Smi67b]. Tew [Che61b]. Text [Ede64]. Textile [Hip69, La 64, Lea67, Out62, Rog60, Wol66]. Textiles [Har69, Rog60]. TFX [Smi69c]. Theater [Con64d]. Theatre [Fer66a]. Theatrum [Alt69a]. Their [Bak67, Bed64c, Big64, Gar65, Gha68a, Hig68a, Les60, New66b, Pru62, Rae61a, She68]. Them [Hig65b, Sca65a]. Theobald [Sch64a]. Theodor [Jen67]. Theodore [Aus63b, Che61a, Emm61, Smi64d, Süs68a, Mal64]. Theodosius [Gre63]. Theophilus [Tho65c, Whi64]. Theories [Jen69]. Theory [Ayr67, Bar67a, Bar61a, Kil63, Hea63]. There [Bri63]. Thinking [Bur66a, Sko66]. Thirsk [Fus67b]. Thomas [Aus68, Chr61, Lon67, Mar69b, Rei65a, Rol67, Emm62b, Hig66c, Nix64]. Thomsen [Hei62]. Thomson [Hug61c, Spe62a]. Thor [Per63]. Thorbecke [Ayr67]. Thorndike [BB66]. Thorne [She65d]. Thornton [Lam61].

Thorstein [Lay69, Lyn67]. Those [Hig66a, Sca65a]. Thread [Sin69a]. Three [Goi69, Hei62, Woo68]. Three-Age [Hei62]. Throwing [Per69]. Tibor [Ble63, Gab66]. T’ien [Yab67]. T’ien-kung [Yab67]. Tillage [Arn69]. Time [Kur63, Bri68, Hoo64]. Times [Bar65b, Bea66, Cha69a, Kou61, Ste67, Ste68b, Bar65b, Rie67, Smi69a]. Timm [Har65a]. Tin [Pur62, Smi68a, Pea66a]. Tin-Plate [Pur62]. Titan [Per63]. Today [Fin65b]. Tomato [Ras68a]. Ton [Dav61]. Tönnies [Ben63]. Tool [Bur61a, Wel66, Woo61b]. Tools [Gha68a, Mum60, She61b, She65b, Woo69b, Zvo61a, Alt69a, Fer66b, Wel65, Woo66a, Arn69]. Toronto [Ano68g]. Torroja [Con60a]. Torsten [Rog69]. Tower [Kni67, Cal66]. Town [Con64a, Jac60, Pau67b]. Toy [Gha68c]. Toynbee [Bri68]. Toys [Wei63]. Trachtenberg [Kou66]. Trackers [Cha69c]. Traction [Bur61d, Erm64]. Tractor [Dal64]. Trade [K.61, Ver64, Rod67]. Trades [For60]. Tradition [Mor67, Rae60b, Rot60, Whi61, PS68]. Traditional [She68, Smi69a, Wel67]. Traditionalism [DPS66]. Trail [Fin63a]. Transactions [Fer64d, Har62a, Jew64]. Transatlantic [Hun62]. Transfer [Dal64, De 68c, Sch64b]. Transition [Lub65b, Ste68c, Wat63]. Translation [Ede64]. Transport [Smi69c]. Transportation [CB64, Emm69, Mul69, Rae69, Uye68a, Uye68b, Con64c, Owe65]. Transporting [Dav61]. Travel [Emm68b]. Traveling [Kin66]. Traveling-Wave [Kin66]. Treatise [Alt69b, Ret63, Sch62a, Whi69]. Treatment [Chi63]. Treibstoff [May67a]. Trends [Bur65b]. Tretmuhle [Ret61]. Treue [Whi68e, Bat66a]. Trevor [Pay63]. Trinity [Hew66]. Triumph [Hun62]. Truesdell [Mur69]. Tsander [Sut66a]. Tube [Kra68, Kin66]. Tuchel [May69d]. Tudor [Goo69a]. Tula [Kel69b]. Tunnard [Con63b]. Tunnelling [Viv68]. Turn [Sin69a]. Turner [Deu66, Bed68a]. Turning [Con64a]. Turnpike [Hun63b]. Turriano [Ret65, Ret67a, Wul66b]. TVA [Con60b, Swa67a]. Twain [Tra65]. Twelve [Cal68b]. Twentieth [BB62, Ste68c]. Two [Hew64, Mes66, Rin68, Rob65, Tas66, Le 64]. Tybout [Fro66]. Tylecote [For64a]. Tyneside [Bur66d]. Type [Fra64]. Typefounding [Rya67].

U [Mor63a, Wyn60]. U. [Sko61]. U.S [Bur69c, Fer68b, Gil66, Sha67a, Tho65b]. U.S. [Hal63, Rei60, Tas66]. U.S.N [Wil68]. Uhr [dSP61]. Uhrmacher [Bry68b]. Umberto [Len60]. Umformens [Hoo62]. Unaware [Kur61a]. Uncountable [Dib69a]. Underdeveloped [DeG69b]. Underdevelopment [Kam63, Bar65a]. Understanding [Mes65, Mes66]. UNESCO [Fer63a]. Unfinished [Bed69]. Unification [Mul68]. Unified [Zvo61a]. Union [Car61b, Kni65, Hug61b, Nie67c]. United [Goo65b, Gra62, Hut63, Hut66, Kle63a, Kni67, Mon67, Mul62, Mul65a, Pur65b, Rae66, Sha67b, Sin66, Swa67b, Top66a, Zap62, Zim65, Pen63, Bed61b, Spe62a]. Universal [Kni65]. Universities [Gul64, Boe63]. University [Buc68c, Dor68a, Koh67, Liv69, Mes65, Mes69b, Rei67, Tha69, Aus68, May69b].

Unsinkables [dC69]. until [Bis69]. Uplift [Har65b]. Upper [Pur64]. Urban [Con64c, Con68c, Gut64, Meu64, Uye64, Uye68b]. Urbanization [Lew67]. Urbanizing [Lub65a]. Use [Gut64, Net65, Tho66, Wal69, Zim65, She65d]. Uses [Spe62b]. Usher [Ano69a, Smi61c, Smi65c]. Utility [Smi67d]. Utilization [Dry61, Bir62]. Utopian [Bar63c]. Utopians [Rie68b, Lok68]. Utrecht [Bed69].

V [Alt69a, Bed64c, Bla62, Boy69, Bra69, Che61a, Emm62a, Fou64, Hal68, Hoo66a, Jac69, Nie67b, Rie65, Tok68, Ano64b, Dor63, Fer64e]. V-2 [Dor63]. Václav [Smi69a]. Vaizey [Wil61a]. Vale [Fer69b]. Valencia [Gli68]. Valjavec [Hug67a]. Vall [Gli66]. Valley [Gha67, Lea67, Woo61b, Spi69]. Value [McH66a]. Values [Tra64, Rie67, Smi68c]. Vanguard [Hag63]. Vannoccio [Cah67]. various [Sto64a]. Vaughan [dC61a]. Vault [She61a]. Veach [Bul66]. Veblen [Lay69, Lyn67]. Vedecké [Süs65]. veneta [Lor67a]. Venice [Ede66]. Venture [Fer68a, Sch65a]. Venus [Sto64b]. Verbrennungsmotor [Bry69]. Verbrennungsmotorenbaues [Bry64a]. Verbrennungsmotors [Bry69]. Vermont [Woo61b]. Verner [Rei65b]. Vernon [Ger66b, Swe65]. Versuch [vW67]. versus [Wil69a]. VI [Col66, Ano65b, Fer64f]. Vice [Smi67b]. Victor [Bur69c, Jaf68]. Victorian [Bri62a, Fer68a, Ste68b]. Vidonov [Ogd66]. Vie [Smi67a]. View [Gul61a, Kur61a, Ols69, Zvo61b, Wei66b]. Viewed [Bri62a]. Views [New66b]. vii [And65b, Ano66a, Fer65c]. VIII [Bra66]. Viking [Hag63]. Village [Ric68, Fus66b, Nea64]. Villages [Lan68]. Vinci [Ano63e, Ano64c, Ano65c, Ano66b, Ano67c, Ano68c, Ano69e, Ret67b, Ste63, Ret65]. Violence [Mon67]. Virginia [Hun63b, Pat62]. Virginian [Con63e]. Virus [Fly68]. Vision [Cla62, McH66a]. Vitruvius [Kil63]. Vladimir [Esp66b]. Vogel [Fer68b]. Vol [And65b, Bra69, Bra66, Bur65a, Col66, Fus67b, Fus67a, Hal68, Har62a, Hut63, Hut66, Jew64, Kni67, Man65, Mei66a, Sha67a, tZS64]. Volkmann [Kin68]. Volume [Ano60b, Ano60f, Ano60d, Ano60e, Ano61c, Ano61k, Ano61i, Ano61j, Ano61h, Ano62c, Ano62i, Ano62g, Ano62h, Ano62f, Ano63d, Ano63l, Ano63j, Ano63k, Ano63i, Ano64b, Ano64k, Ano64i, Ano64j, Ano64h, Ano65b, Ano65j, Ano65h, Ano65i, Ano65g, Ano66a, Ano66k, Ano66i, Ano66j, Ano66h, Ano67b, Ano67h, Ano67g, Ano67i, Ano67j, Ano68a, Ano68i, Ano68h, Ano68j, Ano68k, Ano69d, Ano69j, Ano69i, Ano69k, Ano69l, Kle63a]. Volumes [Fer64d]. voor [Alt65]. Voraussetzungen [May69d]. Vorgeschichte [Ros68]. Vorzeit [May67b]. Voyce [Esp68]. vs [Hea63]. Vyp [Bra69].

W [And68b, Aus63a, Bar67b, Bar63b, Bar63c, Bar65b, Bar66c, Boe63, Bry66b, Buc68b, Bur65e, BB62, Cal66, DeG69a, Fer66b, Ger66b, Gol66a, Gru68, Gul61a, Gul61b, Hal65a, Hid60, Kan65, Kil64, Lay66, Lon65, Mah61, Pea65, Pen66, Pie66, Rae69, Ras61, Rav66, Rei65b, Ric68, Rin64, Rod65, Rol65, Rot69, Sch62a, Sch63a, Sha63a, Sha65, Shr63, Smi64a, Smi69d, Son63, Tow63, Uye68a, Vat68, Wei66a, Wer64b, Wil65c, Wil61c, Wol67, Mos67].

Wachsmann [Con64a]. Wages [Bar62]. Wagley [PS68]. Wagner [Son63]. Wagon [Pau65, Pau67a]. Waldstein [Ush64]. Wales [Fus67b]. Walsh [Hea61]. Walter [Bra65a, Con61a, Dun63, Joh69b, Kol62, Süs68b, Tow65]. Wanetta [Sha65]. Wanted [Kle65]. War [Ful63, Hac68, Sko61, Spe62a, Wei66a, Hal66, Hig65b, Hig69b, Jaf61, Pat62, Rei62a, San60, Sch63a, Sko61, Emm61, Ros64a]. Ware [Nee69]. Warfare [Sko61, Hun64, Nie69]. Warner [Goi69, Rav66, Smi69d]. Warren [Col61, Pea65, Tei67]. Wars [Mos67]. Was [Süs68b]. Washington [Ano62j, Ano67k]. Watch [Bed61b, Hoo66a]. Watches [Bed65a, Bed67b]. Water [Fer67, Hun67, Kel69b, Pur69a, She65d, Wul66b]. Water-driven [Kel69b]. Water-Mills [Hun67]. Waters [Kou61, Swa67a]. Watertown [Gue61]. Waterwheels [Ret65, Ret67c]. Watkins [Vog69]. Watney [Wil64a]. Watson [Tho60, Tho60]. Watson-Watt [Tho60, Tho60]. Watt [Sch65a, Tho60, Tho60]. Wave [Kil63, Kin66]. Wayne [Bri61, Woo61b, Liv69]. Wealth [Sha67b, Bou61, Sha67b]. Weapons [Gra61, Hig65b, Hol65a, Nie69, Sto63, Top65a, Hoo65]. Weather [Mul62]. Weavers [Moe69]. Webb [Esp66a]. Wedel [Hoo62]. Wedgwood [Sch62b]. Weeks [Kol62]. Weg [Ret61]. Wegbereiter [Hug69]. Weidner [Boe63]. Weights [Kle67, Fou64]. Weiher [Hug69]. Weiner [Cha69b, Spe69]. Wells [Lok68]. Welsh [Pur65b]. Welt [Bry66a]. Wentzel [Bed68b]. Werk [Kin68]. Werner [Con66d, Hug69]. Wernher [Emm68b, Sto68]. Wertime [Aus63b, Smi64d]. Wesley [Pie66]. West [Mar69b, Pau67b, Smi64a, Wil67]. Westcott [Con63c]. Western [Bru68, Buc68b, Jaf67, McG65, Nie67c, Oli60, She68, Mil62]. Westin [Jac68]. Westward [Pau65, Pau67a]. Wheel [Fer64c, Whi68d, Cha64]. Wheeler [Hig65b, Gha67, Gha69b, Hig66a]. Wheelock [San61b]. Wheels [Rae61b, Kol66]. Wheelwright [Gil64]. Where [Tyl62, Bri63]. Whisler [Chr61]. White [Sto69b, Str63, Con66b, Fer61, Fus68, Gra60, Her62]. Whitehead [Bur66c]. Whitfield [Svo67]. Whitnah [Mul62, Pur68a]. Whitney [Fed60, Woo60b]. Who [And68b, Sca65a, Mic64]. Wide [Wat63]. Wiejski [Mos67]. Wiener [Bro65b]. Wilburn [Jaf68]. Wilderness [Cal68d]. Wilfred [Edw67, Wol65]. Wilhelm [Bat66a, Whi68e]. Wilkes [Bur66d]. William [Bee66, Bor65b, Bru68, Bur61a, Car61b, Con63d, Fin64b, Flo69a, Gue67, IU65, Kou61, Mei66b, New66b, Pur64, Rae61b, Rae69, Rob69, Rub67, Sel69, Whi68c, Wis61, Sin69a, Sil69]. Williams [Pay63, Wil61b]. Wilmon [Swa67a]. Wilson [Wei66b, Wol67, Bry66b, Hig66e, Lew69a]. Wind [Süs68a]. Windmills [Wai68]. Wing [Hig66c]. Wingo [Gut64]. Winslow [Qui65]. Winter [Ano61h, Ano62f, Ano63i, Ano64h, Ano65g, Ano66h]. Winterthur [Hec67, Cal67, Cal69b, Kou65]. Wire [Wil67]. Wireless [Hut66]. Wirkungen [May69d]. Wirtschafts [May67a]. Wirtschafts- [May67a]. Wirtschaftspolitik [Fre69a]. Wirtschaftsraumen [For61a]. Wisdom [Aga67]. within [Jaf68]. Witt [Min65a]. Wittek [Rot65]. Wittenberg [Ras66]. Wittmann [Woo61a]. Wolf [Bur64a]. Wolfe [Owe65]. Wolfgang [May67a, New66a]. Woman [Ber65a]. Women [Bak66, Bak67]. Wonders [For61b]. Wood [Shr63, Haw66]. Woodbury [For62, Har60, Hug61c, Ush61]. Woodman [Bla62]. Woodruff [Bru68]. Woodson [Pie66]. Woodworking [Fer66b, Wel66]. Woolen [Ush64]. Woolf [Ces65]. Woolley [Spi64b]. Woolrich [And68b, Cal66]. Worcester [Smi67e]. Words [Tho65c]. Work [Jac64, Kur63, Smi69a, Wat65, Wil65c, Zvo61a, Ber65a, McC66a]. Worker [Deu65, Deu66, Gue62a, Gue65]. Workers [Bur65e]. Workmanship [Con69c]. Works [Hal68, McH65, Pur69a, She65a]. Workshops [Bul62, Cal68d]. World [Boe63, Bru68, Bur65b, Cle61, Hac68, Har62b, Hig65a, Hin62, Kle63a, Pil66, San61a, San64, Sko61, Spe62a, Ste63, Ver64, WW62, Wat63, Wei66b, Ful63, Dow65, Emm61, Emm62a, For61b, Fus68, Hoo66b, Lub65a, Mei66a, Bur65b, Bur66a, Le 65]. World-Wide [Wat63]. Worlds [Rin68]. Woroniak [De 68c]. wozu [Süs68b]. Writings [Ret65]. Wrought [Con62b]. wu [Yab67, tZS65]. Wulff [She68, Wer68b]. Wyld [McN64]. Wynn [She65d].

XII [Mar63]. XIX [Dow64]. XVII [Dow64]. XVIIIth [Alt69a]. XXXI [Har62a]. XXXII [Fer64d]. XXXIII [Jew64].

Yamada [Yam68]. Yamey [Min65b]. Yankee [Tra65, Fer61]. Year [Bur66a, Sch63b, Tho65b, Whi68c]. Years [Lea69, Mes66, Pau65, Pau67a, Sto65a, PSD67]. Yeh [hC66]. Ying [Yab67]. Ying-hsing [Yab67]. York [And65b, Ano61d, Joh62, Pur69a, Uye68a]. Yorkshire [Phi63]. Young [Fer65a, Nie67c]. Youngson [Mor61]. Youth [McC67]. Yudelman [DeG65a].

Z [Jac69]. Zagorskii [Esp66b]. Zauberman [Ros65a]. Zeichnens [Har61, She63]. Zeiss [Kin68]. Zeit [May67b, Ret61]. Zemedelského [Süs65]. zum [Woo61a]. zur [May67a, New66a, dSP61]. Zvorikine [Dru61a]. Zvorkine [Zvo61a]. zwischen [Bry69]. Zwölfbrüderstiftung [Whi68e].

References

Agassi:1966:CBS

[Aga66] Joseph Agassi. The confusion between science and technology in the standard philosophies of science. Technology and Culture, 7 (3):348–366, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/ 1/article/894580/pdf. See [Wis66] and replies [Aga67, Wis67].

Agassi:1967:PSR

[Aga67] Joseph Agassi. Planning for success: a reply to Professor Wisdom. Technology and Culture, 8(1):78–81, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894492/pdf. See [Wis66].

Agnew:1967:CMS

[Agn67] Donald C. Agnew. Classrooms on Main Street by Harold F. Clark, Harold S. Sloan (review). Technology and Culture, 8(4):535–537, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894425/
pdf.

Allen:1962:PSI

[All62] Francis R. Allen. Proposed Studies on the Implications of Peaceful Space Activities for Human Affairs by Donald N. Michael (review). Technology and Culture, 3(2):216–219, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895401/pdf.

Althin:1965:EAV

[Alt65] Torsten Althin. Economische Aspecten van bet “Directe Procede” voor de Bereiding van Ijzer en Staal by R. B. Harteveld (review). Technology and Culture, 6(3):454–455, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894686/pdf.

Althin:1967:HH

[Alt67] Torsten Althin. Hargraves, not Hargreaves. Technology and Culture, 8(4):491, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894402/pdf.

Althin:1969:NHT

[Alt69a] Torsten Althin. Nartov and His Theatrum Machinarum, Russian XVIIIth Century Machine Tools by V. V. Danilevskii (review). Technology and Culture, 10(1):96–99, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892325/pdf.

Althin:1969:RTM

[Alt69b] Torsten Althin. Rinman’s treatise on mechanics. Technology and Culture, 10(2):185–186, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892359/pdf.

Anderson:1962:SSB

[And62a] David L. Anderson. Science Since Babylon by Derek J. de Solla Price (review). Technology and Culture, 3(2):175–177, Spring 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895380/pdf.

Anderson:1962:IAE

[And62b] David L. Anderson. The Impact of Atomic Energy: a History of Responses by Governments, Scientists, and Religious Groups by Erwin N. Hiebert (review). Technology and Culture, 3(3):355–358, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895153/pdf.

Anderson:1964:ESE

[And64] David L. Anderson. The Evolution of Science ed. by Guy S. Metraux and François Crouzet (review). Technology and Culture, 5(4):598–600, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894824/
pdf.

Anderson:1965:NPEb

[And65a] David L. Anderson. The Natural Philosopher ed. by Daniel E. Gershenson and Daniel A. Greenberg (review). Technology and Culture, 6(3):536–537, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894727/pdf.

Anderson:1965:NPV

[And65b] David L. Anderson. The Natural Philosopher, Vol. III, edited by Daniel E. Gershenson and Daniel A. Greenberg. New York: Blaisdell, 1964, Pp. vii + 111 (review). Technology and Culture, 6(1):132–134, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894795/
pdf.

Anderson:1968:OSP

[And68a] Robert T. Anderson. Oil and Steel: Processes of Karinya Culture Change in Response to Industrial Development by Karl H. Schwerin (review). Technology and Culture, 9(2):267–268, April 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894331/pdf.

Anderson:1968:MWC

[And68b] Oscar E. Anderson, Jr. The Men Who Created Cold: a History of Refrigeration by W. R. Woolrich (review). Technology and Culture, 9(1):140–141, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894375/pdf.

Angel:1967:MAR

[Ang67] J. Lawrence Angel. Man Adapting by René dubos (review). Technology and Culture, 8(1):148–150, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894525/pdf.

Anonymous:1960:CMD

[Ano60a] Anonymous. The Chicago Meeting, December 28–30, 1959. Technology and Culture, 1(2):195–200, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895262/pdf.

Anonymous:1960:IVT

[Ano60b] Anonymous. Index to volume I of Technology and Culture. Technology and Culture, 1(4):431–433, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895627/pdf.

Anonymous:1960:ON

[Ano60c] Anonymous. Organizational notes. Technology and Culture, 1(4):426– 429, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895626/
pdf.

Anonymous:1960:VNSa

[Ano60d] Anonymous. Volume 1, number 2, spring 1960. Technology and Culture, 1(2):??, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1960:VNSb

[Ano60e] Anonymous. Volume 1, number 3, summer 1960. Technology and Culture, 1(3):??, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1960:VNF

[Ano60f] Anonymous. Volume 1, number 4, fall 1960. Technology and Culture, 1(4):??, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1961:Aa

[Ano61a] Anonymous. Announcements. Technology and Culture, 2(2):217–218, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895568/
pdf.

Anonymous:1961:Ab

[Ano61b] Anonymous. Announcements. Technology and Culture, 2(4):437–439, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895456/
pdf.

Anonymous:1961:IVI

[Ano61c] Anonymous. Index to volume II of Technology and Culture. Technology and Culture, 2(4):440–446, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895457/pdf.

Anonymous:1961:NYM

[Ano61d] Anonymous. The New York Meeting, December 27–30, 1960. Technology and Culture, 2(2):206–216, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895567/pdf.

Anonymous:1961:NA

[Ano61e] Anonymous. Notes and announcements. Technology and Culture, 2(3):302–304, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895488/pdf.

Anonymous:1961:ONa

[Ano61f] Anonymous. Organizational note. Technology and Culture, 2(1):96, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895604/
pdf.

Anonymous:1961:ONb

[Ano61g] Anonymous. Organizational notes. Technology and Culture, 2(4):435– 436, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895455/
pdf.

Anonymous:1961:VNW

[Ano61h] Anonymous. Volume 2, number 1, winter 1961. Technology and Culture, 2(1):??, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1961:VNSa

[Ano61i] Anonymous. Volume 2, number 2, spring 1961. Technology and Culture, 2(2):??, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1961:VNSb

[Ano61j] Anonymous. Volume 2, number 3, summer 1961. Technology and Culture, 2(3):??, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1961:VNF

[Ano61k] Anonymous. Volume 2, number 4, fall 1961. Technology and Culture, 2(4):??, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1962:Aa

[Ano62a] Anonymous. Announcements. Technology and Culture, 3(2):245–246, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895410/
pdf.

Anonymous:1962:Ab

[Ano62b] Anonymous. Announcements. Technology and Culture, 3(3):373–375, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895161/
pdf.

Anonymous:1962:IVI

[Ano62c] Anonymous. Index to volume III (1962) of Technology and Culture. Technology and Culture, 3(4):662–667, Fall 1962. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895125/pdf.

Anonymous:1962:NA

[Ano62d] Anonymous. Notes and announcements. Technology and Culture, 3(1):110–114, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895428/pdf.

Anonymous:1962:ON

[Ano62e] Anonymous. Organizational notes. Technology and Culture, 3(4):659– 661, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895124/
pdf.

Anonymous:1962:VNW

[Ano62f] Anonymous. Volume 3, number 1, winter, 1962. Technology and Culture, 3(1):??, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1962:VNSa

[Ano62g] Anonymous. Volume 3, number 2, spring 1962. Technology and Culture, 3(2):??, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1962:VNSb

[Ano62h] Anonymous. Volume 3, number 3, summer 1962. Technology and Culture, 3(3):??, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1962:VNF

[Ano62i] Anonymous. Volume 3, number 4, fall 1962. Technology and Culture, 3(4):??, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1962:WMD

[Ano62j] Anonymous. The Washington Meeting, December 28–30, 1961. Technology and Culture, 3(2):235–244, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895409/pdf.

Anonymous:1963:Aa

[Ano63a] Anonymous. Announcements. Technology and Culture, 4(1):128–129, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895066/
pdf.

Anonymous:1963:Ab

[Ano63b] Anonymous. Announcements. Technology and Culture, 4(2):274–276, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895027/
pdf.

Anonymous:1963:EN

[Ano63c] Anonymous. Editor’s note. Technology and Culture, 4(4):??, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894957/pdf.

Anonymous:1963:IVI

[Ano63d] Anonymous. Index to volume IV (1963) of Technology and Culture. Technology and Culture, 4(4):531–536, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894970/pdf.

Anonymous:1963:LVM

[Ano63e] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 4(2):257–258, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895025/pdf.

Anonymous:1963:NA

[Ano63f] Anonymous. Notes and announcements. Technology and Culture, 4(3):374–375, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894997/pdf.

Anonymous:1963:PMD

[Ano63g] Anonymous. The Philadelphia Meeting, December 27–29, 1962. Technology and Culture, 4(2):259–273, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895026/pdf.

Anonymous:1963:PAM

[Ano63h] Anonymous. Program for 1963 annual meeting. Technology and Culture, 4(4):529–530, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894969/pdf.

Anonymous:1963:VNW

[Ano63i] Anonymous. Volume 4, number 1, winter 1963. Technology and Culture, 4(1):??, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1963:VNSa

[Ano63j] Anonymous. Volume 4, number 2, spring 1963. Technology and Culture, 4(2):??, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1963:VNSb

[Ano63k] Anonymous. Volume 4, number 3, summer 1963. Technology and Culture, 4(3):??, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1963:VNF

[Ano63l] Anonymous. Volume 4, number 4, fall 1963. Technology and Culture, 4(4):??, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1964:CBH

[Ano64a] Anonymous. Current bibliography in the history of technology (1962). Technology and Culture, 5(1):138–148, Winter 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894956/pdf.

Anonymous:1964:IVV

[Ano64b] Anonymous. Index to volume V (1964) of Technology and Culture. Technology and Culture, 5(4):648–654, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894848/pdf.

Anonymous:1964:LVM

[Ano64c] Anonymous. The Leonardo de Vinci Medal. Technology and Culture, 5(2):280–281, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894915/pdf.

Anonymous:1964:NA

[Ano64d] Anonymous. Notes and announcements. Technology and Culture, 5(3):484–488, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894889/pdf.

Anonymous:1964:PMD

[Ano64e] Anonymous. The Philadelphia Meeting, December 27–29, 1963. Technology and Culture, 5(2):282–295, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894916/pdf.

Anonymous:1964:PAM

[Ano64f] Anonymous. Program for 1964 annual meeting. Technology and Culture, 5(4):645–647, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894847/pdf.

Anonymous:1964:SCS

[Ano64g] Anonymous. Sorby centennial symposium on the history of metallurgy. Technology and Culture, 5(2):295–296, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894917/pdf.

Anonymous:1964:VNW

[Ano64h] Anonymous. Volume 5, number 1, winter 1964. Technology and Culture, 5(1):??, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1964:VNSa

[Ano64i] Anonymous. Volume 5, number 2, spring 1964. Technology and Culture, 5(2):??, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1964:VNSb

[Ano64j] Anonymous. Volume 5, number 3, summer 1964. Technology and Culture, 5(3):??, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1964:VNF

[Ano64k] Anonymous. Volume 5, number 4, fall 1964. Technology and Culture, 5(4):??, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1965:EN

[Ano65a] Anonymous. Editor’s notes. Technology and Culture, 6(1):??, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894775/pdf.

Anonymous:1965:IVV

[Ano65b] Anonymous. Index to volume VI (1965) of Technology and Culture. Technology and Culture, 6(4):697–711, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894671/pdf.

Anonymous:1965:LVM

[Ano65c] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 6(2):324–326, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894772/pdf.

Anonymous:1965:MMD

[Ano65d] Anonymous. The Montreal Meeting, December 27–30, 1964. Technology and Culture, 6(2):327–345, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894773/pdf.

Anonymous:1965:NA

[Ano65e] Anonymous. Notes and announcements. Technology and Culture, 6(4):692–695, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894670/pdf.

Anonymous:1965:ONA

[Ano65f] Anonymous. Organizational notes and announcements. Technology and Culture, 6(3):541–546, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894730/pdf.

Anonymous:1965:VNW

[Ano65g] Anonymous. Volume 6, number 1, winter 1965. Technology and Culture, 6(1):??, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1965:VNSa

[Ano65h] Anonymous. Volume 6, number 2, spring 1965. Technology and Culture, 6(2):??, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1965:VNSb

[Ano65i] Anonymous. Volume 6, number 3, summer 1965. Technology and Culture, 6(3):??, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1965:VNF

[Ano65j] Anonymous. Volume 6, number 4, fall 1965. Technology and Culture, 6(4):??, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1966:IVV

[Ano66a] Anonymous. Index to volume VII (1966) of Technology and Culture. Technology and Culture, 7(4):583–594, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894575/pdf.

Anonymous:1966:LVM

[Ano66b] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 7(2):206–208, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895175/pdf.

Anonymous:1966:NAa

[Ano66c] Anonymous. Notes and announcements. Technology and Culture, 7(1):136–142, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893848/pdf.

Anonymous:1966:NAb

[Ano66d] Anonymous. Notes and announcements. Technology and Culture, 7(3):448–451, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894610/pdf.

Anonymous:1966:NAc

[Ano66e] Anonymous. Notes and announcements. Technology and Culture, 7(4):580–582, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894574/pdf.

Anonymous:1966:PN

[Ano66f] Anonymous. Prefatory note. Technology and Culture, 7(3):301–302, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894576/
pdf.

Anonymous:1966:SFM

[Ano66g] Anonymous. The San Francisco Meeting December 27–30, 1965. Technology and Culture, 7(2):209–224, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895176/pdf.

Anonymous:1966:VNW

[Ano66h] Anonymous. Volume 7, number 1, winter 1966. Technology and Culture, 7(1):??, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1966:VNSa

[Ano66i] Anonymous. Volume 7, number 2, spring 1966. Technology and Culture, 7(2):??, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1966:VNSb

[Ano66j] Anonymous. Volume 7, number 3, summer 1966. Technology and Culture, 7(3):??, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1966:VNF

[Ano66k] Anonymous. Volume 7, number 4, fall 1966. Technology and Culture, 7(4):??, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1967:A

[Ano67a] Anonymous. Announcements. Technology and Culture, 8(2):332–333, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894485/
pdf.

Anonymous:1967:IVT

[Ano67b] Anonymous. Index to volume 8 (1967) of Technology and Culture. Technology and Culture, 8(4):549–561, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894430/pdf.

Anonymous:1967:LVM

[Ano67c] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 8(2):310–313, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894483/pdf.

Anonymous:1967:NAa

[Ano67d] Anonymous. Notes and announcements. Technology and Culture, 8(1):156–157, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894529/pdf.

Anonymous:1967:NAb

[Ano67e] Anonymous. Notes and announcements. Technology and Culture, 8(3):434–436, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894456/pdf.

Anonymous:1967:NAc

[Ano67f] Anonymous. Notes and announcements. Technology and Culture, 8(4):543–548, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894429/pdf.

Anonymous:1967:VNJa

[Ano67g] Anonymous. Volume 8, number 1, January 1967. Technology and Culture, 8(1):??, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1967:VNA

[Ano67h] Anonymous. Volume 8, number 2, April 1967. Technology and Culture, 8(2):??, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1967:VNJb

[Ano67i] Anonymous. Volume 8, number 3, July 1967. Technology and Culture, 8(3):??, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1967:VNO

[Ano67j] Anonymous. Volume 8, number 4, October 1967. Technology and Culture, 8(4):??, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1967:WMD

[Ano67k] Anonymous. The Washington Meeting, December 27–30, 1966. Technology and Culture, 8(2):314–331, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894484/pdf.

Anonymous:1968:IVT

[Ano68a] Anonymous. Index to volume 9 (1968) of Technology and Culture. Technology and Culture, 9(4):642–656, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894245/pdf.

Anonymous:1968:ICH

[Ano68b] Anonymous. International Cooperation in History of Technology Committee (ICOHTEC). Technology and Culture, 9(1):142, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894376/pdf.

Anonymous:1968:LVM

[Ano68c] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 9(2):328–331, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894338/pdf.

Anonymous:1968:NAa

[Ano68d] Anonymous. Notes and announcements. Technology and Culture, 9(2):354–355, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894340/pdf.

Anonymous:1968:NAb

[Ano68e] Anonymous. Notes and announcements. Technology and Culture, 9(3):527–530, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894293/pdf.

Anonymous:1968:NAc

[Ano68f] Anonymous. Notes and announcements. Technology and Culture, 9(4):637–641, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894244/pdf.

Anonymous:1968:TMD

[Ano68g] Anonymous. The Toronto Meeting December 27–29, 1967. Technology and Culture, 9(2):332–353, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894339/pdf.

Anonymous:1968:VNJa

[Ano68h] Anonymous. Volume 9, number 1, January 1968. Technology and Culture, 9(1):??, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1968:VNA

[Ano68i] Anonymous. Volume 9, number 2, April 1968. Technology and Culture, 9(2):??, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1968:VNJb

[Ano68j] Anonymous. Volume 9, number 3, July 1968. Technology and Culture, 9(3):??, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1968:VNO

[Ano68k] Anonymous. Volume 9, number 4, October 1968. Technology and Culture, 9(4):??, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1969:APU

[Ano69a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 10(2):187–188, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892360/pdf.

Anonymous:1969:DMD

[Ano69b] Anonymous. The Dallas Meeting, December 26–29, 1968. Technology and Culture, 10(2):196–227, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892363/pdf.

Anonymous:1969:DP

[Ano69c] Anonymous. The Dexter prize. Technology and Culture, 10(2):189– 190, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892361/pdf.

Anonymous:1969:IVT

[Ano69d] Anonymous. Index to volume 10 (1969) of Technology and Culture. Technology and Culture, 10(4):631–644, October 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892422/pdf.

Anonymous:1969:LVM

[Ano69e] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 10(2):191–195, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892362/pdf.

Anonymous:1969:NAa

[Ano69f] Anonymous. Notes and announcements. Technology and Culture, 10(3):487–488, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901647/pdf.

Anonymous:1969:NAb

[Ano69g] Anonymous. Notes and announcements. Technology and Culture, 10(4):625–630, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892421/pdf.

Anonymous:1969:RR

[Ano69h] Anonymous. Rejoinder and refutation. Technology and Culture, 10(2):183–184, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892358/pdf.

Anonymous:1969:VNJa

[Ano69i] Anonymous. Volume 10, number 1, January 1969. Technology and Culture, 10(1):??, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1969:VNA

[Ano69j] Anonymous. Volume 10, number 2, April 1969. Technology and Culture, 10(2):??, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1969:VNJb

[Ano69k] Anonymous. Volume 10, number 3, July 1969. Technology and Culture, 10(3):??, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1969:VNO

[Ano69l] Anonymous. Volume 10, number 4, October 1969. Technology and Culture, 10(4):??, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Armstrong:1962:CMM

[Arm62] George Armstrong. Capital in Manufacturing and Mining: Its Formation and Financing by Daniel Creamer, et al. (review). Technology and Culture, 3(2):197–199, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895392/pdf.

Armstrong:1967:IER

[Arm67] Frederick H. Armstrong. The Intercity Electric Railway Industry in Canada by John F. Due (review). Technology and Culture, 8 (4):508–510, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894412/pdf.

Arnott:1965:CCE

[Arn65] Margaret L. Arnott. Continuities in Cultural Evolution by Margaret Mead (review). Technology and Culture, 6(2):319–320, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894769/pdf.

Arnott:1969:TTE

[Arn69] Margaret L. Arnott. Tools and Tillage ed. by Axel Steensberg, et al. (review). Technology and Culture, 10(2):344–346, April 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892390/pdf.

Austin:1963:HBS

[Aus63a] J. B. Austin. A History of the British Steel Industry by J. C. Carr and W. Taplin (review). Technology and Culture, 4(3):352–353, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894984/
pdf.

Austin:1963:CAS

[Aus63b] James B. Austin. The Coming of the Age of Steel by Theodore A. Wertime (review). Technology and Culture, 4(1):71–72, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895042/pdf.

Austin:1965:ISN

[Aus65] James B. Austin. Iron and Steel in Nineteenth Century America: an Economic Inquiry by Peter Temin (review). Technology and Culture, 6(3):471–472, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894695/pdf.

Austin:1968:MIA

[Aus68] James B. Austin. Made of Iron by the Art Department of St. Thomas University (review). Technology and Culture, 9(2):223–224, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894304/pdf.

Ayres:1967:TDE

[Ayr67] C. E. Ayres. The Theory and Design of Economic Development ed. by Irma Adelman, Erik Thorbecke (review). Technology and Culture, 8(4):521–523, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894420/pdf.

Bachrach:1969:OAC

[Bac69] Bernard S. Bachrach. The origin of Armorican chivalry. Technology and Culture, 10(2):166–171, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892352/pdf.

Baker:1965:RAC

[Bak65] A. Albert Baker, Jr. The Rise of the American Chemistry Profession, 1850–1900 by Edward H. Beardsley (review). Technology and Culture, 6(4):659–660, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894653/pdf.

Baker:1966:WSP

[Bak66] Elizabeth Faulkner Baker. Women and the Scientific Professions ed. by Jacquelyn A. Mattfeld and Carol G. Van Aken (review). Technology and Culture, 7(4):571–574, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894570/pdf.

Baker:1967:WSE

[Bak67] Elizabeth Faulkner Baker. Women in the Soviet Economy, Their Role in Economic, Scientific, and Technological Developments by Norton T. Dodge (review). Technology and Culture, 8(2):242–245, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894475/pdf.

Barber:1961:SCI

[Bar61a] Bernard Barber. Social Change in the Industrial Revolution: an Application of Theory to the British Cotton Industry, 1770–1840 by Neil J. Smelser (review). Technology and Culture, 2(1):57–59, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895585/pdf.

Barber:1961:SAE

[Bar61b] Bernard Barber. Sociological Aspects of Economic Growth by Bert F. Hoselitz (review). Technology and Culture, 2(3):264–266, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895468/pdf.

Barkin:1962:RWM

[Bar62] Solomon Barkin. Real Wages in Manufacturing, 1890–1914 by Albert Rees (review). Technology and Culture, 3(3):343–346, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895148/pdf.

Baranson:1963:ESC

[Bar63a] Jack Baranson. Economic and social considerations in adapting technologies for developing countries. Technology and Culture, 4(1):22–29, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895030/
pdf.

Barnes:1963:BEG

[Bar63b] Donald Grove Barnes. British Economic Growth 1688–1959 by Phyllis Deane and W. A. Cole (review). Technology and Culture, 4 (3):350–352, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894983/pdf.

Barnes:1963:HBU

[Bar63c] Donald Grove Barnes. Heavens Below: Utopian Experiments in England, 1560–1960 by W. H. G. Armytage (review). Technology and Culture, 4(1):69–71, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895041/pdf.

Barnes:1964:BER

[Bar64] Donald Grove Barnes. The British Economy by Roy Harrod (review). Technology and Culture, 5(3):462–464, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894880/pdf.

Baranson:1965:DUC

[Bar65a] Jack Baranson. Development and Underdevelopment by Celso Furtado (review). Technology and Culture, 6(3):510–512, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894714/pdf.

Barnes:1965:CEH

[Bar65b] Donald Grove Barnes. A Concise Economic History of Britain: From the Earliest Times to 1750 by Sir John Clapham, and: A Concise Economic History of Britain: From 1750 to Recent Times by W. H. B. Court (review). Technology and Culture, 6(3):455–457, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894687/pdf.

Barber:1966:GCS

[Bar66a] Bernard Barber. A Guide to Case Studies of Scientific Activity by Ben-Ami Lipetz (review). Technology and Culture, 7(3):440–441, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894604/
pdf.

Barber:1966:SSE

[Bar66b] Bernard Barber. Science and Society ed. by Norman Kaplan (review). Technology and Culture, 7(4):570–571, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894569/pdf.

Barnes:1966:BEH

[Bar66c] Donald Grove Barnes. British Economic History, 1870–1914: Commentary and Documents by W. H. B. Court (review). Technology and Culture, 7(4):530–532, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894549/pdf.

Barnes:1966:RIS

[Bar66d] Donald Grove Barnes. The Rise of Industrial Society in England, 1815–1885 by Sidney G. Checkland (review). Technology and Culture, 7(1):84–86, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893866/pdf.

Baranson:1967:TAT

[Bar67a] Jack Baranson. Technical Assistance — Theory and Guidelines by Sidney C. Sufrin (review). Technology and Culture, 8(4):519–520, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894419/
pdf.

Barber:1967:DAD

[Bar67b] Bernard Barber. Drug Adulteration: Detection and Control in Nineteenth-Century Britain by Ernst W. Stieb (review). Technology and Culture, 8(3):431–432, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894454/pdf.

Barber:1967:SFS

[Bar67c] Bernard Barber. The Social Function of Science by J. D. Bernal (review). Technology and Culture, 8(4):532–533, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894423/pdf.

Baranson:1969:EPT

[Bar69a] Jack Baranson. Economic Policies toward Less Developed Countries by Harry G. Johnson (review). Technology and Culture, 10 (1):133–135, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892347/pdf.

Baranson:1969:ACD

[Bar69b] Jack Baranson. The American Challenge (“Le Défi Américain”) by J.- J. Servan-Schreiber (review). Technology and Culture, 10(3):484–486, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/901646/
pdf.

Battison:1962:DCI

[Bat62] Edwin A. Battison. Decorative Cast Ironwork in Great Britain by Raymond Lister (review). Technology and Culture, 3(3):316–317, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895135/
pdf.

Battison:1966:KSR

[Bat66a] Edwin A. Battison. Die Kulturgeschichte der Schraube by Rudolf Kellermann and Wilhelm Treue (review). Technology and Culture, 7(1):76, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893860/pdf.

Battison:1966:KGT

[Bat66b] Edwin A. Battison. Knaurs Geschichte der Technik by Carl Graf von Klinckowstroem (review). Technology and Culture, 7(2):229–230, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895179/
pdf.

Battison:1966:SCP

[Bat66c] Edwin A. Battison. Stone-Cutting and Polishing Lathe, by Jacques Besson. Technology and Culture, 7(2):202–205, Spring 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895174/pdf.

Battison:1967:BSF

[Bat67] Edwin A. Battison. The British Soldier’s Firearm, 1850–1864 by C. H. Roads (review). Technology and Culture, 8(2):223–224, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894468/pdf.

Baumgarten:1964:SRB

[Bau64] Karl Baumgarten. Studies of rural buildings in Mecklenburg. Technology and Culture, 5(2):234–240, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894895/pdf.

Bush-Brown:1962:ABA

[BB62] Albert Bush-Brown. American Building Art: The Twentieth Century by Carl W. Condit (review). Technology and Culture, 3(2):190–192, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895389/
pdf.

Boyer:1966:LT

[BB66] Carl B. Boyer and Marjorie N. Boyer. Lynn Thorndike (1882– 1965). Technology and Culture, 7(3):391–394, Summer 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894584/pdf.

Beatty:1966:LTE

[Bea66] William K. Beatty. The Life and Times of Emil H. Grubbe by Paul C. Hodges (review). Technology and Culture, 7(2):241–242, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895186/pdf.

Bedini:1961:SRD

[Bed61a] Silvio A. Bedini. Shop Records of Daniel Burnap, Clockmaker by Penrose R. Hoopes (review). Technology and Culture, 2(3):280–282, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895477/
pdf.

Bedini:1961:USC

[Bed61b] Silvio A. Bedini. United States Clock and Watch Patents, 1790– 1890 by George H. Eckhardt (review). Technology and Culture, 2 (3):278–280, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895476/pdf.

Bedini:1962:CCC

[Bed62] Silvio A. Bedini. The compartmented cylindrical Clepsydra. Technology and Culture, 3(2):115–141, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895376/pdf.

Bedini:1964:KFC

[Bed64a] Silvio A. Bedini. The Knibb Family — Clockmakers by Ronald A. Lee (review). Technology and Culture, 5(4):608–610, Fall 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894829/pdf.

Bedini:1964:MSI

[Bed64b] Silvio A. Bedini. The Makers of Surveying Instruments in America since 1700 by Charles E. Smart (review). Technology and Culture, 5(2):260–261, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894902/pdf.

Bedini:1964:PMC

[Bed64c] Silvio A. Bedini. The Peabody Museum Collection of Navigating Instruments, with Notes on Their Makers by Marion V. Brewington (review). Technology and Culture, 5(4):640–641, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894845/pdf.

Bedini:1964:RAH

[Bed64d] Silvio A. Bedini. The role of automata in the history of technology. Technology and Culture, 5(1):24–42, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894920/pdf.

Bedini:1965:BOC

[Bed65a] Silvio A. Bedini. The Book of Old Clocks and Watches by Ernst von Bassermann-Jordan, and: The Collector’s Dictionary of Clocks by H. Alan Lloyd (review). Technology and Culture, 6(3):451–454, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894685/
pdf.

Bedini:1965:ESM

[Bed65b] Silvio A. Bedini. The evolution of science museums. Technology and Culture, 6(1):1–29, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894776/pdf.

Bedini:1965:RLL

[Bed65c] Silvio A. Bedini. A Renaissance lapidary lathe. Technology and Culture, 6(3):407–415, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894675/pdf.

Bedini:1967:AT

[Bed67a] Silvio A. Bedini. The aerial telescope. Technology and Culture, 8(3):395–401, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894437/pdf.

Bedini:1967:WCC

[Bed67b] Silvio A. Bedini. Watches by Cecil Clutton, George Daniels (review). Technology and Culture, 8(1):94–95, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894498/pdf.

Bedini:1967:EOL

[Bed67c] Silvio A. Bedini. An early optical lens-grinding lathe. Technology and Culture, 8(1):74–77, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894491/pdf.

Bedini:1968:HAM

[Bed68a] Silvio A. Bedini. Historical Aspects of Microscopy ed. by S. Bradbury, G. L’E. Turner (review). Technology and Culture, 9(4):603–604, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894224/
pdf.

Bedini:1968:PMW

[Bed68b] Silvio A. Bedini. The perspective machine of Wentzel jamnitzer. Technology and Culture, 9(2):197–202, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894299/pdf.

Bedini:1969:UUQ

[Bed69] Silvio A. Bedini. The unfinished Utrecht quadrant. Technology and Culture, 10(3):432–435, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/901625/pdf.

Beer:1965:HRS

[Bee65] John J. Beer. The historical relations of science and technology. Technology and Culture, 6(4):547–552, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894637/pdf.

Beer:1966:SRH

[Bee66] John J. Beer. Selected Readings in the History of Chemistry comp. by Aaron J. Ihde and William F. Kieffer (review). Technology and Culture, 7(4):535–538, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894552/pdf.

Bennett:1962:HMC

[Ben62a] John W. Bennett. The History of Modern Culture by Maurice Parmelee (review). Technology and Culture, 3(2):227–228, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895406/pdf.

Benton:1962:F

[Ben62b] William Benton. Foreword. Technology and Culture, 3(4):381, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895096/pdf.

Bennett:1963:CES

[Ben63] John W. Bennett. Custom: an Essay on Social Codes by Ferdinand Tönnies (review). Technology and Culture, 4(1):115, Winter 1963.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895059/pdf.

Benfey:1966:DMC

[Ben66] O. T. Benfey. The Development of Modern Chemistry by Aaron J. Ihde (review). Technology and Culture, 7(2):235–237, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895183/pdf.

Bennett:1968:BRA

[Ben68] John W. Bennett. Biennial Review of Anthropology 1967 ed. by Bernard J. Siegel and Alan R. Beals (review). Technology and Culture, 9(3):518–520, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894288/pdf.

Bernard:1965:TWW

[Ber65a] Jessie Bernard. Technology and Woman’s Work by Elizabeth Faulkner Baker (review). Technology and Culture, 6(3):521–523, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894720/pdf.

Bernstein:1965:MCS

[Ber65b] Herbert M. Bernstein. The Mass Consumption Society by George Katona (review). Technology and Culture, 6(3):525–527, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894722/pdf.

Bernstein:1967:EHE

[Ber67] Herbert M. Bernstein. Essays in the History of Economics by George J. Stigler (review). Technology and Culture, 8(1):136–138, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894521/pdf.

Biggs:1964:STH

[Big64] Robert D. Biggs. The Sumerians: Their History, Culture, and Character by Samuel Noah Kramer (review). Technology and Culture, 5(3):445–446, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894869/pdf.

Billington:1968:HCC

[Bil68] William A. Billington. A History of the Cornish China-Clay Industry by R. M. Barton (review). Technology and Culture, 9(1):118–120, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894361/
pdf.

Bilstein:1969:TCA

[Bil69] Roger E. Bilstein. Technology and commerce: Aviation in the conduct of American business 1918–29. Technology and Culture, 10(3):392– 411, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901620/pdf.

Birr:1962:SAI

[Bir62] Kendall Birr. The Scientist in American Industry: Some Organizational Determinants in Manpower Utilization by Simon Marcson (review). Technology and Culture, 3(1):101–102, Winter 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895423/pdf.

Birr:1966:RAE

[Bir66] Kendall A. Birr. The Roots of American Economic Growth, 1607– 1861: an Essay in Social Causation by Stuart Bruchey (review). Technology and Culture, 7(1):103–104, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893876/pdf.

Biswas:1969:EAE

[Bis69] Asit K. Biswas. Experiments on atmospheric evaporation until the end of the Eighteenth Century. Technology and Culture, 10(1):49–58, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892314/
pdf.

Blair:1962:CAJ

[Bla62] C. Blair. Catalogue of Armor: The John Woodman Higgins Armory by Stephen V. Grancsay (review). Technology and Culture, 3(2):182–183, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895383/pdf.

Black:1963:SMN

[Bla63] Cyril E. Black. Soviet Marxism and Natural Science, 1917–1932 by David Joravsky (review). Technology and Culture, 4(1):123–124, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895063/
pdf.

Black:1964:SID

[Bla64] Cyril E. Black. The Soviet Industrialization Debate, 1924–1928 by Alexander Erlich (review). Technology and Culture, 5(3):461–462, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894879/
pdf.

Bleich:1961:CPC

[Ble61] Edmund M. Bleich. Central Planning in Czechoslovakia: Organization for Growth in a Mature Economy by Jan M. Michal (review). Technology and Culture, 2(4):426–428, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895451/pdf.

Bleich:1963:IGP

[Ble63] Edmund M. Bleich. Investment and Growth Policies in British Industrial Firms by Tibor Barna (review). Technology and Culture, 4(3):354–355, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894985/pdf.

Boehm:1962:PMP

[Boe62] Dwight Boehm. The Poet and the Machine by Paul Ginestier (review). Technology and Culture, 3(2):230–234, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895408/pdf.

Boelter:1963:WRU

[Boe63] L. M. K. Boelter. The World Role of Universities by Edward W. Weidner (review). Technology and Culture, 4(2):239–241, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895019/pdf.

Bonelli:1969:RIM

[Bon69] Maria Luisa Righini Bonelli. Rehabilitation of the istituto e museo di storia della scienza in Florence. Technology and Culture, 10(1):62–64, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892316/
pdf.

Borgstrom:1965:FST

[Bor65a] Georg Borgstrom. Food Science and Technology by Magnus Pyke (review). Technology and Culture, 6(4):665–666, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894656/pdf.

Borgstrom:1965:MME

[Bor65b] Georg Borgstrom. Man Must Eat by Sir William Slater (review). Technology and Culture, 6(3):494–496, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894706/pdf.

Boulding:1961:CWJ

[Bou61] Kenneth E. Boulding. The Causes of Wealth by Jean Fourastié (review). Technology and Culture, 2(3):262–264, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895467/pdf.

Boulding:1966:RMP

[Bou66] Kenneth E. Boulding. The role of the museum in the propagation of developed images. Technology and Culture, 7(1):64–66, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893853/pdf.

Boulding:1967:IDC

[Bou67] Kenneth E. Boulding. The Institutional Drive by Carl G. Gustavson (review). Technology and Culture, 8(4):534–535, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894424/pdf.

Bowles:1966:OKM

[Bow66] Edmund A. Bowles. On the origin of the keyboard mechanism in the Late Middle Ages. Technology and Culture, 7(2):152–162, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895170/pdf.

Boyer:1960:MPA

[Boy60] Marjorie Nice Boyer. Medieval pivoted axles. Technology and Culture, 1(2):128–138, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895248/pdf.

Boyd:1964:CFK

[Boy64] T. A. Boyd. The Charles F. Kettering Archives. Technology and Culture, 5(3):412–415, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894862/pdf.

Boyer:1966:RBA

[Boy66] Marjorie N. Boyer. Rebuilding the bridge at albi, 1408–1410. Technology and Culture, 7(1):24–37, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893850/pdf.

Boyer:1967:PPM

[Boy67] Ronald K. Boyer. Personality and Productivity in Management by James H. Mullen (review). Technology and Culture, 8(2):255–257, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894481/
pdf.

Boyd:1968:SS

[Boy68] T. A. Boyd. The self-starter. Technology and Culture, 9(4):585–591, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894218/
pdf.

Boyle:1969:EMG

[Boy69] Leonard E. Boyle. English Medieval Graffiti by V. Pritchard (review). Technology and Culture, 10(4):607, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892410/pdf.

Brauer:1965:NEI

[Bra65a] Kinley J. Brauer. The New Empire: an Interpretation of American Expansion, 1860–1898 by Walter LaFeber (review). Technology and Culture, 6(4):657–659, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894652/pdf.

Braugh:1965:HBN

[Bra65b] Frank C. Braugh. History and Bibliography of The New American Practical Navigator and The American Coast Pilot by John F. Campbell (review). Technology and Culture, 6(1):131–132, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894794/pdf.

Brant:1966:PAH

[Bra66] Irving Brant. The Papers of Alexander Hamilton, Vol. VIII: February 1791–July 1791; Vol. IX: July 1791–December 1791 ed. by Harold C. Syrett and Jacob E. Cooke (review). Technology and Culture, 7(4): 539–542, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894554/pdf.

Bramwell:1969:HAC

[Bra69] A. R. S. Bramwell. History of Aviation and Cosmonautics. Vol. V (Iz istorii aviatsii i kosmonautiki, Vyp. 5) ed. by N. D. Anoshchenko (review). Technology and Culture, 10(3):468–470, July 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901640/pdf.

Brittain:1961:RHA

[Bri61] Robert Brittain. Readings in the History of American Agriculture ed. by Wayne D. Rasmussen (review). Technology and Culture, 2 (2):183–185, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895557/pdf.

Brightfield:1962:CRE

[Bri62a] Myron F. Brightfield. The coming of the railroad to Early Victorian England, as viewed by novels of the period (1840–1870). Technology and Culture, 3(1):45–72, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895413/pdf.

Brittain:1962:FDS

[Bri62b] Robert Brittain. The Farmer’s Dilemma by Stanley Andrews (review). Technology and Culture, 3(3):353–355, Summer 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895152/pdf.

Brittain:1963:WTL

[Bri63] Robert Brittain. Where There Is Life by Paul B. Sears (review). Technology and Culture, 4(1):113–114, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895058/pdf.

Brinton:1965:MHE

[Bri65] Crane Brinton. The Meaning of History by Erich Kahler (review). Technology and Culture, 6(3):535–536, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894726/pdf.

Brinton:1968:CHC

[Bri68] Crane Brinton. Change and Habit: The Challenge of Our Time by Arnold J. Toynbee (review). Technology and Culture, 9(1):102–103, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894353/
pdf.

Brown:1962:TCC

[Bro62] Robert McAfee Brown. Technology and Christian Culture ed. by Robert Paul Mohan (review). Technology and Culture, 3(3):371–372, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895160/
pdf.

Brown:1965:STC

[Bro65a] John J. Brown. A survey of technology in Canadian museums. Technology and Culture, 6(1):83–98, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894781/pdf.

Brown:1965:GGI

[Bro65b] Robert McAfee Brown. God & Golem, Inc. by Norbert Wiener (review). Technology and Culture, 6(3):531–532, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894724/pdf.

Brugmans:1968:IWM

[Bru68] H. Brugmans. Impact of Western Man. A Study of Europe’s Role in the World Economy 1750–1960 by William Woodruff (review). Technology and Culture, 9(1):98–100, January 1968. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894351/pdf.

Bryant:1964:GDV

[Bry64a] Lynwood Bryant. Geschichte des deutschen Verbrennungsmotorenbaues von 1860 bis 1918 by Friedrich Sass (review). Technology and Culture, 5(1):82–86, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894928/pdf.

Bryant:1964:NEP

[Bry64b] Paul T. Bryant. The New Exploration: a Philosophy of Regional Planning by Benton MacKaye (review). Technology and Culture, 5 (3):469–471, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894883/pdf.

Bryant:1966:KWK

[Bry66a] Lynwood Bryant. Kraft für die Welt: 1864–1964; Klöckner– Humboldt–Deutz AG by Gustav Goldbeck (review). Technology and Culture, 7(3):429–432, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894599/pdf.

Bryant:1966:RDP

[Bry66b] Lynwood Bryant. Rudolf Diesel: Pioneer of the Age of Power by W. Robert Nitske and Charles Morrow Wilson (review). Technology and Culture, 7(1):90–93, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893869/pdf.

Bryant:1966:SO

[Bry66c] Lynwood Bryant. The silent Otto. Technology and Culture, 7(2):184– 200, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895172/pdf.

Bryant:1967:GKG

[Bry67a] Lynwood Bryant. Gebändigte Kraft: Die Geschichte der Erfindung des Otto-motors by Gustav Goldbeck (review). Technology and Culture, 8(1):105–107, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894504/pdf.

Bryant:1967:OFS

[Bry67b] Lynwood Bryant. The origin of the four-stroke cycle. Technology and Culture, 8(2):178–198, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894458/pdf.

Bryant:1968:ANA

[Bry68a] Lynwood Bryant. Ainsi naquit l’automobile by Jacques Ickx (review). Technology and Culture, 9(2):237–239, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894313/pdf.

Bryant:1968:CRU

[Bry68b] Lynwood Bryant. Christian Reithmann: Uhrmacher und Motorenerfinder, 1818–1909 by Gustav Goldbeck (review). Technology and Culture, 9(3):489–491, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894271/pdf.

Bryant:1969:RZH

[Bry69] Lynwood Bryant. Die Rivalität zwischen Heissluftmaschine und Verbrennungsmotor als Kleingewerbemaschinen zwischen 1860 und 1890. Der Sieg des Verbrennungsmotors und seine Gründe by Kurt Mauel (review). Technology and Culture, 10(1):104–106, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892330/pdf.

Bramwell:1969:RRH

[BT69] A. R. S. Bramwell and G. A. Tokaty. Russian Rocketry: a Historical Survey by I. A. Slukhai (review). Technology and Culture, 10 (1):124–125, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892340/pdf.

Buchanan:1962:TSE

[Buc62] Scott Buchanan. Technology as a system of exploitation. Technology and Culture, 3(4):535–543, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895110/pdf.

Buchanan:1968:TIA

[Buc68a] R. A. Buchanan. Techniques of Industrial Archaeology by J. P. M. Pannell (review). Technology and Culture, 9(1):109–111, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894357/pdf.

Buchanan:1968:TWC

[Buc68b] R. A. Buchanan. Technology in Western Civilization ed. by Melvin Kranzberg and Carroll W. Pursell, Jr. (review). Technology and Culture, 9(3):468–476, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894265/pdf.

Buchanan:1968:CSH

[Buc68c] R. A. Buchanan. The Centre for the Study of the History of Technology, Bath University of Technology, England. Technology and Culture, 9(3):430–435, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894260/pdf.

Bullough:1962:AGW

[Bul62] Vern L. Bullough. Attic Geometric Workshops by Jean M. Davidson (review). Technology and Culture, 3(2):181–182, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895382/pdf.

Bullough:1966:TPA

[Bul66] Vern L. Bullough. The Techniques of Painted Attic Pottery by Joseph Veach Noble (review). Technology and Culture, 7(4):521–522, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894542/pdf.

Bunge:1966:FMP

[Bun66a] Mario Bunge. I filosofi e le macchine (1400–1700) by Paolo Rossi (review). Technology and Culture, 7(1):73–74, Winter 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL http://www.jstor.org/stable/3101607;   https://muse.
jhu.edu/pub/1/article/893858/pdf.

Bunge:1966:TAS

[Bun66b] Mario Bunge. Technology as applied science. Technology and Culture, 7(3):329–347, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894579/pdf.

Burgess:1960:RCS

[Bur60a] E. Martin Burgess. A reply to cyril Stanley Smith on Mail making methods. Technology and Culture, 1(2):151–155, Spring 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895250/pdf.

Burstall:1960:M

[Bur60b] Aubrey F. Burstall. Machinery. Technology and Culture, 1(4):401– 407, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895620/
pdf.

Burger:1961:IET

[Bur61a] Henry G. Burger. Industrial Estates: Tool for Industrialization by William Bredo (review). Technology and Culture, 2(3):295–298, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895485/
pdf.

Burger:1961:T

[Bur61b] Henry G. Burger. “Technoculture”. Technology and Culture, 2 (3):260–261, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895466/pdf.

Burlingame:1961:TNC

[Bur61c] Roger Burlingame. Technology: Neglected clue to historical change. Technology and Culture, 2(3):219–229, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895458/pdf.

Burstall:1961:DET

[Bur61d] Aubrey F. Burstall. The Development of the English Traction Engine by Ronald H. Clark (review). Technology and Culture, 2(3):288–289, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895481/
pdf.

Burchard:1962:AEP

[Bur62a] John E. Burchard. Architecture and the Esthetics of Plenty by James Marston Fitch (review). Technology and Culture, 3(3):327–332, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895141/
pdf.

Burger:1962:MTC

[Bur62b] David H. Burger. Man Takes Control: Cultural Development and American Aid by Charles J. Erasmus (review). Technology and Culture, 3(2):219–222, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895402/pdf.

Burke:1963:PPP

[Bur63] John G. Burke. Petroleum Progress and Profits: a History of Process Innovation by John Lawrence Enos (review). Technology and Culture, 4(1):82–84, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895047/pdf.

Burchard:1964:BEC

[Bur64a] John E. Burchard. Babylon is Everywhere: The City as Man’s Fate by Wolf Schneider (review). Technology and Culture, 5(3):472–473, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894884/
pdf.

Burchard:1964:EPB

[Bur64b] John E. Burchard. The Eternal Present: The Beginnings of Art by S. Giedion (review). Technology and Culture, 5(1):120–122, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894946/pdf.

Burchard:1965:BAV

[Bur65a] John E. Burchard. The Beginnings of Architecture. Vol. II of The Eternal Present by Sigfried Giedion (review). Technology and Culture, 6(2):267–269, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894741/pdf.

Burger:1965:WDS

[Bur65b] David H. Burger. World Design Science Decade, 1965–1975: Phase I (1963) Document 1, Inventory of World Resources, Human Trends, and Needs; Phase I (1964) Document 2, The Design Initiative by R. Buckminster Fuller and John McHale (review). Technology and Culture, 6(3):504–506, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894711/pdf.

Burnham:1965:BBH

[Bur65c] Alan Burnham. Bulfinch’s Boston, 1787–1817 by Harold Kirker and James Kirker (review). Technology and Culture, 6(3):468–469, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894693/
pdf.

Burstall:1965:BLG

[Bur65d] Aubrey F. Burstall. Bulleid: Last Giant of Steam by Sean Day Lewis (review). Technology and Culture, 6(3):457–459, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894688/pdf.

Burstall:1965:EIW

[Bur65e] Aubrey F. Burstall. Engineers, Inventors and Workers by P. W. Kingsford (review). Technology and Culture, 6(2):276–277, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894745/pdf.

Burger:1966:WDS

[Bur66a] David H. Burger. World Design Science Decade, 1965–1975: Phase 1 (1965), Document III: Comprehensive Thinking; Phase 1 (1965); Document IV: The Ten Year Program by R. Buckminster Fuller and John McHale (review). Technology and Culture, 7(2):262–264, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895196/pdf.

Burke:1966:BBF

[Bur66b] John G. Burke. Bursting boilers and the Federal power. Technology and Culture, 7(1):1–23, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893849/pdf.

Burstall:1966:GLR

[Bur66c] Aubrey F. Burstall. Garretts of Leiston by R. A. Whitehead (review). Technology and Culture, 7(1):88–90, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893868/pdf.

Burstall:1966:TCN

[Bur66d] Aubrey F. Burstall. Tyneside Classical, the Newcastle of Grainger, Dobson and Clayton by Lyall Wilkes and Gordon Dodds (review). Technology and Culture, 7(2):242–244, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895187/pdf.

Burke:1967:SCS

[Bur67] John G. Burke. The Sorby Centennial Symposium on the History of Metallurgy ed. by Cyril Stanley Smith (review). Technology and Culture, 8(1):82–86, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894493/pdf.

Burke:1968:ERF

[Bur68] John G. Burke. Early Research at the Franklin Institute: The Investigation into the Causes of Steam Boiler Explosions: 1830–1837 by Bruce Sinclair (review). Technology and Culture, 9(2):230–232, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894309/pdf.

Burke:1969:SHS

[Bur69a] John G. Burke. Sources for the History of the Science of Steel, 1532– 1786 ed. by Cyril Stanley Smith (review). Technology and Culture, 10(1):94–96, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892324/pdf.

Burke:1969:CLS

[Bur69b] John G. Burke. Comment: Let’s be sure technology is for man. Technology and Culture, 10(1):11–13, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892309/pdf.

Burstyn:1969:BPU

[Bur69c] Harold L. Burstyn. Benjamin Peirce and the U.S. Coast Survey by Victor F. Lenzen (review). Technology and Culture, 10(3):452–453, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/901632/
pdf.

Cahn:1967:PVB

[Cah67] R. W. Cahn. Pirotechnia by Vannoccio Biringuccio (review). Technology and Culture, 8(1):88–90, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894495/pdf.

Calder:1962:C

[Cal62a] Ritchie Calder. Commentary. Technology and Culture, 3(4):516–518, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895107/
pdf.

Calder:1962:TF

[Cal62b] Ritchie Calder. Technology in focus. Technology and Culture, 3(4): 563–580, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895114/pdf.

Calhoun:1966:MIB

[Cal66] John C. Calhoun, Jr. Men of Ingenuity from beneath the Orange Tower by W. R. Woolrich (review). Technology and Culture, 7(1):111– 112, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893881/pdf.

Calvert:1967:WPI

[Cal67] Monte A. Calvert. Winterthur Portfolio II ed. by Milo M. Naeve (review). Technology and Culture, 8(3):406–407, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894442/pdf.

Calvert:1968:IGL

[Cal68a] H. R. Calvert. Irascible Genius: a Life of Charles Babbage, Inventor by Maboth Moseley (review). Technology and Culture, 9(1):122–123, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL http://www.jstor.org/stable/3102064;
 https://muse.jhu.edu/pub/1/article/894364/pdf.

Calvert:1968:MIE

[Cal68b] Monte A. Calvert. Men and Ideas in Engineering: Twelve Histones from Illinois by R. A. Kingery, et al. (review). Technology and Culture, 9(2):239–240, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894314/pdf.

Calvert:1968:MPP

[Cal68c] Monte A. Calvert. The Mechanicals: Progress of a Profession by L. T. C. Rolt (review). Technology and Culture, 9(4):611–612, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894229/pdf.

Calvert:1968:WWE

[Cal68d] Monte A. Calvert. Workshops in the Wilderness: The European Response to American Industrialization, 1830–1860 by Marvin Fisher (review). Technology and Culture, 9(1):125–126, January 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894366/pdf.

Calvert:1969:ETS

[Cal69a] Monte A. Calvert. Education for a Technological Society: a Sesquicentennial History of Rensselaer Polytechnic Institute by Samuel Rezneck (review). Technology and Culture, 10(2):319–320, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892374/pdf.

Calvert:1969:WPI

[Cal69b] Monte A. Calvert. Winterthur Portfolio III ed. by Milo M. Naeve (review). Technology and Culture, 10(3):448–449, July 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901630/pdf.

Carmichael:1960:HCH

[Car60] Leonard Carmichael. The Human Condition by Hannah Arendt (review). Technology and Culture, 1(2):177–178, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895253/pdf.

Carlson:1961:PAS

[Car61a] Eric T. Carlson. Psychophysiological Aspects of Space Flight ed. by Bernard E. Flaherty (review). Technology and Culture, 2(4):433–434, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895454/
pdf.

Carlson:1961:UPR

[Car61b] Robert E. Carlson. The Union Pacific Railroad: a Case in Premature Enterprise by Robert William Fogel (review). Technology and Culture, 2(4):405–407, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895441/pdf.

Carmichael:1961:CCE

[Car61c] Leonard Carmichael. Creativity and Its Cultivation ed. by Harold H. Anderson (review). Technology and Culture, 2(1):89–92, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895602/pdf.

Carlson:1962:MBA

[Car62a] Eric T. Carlson. Medical and Biological Aspects of the Energies of Space ed. by Paul A. Campbell (review). Technology and Culture, 3(2):214–216, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895400/pdf.

Carlson:1962:RBH

[Car62b] Robert E. Carlson. The Railways of Britain: an Historical Introduction by Jack Simmons (review). Technology and Culture, 3 (3):317–319, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895136/pdf.

Carmichael:1962:RMS

[Car62c] Leonard Carmichael. Ressentiment by Max Scheler (review). Technology and Culture, 3(2):228–230, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895407/pdf.

Cardwell:1965:PTA

[Car65] D. S. L. Cardwell. Power technologies and the advance of science, 1700–1825. Technology and Culture, 6(2):188–207, Spring 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894732/pdf.

Carlson:1967:GRN

[Car67] Robert E. Carlson. Government and the Railways in Nineteenth-Century Britain by Henry Parris (review). Technology and Culture, 8(1):99–102, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894501/pdf.

Carlson:1968:SHM

[Car68] Robert E. Carlson. A Short History of the Midland & Great Northern Joint Railway by Ronald H. Clark (review). Technology and Culture, 9(2):233–235, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894310/pdf.

Careless:1969:CBR

[Car69] J. M. S. Careless. Canada Builds, 1861–1961 by T. Ritchie (review). Technology and Culture, 10(1):113–114, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892334/pdf.

Chase:1964:ICA

[CB64] Edward T. Chase and Emile Benoit. The issues and challenges of air transportation: The impact of new technology. Technology and Culture, 5(1):57–67, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894922/pdf.

Cesari:1965:SCF

[Ces65] Gene S. Cesari. Science as a Cultural Force ed. by Harry Woolf (review). Technology and Culture, 6(4):683–686, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894666/pdf.

Cesari:1967:EQG

[Ces67] Gene S. Cesari. Environmental Quality in a Growing Economy ed. by Henry Jarrett (review). Technology and Culture, 8(4):523–524, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894421/
pdf.

Cesari:1969:PDD

[Ces69a] G. S. Cesari. Progress and Disillusion: The Dialectics of Modern Society by Raymond Aron (review). Technology and Culture, 10(3):478– 480, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901643/pdf.

Cesari:1969:IEG

[Ces69b] Gene S. Cesari. Invention and Economic Growth by Jacob Schmookler (review). Technology and Culture, 10(1):127–128, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892342/pdf.

Chang:1961:DIS

[Cha61] Tien S. Chang. The Development of Iron and Steel Technology in China by Joseph Needham (review). Technology and Culture, 2(2):170–171, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895551/pdf.

Chapelle:1964:RDP

[Cha64] Howard I. Chapelle. The Rise and Decline of the Paddle-Wheel by Greville Bathe (review). Technology and Culture, 5(4):611–612, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894831/pdf.

Chandler:1966:EEE

[Cha66] Alfred D. Chandler, Jr. Explorations in Enterprise ed. by Hugh G. J. Aitken (review). Technology and Culture, 7(3):434–436, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894601/pdf.

Chapin:1968:HPL

[Cha68] Seymour L. Chapin. Histoire et prestige de l’Académie des Sciences, 1666–1966 by the Musée du Conservatoire National des Arts et Métiers (review). Technology and Culture, 9(1):126–128, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894367/pdf.

Chapelle:1969:RHT

[Cha69a] Howard I. Chapelle. Rushton and His Times in American Canoeing by Atwood Manley, Paul F. Jamieson (review). Technology and Culture, 10(1):107–109, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892331/pdf.

Chapin:1969:RSR

[Cha69b] Seymour L. Chapin. Raspail, Scientist and Reformer by Dora B. Weiner (review). Technology and Culture, 10(3):457–459, July 1969.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901635/pdf.

Chapin:1969:TSN

[Cha69c] Seymour L. Chapin. Trackers of the Skies by E. Nelson Hayes (review). Technology and Culture, 10(4):620–622, October 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892418/pdf.

Chenault:1961:BCM

[Che61a] Lawrence R. Chenault. Blue Collar Man: Patterns of Dual Allegiance in Industry by Theodore V. Purcell (review). Technology and Culture, 2(4):411–414, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895445/pdf.

Chenault:1961:SCF

[Che61b] Lawrence R. Chenault. Studies in Company Finance: a Symposium on the Economic Analysis and Interpretation of British Company Accounts ed. by Brian Tew, R. F. Henderson (review). Technology and Culture, 2(1):77–79, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895595/pdf.

Childs:1963:DTA

[Chi63] Donald S. Childs, Jr. Diagnosis and Treatment of Acute Radiation Injury (review). Technology and Culture, 4(1):107–108, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895055/pdf.

Chilton:1965:HED

[Chi65] D. Chilton. A History of Engineering Drawing by P. J. Booker (review). Technology and Culture, 6(1):128–130, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894792/pdf.

Christian:1961:MOC

[Chr61] W. C. Christian. Management Organization and the Computer ed. by George P. Shultz, Thomas L. Whisler (review). Technology and Culture, 2(2):193–196, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895562/pdf.

Christian:1962:AOI

[Chr62] W. C. Christian. Automation in the Office by Ida Russakoff Hoos (review). Technology and Culture, 3(2):207–209, Spring 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895396/pdf.

Clarke:1962:TMC

[Cla62] W. Norris Clarke, S.J. Technology and man: a Christian vision. Technology and Culture, 3(4):422–442, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895100/pdf.

Clagett:1966:PMH

[Cla66] Marshall Clagett. Perspectives in Medieval History ed. by K. F. Drew and F. S. Lear (review). Technology and Culture, 7(2):225–228, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895177/pdf.

Clemen:1961:RAW

[Cle61] Rudolf A. Clemen. Rings Around the World: Man’s Progress from Steam Engine to Satellite by Heinz Gartmann (review). Technology and Culture, 2(1):37–41, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895577/pdf.

Cohen:1969:PIL

[Coh69] Sanford Cohen. The Peruvian Industrial Labor Force by David Chaplin (review). Technology and Culture, 10(1):131–133, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892345/pdf.

Cole:1961:PHF

[Col61] Charles W. Cole. The Persecution of the Huguenots and French Economic Development, 1680–1720 by Warren Candler Scoville (review). Technology and Culture, 2(2):174–176, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895553/pdf.

Cole:1965:ECS

[Col65] Stephen Cole. Essays on Creativity in the Sciences ed. by Myron A. Coler (review). Technology and Culture, 6(1):158–161, Winter 1965.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894812/pdf.

Cole:1966:CEH

[Col66] Charles W. Cole. The Cambridge Economic History of Europe, Vol. VI, Parts I and II: The Industrial Revolutions and After: Population and Technological Change ed. by H. J. Habakkuk and M. Postan (review). Technology and Culture, 7(4):529–530, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894548/pdf.

Condit:1960:PSE

[Con60a] Carl W. Condit. Philosophy of Structures by Eduardo Torroja (review). Technology and Culture, 1(3):280–283, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895644/pdf.

Condit:1960:BTI

[Con60b] Carl W. Condit. The Building of TVA; An Illustrated History by John H. Kyle (review). Technology and Culture, 1(2):189–191, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895259/pdf.

Condit:1960:BCE

[Con60c] Carl W. Condit. Building and civil engineering. Technology and Culture, 1(4):349–359, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895612/pdf.

Condit:1961:OSE

[Con61a] Carl W. Condit. Offices in the Sky by Earle Schultz, Walter Simmons (review). Technology and Culture, 2(1):71–73, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895592/pdf.

Condit:1961:AAS

[Con61b] Carl W. Condit. The Architecture of America: a Social and Cultural History by John Burchard and Albert Bush-Brown (review). Technology and Culture, 2(4):98–101, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895442/pdf.

Condit:1962:HBM

[Con62a] Carl W. Condit. A History of Building Materials by Norman Davey (review). Technology and Culture, 3(3):323–325, Summer 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895139/pdf.

Condit:1962:PCW

[Con62b] Carl W. Condit. Peter Cooper and the Wrought Iron Beam by Esmond Shaw (review). Technology and Culture, 3(3):325–326, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895140/pdf.

Condit:1962:AAS

[Con62c] Carl W. Condit. The Architecture of America: a Social and Cultural History by John Burchard and Albert Bush-Brown (review). Technology and Culture, 3(1):98–101, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895422/pdf.

Condit:1963:EEM

[Con63a] Carl W. Condit. Environmental Engineering and Metropolitan Planning ed. by John A. Logan, et al. (review). Technology and Culture, 4(3):360–362, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894989/pdf.

Condit:1963:MMA

[Con63b] Carl W. Condit. Man-Made America: Chaos or Control? by Christopher Tunnard and Boris Pushkarev (review). Technology and Culture, 4(3):362–365, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894990/pdf.

Condit:1963:SLE

[Con63c] Carl W. Condit. Steam Locomotives ed. by Linn H. Westcott (review). Technology and Culture, 4(1):75–80, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895045/pdf.

Condit:1963:IEW

[Con63d] Carl W. Condit. The Interurban Era by William D. Middleton (review). Technology and Culture, 4(2):223–224, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895012/pdf.

Condit:1963:VRH

[Con63e] Carl W. Condit. The Virginian Railway by H. Reid, and: Canadian Steam ed. by David P. Morgan (review). Technology and Culture, 4(1):80–82, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895046/pdf.

Condit:1964:CAT

[Con64a] Carl W. Condit. Chinese Architecture and Town Planning, 1500 B.C.–A.D. 1911 by Andrew Boyd, and: The Turning Point of Building: Structure and Design by Konrad Wachsmann, and: Structure and Form in Modern Architecture by Curt Siegel (review). Technology and Culture, 5(1):92–96, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894931/pdf.

Condit:1964:SAM

[Con64b] Carl W. Condit. Structure in Architecture by Mario Salvadori and Robert Heller, and: Candela: The Shell Builder by Colin Faber (review). Technology and Culture, 5(4):618–620, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894836/pdf.

Condit:1964:TUT

[Con64c] Carl W. Condit. The Technology of Urban Transportation by Donald S. Berry, et al. (review). Technology and Culture, 5(2):262–263, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894904/
pdf.

Condit:1964:SSA

[Con64d] Carl W. Condit. The structural system of Adler and Sullivan’s Garrick Theater building. Technology and Culture, 5(4):523–540, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894817/pdf.

Condit:1965:SAJ

[Con65a] Carl W. Condit. Shell Architecture by Jürgen Joedicke, and: The People’s Architects ed. by Harry S. Ransom (review). Technology and Culture, 6(3):476–478, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894697/pdf.

Condit:1965:CSR

[Con65b] Carl W. Condit. Comment: Stages in the relationships between science and technology. Technology and Culture, 6(4):587–590, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894640/pdf.

Condit:1966:ATB

[Con66a] Carl W. Condit. Aesthetics and Technology in Building by Pier Luigi Nervi (review). Technology and Culture, 7(3):432–434, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894600/pdf.

Condit:1966:CLB

[Con66b] Carl W. Condit. Cincinnati Locomotive Builders, 1845–1868 by John H. White (review). Technology and Culture, 7(3):428–429, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894598/pdf.

Condit:1966:IAD

[Con66c] Carl W. Condit. Islamic Architecture and Its Decoration by Derek Hill and Oleg Grabar (review). Technology and Culture, 7(2):228–229, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895178/
pdf.

Condit:1966:MVR

[Con66d] Carl W. Condit. Mies van der Rohe: The Art of Structure by Werner Blaser (review). Technology and Culture, 7(4):552–554, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894561/pdf.

Condit:1967:SAA

[Con67a] Carl W. Condit. Structure: an Architect’s Approach by H. Seymour Howard, Jr. (review). Technology and Culture, 8(4):507–508, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894411/pdf.

Condit:1967:CAR

[Con67b] Carl W. Condit. The Concrete Architecture of Riccardo Morandi by Giorgio Boaga, Benito Boni (review). Technology and Culture, 8(1):111–113, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894509/pdf.

Condit:1968:HOA

[Con68a] Carl W. Condit. A Historical Outline of Architectural Science by Henry J. Cowan (review). Technology and Culture, 9(2):226–227, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894306/
pdf.

Condit:1968:ABH

[Con68b] Carl W. Condit. American Building: The Historical Forces That Shaped It by James Marston Fitch (review). Technology and Culture, 9(2):227–229, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894307/pdf.

Condit:1968:HBU

[Con68c] Carl W. Condit. Highrise Building and Urban Design by Hans Aregger and Otto Glaus (review). Technology and Culture, 9(3):502–503, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894279/
pdf.

Condit:1968:FRC

[Con68d] Carl W. Condit. The first reinforced-concrete skyscraper: The ingalls building in Cincinnati and its place in structural history. Technology and Culture, 9(1):1–33, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894341/pdf.

Condit:1969:HEA

[Con69a] Carl W. Condit. A History of Egyptian Architecture: The Empire (The New Kingdom) by Alexander Badawy (review). Technology and Culture, 10(4):603–604, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892407/pdf.

Condit:1969:AAE

[Con69b] Carl W. Condit. Architecture in Ancient Egypt and the Near East by Alexander Badawy (review). Technology and Culture, 10(1):86–87, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892320/
pdf.

Condit:1969:NAW

[Con69c] Carl W. Condit. The Nature and Art of Workmanship by David Pye (review). Technology and Culture, 10(2):307–309, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892366/pdf.

dArge:1968:EIC

[d’A68] Ralph d’Arge. The Economics of Irrigation by Colin Clark (review). Technology and Culture, 9(2):269–271, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894333/pdf.

Dalrymple:1964:ATC

[Dal64] Dana G. Dalrymple. The American tractor comes to Soviet agriculture: The transfer of a technology. Technology and Culture, 5 (2):191–214, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894892.

Dalrymple:1969:DAT

[Dal69] Dana G. Dalrymple. The development of an agricultural technology: Controlled-atmosphere storage of fruit. Technology and Culture, 10(1):35–48, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892313/pdf.

Daum:1960:OSH

[Dau60a] Arnold R. Daum. Oil on Stream! A History of Interstate Oil Pipe Line Company, 1909–1959 by John L. Loos (review). Technology and Culture, 1(3):276–278, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895642/pdf.

Daumas:1960:LGT

[Dau60b] Maurice Daumas. L’Histoire generale des techniques. Technology and Culture, 1(4):415–418, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/895623/pdf.

Daum:1962:POS

[Dau62] Arnold R. Daum. The Politics of Oil: a Study of Private Power and Democratic Directions by Robert Engler (review). Technology and Culture, 3(3):365–368, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895158/pdf.

Daum:1964:COG

[Dau64] Arnold R. Daum. A Century of Oil and Gas in Books: a Descriptive Bibliography comp. by E. B. Swanson (review). Technology and Culture, 5(1):134–136, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894954/pdf.

Davison:1961:TST

[Dav61] C. St. C. Davison. Transporting sixty-ton statues in Early Assyria and Egypt. Technology and Culture, 2(1):11–16, Winter 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895571/pdf.

Davis:1962:ASD

[Dav62] Robert C. Davis. The Achieving Society by David C. McClelland (review). Technology and Culture, 3(3):351–353, Summer 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895151/pdf.

Davis:1964:DAS

[Dav64] Robert C. Davis. A Dictionary of American Social Reform by Louis Filler (review). Technology and Culture, 5(1):125–126, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894949/pdf.

Davis:1966:PJB

[Dav66] Natalie Zemon Davis. The Protestantism of Jacques Besson. Technology and Culture, 7(4):513, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894536/pdf.

Davenport:1968:TCP

[Dav68] William H. Davenport. Technology and Culture in Perspective by Harvey Cox, et al. (review). Technology and Culture, 9(1):103–105, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894354/
pdf.

deCamp:1961:HDA

[dC61a] L. Sprague de Camp. The House of the Double Axe: The Palace at Knossos by Agnes Carr Vaughan (review). Technology and Culture, 2(1):53–54, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895583/pdf.

deCamp:1961:MGA

[dC61b] L. Sprague de Camp. Master Gunner apollonios. Technology and Culture, 2(3):240–244, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895461/pdf.

deCamp:1963:PPE

[dC63a] L. Sprague de Camp. Pharaoh’s Pump by Edward J. Kunkel (review). Technology and Culture, 4(1):56–57, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895036/pdf.

deCamp:1963:PAF

[dC63b] L. Sprague de Camp. The Pyramids by Ahmed Fakhry (review). Technology and Culture, 4(1):54–56, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895035/pdf.

deCamp:1964:CI

[dC64a] L. Sprague de Camp. Custom and innovation. Technology and Culture, 5(4):595, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894822/pdf.

deCamp:1964:EPL

[dC64b] L. Sprague de Camp. An early patent law proposal? Technology and Culture, 5(3):377–379, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894854/pdf.

deCamp:1965:DA

[dC65] L. Sprague de Camp. The “Darkhouse” of Alexandria. Technology and Culture, 6(3):423–427, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894677/pdf.

deCamp:1969:U

[dC69] L. Sprague de Camp. The unsinkables. Technology and Culture, 10(3):422–423, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901622/pdf.

DeGregori:1964:SSE

[De 64] Thomas De Gregori. From Stone to Steel: Economic Consequences of a Technological Change in New Guinea by R. F. Salisbury (review). Technology and Culture, 5(2):272–274, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894911/pdf.

DeGregori:1965:STD

[De 65] Thomas De Gregori. Science and technology for development. Technology and Culture, 6(4):636–648, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894647/pdf.

DeGregori:1967:FIC

[De 67a] Thomas De Gregori. First International Conference of Economic History, Stockholm, 1960, and: Second International Conference of Economic History, Aix-en-Provence, 1962 (review). Technology and Culture, 8(2):219–220, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894465/pdf.

DeGregori:1967:MEG

[De 67b] Thomas R. De Gregori. Modern Economic Growth: Rate, Structure, and Spread by Simon Kuznets (review). Technology and Culture, 8(3):419–420, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894449/pdf.

DeFrance:1968:HAC

[De 68a] Smith J. De France. Helicopters and Autogiros by Charles Gablehouse (review). Technology and Culture, 9(2):249–250, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894321/pdf.

DeGregori:1968:MSD

[De 68b] Thomas De Gregori. Manpower Strategy for Developing Countries: Lessons from Ethiopia by Eli Ginzberg, Herbert A. Smith (review). Technology and Culture, 9(2):269, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894332/pdf.

DeGregori:1968:TTD

[De 68c] Thomas R. De Gregori. The Transfer of Technology to Developing Countries ed. by Daniel L. Spencer, Alexander Woroniak (review). Technology and Culture, 9(4):631–633, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894241/pdf.

DeVore:1968:IEA

[De 68d] Paul W. De Vore. Industrial Education: American Ideals and Institutions by Berenice M. Fisher (review). Technology and Culture, 9(3):516–518, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894287/pdf.

Dean:1969:OAR

[Dea69] Burton V. Dean. Organization and Administration of Research and Development ed. by N. H. Atthreya (review). Technology and Culture, 10(2):335–336, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892384/pdf.

DeGolyer:1964:LDF

[DeG64] E. DeGolyer, Jr. The library of the DeGolyer Foundation. Technology and Culture, 5(3):408–411, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894861/pdf.

DeGregori:1965:ALE

[DeG65a] Thomas DeGregori. Africans on the Land: Economic Problems of African Agricultural Development in Southern, Central, and East Africa, with Special Reference to Southern Rhodesia by Montague Yudelman (review). Technology and Culture, 6(3):509, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894713/pdf.

DeGregori:1965:SGE

[DeG65b] Thomas DeGregori. Scarcity and Growth: The Economics of Natural Resource Availability by Harold J. Barnett and Chandler Morse (review). Technology and Culture, 6(3):506–509, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894712/pdf.

DeGregori:1969:TCE

[DeG69a] Thomas R. DeGregori. Technological Change and Economic Development: The Manufacturing Experience of Mexico and Puerto Rico by W. Paul Strassmann (review). Technology and Culture, 10 (1):130–131, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892344/pdf.

DeGregori:1969:TUA

[DeG69b] Thomas R. DeGregori. Technology for Underdeveloped Areas: an Annotated Bibliography by Jack Baranson (review). Technology and Culture, 10(1):133, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892346/pdf.

Deutsch:1965:MHI

[Deu65] Steven E. Deutsch. Mental Health of the Industrial Worker: a Detroit Study by Arthur Kornhauser (review). Technology and Culture, 6(4):670–672, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894660/pdf.

Deutsch:1966:IJW

[Deu66] Steven E. Deutsch. Industrial Jobs and the Worker: an Investigation of Response to Task Attributes by Arthur N. Turner and Paul R. Lawrence (review). Technology and Culture, 7(3):436–438, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894602/pdf.

Dibner:1964:MEA

[Dib64] Bern Dibner. The Making of the Electrical Age: From the Telegraph to Automation by Harold I. Sharlin (review). Technology and Culture, 5(3):441–443, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894867/pdf.

Dibner:1965:AGL

[Dib65] Bern Dibner. “Assaying Gold” by Lazarus Ercker. Technology and Culture, 6(3):445–446, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894681/pdf.

Dibner:1966:EEC

[Dib66] Bern Dibner. Early Electrical Communication by E. A. Marland (review). Technology and Culture, 7(1):96–98, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893873/pdf.

Dibbits:1969:MMU

[Dib69a] H. A. M. C. Dibbits. Machines, Mills and Uncountable Costly Necessities: a Short History of the Drainage of the Fens by Richard L. Hills (review). Technology and Culture, 10(2):312–313, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892369/pdf.

Dibner:1969:CSC

[Dib69b] Bern Dibner. Cableships and Submarine Cables by K. R. Haigh (review). Technology and Culture, 10(1):102–104, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892329/pdf.

Dibner:1969:ESI

[Dib69c] Bern Dibner. Elektrische Spielereien im Barock und Rokoko by Fritz Fraunberger (review). Technology and Culture, 10(1):100–101, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892327/
pdf.

Didsbury:1962:PME

[Did62] Howard F. Didsbury, Jr. Professional Manpower and Education in Communist China by Leo A. Orleans (review). Technology and Culture, 3(2):226–227, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895405/pdf.

delaCroix:1963:LFR

[dlC63] Horst de la Croix. The literature on fortification in Renaissance Italy. Technology and Culture, 4(1):30–50, Winter 1963. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895031/pdf.

deMedina:1967:HPT

[dM67] Carlos Alberto de Medina. Human Problems in Technological Change ed. by Edward H. Spicer (review). Technology and Culture, 8(1):138– 143, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894522/pdf.

Dobrovolny:1962:SSF

[Dob62] Jerry S. Dobrovolny. Samuel Slater, Father of American Manufactures by E. H. Cameron (review). Technology and Culture, 3(2):188– 189, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895387/pdf.

Dornberger:1963:GV

[Dor63] Walter R. Dornberger. The German V-2. Technology and Culture, 4(4):393–409, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894960/pdf.

Dorf:1968:TMF

[Dor68a] Richard C. Dorf. Technology and Man’s Future: University of Santa Clara April 1968. Technology and Culture, 9(4):580–584, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894217/pdf.

Dorn:1968:NSA

[Dor68b] Harold Dorn. A note on “Structural Antecedents of the I-Beam”. Technology and Culture, 9(3):415–418, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894257/pdf. See [Jew67].

Dorn:1968:R

[Dor68c] Harold Dorn. The rejoinder. Technology and Culture, 9(3):427–429, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894259/
pdf.

Dorwart:1968:FEO

[Dor68d] Reinhold A. Dorwart. Fire Engines and Other Fire-fighting Appliances by K. R. Gilbert (review). Technology and Culture, 9(1):120, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894362/
pdf.

Dowd:1964:ALA

[Dow64] Douglas F. Dowd. Gli Aratri e l’Economia Agraria Nel Bolognese, dal XVII al XIX Secolo by Carlo Poni (review). Technology and Culture, 5(4):604–605, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894827/pdf.

Dowd:1965:CMW

[Dow65] Douglas F. Dowd. The Conquest of the Material World by John Nef (review). Technology and Culture, 6(4):651–653, Fall 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894650/pdf.

DeGregori:1966:TTM

[DPS66] Thomas De Gregori and Oriol Pi-Sunyer. Technology, traditionalism, and military establishments. Technology and Culture, 7(3):402–407, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894587/
pdf.

Dresbeck:1967:SHH

[Dre67] LeRoy J. Dresbeck. The ski: Its history and historiography. Technology and Culture, 8(4):467–479, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894398/pdf.

Drucker:1961:DCZ

[Dru61a] Peter F. Drucker. Drucker’s comments on Zvorikine. Technology and Culture, 2(3):253–254, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895464/pdf. See [Zvo61a].

Drucker:1961:TRN

[Dru61b] Peter F. Drucker. The technological revolution: Notes on the relationship of technology, science, and culture. Technology and Culture, 2(4):342–351, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895432/pdf.

Drucker:1963:MTA

[Dru63] Peter F. Drucker. Modern technology and ancient jobs. Technology and Culture, 4(3):277–281, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894971/pdf.

Drucker:1966:FTR

[Dru66] Peter F. Drucker. The first technological revolution and its lessons. Technology and Culture, 7(2):143–151, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895169/pdf.

Drucker:1968:MMT

[Dru68] Peter F. Drucker. The Myth of the Machine: Technics and Human Development by Lewis Mumford (review). Technology and Culture, 9(1):94–98, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894350/pdf.

Drucker:1969:CTP

[Dru69] Peter F. Drucker. Comment: Is technology predictable? Technology and Culture, 10(4):522–527, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892397/pdf.

Dryden:1961:FEU

[Dry61] Hugh L. Dryden. Future exploration and utilization of outer space. Technology and Culture, 2(2):112–126, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895546/pdf.

Price:1961:USA

[dSP61] Derek J. de Solla Price. Die Uhr, von der Sonnenuhr zur Atomuhr by Anton Lübke (review). Technology and Culture, 2(1):49–50, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895581/pdf.

Price:1964:AOM

[dSP64] Derek J. de Solla Price. Automata and the origins of mechanism and mechanistic philosophy. Technology and Culture, 5(1):9–23, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894919/pdf.

Price:1965:THI

[dSP65] Derek J. de Solla Price. Is technology historically independent of science? A study in statistical historiography. Technology and Culture, 6(4):553–568, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894638/pdf.

Duncan:1963:SSE

[Dun63] Otis Dudley Duncan. The Sociology of Science ed. by Bernard Barber and Walter Hirsch (review). Technology and Culture, 4(3):372–373, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894996/
pdf.

Duncan:1968:NOH

[Dun68] Francis Duncan. This New Ocean: a History of Project Mercury by Loyd S. Swenson, Jr., et al. (review). Technology and Culture, 9(2):254–256, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894325/pdf.

Dunbar:1969:BBH

[Dun69] Robert G. Dunbar. The Black Book of Hunger by Josué de Castro (review). Technology and Culture, 10(2):342–343, April 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892388/pdf.

Dupree:1969:CRT

[Dup69] A. Hunter Dupree. Comment: The role of technology in society and the need for historical perspective. Technology and Culture, 10 (4):528–534, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892398/pdf.

Duveen:1964:NPC

[Duv64] Denis Duveen. Nobel Prizewinners in Chemistry 1901–1961 by Eduard Farber (review). Technology and Culture, 5(1):114, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894942/pdf.

Duveen:1966:ESC

[Duv66] Denis Duveen. Enterprise in Soap and Chemicals: Joseph Crosfield and Sons, Limited, 1815–1965 by A. E. Musson (review). Technology and Culture, 7(4):538–539, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894553/pdf.

Edelstein:1964:AMF

[Ede64] Sidney M. Edelstein. The allerley matkel (1532): Facsimile text, translation, and critical study of the earliest printed book on spot removing and dyeing. Technology and Culture, 5(3):297–321, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894849/pdf.

Edelstein:1965:MM

[Ede65] Sidney M. Edelstein. “Matkel” or “Mackel”. Technology and Culture, 6(3):444, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894680/pdf.

Edelstein:1966:DFS

[Ede66] Sidney M. Edelstein. Dyeing fabrics in Sixteenth-Century Venice. Technology and Culture, 7(3):395–397, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894585/pdf.

Edwards:1967:MPH

[Edw67] Charles E. Edwards. Master of Precision: Henry M. Leland by Mrs. Wilfred C. Leland (review). Technology and Culture, 8(4):503–505, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894409/
pdf.

Edwards:1968:BKL

[Edw68] Charles E. Edwards. Boyd on kettering and Leland. Technology and Culture, 9(4):592–593, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894219/pdf.

Eichner:1964:CPF

[Eic64] L. C. Eichner. The cooling pit of Filarete’s smelter. Technology and Culture, 5(3):404–405, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894859/pdf. See [Spe63, Spe64].

Eisenman:1966:BDA

[Eis66] Harry J. Eisenman. The brush double-arc lamp. Technology and Culture, 7(4):511–512, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894535/pdf.

Emme:1961:WMW

[Emm61] Eugene M. Emme. War in the Modern World by Theodore Ropp (review). Technology and Culture, 2(3):275–277, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895474/pdf.

Emme:1962:AMW

[Emm62a] Eugene M. Emme. Aviation in the Modern World by James V. Bernardo (review). Technology and Culture, 3(1):105, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895425/pdf.

Emme:1962:MSP

[Emm62b] Eugene M. Emme. Men of Space: Profiles of the Leaders of Space Research, Development and Exploration by Shirley Thomas (review). Technology and Culture, 3(2):212–214, Spring 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895399/pdf.

Emme:1962:RDL

[Emm62c] Eugene M. Emme. Rocket Development: Liquid Fuel Rocket Research, 1929–1941 by Robert H. Goddard (review). Technology and Culture, 3(3):364–365, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895157/pdf.

Emme:1963:NAS

[Emm63a] Eugene M. Emme. The National Aeronautics and Space Act: a Study of the Development of Public Policy by Alison Griffith (review). Technology and Culture, 4(2):231–232, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895015/pdf.

Emme:1963:IHR

[Emm63b] Eugene M. Emme. Introduction to the history of rocket technology. Technology and Culture, 4(4):377–383, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894958/pdf.

Emme:1964:CSE

[Emm64] Eugene M. Emme. The Challenges of Space ed. by Hugh Odishaw (review). Technology and Culture, 5(2):265–267, Spring 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894907/pdf.

Emme:1967:DBH

[Emm67] Eugene M. Emme. Deterrence before Hiroshima: The Airpower Background of Modern Strategy by George H. Quester (review). Technology and Culture, 8(2):245–247, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894476/pdf.

Emme:1968:ARA

[Emm68a] Eugene M. Emme. Aeronautics, rocketry, and astronautics. Technology and Culture, 9(3):436–455, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894261/pdf.

Emme:1968:HRS

[Emm68b] Eugene M. Emme. History of Rocketry and Space Travel by Wernher von Braun, Frederick I. Ordway III (review). Technology and Culture, 9(2):250–251, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894322/pdf.

Emme:1969:HCA

[Emm69] Eugene M. Emme. Historical considerations in aviation development: Transportation engineering conference, October 1968. Technology and Culture, 10(3):429–431, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/901624/pdf.

Endrei:1968:FTE

[End68] Walter G. Endrei. The first technical exhibition. Technology and Culture, 9(2):181–183, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894296/pdf.

Ermenc:1964:DET

[Erm64] J. J. Ermenc. The Development of the English Traction Engine by Ronald H. Clark (review). Technology and Culture, 5(2):258–259, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894901/
pdf.

Esper:1965:RLF

[Esp65] Thomas Esper. The replacement of the longbow by firearms in the English Army. Technology and Culture, 6(3):382–393, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894673/pdf.

Esper:1966:EMS

[Esp66a] Thomas Esper. Elizabethan Military Science: The Books and the Practice by Henry J. Webb (review). Technology and Culture, 7 (3):420–421, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894594/pdf.

Esper:1966:VSK

[Esp66b] Thomas Esper. Vladimir Sergeevich Knabbe by F. N. Zagorskii (review). Technology and Culture, 7(1):93–94, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893870/pdf.

Esper:1967:GSE

[Esp67] Thomas Esper. Guns, Sails, & Empires: Technological Innovation and the Early Phases of European Expansion, 1400–1700 by Carlo M. Cipolla (review). Technology and Culture, 8(2):220–222, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894466/pdf.

Esper:1968:AAM

[Esp68] Thomas Esper. The Art and Architecture of Medieval Russia by Arthur Voyce (review). Technology and Culture, 9(1):133–134, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894370/
pdf.

Etz:1966:FTR

[Etz66] Donald V. Etz. The first technological revolution. Technology and Culture, 7(4):515–516, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894538/pdf.

Evans:1965:EME

[Eva65] Luther H. Evans. Education, Manpower, and Economic Growth: Strategies of Human Resource Development by Frederick Harbison and Charles Myers (review). Technology and Culture, 6(1):157–158, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894811/
pdf.

Ellul:1962:TO

[EW62] Jacques Ellul and John Wilkinson. The technological order. Technology and Culture, 3(4):394–421, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895099/pdf.

Fairbank:1962:BCB

[Fai62] Wilma Fairbank. Bronze Casting and Bronze Alloys in Ancient China by Noel Barnard (review). Technology and Culture, 3(2):178–180, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895381/
pdf.

Falk:1965:TSJ

[Fal65] Howard Falk. The Technological Society by Jacques Ellul (review). Technology and Culture, 6(3):532–535, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894725/pdf.

Farber:1961:SCG

[Far61] Eduard Farber. The Story of Chemistry by Georg Lockemann (review). Technology and Culture, 2(3):270–272, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895471/pdf.

Farber:1963:HCJ

[Far63a] Eduard Farber. A History of Chemistry by J. R. Partington (review). Technology and Culture, 4(3):338–340, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894978/pdf.

Farber:1963:MEC

[Far63b] Eduard Farber. Michel Eugène Chevreul, Pioneer of Organic Chemistry by Albert B. Costa (review). Technology and Culture, 4(1):74–75, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895044/
pdf.

Farber:1964:HCJ

[Far64a] Eduard Farber. A History of Chemistry by J. R. Partington (review). Technology and Culture, 5(4):602–603, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894826/pdf.

Farber:1964:JCE

[Far64b] Eduard Farber. Ein Jahrhundert Chemie by Ernst Bäumler (review). Technology and Culture, 5(3):440–441, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894866/pdf.

Farber:1965:HCJ

[Far65] Eduard Farber. A History of Chemistry by J. R. Partington (review). Technology and Culture, 6(2):280–282, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894748/pdf.

Farber:1968:OCR

[Far68a] Eduard Farber. The Origins of Chemistry by Robert P. Multhauf (review). Technology and Culture, 9(2):224–226, April 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894305/pdf.

Farber:1968:SFH

[Far68b] Eduard Farber. The Story of Fritz Haber by Morris Goran (review). Technology and Culture, 9(3):499–500, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894277/pdf.

Farber:1969:RR

[Far69] Eduard Farber. The reviewer replies. Technology and Culture, 10(2): 183, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892357/pdf.

Federico:1960:REW

[Fed60] P. J. Federico. Records of Eli Whitney’s cotton gin patent. Technology and Culture, 1(2):168–176, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895252/pdf.

Feibleman:1961:PSA

[Fei61] James K. Feibleman. Pure science, applied science, technology, engineering: an attempt at definitions. Technology and Culture, 2(4):305– 317, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895429/
pdf.

Feibleman:1966:TS

[Fei66] James K. Feibleman. Technology as skills. Technology and Culture, 7(3):318–328, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894578/pdf.

Ferguson:1960:N

[Fer60] Eugene S. Ferguson. Notes. Technology and Culture, 1(3):292–297, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895648/
pdf.

Ferguson:1961:YSD

[Fer61] Eugene S. Ferguson. Yankee from Sweden: The Dream and the Reality in the Days of John Ericsson by Ruth White (review). Technology and Culture, 2(1):65–67, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895589/pdf.

Ferguson:1962:CBHa

[Fer62a] Eugene S. Ferguson. Contributions to bibliography in the history of technology (part i). Technology and Culture, 3(1):73–84, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895414/pdf.

Ferguson:1962:CBHb

[Fer62b] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part II). Technology and Culture, 3(2):167–174, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895379/pdf.

Ferguson:1962:CBHc

[Fer62c] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part III). Technology and Culture, 3(3):298–306, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895130/pdf.

Ferguson:1963:BIS

[Fer63a] Eugene S. Ferguson. Bibliography of Interlingual Scientific and Technical Dictionaries by UNESCO (review). Technology and Culture, 4(1):125–126, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895064/pdf.

Ferguson:1963:CBH

[Fer63b] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part IV). Technology and Culture, 4(3):318–330, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894975/pdf.

Ferguson:1964:HME

[Fer64a] Eugene S. Ferguson. A History of Mechanical Engineering by Aubrey F. Burstall (review). Technology and Culture, 5(3):438–440, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894865/pdf.

Ferguson:1964:BAP

[Fer64b] Eugene S. Ferguson. Bibliographia Antiqua: Philosophia Naturalis, Supplement II, 1950–1960 by R. J. Forbes (review). Technology and Culture, 5(4):639–640, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894844/pdf.

Ferguson:1964:CWO

[Fer64c] Eugene S. Ferguson. Cogs in the Wheel. One Hundred Books from the Llewellyn Powell Collection of Early Technology (review). Technology and Culture, 5(1):133–134, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894953/pdf.

Ferguson:1964:GIT

[Fer64d] Eugene S. Ferguson. General Index to Transactions, Volumes I– XXXII 1920–1960 by the Newcomen Society for the Study of the History of Engineering and Technology (review). Technology and Culture, 5(1):136–137, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894955/pdf.

Ferguson:1964:CBHa

[Fer64e] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part V). Technology and Culture, 5(3):416–434, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894863/pdf.

Ferguson:1964:CBHb

[Fer64f] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part VI). Technology and Culture, 5(4):578–594, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894821/pdf.

Ferguson:1965:OCB

[Fer65a] Eugene S. Ferguson. Octave Chanute, 1832–1910: a Bibliography by Pearl I. Young (review). Technology and Culture, 6(1):147–148, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894804/
pdf.

Ferguson:1965:BPA

[Fer65b] Eugene S. Ferguson. The Business Press in America, 1750–1865 by David P. Forsyth (review). Technology and Culture, 6(3):538–540, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894729/
pdf.

Ferguson:1965:CBH

[Fer65c] Eugene S. Ferguson. Contributions to bibliography in the history of technology (Part VII). Technology and Culture, 6(1):99–107, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894782/pdf.

Ferguson:1965:TMI

[Fer65d] Eugene S. Ferguson. Technical museums and international exhibitions. Technology and Culture, 6(1):30–46, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894777/pdf.

Ferguson:1966:TMG

[Fer66a] Eugene S. Ferguson. A Theatre of Machines by A. G. Keller (review). Technology and Culture, 7(4):527–528, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894546/pdf.

Ferguson:1966:HWT

[Fer66b] Eugene S. Ferguson. The History of Woodworking Tools by W. L. Goodman (review). Technology and Culture, 7(4):525–527, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894545/pdf.

Ferguson:1967:BCS

[Fer67] Eugene S. Ferguson. The Beginnings of a Century of Steam and Water Heating by the H. B. Smith Company by Susan Reed Stifler (review). Technology and Culture, 8(4):502–503, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894408/pdf.

Ferguson:1968:ARV

[Fer68a] Eugene S. Ferguson. Atmospheric Railways: a Victorian Venture in Silent Speed by Charles Hadfield (review). Technology and Culture, 9(4):612–613, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894230/pdf.

Ferguson:1968:HPM

[Fer68b] Eugene S. Ferguson. Hall of Power Machinery, Museum of History and Technology, U.S. National Museum by Robert M. Vogel (review). Technology and Culture, 9(1):75–85, January 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894347/pdf.

Ferguson:1968:CIA

[Fer68c] Eugene S. Ferguson. Cast-iron aqueduct in Rees’s “Cyclopaedia”. Technology and Culture, 9(4):597–600, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894221/pdf.

Ferguson:1969:ECM

[Fer69a] Eugene S. Ferguson. Essays in Cornish Mining History by D. B. Barton (review). Technology and Culture, 10(3):449–452, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901631/pdf.

Ferguson:1969:HHE

[Fer69b] Eugene S. Ferguson. The Harveys of Hayle. Engine Builders, Shipwrights and Merchants of Cornwall by Edmund Vale (review). Technology and Culture, 10(2):313–314, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892370/pdf.

Finch:1961:ESH

[Fin61] James Kip Finch. Engineering and science: a historical review and appraisal. Technology and Culture, 2(4):318–332, Fall 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895430/pdf.

Finch:1962:HAA

[Fin62a] James Kip Finch. The Heroic Age of American Invention by L. Sprague de Camp (review). Technology and Culture, 3(1):95–98, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895421/
pdf.

Finch:1962:RPT

[Fin62b] James Kip Finch. The role of purely technical innovation in the history of engineering. Technology and Culture, 3(3):307–309, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895131/pdf.

Finkelstein:1962:IIM

[Fin62c] N. A. Finkelstein. Industrialism and Industrial Man by Clark Kerr, et al. (review). Technology and Culture, 3(3):346–348, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895149/pdf.

Finch:1963:TBT

[Fin63a] James K. Finch. Trail Blazers of Technology by Harland Manchester (review). Technology and Culture, 4(3):344–347, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894981/pdf.

Finch:1963:ECR

[Fin63b] James Kip Finch. Engineering as a Career by Ralph J. Smith (review). Technology and Culture, 4(2):212–214, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895005/pdf.

Finch:1964:CET

[Fin64a] James Kip Finch. Civil Engineering through the Ages by Charles J. Merdinger (review). Technology and Culture, 5(3):435–438, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894864/pdf.

Finkelstein:1964:SIW

[Fin64b] Nisson A. Finkelstein. Scientists in Industry by William Kornhauser (review). Technology and Culture, 5(1):115–117, Winter 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894943/pdf.

Finch:1965:FEP

[Fin65a] James Kip Finch. Foundations of Engineering by Philip Sporn (review). Technology and Culture, 6(2):263–266, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894739/pdf.

Finn:1965:SMT

[Fin65b] Bernard S. Finn. The science museum today. Technology and Culture, 6(1):74–82, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894780/pdf.

Finn:1968:CSP

[Fin68] Bernard S. Finn. Contact at Sea by Peter B. Schroeder (review). Technology and Culture, 9(3):498, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894276/pdf.

FitzSimons:1968:JKF

[Fit68] Neal FitzSimons. James Kip Finch (1883–1967). Technology and Culture, 9(4):594–596, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894220/pdf.

Fletcher:1965:EGF

[Fle65] Max E. Fletcher. Economic Growth in France and Britain, 1851– 1950 by Charles P. Kindleberger (review). Technology and Culture, 6(2):285–287, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894751/pdf.

Florman:1969:CDM

[Flo69a] Samuel C. Florman. Cultural Development of Managers, Executives, and Professionals by Joseph Basile, and: Engineering: Its Role and Function in Human Society ed. by William H. Davenport, Daniel Rosenthal (review). Technology and Culture, 10(1):136–138, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892349/pdf.

Florman:1969:CEE

[Flo69b] Samuel C. Florman. Comment: Engineers and the end of innocence. Technology and Culture, 10(1):14–16, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892310/pdf.

Flynn:1968:VHG

[Fly68] John V. Flynn. The Virus House: Germany’s Atomic Research and Allied Countermeasures by David Irving (review). Technology and Culture, 9(2):243–245, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894317/pdf.

Fryling:1961:C

[FMG61] Glenn R. Fryling, Frank Makara, and S. Colum Gilfillan. Communications. Technology and Culture, 2(1):93–95, Winter 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895603/pdf.

Forbes:1960:DPE

[For60] R. J. Forbes. A Diderot Pictorial Encyclopedia of Trades and Industry: Manufacturing and the Technical Arts in Plates Selected from “L’Encyclopedie, ou Dictionnaire Raisonné des Sciences, des Arts et des Métiers” of Denis Diderot ed. by Charles Coulston Gillispie (review). Technology and Culture, 1(3):267–268, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895638/pdf.

Forbes:1961:ISA

[For61a] R. J. Forbes. Der Ingenieur und Seine Aufgaben in Neuen Wirtschaftsraumen (review). Technology and Culture, 2(1):82–83, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895597/pdf.

Forbes:1961:FWW

[For61b] R. J. Forbes. The Fifteen Wonders of the World by René Poirier (review). Technology and Culture, 2(3):269–270, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895470/pdf.

Forbes:1961:SEJ

[For61c] R. J. Forbes. The Story of Engineering by James Kip Finch (review). Technology and Culture, 2(1):33–34, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895575/pdf.

Forbes:1962:HLR

[For62] R. J. Forbes. History of the Lathe to 1850 by Robert S. Woodbury (review). Technology and Culture, 3(3):312–313, Summer 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895133/pdf.

Forbes:1963:HGT

[For63a] R. J. Forbes. Histoire générale des techniques ed. by Maurice Daumas (review). Technology and Culture, 4(3):336–337, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894977/pdf.

Forbes:1963:KGT

[For63b] R. J. Forbes. Kurze Geschichte der Technik by Friederich Klemm (review). Technology and Culture, 4(1):52, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895033/pdf.

Forbes:1964:MAP

[For64a] R. J. Forbes. Metallurgy in Archaeology: a Prehistory of Metallurgy in the British Isles by R. F. Tylecote (review). Technology and Culture, 5(2):254–255, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894898/pdf.

Forbes:1964:AES

[For64b] R. J. Forbes. The Ancient Engineers by L. Sprague de Camp, and: The Mechanical Technology of Greek and Roman Antiquity by A. G. Drachmann, and: Dictionnaire Archéologique des Techniques (review). Technology and Culture, 5(1):69–74, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894924/pdf.

Forbes:1965:AMP

[For65a] R. J. Forbes. Ancient Mesopotamia, Portrait of a Dead Civilization by A. Leo Oppenheim (review). Technology and Culture, 6(3):449–450, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894684/
pdf.

Forbes:1965:DAT

[For65b] R. J. Forbes. Dictionnaire archéologiques des techniques (review). Technology and Culture, 6(2):267, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894740/pdf.

Forbes:1966:GDG

[For66] R. J. Forbes. Geschichte der deutschen Gasindustrie by Johannes Körting (review). Technology and Culture, 7(1):82, Winter 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893864/pdf.

Fountain:1964:WMI

[Fou64] Alvin M. Fountain. Weights and Measures: an Informal Guide by Stacy V. Jones (review). Technology and Culture, 5(2):276–277, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894913/
pdf.

Fox:1969:FPO

[Fox69] Robert Fox. The fire piston and its origins in Europe. Technology and Culture, 10(3):355–370, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901618/pdf.

Frankel:1963:OIP

[Fra63] J. P. Frankel. The origin of Indonesian “Pamor”. Technology and Culture, 4(1):14–21, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895029/pdf.

Frazier:1964:DFH

[Fra64] Arthur H. Frazier. Daniel Farrand Henry’s cup type “Telegraphic” river current meter. Technology and Culture, 5(4):541–565, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894818/pdf.

Freudenberger:1969:KWI

[Fre69a] Herman Freudenberger. Die Kaiserliche Wirtschaftspolitik im Banat im 18. Jahrhundert by Sonja Jordan (review). Technology and Culture, 10(2):314–315, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892371/pdf.

Freudenberger:1969:GMM

[Fre69b] Herman Freudenberger. On the “ Genesis of modern management ”. Technology and Culture, 10(4):593–602, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892406/pdf.

Froomkin:1966:ERD

[Fro66] Joseph N. Froomkin. Economics of Research and Development ed. by Richard A. Tybout (review). Technology and Culture, 7(4):564–566, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894566/
pdf.

Fryling:1961:EHD

[Fry61] Glenn R. Fryling. Die Entwicklung des Hochdruckdampfes in Deutschland (The Development of High Pressure Steam in Germany) by Ernst-Otto Jochmann (review). Technology and Culture, 2(2):178– 180, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895555/pdf.

Fulton:1963:GCT

[Ful63] Garland Fulton. The general climate for technological developments in Naval aeronautics on the eve of World War I. Technology and Culture, 4(2):154–165, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894999/pdf.

Fussell:1966:SSS

[Fus66a] G. E. Fussell. Settlement and Society: a Study of the Early Agrarian History of South Lincolnshire by H. E. Hallam (review). Technology and Culture, 7(3):418–420, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894593/pdf.

Fussell:1966:SFF

[Fus66b] G. E. Fussell. Sherington: Fiefs and Fields of a Buckinghamshire Village by A. C. Chibnall (review). Technology and Culture, 7(4):524– 525, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894544/
pdf.

Fussell:1966:CAG

[Fus66c] G. E. Fussell. The Conditions of Agricultural Growth: The Economics of Agricultural Change under Population Pressure by Ester Boserup (review). Technology and Culture, 7(3):415–418, Summer 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894592/pdf.

Fussell:1967:CEH

[Fus67a] G. E. Fussell. Cambridge Economic History of Europe, Vol. I: The Agrarian Life of the Middle Ages ed. by M. M. Postan (review). Technology and Culture, 8(3):403–405, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894439/pdf.

Fussell:1967:AHE

[Fus67b] G. E. Fussell. The Agrarian History of England and Wales. General editor, H. P. R. Finberg. Vol. IV: 1500–1640 ed. by Joan Thirsk (review). Technology and Culture, 8(4):492–495, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894403/pdf.

Fussell:1967:FSC

[Fus67c] G. E. Fussell. Farming systems of the Classical Era. Technology and Culture, 8(1):16–44, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894487/pdf.

Fussell:1968:AIR

[Fus68] G. E. Fussell. Agricultural Implements of the Roman World by K. D. White (review). Technology and Culture, 9(4):602–603, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894223/pdf.

Gabor:1966:CSH

[Gab66] Dennis Gabor. A Case Study of the History of the Incandescent Lamp, 1830–1940: From Scientific Research to Industrial Production by Tibor Balázs (review). Technology and Culture, 7(4):534–535, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894551/pdf.

Gaines:1967:RSC

[Gai67] David I. Gaines. Report on the Sanitary Conditions of the Labouring Population of Great Britain by Edwin Chadwick, and: Doctors and the State: The British Medical Profession and Government Action in Public Health, 1870–1912 by Jeanne L. Brand (review). Technology and Culture, 8(2):225–228, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894469/pdf.

Garvan:1965:EAS

[Gar65] Anthony N. B. Garvan. Early American Scientific Instruments and Their Makers by Silvio A. Bedini (review). Technology and Culture, 6(3):464–465, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894691/pdf.

Garber:1967:JMP

[Gar67] Paul Edward Garber. James Means and the Problem of Manflight: During the Period 1882–1920 by James Howard Means (review). Technology and Culture, 8(1):115–117, January 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894511/pdf.

Gerstner:1966:SIB

[Ger66a] Patsy A. Gerstner. Surgical instruments, from Brunschwig’s Cirurgia. Technology and Culture, 7(1):70–71, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893856/pdf.

Gertel:1966:EDI

[Ger66b] Karl Gertel. The Economic Demand for Irrigated Acreage: New Methodology and Some Preliminary Projections, 1954–1980 by Vernon W. Ruttan (review). Technology and Culture, 7(2):245–248, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895189/
pdf.

Ghaswala:1961:IME

[Gha61] S. K. Ghaswala. Industrialization in the Middle East by Kurt Grunwald, Joachim O. Ronall (review). Technology and Culture, 2(1):79– 82, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895596/pdf.

Ghaswala:1965:HRD

[Gha65] S. K. Ghaswala. History of Road Development in India (review). Technology and Culture, 6(2):301–303, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894760/pdf.

Ghaswala:1967:CIV

[Gha67] S. K. Ghaswala. Civilizations of the Indus Valley and Beyond by Sir Mortimer Wheeler (review). Technology and Culture, 8(1):90–91, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894496/
pdf.

Ghaswala:1968:SAT

[Gha68a] S. K. Ghaswala. Stone Age Tools: Their Techniques, Names and Probable Functions by H. D. Sankalia (review). Technology and Culture, 9(1):136–137, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894372/pdf.

Ghaswala:1968:EIL

[Gha68b] S. K. Ghaswala. The Emergence of an Industrial Labor Force in India: a Study of the Bombay Cotton Mills, 1854–1947 by Morris David Morris (review). Technology and Culture, 9(1):137–139, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894373/pdf.

Ghaswala:1968:FCI

[Gha68c] S. K. Ghaswala. The Fortified Cities of India by Sidney Toy (review). Technology and Culture, 9(1):139–140, January 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894374/pdf.

Ghaswala:1969:BIC

[Gha69a] S. K. Ghaswala. The Birth of Indian Civilization: India and Pakistan before 500 B.C. by Bridget Allchin, Raymond Allchin (review). Technology and Culture, 10(4):606, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892409/pdf.

Ghaswala:1969:ICS

[Gha69b] S. K. Ghaswala. The Indus Civilization by Sir Mortimer Wheeler (review). Technology and Culture, 10(4):604–605, October 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892408/pdf.

Gilfillan:1960:AMR

[Gil60a] S. C. Gilfillan. An attempt to measure the rise of American inventing and the decline of patenting. Technology and Culture, 1(3):201–214, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895628/
pdf.

Gilfillan:1960:DGR

[Gil60b] S. C. Gilfillan. Dr. Gilfillan’s reply. Technology and Culture, 1 (3):227–234, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895631/pdf.

Gilfillan:1962:ILC

[Gil62] S. Colum Gilfillan. The inventive lag in classical Mediterranean society. Technology and Culture, 3(1):85–87, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895415/pdf.

Gilbert:1963:ARM

[Gil63] K. R. Gilbert. The Ames Recessing Machine: a survivor of the original Enfield rifle machinery. Technology and Culture, 4(2):207–211, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895004/
pdf.

Gilbert:1964:WSG

[Gil64] K. R. Gilbert. The Wheelwright’s Shop by George Sturt (review). Technology and Culture, 5(4):610–611, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894830/pdf.

Gilfillan:1966:ASU

[Gil66] S. C. Gilfillan. “175th Anniversary Symposium of the U.S. Patent System” in the Journal of the Patent Office Society (review). Technology and Culture, 7(4):546–547, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894557/pdf.

Glaab:1962:CRI

[Gla62] Charles N. Glaab. Canal or Railroad? Imitation and Innovation in the Response to the Erie Canal in Philadelphia, Baltimore, and Boston by Julius Rubin (review). Technology and Culture, 3(3):321–323, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895138/
pdf.

Glick:1966:FVS

[Gli66] Thomas F. Glick. La Font de la Vall de Segó by José María Cueco Adrián (review). Technology and Culture, 7(4):522–524, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894543/pdf.

Glick:1968:LLS

[Gli68] Thomas F. Glick. Levels and levelers: Surveying irrigation canals in Medieval Valencia. Technology and Culture, 9(2):165–180, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894295/pdf.

Glick:1969:ASM

[Gli69a] Thomas F. Glick. Arturo Soria y Mata y la Ciudad Lineal ed. by George R. Collins and Carlos Flores (review). Technology and Culture, 10(3):460–461, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901636/pdf.

Glick:1969:MNM

[Gli69b] Thomas F. Glick. La minería en el Nuevo Mundo by Carlos Prieto (review). Technology and Culture, 10(4):607–608, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892411/pdf.

Glick:1969:MIC

[Gli69c] Thomas F. Glick. Medieval irrigation clocks. Technology and Culture, 10(3):424–428, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901623/pdf.

Goist:1969:PCP

[Goi69] Park Dixon Goist. The Private City: Philadelphia in Three Periods of Its Growth by Sam Bass Warner, Jr. (review). Technology and Culture, 10(4):618–620, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892417/pdf.

Goldschmidt:1962:TEC

[Gol62] Arthur Goldschmidt. Technology in emerging countries. Technology and Culture, 3(4):581–600, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895115/pdf.

Goldbeck:1963:C

[Gol63] Gustav Goldbeck. Communication. Technology and Culture, 4(1):127, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895065/
pdf.

Goldstein:1965:NGD

[Gol65] Walter Goldstein. The Next Generation by Donald M. Michael (review). Technology and Culture, 6(4):686–688, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894667/pdf.

Goldbeck:1966:GGT

[Gol66a] G. Goldbeck. Grundlagen der Geschichte der Technik by S. W. Schuchardin (review). Technology and Culture, 7(3):412–413, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894590/
pdf.

Goldbeck:1966:LIR

[Gol66b] Gustav Goldbeck. L’inventeur Isaac de Rivaz (1752–1828) by Henry Michelet (review). Technology and Culture, 7(4):548–549, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894558/pdf.

Goodrich:1961:HCG

[Goo61] L. Carrington Goodrich. Heavenly Clockwork: The Great Astronomical Clocks of Medieval China by Joseph Needham, et al. (review). Technology and Culture, 2(1):50–52, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895582/pdf.

Goodrich:1965:LCE

[Goo65a] L. Carrington Goodrich. The Legacy of China ed. by Raymond Dawson (review). Technology and Culture, 6(2):298–299, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894758/pdf.

Goodwin:1965:BHM

[Goo65b] Jack Goodwin. Bibliography of the History of Medicine of the United States and Canada, 1939–1960 ed. by Genevieve Miller (review). Technology and Culture, 6(4):690–691, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894669/pdf.

Goodwin:1965:CBH

[Goo65c] Jack Goodwin. Current bibliography in the history of technology (1963). Technology and Culture, 6(2):346–374, Spring 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894774/pdf.

Goodwin:1966:SBI

[Goo66a] Jack Goodwin. Sources of Business Information by Edwin T. Coman, Jr. (review). Technology and Culture, 7(1):123–124, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893887/pdf.

Goodwin:1966:CBH

[Goo66b] Jack Goodwin. Current bibliography in the history of technology (1964). Technology and Culture, 7(2):268–300, Spring 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895198/pdf.

Goodrich:1967:PPR

[Goo67a] L. Carrington Goodrich. Printing: Preliminary report on a new discovery. Technology and Culture, 8(3):376–378, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894434/pdf.

Goodwin:1967:CBH

[Goo67b] Jack Goodwin. Current bibliography in the history of technology (1965). Technology and Culture, 8(2):258–309, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894482/pdf.

Goodwin:1968:SMB

[Goo68a] Jack Goodwin. Source Materials for Business and Economic History ed. by Laurence J. Kipp (review). Technology and Culture, 9(2):275– 276, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894336/pdf.

Goodwin:1968:HST

[Goo68b] Jack Goodwin. The History of Science and Technology, a Select Bibliography for Students by K. J. Rider (review). Technology and Culture, 9(2):274–275, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894335/pdf.

Goodwin:1968:CBH

[Goo68c] Jack Goodwin. Current bibliography in the history of technology (1966). Technology and Culture, 9(2):277–327, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894337/pdf.

Goodwin:1969:TEM

[Goo69a] Jack Goodwin. Tudor England, 1485–1603 by Mortimer Levine (review). Technology and Culture, 10(2):352–353, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892395/pdf.

Goodwin:1969:CBH

[Goo69b] Jack Goodwin. Current bibliography in the history of technology (1967). Technology and Culture, 10(2):228–303, April 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892364/pdf.

Gorman:1968:EEM

[Gor68] Mel Gorman. An early electric motor in India. Technology and Culture, 9(2):184–190, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894297/pdf.

Goran:1969:SFH

[Gor69] Morris Goran. The story of Fritz Haber. Technology and Culture, 10(2):182, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892356/pdf.

Graham:1960:FAB

[Gra60] W. Earl Graham. Food: America’s Biggest Business by Pauline Arnold and Percival White (review). Technology and Culture, 1(2): 194, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895261/pdf.

Grancsay:1961:AWA

[Gra61] Stephen V. Grancsay. The Archaeology of Weapons: Arms and Armour from Prehistory to the Age of Chivalry by R. Ewart Oakeshott (review). Technology and Culture, 2(2):171–174, Spring 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895552/pdf.

Graham:1962:MDU

[Gra62] W. Earl Graham. Museums Directory of the United States and Canada ed. by Erwin O. Christensen (review). Technology and Culture, 3(1):108–109, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/895427/pdf.

Grancsay:1965:MAH

[Gra65] Stephen V. Grancsay. The Manufacture of Armour and Helmets in Sixteenth Century Japan by Sakakibara Kōzan (review). Technology and Culture, 6(2):299–301, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894759/pdf.

Greene:1963:MEE

[Gre63] John C. Greene. Mankind Evolving: The Evolution of the Human Species by Theodosius Dobzhansky (review). Technology and Culture, 4(1):111–112, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/895057/pdf.

Green:1965:IAK

[Gre65] E. R. R. Green. Industrial Archaeology by Kenneth Hudson (review). Technology and Culture, 6(4):688–690, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894668/pdf.

Grundstein:1967:AHN

[Gru67] Nathan D. Grundstein. An Administrative History of NASA, 1958– 1963 by Robert L. Rosholt (review). Technology and Culture, 8 (4):511–512, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894414/pdf.

Grundstein:1968:SMM

[Gru68] Nathan D. Grundstein. Scientific and Managerial Manpower in Nuclear Industry by James W. Kuhn (review). Technology and Culture, 9(2):260–262, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894328/pdf.

Guenther:1960:SFH

[Gue60] Paul E. Guenther. Steam’s Finest Hour ed. by David P. Morgan (review). Technology and Culture, 1(3):273–276, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895641/pdf.

Guest:1961:TWA

[Gue61] R. H. Guest. Taylorism at Watertown Arsenal: Scientific Management in Action 1908–1915 by Hugh G. J. Aitken (review). Technology and Culture, 2(2):191–193, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895561/pdf.

Guest:1962:PFC

[Gue62a] R. H. Guest. A Polish Factory: a Case Study of Workers’ Participation in Decision Making by Jiri Kolaja (review). Technology and Culture, 3(1):103–104, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895424/pdf.

Guest:1962:GOE

[Gue62b] Robert H. Guest. The Great Organizers by Ernest Dale (review). Technology and Culture, 3(3):349–351, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895150/pdf.

Guest:1965:AFF

[Gue65] Robert H. Guest. Alienation and Freedom: The Factory Worker and His Industry by Robert Blauner (review). Technology and Culture, 6(3):518–519, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894718/pdf.

Guest:1967:COW

[Gue67] Robert H. Guest. Counseling in an Organization by William J. Dickson, Fritz J. Roethlisberger (review). Technology and Culture, 8 (4):539–540, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894427/pdf.

Guest:1968:IRB

[Gue68] Robert H. Guest. Industrial Relations in a British Car Factory by Garfield Clack (review). Technology and Culture, 9(2):242–243, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894316/pdf.

Gullette:1961:SEC

[Gul61a] George A. Gullette. Science and Education at the Crossroads: a View from the Laboratory by Joseph W. Still (review). Technology and Culture, 2(3):300–301, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895487/pdf.

Gullette:1961:SRC

[Gul61b] George A. Gullette. The Scientific Revolution: Challenge and Promise ed. by Gerald W. Elbers, Paul Duncan (review). Technology and Culture, 2(1):88–89, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895601/pdf.

Gullette:1964:TAE

[Gul64] George A. Gullette. Technology and the Academics: an Essay on Universities and the Scientific Revolution by Eric Ashby (review). Technology and Culture, 5(4):600–601, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894825/pdf.

Gullette:1969:ELA

[Gul69] George A. Gullette. Engineering and the Liberal Arts: a Technologist’s Guide to History, Literature, Philosophy, Art, and Music by Samuel C. Florman (review). Technology and Culture, 10(1):135–136, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892348/
pdf.

Gutman:1964:CSF

[Gut64] Robert Gutman. Cities and Space: The Future Use of Urban Land ed. by Lowdon Wingo, Jr. (review). Technology and Culture, 5(4):628– 631, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894840/
pdf.

Hacker:1968:GCC

[Hac68] Barton C. Hacker. Greek catapults and catapult technology: Science, technology, and war in the ancient world. Technology and Culture, 9(1):34–50, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894342/pdf.

Haden:1960:CG

[Had60] H. J. Haden. Ceramics and glass. Technology and Culture, 1(4):364– 371, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895614/
pdf.

Haden:1961:NCG

[Had61] H. J. Haden. Nineteenth Century Glass: Its Genesis and Development by Albert Christian Revi (review). Technology and Culture, 2(1):68–70, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895590/pdf.

Haden:1963:AAG

[Had63] H. J. Haden. Analysis of Ancient Glasses, 1790–1957: a Comprehensive and Critical Survey by Earle R. Caley (review). Technology and Culture, 4(1):58–61, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895037/pdf.

Haden:1964:APG

[Had64a] H. J. Haden. American Pressed Glass and Figure Bottles by Albert Christian Revi (review). Technology and Culture, 5(4):605–608, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894828/pdf.

Haden:1964:AGF

[Had64b] H. J. Haden. Ancient Glass by Frederic Neuburg (review). Technology and Culture, 5(1):74–76, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894925/pdf.

Haden:1967:JGS

[Had67] H. Jack Haden. Journal of Glass Studies (review). Technology and Culture, 8(2):215–219, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894464/pdf.

Haden:1968:JGS

[Had68] H. Jack Haden. Journal of Glass Studies by Corning Museum of Glass (review). Technology and Culture, 9(1):113–117, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894359/pdf.

Haden:1969:JGS

[Had69] H. Jack Haden. Journal of Glass Studies ed. by Paul N. Perrot (review). Technology and Culture, 10(2):346–349, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892391/pdf.

Hagen:1963:VV

[Hag63] John P. Hagen. The Viking and the Vanguard. Technology and Culture, 4(4):435–451, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894962/pdf.

Hall:1960:HTS

[Hal60a] A. Rupert Hall. “A history of technology”: Some editorial reflections. Technology and Culture, 1(4):311–319, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895607/pdf.

Hall:1960:SMM

[Hal60b] A. Rupert Hall. The Science of Mechanics in the Middle Ages by Marshall Clagett (review). Technology and Culture, 1(3):259–262, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895635/
pdf.

Hall:1961:ESR

[Hal61a] A. Rupert Hall. Engineering and the scientific revolution. Technology and Culture, 2(4):333–341, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895431/pdf.

Hall:1961:MMP

[Hal61b] A. Rupert Hall. More on medieval pivoted axles. Technology and Culture, 2(1):17–22, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895572/pdf.

Hall:1962:CTA

[Hal62] A. Rupert Hall. The changing technical act. Technology and Culture, 3(4):501–515, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895106/pdf.

Hall:1963:EUS

[Hal63] R. Cargill Hall. Early U.S. Satellite Proposals. Technology and Culture, 4(4):410–434, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894961/pdf.

Hall:1965:HBW

[Hal65a] A. Rupert Hall. The History of the Barometer by W. E. Knowles Middleton (review). Technology and Culture, 6(1):130–131, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894793/pdf.

Halsey:1965:PKR

[Hal65b] Ashley Halsey, Jr. The Pennsylvania–Kentucky Rifle by Henry J. Kauffman (review). Technology and Culture, 6(1):126–128, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894791/pdf.

Hall:1966:AWS

[Hal66] A. Rupert Hall. The Art of War by Simon Stevin (review). Technology and Culture, 7(1):72–73, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893857/pdf.

Hall:1968:PWS

[Hal68] A. Rupert Hall. The Principal Works of Simon Stevin. Vol. V: Engineering ed. by R. J. Forbes, et al. (review). Technology and Culture, 9(3):481–483, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894268/pdf.

Hamill:1968:MMI

[Ham68] Hugh M. Hamill, Jr. The Mexican Mining Industry, 1890–1950: a Study of the Interaction of Politics, Economics, and Technology by Marvin D. Bernstein (review). Technology and Culture, 9(1):134–136, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894371/
pdf.

Hammond:1969:PPU

[Ham69] Thomas T. Hammond. Planning and Productivity under Soviet Socialism by Abram Bergson (review). Technology and Culture, 10(3):462–464, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901638/pdf.

Hanford:1961:RRL

[Han61] W. E. Hanford. The Research Revolution by Leonard S. Silk (review). Technology and Culture, 2(3):298–300, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895486/pdf.

Hannah:1967:IRI

[Han67] Theodore M. Hannah. Izobretenie Radio (“The Invention of Radio”) ed. by A. I. Berg (review). Technology and Culture, 8(3):411–413, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894445/
pdf.

Hartenberg:1960:HGC

[Har60] R. S. Hartenberg. History of the Gear-Cutting Machine: a Historical Study in Geometry and Machines by Robert S. Woodbury (review). Technology and Culture, 1(2):186–187, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895257/pdf.

Hartenberg:1961:GTZ

[Har61] R. S. Hartenberg. Geschichte des Technischen Zeichnens by Franz Maria Feldhaus (review). Technology and Culture, 2(1):45–49, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895580/
pdf.

Hartenberg:1962:TNS

[Har62a] R. S. Hartenberg. Transactions of the Newcomen Society, Vol. XXXI, 1957–58, 1958–59 (review). Technology and Culture, 3(3):313–316, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895134/
pdf.

Hartner:1962:PHT

[Har62b] Willy Hartner. The place of humanism in a technological world. Technology and Culture, 3(4):544–553, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895111/pdf.

Hartenberg:1963:BTR

[Har63] R. S. Hartenberg. Blätter für Technikgeschichte (review). Technology and Culture, 4(3):347–350, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894982/pdf.

Hartenberg:1964:BTR

[Har64] R. S. Hartenberg. Blätter für Technikgeschichte (review). Technology and Culture, 5(4):641–644, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894846/pdf.

Hartenberg:1965:KGT

[Har65a] R. S. Hartenberg. Kleine Geschichte der Technologie by Albrecht Timm (review). Technology and Culture, 6(3):448–449, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894683/pdf.

Hartman:1965:EUS

[Har65b] Lawton M. Hartman. Efficiency and Uplift: Scientific Management in the Progressive Era, 1890–1920 by Samuel Haber (review). Technology and Culture, 6(2):307–309, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894763/pdf.

Hartenberg:1966:TR

[Har66] Richard S. Hartenberg. Technikgeschichte (review). Technology and Culture, 7(3):410–411, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894589/pdf.

Hartenberg:1967:CEE

[Har67] R. S. Hartenberg. Combustion Engineering ed. by Glenn R. Fryling (review). Technology and Culture, 8(4):506–507, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894410/pdf.

Harrison:1969:LTC

[Har69] James Harrison. Lancashire Textiles — a Case Study of Industrial Change by Caroline Miles (review). Technology and Culture, 10(3):461–462, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901637/pdf.

Hattery:1962:DED

[Hat62] Lowell H. Hattery. The Diplomacy of Economic Development by Eugene R. Black (review). Technology and Culture, 3(2):222–223, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895403/pdf.

Hathway:1966:NSS

[Hat66] Jack F. Hathway. Neville Shute’s sliderule as a historical source. Technology and Culture, 7(1):69, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893855/pdf.

Hawkins:1966:RWE

[Haw66] Leslie V. Hawkins. A Reverence for Wood by Eric Sloane (review). Technology and Culture, 7(3):424–426, Summer 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894596/pdf.

Chou:1966:ECM

[hC66] Shun hsin Chou. The Economy of the Chinese Mainland: National Income and Economic Development, 1933–1959 by Ta-chung Liu and Kung-chia Yeh (review). Technology and Culture, 7(2):264–267, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895197/
pdf.

Heald:1961:CSL

[Hea61] Morrell Heald. Charleston’s Sons of Liberty: a Study of the Artisans, 1763–1789 by Richard Walsh (review). Technology and Culture, 2(1):59–61, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895586/pdf.

Heald:1963:CRA

[Hea63] Morrell Heald. The Corporate Revolution in America, Economic Reality vs. Economic Theory by Gardiner C. Means (review). Technology and Culture, 4(3):359–360, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894988/pdf.

Heckscher:1967:AFF

[Hec67] Morrison H. Heckscher. American Furniture — the Federal Period in the Henry Francis du Pont Winterthur Museum by Charles F. Montgomery (review). Technology and Culture, 8(3):408–410, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894443/pdf.

Heizer:1962:BTT

[Hei62] Robert F. Heizer. The background of Thomsen’s three-age system. Technology and Culture, 3(3):259–266, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895127/pdf.

Heilbroner:1967:DMM

[Hei67] Robert L. Heilbroner. Do machines make history? Technology and Culture, 8(3):335–345, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894431/pdf.

Heilbroner:1968:ETI

[Hei68] Robert L. Heilbroner. The economic and technological implications of 3 per cent growth: The SHOT-AEA Program December 1961. Technology and Culture, 9(4):570–575, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894215.

Hertz:1962:AIR

[Her62] David B. Hertz. American Industrial Research Laboratories by Frederick White (review). Technology and Culture, 3(2):199–201, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895393/pdf.

Hewlett:1964:PNF

[Hew64] Richard G. Hewlett. Pioneering on nuclear frontiers: Two early landmarks in reactor technology. Technology and Culture, 5(4):512–522, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894816/
pdf.

Hewlett:1966:PHS

[Hew66] Richard G. Hewlett. A Peril and a Hope: The Scientists’ Movement in America, 1945–47 by Alice Kimball Smith, and: Day of Trinity by Lansing Lamont (review). Technology and Culture, 7(2):250–252, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895191/
pdf.

Hewlett:1967:PAS

[Hew67] Richard G. Hewlett. The Politics of American Science, 1939 to the Present ed. by James L. Penick, Jr., et al. (review). Technology and Culture, 8(1):125–126, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894517/pdf.

Hidy:1960:RTI

[Hid60] Ralph W. Hidy. Risk and Technological Innovation: American Manufacturing Methods during the Nineteenth Century by Paul W. Strassman (review). Technology and Culture, 1(3):272–273, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895640/pdf.

Hidy:1962:MDH

[Hid62] Ralph W. Hidy. Mill on the Dan: a History of Dan River Mills, 1882– 1950 by Robert Sidney Smith (review). Technology and Culture, 3 (3):341–343, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895147/pdf.

Hidy:1968:ABC

[Hid68] Ralph W. Hidy. Andrew Brown and Cypress Lumbering in the Old Southwest by John Hebron Moore (review). Technology and Culture, 9(4):610–611, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894228/pdf.

Hiebert:1963:LAR

[Hie63] Erwin N. Hiebert. Living with the Atom by Ritchie Calder (review). Technology and Culture, 4(1):104–107, Winter 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895053/pdf.

Higgins:1960:EE

[Hig60] Thomas James Higgins. Electrical engineering. Technology and Culture, 1(4):408–410, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895621/pdf.

Higgins:1961:BBEa

[Hig61a] Thomas James Higgins. A biographical bibliography of electrical engineers and electrophysicists (Part I). Technology and Culture, 2(1):28– 32, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895574/pdf.

Higgins:1961:BBEb

[Hig61b] Thomas James Higgins. A biographical bibliography of electrical engineers and electrophysicists (Part II). Technology and Culture, 2(2):146–165, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895549/pdf.

Higham:1963:SFA

[Hig63] Robin Higham. Sky Fever: The Autobiography of Sir Geoffrey de Havilland (review). Technology and Culture, 4(1):99–100, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895051/pdf.

Higham:1964:RAB

[Hig64] Robin Higham. The Royal Arsenal: Its Background, Origin, and Subsequent History by Brigadier O. F. G. Hogg (review). Technology and Culture, 5(2):255–257, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894899/pdf.

Higham:1965:HWA

[Hig65a] Robin Higham. A History of the World’s Airlines by R. E. G. Davies (review). Technology and Culture, 6(3):481–482, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894700/pdf.

Higham:1965:LNF

[Hig65b] Robin Higham. … That Nothing Failed Them: Testing Aeroplanes in War by A. H. Wheeler, and: Operational Research in the Royal Air Force by the Air Ministry, and: The Design and Development of Weapons: Studies in Government and Industrial Organization by M. M. Postan, et al. (review). Technology and Culture, 6(3):478–480, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894698/
pdf.

Higham:1965:BSD

[Hig65c] Robin Higham. The Blunted Sword by David Divine (review). Technology and Culture, 6(1):151–152, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894807/pdf.

Higham:1966:BAMa

[Hig66a] Robin Higham. Building Aeroplanes for “Those Magnificent Men” by Allen H. Wheeler (review). Technology and Culture, 7(1):112–113, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893882/
pdf.

Higham:1966:BAMb

[Hig66b] Robin Higham. Buying Aircraft: Matériel Procurement for the Army Air Forces by Irving Brinton Holley, Jr. (review). Technology and Culture, 7(4):558–560, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894563/pdf.

Higham:1966:WAS

[Hig66c] Robin Higham. Out on a Wing: an Autobiography by Sir Miles Thomas (review). Technology and Culture, 7(3):443–444, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894606/pdf.

Higham:1966:QRA

[Hig66d] Robin Higham. Qantas Rising: The Autobiography of the Flying Fysh by Hudson Fysh (review). Technology and Culture, 7(4):560–561, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894564/pdf.

Higham:1966:SAA

[Hig66e] Robin Higham. Slipstream: The Autobiography of an Air Craftsman by Eugene E. Wilson (review). Technology and Culture, 7(3):441–443, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894605/
pdf.

Higham:1968:BMF

[Hig68a] Robin Higham. Bold Men, Far Horizons: Pioneer Pilots and Their Flights by Herbert Molloy Mason, Jr. (review). Technology and Culture, 9(2):247–248, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894319/pdf.

Higham:1968:DRB

[Hig68b] Robin Higham. Defense of the Realm: British Strategy in the Nuclear Epoch by R. N. Rosecrance (review). Technology and Culture, 9(4):620–621, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894236/pdf.

Higham:1968:GIT

[Hig68c] Robin Higham. Government, industry, and technological change: The SHOT-OAH program April 1967. Technology and Culture, 9(1):63–69, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894345/
pdf.

Higham:1969:CBA

[Hig69a] Robin Higham. A Century of British Aeronautics: The Centenary Journal of the Royal Aeronautical Society, 1866–1966 by Royal Aeronautical Society (review). Technology and Culture, 10(2):323–324, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892377/
pdf.

Higham:1969:QWS

[Hig69b] Robin Higham. Qantas at War by Sir Hudson Fysh (review). Technology and Culture, 10(1):119–120, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892338/pdf.

Hindle:1962:AWS

[Hin62] Brooke Hindle. Adventures in the World of Science by Charles Greeley Abbot (review). Technology and Culture, 3(3):334–335, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895143/pdf.

Hinkle:1964:SHS

[Hin64] Roscoe C. Hinkle. A Short History of Sociology by Heinz Maus (review). Technology and Culture, 5(1):123–125, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894948/pdf.

Hine:1966:EAM

[Hin66] Robert V. Hine. Exploration of Alaska, 1865–1900 by Morgan B. Sherwood (review). Technology and Culture, 7(1):108, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893879/pdf.

Hippen:1967:JHS

[Hip67] James C. Hippen. James Hargreaves and the Spinning Jenny by C. Aspin, S. D. Chapman (review). Technology and Culture, 8(1):95–97, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894499/
pdf.

Hippen:1969:ITM

[Hip69] James C. Hippen. Industrial textile machinery: Five North American museums. Technology and Culture, 10(4):570–586, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892404/pdf.

Hirschfeld:1964:TIM

[Hir64] Gerhard Hirschfeld. Technology and the idea of mankind. Technology and Culture, 5(1):43–56, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894921/pdf.

Hoffman:1964:AES

[Hof64] Myron A. Hoffman. Astronautical Engineering and Science ed. by E. Stuhlinger, et al. (review). Technology and Culture, 5(2):264–265, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894906/
pdf.

Hoffman:1965:AAN

[Hof65] Nathan J. Hoffman. Astronautics and Aeronautics, 1963 by National Aeronautics and Space Administration (review). Technology and Culture, 6(4):669–670, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894659/pdf.

Holley:1961:AHM

[Hol61a] I. B. Holley, Jr. Armor: a History of Mechanized Forces by Richard M. Ogorkiewicz (review). Technology and Culture, 2(4):415–416, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895446/pdf.

Hollon:1961:CMA

[Hol61b] W. Eugene Hollon. Colt: The Man, The Arms, The Company by James L. Mitchell (review). Technology and Culture, 2(1):64–65, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895588/
pdf.

Holley:1965:WAP

[Hol65a] I. B. Holley, Jr. The Weapons Acquisition Process: Economic Incentives by Frederic M. Scherer (review). Technology and Culture, 6(1):152–154, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894808/pdf.

Holley:1965:HML

[Hol65b] I. B. Holley, Jr. This High Man: The Life of Robert H. Goddard by Milton Lehman (review). Technology and Culture, 6(3):482–484, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894701/
pdf.

Holmes:1965:LSB

[Hol65c] Frederic L. Holmes. The Lunar Society of Birmingham: a Social History of Provincial Science and Industry in 18th Century England by Robert E. Schofield (review). Technology and Culture, 6(1):112–115, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894785/
pdf.

Holley:1969:HAA

[Hol69] I. B. Holley. History of the American Aircraft Industry: an Anthology ed. by G. R. Simonson (review). Technology and Culture, 10(2):324–326, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892378/pdf.

Hoopes:1962:GEU

[Hoo62] Thomas T. Hoopes. Die geschichtliche Entwicklung des Umformens in Gesenken by Ernst von Wedel (review). Technology and Culture, 3(1):89–91, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895417/pdf.

Hoopes:1964:STS

[Hoo64] Thomas T. Hoopes. The Scent of Time by Silvio A. Bedini (review). Technology and Culture, 5(3):446–449, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894870/pdf.

Hoopes:1965:EGA

[Hoo65] Thomas T. Hoopes. Early Greek Armour and Weapons by Anthony Snodgrass (review). Technology and Culture, 6(4):649–650, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894648/pdf.

Hoopes:1966:TWP

[Hoo66a] Penrose R. Hoopes. Technology of Watch Production by S. V. Tarasov (review). Technology and Culture, 7(1):95–96, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893872/pdf.

Hoopes:1966:GRW

[Hoo66b] Thomas T. Hoopes. Guns and Rifles of the World by Howard L. Blackmore (review). Technology and Culture, 7(3):421–424, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894595/pdf.

Houston:1961:BRM

[Hou61] Charles O. Houston, Jr. Black Rock: Mining Folklore of the Pennsylvania Dutch by George Korson (review). Technology and Culture, 2(4):408–409, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895443/pdf.

Houston:1963:MDS

[Hou63] Charles O. Houston, Jr. Metal Decorating from Start to Finishes by Charles R. Bragdon, and: Craft and Contemporary Culture by Seonaid Mairi Robertson (review). Technology and Culture, 4(2):219–221, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895010/
pdf.

Howland:1969:AEE

[How69] W. E. Howland. The argument: Engineering education for social leadership. Technology and Culture, 10(1):1–10, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892308/pdf.

Hartenberg:1969:EDO

[HS69] Richard S. Hartenberg and John Schmidt, Jr. The Egyptian drill and the origin of the Crank. Technology and Culture, 10(2):155–165, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892351/
pdf.

Hubbell:1964:EAR

[Hub64] Charles H. Hubbell. Early Aeroplanes, 1907–18 by Roy Cross (review). Technology and Culture, 5(1):98–99, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894933/pdf.

Hudson:1965:IAC

[Hud65a] Kenneth Hudson. The Industrial Archaeology of County Down by E. R. R. Green (review). Technology and Culture, 6(1):123–125, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894789/pdf.

Hudson:1965:GPI

[Hud65b] Kenneth Hudson. The growing pains of industrial archaeology. Technology and Culture, 6(4):621–626, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894643/pdf.

Hughes:1960:EBM

[Hug60] Thomas P. Hughes. Edison: a Biography by Matthew Josephson (review). Technology and Culture, 1(3):284–286, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895645/pdf.

Hughes:1961:CTH

[Hug61a] H. Stuart Hughes. Commentary: Technology and the history of ideas. Technology and Culture, 2(3):237–239, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895460/pdf.

Hughes:1961:AES

[Hug61b] Thomas P. Hughes. Atomic Energy in the Soviet Union by Arnold Kramish (review). Technology and Culture, 2(1):85–86, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895599/pdf.

Hughes:1961:ETB

[Hug61c] Thomas P. Hughes. Elihu Thomson: Beloved Scientist by David O. Woodbury (review). Technology and Culture, 2(4):419–420, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895448/pdf.

Hughes:1962:EMP

[Hug62a] Thomas Hughes. The Edison Motion Picture Myth by Gordon Hendricks (review). Technology and Culture, 3(3):335–337, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895144/pdf.

Hughes:1962:ODE

[Hug62b] Thomas P. Hughes. Oersted and the Discovery of Electromagnetism by Bern Dibner (review). Technology and Culture, 3(1):92–93, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895419/pdf.

Hughes:1962:BEI

[Hug62c] Thomas Parke Hughes. British electrical industry lag: 1882–1888. Technology and Culture, 3(1):27–44, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895412/pdf.

Hughes:1963:ABT

[Hug63] Thomas P. Hughes. American and British Technology in the Nineteenth Century: The Search for Labour-Saving Inventions by H. J.

Habakkuk (review). Technology and Culture, 4(1):67–69, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895040/pdf.

Hughes:1966:HGF

[Hug66] Thomas P. Hughes. Hitler, I. G. Farben, and synthetic gasoline. Technology and Culture, 7(4):514, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894537/pdf.

Hughes:1967:HMJ

[Hug67a] Thomas Parke Hughes. Historia mundi: Das 19. und 20. Jahrhundert ed. by Fritz Valjavec (review). Technology and Culture, 8(1):153–155, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894528/
pdf.

Hughes:1967:SRD

[Hug67b] Thomas Parke Hughes. Soviet Research and Development: Its Organization, Personnel, and Funds by Alexander G. Korol (review). Technology and Culture, 8(1):132–135, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894519/pdf.

Hughes:1968:AIL

[Hug68] Thomas P. Hughes. Army, Industry, and Labor in Germany, 1914– 1918 by Gerald D. Feldman (review). Technology and Culture, 9 (1):128–129, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894368/pdf.

Hughes:1969:WSW

[Hug69] Thomas P. Hughes. Werner von Siemens: Ein Wegbereiter der deutschen Industrie by Sigfrid von Weiher (review). Technology and Culture, 10(4):614–615, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892414/pdf.

Hunsaker:1962:TIS

[Hun62] James C. Hunsaker. Triumph: The Incredible Saga of the First Transatlantic Flight by Hy Steirman and Glenn D. Kittler (review). Technology and Culture, 3(3):363–364, Summer 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895156/pdf.

Hunsaker:1963:SGC

[Hun63a] J. C. Hunsaker. Sir George Cayley: The Inventor of the Aeroplane by J. Laurence Pritchard (review). Technology and Culture, 4(1):88–94, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895049/
pdf.

Hunter:1963:TCA

[Hun63b] Robert F. Hunter. Turnpike construction in Antebellum Virginia. Technology and Culture, 4(2):177–200, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895002/pdf.

Hunter:1964:MRW

[Hun64] Robert F. Hunter. Missiles and the Revolution in Warfare by Nels A. Parson, Jr. (review). Technology and Culture, 5(2):263–264, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894905/pdf.

Hunter:1967:LPA

[Hun67] Louis C. Hunter. The living past in the Appalachias of Europe: Water-mills in Southern Europe. Technology and Culture, 8(4):446– 466, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894397/pdf.

Hutchins:1962:C

[Hut62] Robert M. Hutchins. Commentary. Technology and Culture, 3(4):643– 646, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895121/
pdf.

Hutcheson:1963:BBV

[Hut63] J. A. Hutcheson. The Birth of Broadcasting. Vol. I of The History of Broadcasting in the United Kingdom by Asa Briggs (review). Technology and Culture, 4(1):103–104, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895054/pdf.

Hutcheson:1966:GAW

[Hut66] J. A. Hutcheson. The Golden Age of Wireless. Vol. II of The History of Broadcasting in the United Kingdom by Asa Briggs (review). Technology and Culture, 7(1):102–103, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893875/pdf.

Huxley:1962:C

[Hux62] Aldous Huxley. Commentary. Technology and Culture, 3(4):636–642, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895120/
pdf.

Ihde:1964:BRE

[Ihd64] Aaron J. Ihde. The Beginning of Rayon by Edwin J. Beer (review). Technology and Culture, 5(3):453–454, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894874/pdf.

Ihde:1965:PLL

[Ihd65] Aaron J. Ihde. Pesticides and the Living Landscape by Robert L. Rudd (review). Technology and Culture, 6(3):490–492, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894704/pdf.

Ihde:1967:EIS

[Ihd67] Aaron J. Ihde. Ecology and the Industrial Society ed. by Gordon T. Goodman, et al. (review). Technology and Culture, 8(1):150–151, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894526/
pdf.

Ihde:1969:SBC

[Ihd69] Aaron J. Ihde. Source Book in Chemistry, 1900–1950 ed. by Henry M. Leicester (review). Technology and Culture, 10(1):115–116, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892336/
pdf.

Isaac:1965:GMA

[IU65] Erich Isaac and Alvin W. Urquhart. The Geography of Modern Africa by William A. Hance (review). Technology and Culture, 6(2):296–298, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894757/
pdf.

Jacobs:1960:TP

[Jac60] Stephen W. Jacobs. Town planning. Technology and Culture, 1(4): 359–364, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895613/pdf.

Jacobeit:1964:RPW

[Jac64] Wolfgang Jacobeit. Research into peasant work and economy (agrarian ethnography) in the German democratic Republic. Technology and Culture, 5(3):379–385, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894855/pdf.

Jackson:1968:PFA

[Jac68] Joseph Gray Jackson. Privacy and Freedom by Alan F. Westin (review). Technology and Culture, 9(2):271–274, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894334/pdf.

Jacobeit:1969:AMA

[Jac69] Wolfgang Jacobeit. Acta Musaeorum Agriculturae ed. by V. Smelhaus and Z. Tempir (review). Technology and Culture, 10(3):482–484, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/901645/pdf.

Jaffe:1961:CWF

[Jaf61] A. J. Jaffe. The Crisis We Face; Automation and the Cold War by George Steele and Paul Kircher (review). Technology and Culture, 2(4):428–430, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895452/pdf.

Jaffe:1963:OCE

[Jaf63] William J. Jaffe. Organization Change: The Effect of Successful Leadership by Robert Henry Guest (review). Technology and Culture, 4(2):234–237, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895017/pdf.

Jaffe:1964:PPR

[Jaf64] William J. Jaffe. People and Productivity by Robert A. Sutermeister (review). Technology and Culture, 5(3):475–477, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894886/pdf.

Jaffe:1965:MEG

[Jaf65a] A. J. Jaffe. Manpower in Economic Growth: The American Record since 1800 by Stanley Lebergott (review). Technology and Culture, 6(2):282–284, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894749/pdf.

Jaffe:1965:OPM

[Jaf65b] William J. Jaffe. Organization for Profit: Management for the Age of Technology by Gerald G. Fisch (review). Technology and Culture, 6(3):513–515, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894716/pdf.

Jaffe:1967:RLM

[Jaf67] A. J. Jaffe. Retraining and Labor Market Adjustment in Western Europe by Margaret S. Gordon (review). Technology and Culture, 8(1):146–148, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894524/pdf.

Jaffe:1968:PDW

[Jaf68] A. J. Jaffe. Productivity Differences within the Service Sector by Victor R. Fuchs, Jean Alexander Wilburn (review). Technology and Culture, 9(2):262–263, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894329/pdf.

Jarvie:1966:SCT

[Jar66] I. C. Jarvie. The social character of technological problems: Comments on Skolimowski’s paper. Technology and Culture, 7(3):384–390, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894583/
pdf.

Jenkins:1967:KCT

[Jen67] Reese Jenkins. Kekulé Centennial by O. Theodor Benfey (review). Technology and Culture, 8(4):500–502, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894407/pdf.

Jenkins:1969:OOT

[Jen69] Reese V. Jenkins. Oxygen and Oxidation: Theories and Techniques in the 19th Century and the First Part of the 20th by Eduard Farber (review). Technology and Culture, 10(1):114–115, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892335/pdf.

Jewett:1964:TNS

[Jew64] Richard L. Jewett. Transactions of the Newcomen Society, Vol. XXXIII, 1960–1961 (review). Technology and Culture, 5(2):277–279, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894914/
pdf.

Jewett:1967:SAB

[Jew67] Robert A. Jewett. Structural antecedents of the I-beam, 1800–1850. Technology and Culture, 8(3):346–362, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894432/pdf. See note [Dor68b].

Jewett:1968:R

[Jew68] Robert A. Jewett. The response. Technology and Culture, 9(3):419– 426, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894258/pdf.

Jewett:1969:SPF

[Jew69] Robert A. Jewett. Solving the puzzle of the first American structural rail-beam. Technology and Culture, 10(3):371–391, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901619/pdf.

Johnson:1962:MNY

[Joh62] Arthur M. Johnson. Made in New York, Case Studies in Metropolitan Manufacturing by Roy B. Helfgott, et al. (review). Technology and Culture, 3(2):194–197, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895391/pdf.

Johnson:1965:IAE

[Joh65a] Arthur M. Johnson. Investment in Australian Economic Development, 1861–1900 by N. G. Butlin (review). Technology and Culture, 6(2):287–289, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894752/pdf.

Johnston:1965:NSF

[Joh65b] L. D. Johnston. Nuclear Secrecy and Foreign Policy by Harold L. Nieburg (review). Technology and Culture, 6(1):154–156, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894809/pdf.

Johnson:1969:SIR

[Joh69a] Arthur M. Johnson. The State, the Investor, and the Railroad: The Boston & Albany, 1825–1867 by Stephen Salsbury (review). Technology and Culture, 10(2):317–319, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892373/pdf.

Johnston:1969:LSS

[Joh69b] S. Paul Johnston. Over Land and Sea: The Story of Glenn H. Curtiss by Robert Scharff, Walter S. Taylor (review). Technology and Culture, 10(2):326–327, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892379/pdf.

Jolliffe:1968:BMA

[Jol68] A. W. Jolliffe. Book of Minerals by Albertus Magnus (review). Technology and Culture, 9(2):219–223, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894303/pdf.

Joravsky:1961:HTS

[Jor61] David Joravsky. The history of technology in Soviet Russia and Marxist doctrine. Technology and Culture, 2(1):5–10, Winter 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895570/pdf.

Jorgensen:1962:CCS

[Jor62] Paul A. Jorgensen. Castles and Cannon: a Study of Early Artillery Fortifications in England by B. H. St. J. O’Neil (review). Technology and Culture, 3(1):91–92, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895418/pdf.

Junker:1968:IAS

[Jun68] Louis J. Junker. International Aid: a Summary by Hugh L. Keenleyside (review). Technology and Culture, 9(4):629–631, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894240/pdf.

K:1961:GAT

[K.61] M. K. A Guide to American Trade Catalogs, 1744–1900 by Lawrence B. Romaine (review). Technology and Culture, 2(3):282–284, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895478/pdf.

Kamerschen:1963:ELU

[Kam63] David R. Kamerschen. Economic Liberalism and Underdevelopment: Studies in the Disintegration of an Idea by Frederick Clairmonte (review). Technology and Culture, 4(3):366–368, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894992/pdf.

Kann:1965:CSA

[Kan65] Henry Robert Kann. The Chicago School of Architecture by Carl W. Condit (review). Technology and Culture, 6(3):472–475, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894696/pdf.

Kannappan:1966:EIM

[Kan66] Subbiah Kannappan. Economics and the Idea of Mankind ed. by Bert F. Hoselitz (review). Technology and Culture, 7(1):129–131, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/893844/pdf.

Kato:1963:BCH

[Kat63] Hidetoshi Kato. The Beer Can by the Highway by John A. Kouwenhoven (review). Technology and Culture, 4(1):116–117, Winter 1963.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895060/pdf.

Keachie:1968:TEG

[Kea68] E. C. Keachie. Technology, Economic Growth, and Public Policy by Richard R. Nelson, et al. (review). Technology and Culture, 9(2):263– 267, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894330/pdf.

Keller:1965:IRB

[Kel65] Alex Keller. Les Ingénieurs de la Renaissance by Bertrand Gille (review). Technology and Culture, 6(2):271–274, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894743/pdf.

Keller:1968:RSC

[Kel68] A. G. Keller. The Royal Society: Concept and Creation by Margery Purver (review). Technology and Culture, 9(3):484–488, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894269/pdf.

Keller:1969:MBD

[Kel69a] Alex Keller. Le Machine: Bollettino dell’Istituto Italiano per la Storia della Tecnica ed. by G. Polvani (review). Technology and Culture, 10(1):84–86, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892319/pdf.

Keller:1969:CTP

[Kel69b] Alex Keller. The Craftsmen of Tula: Pioneer Builders of Waterdriven Machinery by A. S. Britkin (review). Technology and Culture, 10(4):608–610, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892412/pdf.

Keller:1969:IPC

[Kel69c] Alex Keller. The Idea of Progress in Classical Antiquity by Ludwig Edelstein (review). Technology and Culture, 10(1):87–90, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892321/pdf.

Kerker:1961:SSE

[Ker61] Milton Kerker. Science and the steam engine. Technology and Culture, 2(4):381–390, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895436/pdf.

Kohlmeyer:1961:SEA

[KH61] Fred W. Kohlmeyer and Floyd L. Herum. Science and engineering in agriculture: a historical perspective. Technology and Culture, 2(4):368–380, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895435/pdf.

Kilgour:1962:CIM

[Kil62] Frederick G. Kilgour. A Catalogue of Instruments and Models in the Possession of the American Philosophical Society by Robert P. Multhauf (review). Technology and Culture, 3(2):184–185, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895385/pdf.

Kilgour:1963:VEH

[Kil63] Frederick G. Kilgour. Vitruvius and the early history of wave theory. Technology and Culture, 4(3):282–286, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894972/pdf.

Kilgour:1964:SHE

[Kil64] Frederick G. Kilgour. A Social History of Engineering by W. H. G. Armytage (review). Technology and Culture, 5(1):68, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894923/pdf.

Kimball:1964:E

[Kim64] Charles Kimball. Epilogue. Technology and Culture, 5(3):358, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894852/
pdf.

King:1966:PBR

[Kin66] W. James King. Popov and the Beginnings of Radiotelegraphy by Charles Susskind, and: The Invention of the Traveling-Wave Tube by Rudolf Kompfner, and: On the Early History of Radio Guidance by Benjamin Franklin Miessner (review). Technology and Culture, 7(1):98–102, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893874/pdf.

Kingslake:1968:CZE

[Kin68] R. Kingslake. Carl Zeiss und Ernst Abbe: Ihr Leben und ihr Werk by Harald Volkmann (review). Technology and Culture, 9(2):235–236, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894311/
pdf.

Klein:1962:LHE

[Kle62] Stanley Klein. The Legacy of Hiroshima by Edward Teller with Allen Brown (review). Technology and Culture, 3(3):359–362, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/3100847.
pdf;   https://muse.jhu.edu/pub/1/article/895155/pdf.

Klein:1963:NWV

[Kle63a] Stanley Klein. The New World, 1939–1946. Volume I of A History of the United States Atomic Energy Commission by Richard G. Hewlett and Oscar E. Anderson, Jr. (review). Technology and Culture, 4(2):227–231, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895014/pdf.

Klemperer:1963:OLE

[Kle63b] Wolfgang B. Klemperer. Otto Lilienthal, Der erste Flieger by Gerhard Halle (review). Technology and Culture, 4(1):94–98, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895050/pdf.

Klein:1964:ERM

[Kle64] Stanley Klein. Environmental Radioactivity by Merril Eisenbud (review). Technology and Culture, 5(1):101–103, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894935/pdf.

Klein:1965:WME

[Kle65] Stanley Klein. Wanted: More engineers in public affairs: The EJC symposium on national engineering problems. Technology and Culture, 6(3):416–422, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894676/pdf.

Klemm:1967:SWH

[Kle67] Friedrich Klemm. Scales and Weights: a Historical Outline by Bruno Kisch (review). Technology and Culture, 8(3):402–403, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894438/pdf.

Knights:1965:UPU

[Kni65] Peter R. Knights. The Universal Postal Union: Coordinator of the International Mails by George A. Codding, Jr. (review). Technology and Culture, 6(3):480–481, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894699/pdf.

Knights:1967:TBH

[Kni67] Peter R. Knights. A Tower in Babel: a History of Broadcasting in the United States. Vol. I: To 1933 by Erik Barnouw (review). Technology and Culture, 8(4):510–511, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894413/pdf.

Kohlmeyer:1964:PCB

[Koh64] Fred W. Kohlmeyer. From Prairie to Corn Belt: Farming on the Illinois and Iowa Prairies in the Nineteenth Century by Allan G. Bogue (review). Technology and Culture, 5(3):454–455, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894875/pdf.

Kohlmeyer:1966:HGA

[Koh66] Fred W. Kohlmeyer. A History of Georgia Agriculture, 1732–1860 by James C. Bonner (review). Technology and Culture, 7(1):105–106, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893877/
pdf.

Kohlmeyer:1967:EDA

[Koh67] Fred W. Kohlmeyer. Economic Development of Agriculture: The Modernization of Farming by Iowa State University Center for Agricultural and Economic Development (review). Technology and Culture, 8(3):415–416, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894447/pdf.

Kohlmeyer:1968:EDF

[Koh68] Fred W. Kohlmeyer. The English Dairy Farmer 1500–1900 by G. E. Fussell (review). Technology and Culture, 9(1):117–118, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894360/pdf.

Kohlmeyer:1969:CSS

[Koh69] Fred W. Kohlmeyer. Claus Spreckels: The Sugar King in Hawaii by Jacob Adler (review). Technology and Culture, 10(2):320–321, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892375/pdf.

Kolnicker:1962:MMS

[Kol62] Morris Kolnicker. Machines and the Man: a Sourcebook on Automation ed. by Robert P. Weeks, and: Automation: Its Impact on Business and People by Walter Buckingham (review). Technology and Culture, 3(2):203–207, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895395/pdf.

Kolnicker:1964:ATC

[Kol64] Morris Kolnicker. Automation and Technological Change ed. by John T. Dunlop (review). Technology and Culture, 5(1):103–104, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894936/pdf.

Kolnicker:1966:WNF

[Kol66] Morris Kolnicker. Wheels for a Nation by Frank Donovan (review). Technology and Culture, 7(4):551–552, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894560/pdf.

Kouwenhoven:1961:HBW

[Kou61] John A. Kouwenhoven. Highways Over Broad Waters: Life and Times of David B. Steinman, Bridgebuilder by William Ratigan (review). Technology and Culture, 2(2):204–205, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895566/pdf.

Kouwenhoven:1965:WPO

[Kou65] John A. Kouwenhoven. Winterthur Portfolio One (review). Technology and Culture, 6(2):322–323, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894771/pdf.

Kouwenhoven:1966:BBF

[Kou66] John A. Kouwenhoven. Brooklyn Bridge: Fact and Symbol by Alan Trachtenberg (review). Technology and Culture, 7(1):109–111, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/893880/pdf.

Kranzberg:1960:CSH

[Kra60] Melvin Kranzberg. Charles Singer and “A history of technology”. Technology and Culture, 1(4):299–302, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895605/pdf.

Kranzberg:1962:C

[Kra62] Melvin Kranzberg. Commentary. Technology and Culture, 3(4):519– 523, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895108/
pdf.

Kranzberg:1964:IRE

[Kra64] Melvin Kranzberg. The Industrial Revolution in the Eighteenth Century by Paul Mantoux (review). Technology and Culture, 5(1):81–82, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894927/
pdf.

Kramer:1966:LBB

[Kra66] Samuel Noah Kramer. Land behind Baghdad: a History of Settlement on the Diyala Plains by Robert McC. Adams (review). Technology and Culture, 7(1):74–75, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893859/pdf.

Kraus:1968:BET

[Kra68] Jerome Kraus. The British electron-tube and semiconductor industry, 1935–62. Technology and Culture, 9(4):544–561, October 1968.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894213/pdf.

Kretzmann:1967:STA

[Kre67] Edwin M. J. Kretzmann. Science, Technology, and American Foreign Policy by Eugene B. Skolnikoff (review). Technology and Culture, 8(4):517–519, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894418/pdf.

Kunik:1960:PAT

[Kun60] I. Jordan Kunik. A patent attorney takes issue. Technology and Culture, 1(3):221–227, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895630/pdf.

Kurland:1961:SUM

[Kur61a] Norman Kurland. Slaves Unaware? A Mid-Century View of Applied Science by Magnus Pyke (review). Technology and Culture, 2(2):202–203, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895565/pdf.

Kurland:1961:LAS

[Kur61b] Norman D. Kurland. Leisure in America: a Social Inquiry by Max Kaplan (review). Technology and Culture, 2(4):431–433, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895453/pdf.

Kurland:1963:TWL

[Kur63] Norman D. Kurland. Of Time, Work, and Leisure by Sebastian de Grazia (review). Technology and Culture, 4(3):370–372, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894995/pdf.

Kyropoulos:1964:AQR

[Kyr64] Peter Kyropoulos. Automobile Quarterly (review). Technology and Culture, 5(1):97–98, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894932/pdf.

LaForce:1964:TDC

[La 64] J. Clayburn La Force. Technological diffusion in the 18th century: The Spanish textile industry. Technology and Culture, 5(3):322–343, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894850/
pdf.

Lampman:1961:BSR

[Lam61] L. R. Lampman. British Shipping by R. H. Thornton (review). Technology and Culture, 2(1):73–74, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895593/pdf.

Langdon:1968:HPF

[Lan68] George D. Langdon, Jr. Husbandmen of Plymouth: Farms and Villages in the Old Colony, 1620–1692 by Darrett B. Rutman (review). Technology and Culture, 9(4):605–606, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894225/pdf.

Lasby:1966:COJ

[Las66] Clarence G. Lasby. Crossbow and Overcast by James McGovern (review). Technology and Culture, 7(4):561–564, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894565/pdf.

Lasby:1967:MND

[Las67] Clarence G. Lasby. The Mare’s Nest by David Irving (review). Technology and Culture, 8(3):429–431, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894453/pdf.

Layton:1966:RTS

[Lay66] Edwin Layton. The Rise of the Technocrats: a Social History by W. H. G. Armytage (review). Technology and Culture, 7(4):532–534, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894550/
pdf.

Layton:1967:TEJ

[Lay67] Edwin Layton. The Technical Elite by Jay M. Gould (review). Technology and Culture, 8(2):240–242, April 1967. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894474/pdf.

Layton:1968:TPA

[Lay68] Edwin T. Layton. The Technocrats, Prophets of Automation by Henry Eisner, Jr. (review). Technology and Culture, 9(2):256–259, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894326/pdf.

Layton:1969:TVC

[Lay69] Edwin Layton. Thorstein Veblen: The Carleton College Veblen Seminar Essays ed. by Carlton C. Qualey (review). Technology and Culture, 10(4):616–617, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892416/pdf.

Levey:1960:CT

[LD60] Martin Levey and Denis I. Duveen. Chemical technology. Technology and Culture, 1(4):371–376, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895615/pdf.

LeCorbeiller:1963:SMM

[Le 63] Jean Le Corbeiller. The Science Myth by Magnus Pyke (review). Technology and Culture, 4(2):214–215, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895006/pdf.

LeCorbeiller:1964:TCS

[Le 64] Jean Le Corbeiller. Two Cultures? The Significance of C. P. Snow by F. R. Leavis (review). Technology and Culture, 5(1):119–120, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894945/pdf.

LeCorbeiller:1965:WTH

[Le 65] Jean Le Corbeiller. World Technology and Human Destiny ed. by Raymond Aron (review). Technology and Culture, 6(1):144–145, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894802/pdf.

Leavitt:1966:IAL

[Lea66] Thomas W. Leavitt. Industrial archaeology in Lawrence, Massachusetts. Technology and Culture, 7(2):201, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895173/pdf.

Leavitt:1967:MVT

[Lea67] Thomas W. Leavitt. The Merrimack Valley Textile Museum. Technology and Culture, 8(2):204–206, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894460/pdf.

Leavitt:1968:TSE

[Lea68] Thomas W. Leavitt. Toward a standard of excellence: The nature and purpose of exhibit reviews. Technology and Culture, 9(1):70–75, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894346/
pdf.

Leavitt:1969:FYI

[Lea69] Thomas W. Leavitt. Five years of “Industrial archaeology”. Technology and Culture, 10(4):587–592, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892405/pdf.

Leicester:1961:HGF

[Lei61a] Henry M. Leicester. History of Greek Fire and Gunpowder by J. R. Partington (review). Technology and Culture, 2(3):272–273, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895472/pdf.

Leicester:1961:CCT

[Lei61b] Henry M. Leicester. Chemistry, chemical technology, and scientific progress. Technology and Culture, 2(4):352–356, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895433/pdf.

Leicester:1963:MAB

[Lei63] Henry M. Leicester. Mediaeval Arabic Bookmaking and Its Relation to Early Chemistry and Pharmacology by Martin Levey (review). Technology and Culture, 4(3):343–344, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894980/pdf.

Leighton:1967:MCI

[Lei67] Albert C. Leighton. The mule as a cultural invention. Technology and Culture, 8(1):45–52, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894488/pdf.

Leighton:1969:SCA

[Lei69] Albert C. Leighton. Secret communication among the Greeks and Romans. Technology and Culture, 10(2):139–154, April 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892350/pdf.

Lenzi:1960:SDT

[Len60] Giulio Lenzi. Storia della Tecnica dal Medioevo al Rinascimento by Umberto Forti. Technology and Culture, 1(4):419–420, Fall 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895624/pdf.

Leser:1960:APT

[Les60] Paul Leser. Africa: Its Peoples and Their Culture History by George Peter Murdock (review). Technology and Culture, 1(3):256–259, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895634/
pdf.

Levey:1961:SEP

[Lev61a] Martin Levey. Studies in Early Petroleum History by R. J. Forbes, and: More Studies in Early Petroleum History, 1860–1880 by R. J. Forbes (review). Technology and Culture, 2(1):55–57, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895584/pdf.

Levinson:1961:CMC

[Lev61b] Mark Levinson. Classical Mathematics: a Concise History of the Classical Era in Mathematics by Joseph Ehrenfried Hofmann (review). Technology and Culture, 2(1):86–88, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895600/pdf.

Lewis:1964:MSS

[Lew64] W. David Lewis. Managers and Scientists: Some Human Problems in Industrial Research Organizations by Ralph M. Hower and Charles D. Orth, III (review). Technology and Culture, 5(3):473–475, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894885/pdf.

Lewis:1965:ARD

[Lew65] W. David Lewis. Administering Research and Development: The Behavior of Scientists and Engineers in Organizations by Charles D. Orth III, et al. (review). Technology and Culture, 6(2):310–312, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894765/pdf.

Lewis:1966:ORE

[Lew66] W. David Lewis. The Organization of Research Establishments ed. by Sir John Cockcroft (review). Technology and Culture, 7(4):566–568, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894567/
pdf.

Lewis:1967:TUS

[Lew67] W. David Lewis. Technology and urbanization: The SHOT-OAH program April 1966. Technology and Culture, 8(1):67–73, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894490/pdf.

Lewis:1969:CMG

[Lew69a] W. David Lewis. Corning Museum of Glass by Paul N. Perrot, Kenneth M. Wilson (review). Technology and Culture, 10(1):68–83, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892318/
pdf.

Lewis:1969:PP

[Lew69b] W. David Lewis. Parrying perrot. Technology and Culture, 10(2):178– 181, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892355/pdf.

Livingston:1969:TCH

[Liv69] H. K. Livingston. Technology and the city: Human and social dimensions Wayne State University, May 1968. Technology and Culture, 10(1):59–61, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892315/pdf.

Lockheimer:1969:RNS

[Loc69] F. Roy Lockheimer. Reviews of National Science Policy: Japan by OECD (review). Technology and Culture, 10(2):336–338, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892385/pdf.

Lokke:1968:FNH

[Lok68] Virgil L. Lokke. The Future as Nightmare: H. G. Wells and the Anti-Utopians by Mark R. Hillegas (review). Technology and Culture, 9(3):523–525, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894291/pdf.

Long:1965:PJE

[Lon65] T. D. Long. Paternalism in the Japanese Economy by John W. Bennett and Iwao Ishino (review). Technology and Culture, 6(3):512–513, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894715/
pdf.

Long:1966:OEM

[Lon66] T. D. Long. The Origins of Entrepreneurship in Meiji Japan by Johannes Hirschmeier (review). Technology and Culture, 7(1):86–88, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893867/
pdf.

Long:1967:PCI

[Lon67] T. D. Long. Political Change and Industrial Development in Japan: Government Enterprise, 1868–1880 by Thomas C. Smith (review). Technology and Culture, 8(1):122–124, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894515/pdf.

Long:1969:PSC

[Lon69] T. D. Long. The Politics of Space Cooperation by Don E. Kash (review). Technology and Culture, 10(1):126, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892341/pdf.

Loria:1967:LVD

[Lor67a] Mario Loria. L’agricoltura veneta dalla caduta della Repubblica all’Unità by Marino Berengo (review). Technology and Culture, 8 (4):495–497, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894404/pdf.

Loria:1967:CDF

[Lor67b] Mario Loria. Cavour and the development of the fertilizer industry in Piedmont. Technology and Culture, 8(2):159–177, April 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894457/pdf.

Lubove:1965:MSS

[Lub65a] Roy Lubove. Man’s Struggle for Shelter in an Urbanizing World by Charles Abrams (review). Technology and Culture, 6(3):488–490, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894703/
pdf.

Lubove:1965:RTE

[Lub65b] Roy Lubove. Region in Transition: Economic Study of the Pittsburgh Region (review). Technology and Culture, 6(3):485–488, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894702/
pdf.

Lynd:1967:TVD

[Lyn67] Staughton Lynd. Thorstein Veblen by Douglas F. Dowd (review). Technology and Culture, 8(1):135–136, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894520/pdf.

Mahler:1961:ABA

[Mah61] Harry Mahler. American Building Art: The Nineteenth Century by Carl W. Condit (review). Technology and Culture, 2(4):402–403, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895439/pdf.

Malina:1964:TK

[Mal64] Frank J. Malina. Theodore Von Kármán, 1881–1963. Technology and Culture, 5(2):241–246, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894896/pdf.

Mantel:1965:PGV

[Man65] Samuel J. Mantel, Jr. A Programme for Growth. Vol. I. A Computable Model of Economic Growth ed. by Richard Stone and Alan Brown (review). Technology and Culture, 6(4):677–679, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894663/pdf.

Marshall:1963:PXT

[Mar63] James L. Marshall. Pius XII and Technology comp. by Rev. Leo J. Haigerty (review). Technology and Culture, 4(2):241–243, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895020/pdf.

Marcson:1969:NBS

[Mar69a] Simon Marcson. The New Brahmins by Spencer Klaw (review). Technology and Culture, 10(3):473–478, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/901642/pdf.

Marx:1969:FSL

[Mar69b] Leo Marx. Flesh of Steel: Literature and the Machine in American Culture by Thomas Reed West (review). Technology and Culture, 10(1):111–113, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892333/pdf.

Mawardi:1965:ICP

[Maw65] Betty Hosmer Mawardi. Industrial Creativity: The Psychology of the Inventor by Joseph Rossman (review). Technology and Culture, 6(3):523–525, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894721/pdf.

Mayr:1967:STB

[May67a] Otto Mayr. Der synthetische Treibstoff 1933–1945. Ein Beitrag zur nationalsozialistischen Wirtschafts- und Rüstungspolitik by Wolfgang Birkenfeld (review). Technology and Culture, 8(3):425–429, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894452/pdf.

Mayr:1967:TVG

[May67b] Otto Mayr. Die Technik der Vorzeit, der geschichtlichen Zeit und der Naturvölker by Franz Maria Feldhaus (review). Technology and Culture, 8(2):213–214, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894463/pdf.

Mayr:1968:BTE

[May68] Otto Mayr. Blätter für Technikgeschichte ed. by Josef Nagler (review). Technology and Culture, 9(3):525–526, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894292/pdf.

Mayfield:1969:GGB

[May69a] Robert C. Mayfield. From Geography to Geotechnics by Benton MacKaye (review). Technology and Culture, 10(4):622–623, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892419/pdf.

Mayo:1969:SPU

[May69b] Louis H. Mayo. Science Policy and the University ed. by Harold Orlans (review). Technology and Culture, 10(2):338–341, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892386/pdf.

Mayr:1969:FPJ

[May69c] Otto Mayr. Die französische Pendule des 18. Jahrhunderts. Ein Beitrag zu ihrer Ikonologie by Klaus Maurice (review). Technology and Culture, 10(1):99–100, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892326/pdf.

Mayr:1969:HTG

[May69d] Otto Mayr. Herausforderung der Technik: Gesellschaftliche Voraussetzungen und Wirkungen der technischen Entwicklung by Klaus Tuchel (review). Technology and Culture, 10(2):349–350, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892392/pdf.

Mazlish:1967:FD

[Maz67] Bruce Mazlish. The fourth discontinuity. Technology and Culture, 8(1):1–15, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894486/pdf.

McCarthy:1963:NAD

[McC63] Charles J. McCarthy. Naval aircraft design in the mid-1930’s. Technology and Culture, 4(2):165–174, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895000/pdf.

McCutcheon:1965:INN

[McC65] W. A. McCutcheon. Inland navigations of the North of Ireland. Technology and Culture, 6(4):596–620, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894642/pdf.

McCollum:1966:HWE

[McC66a] Sylvia G. McCollum. Hours of Work ed. by Clyde E. Dankert, et al. (review). Technology and Culture, 7(3):438–440, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894603/pdf.

McCollum:1966:MEC

[McC66b] Sylvia G. McCollum. Manpower and Education: Country Studies in Economic Development ed. by Frederick Harbison and Charles A. Meyers (review). Technology and Culture, 7(1):126–129, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893843/pdf.

McCollum:1967:MPY

[McC67] Sylvia G. McCollum. Manpower Policies for Youth ed. by Eli E. Cohen, Louis Kapp (review). Technology and Culture, 8(4):537–539, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894426/
pdf.

McCutcheon:1969:IAT

[McC69] W. A. McCutcheon. Industrial archaeology and technological conservation in Northern Ireland. Technology and Culture, 10(3):412–421, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/901621/
pdf.

McGuigan:1965:CSW

[McG65] Gerald F. McGuigan. The Changing Structure of the Western Economy by Sir Douglas Copland (review). Technology and Culture, 6 (2):291–293, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894754/pdf.

McHugh:1961:SAN

[McH61] Jeanne McHugh. Steelworkers in America: The Nonunion Era by David Brody (review). Technology and Culture, 2(4):409–411, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895444/pdf.

McHugh:1963:ILG

[McH63] Jeanne McHugh. The Iron Library of the George Fischer Company. Technology and Culture, 4(3):331–335, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894976/pdf.

McHugh:1965:FIM

[McH65] Jeanne McHugh. Frontier Iron: The Maramec Iron Works, 1826– 1876 by James D. Norris (review). Technology and Culture, 6(4):653– 657, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894651/
pdf.

McHale:1966:VVE

[McH66a] John McHale. Vision and Value ed. by Gyorgy Kepes (review). Technology and Culture, 7(1):132–134, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893846/pdf.

McHugh:1966:AIS

[McH66b] Jeanne McHugh. The Australian Iron and Steel Industry, 1848–1962 by Helen Hughes (review). Technology and Culture, 7(1):82–84, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893865/
pdf.

McKelvey:1964:HCE

[McK64a] Blake McKelvey. The Historian and the City ed. by Oscar Handlin and John Burchard (review). Technology and Culture, 5(4):620–624, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894837/
pdf.

McKim:1964:HEG

[McK64b] Robert H. McKim. Human Engineering Guide to Equipment Design ed. by Clifford T. Morgan, et al. (review). Technology and Culture, 5(1):112–114, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894941/pdf.

McKim:1965:HFE

[McK65] R. H. McKim. Human Factors Engineering by Ernest J. McCormick (review). Technology and Culture, 6(3):520–521, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894719/pdf.

McNett:1964:LBF

[McN64] Charles McNett, Jr. Low Bridge!: Folklore and the Erie Canal by Lionel D. Wyld (review). Technology and Culture, 5(4):612–613, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894832/pdf.

Mechanic:1966:JLS

[Mec66] David Mechanic. Journey to Labor: a Study of Births in Hospitals and Technology by Donnell M. Pappenfort (review). Technology and Culture, 7(1):134–135, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893847/pdf.

Mehl:1961:HMD

[Meh61] Robert F. Mehl. A History of Metallography: The Development of Ideas on the Structure of Metals before 1890 by Cyril Stanley Smith (review). Technology and Culture, 2(3):266–269, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895469/pdf.

Meighan:1960:BT

[Mei60] Clement W. Meighan. The beginnings of technology. Technology and Culture, 1(4):334–339, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/895609/pdf.

Meier:1962:C

[Mei62a] Richard L. Meier. Commentary. Technology and Culture, 3(4):554– 557, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895112/
pdf.

Meighan:1962:DCE

[Mei62b] Clement W. Meighan. The Dawn of Civilization ed. by Stuart Piggott (review). Technology and Culture, 3(3):310–311, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895132/pdf.

Meighan:1966:HMV

[Mei66a] Clement W. Meighan. History of Mankind, Vol. II: The Ancient World by Luigi Pareti (review). Technology and Culture, 7(4):519– 520, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894540/
pdf.

Meighan:1966:ARE

[Mei66b] Clement W. Meighan. The Architecture of the Roman Empire. I. An Introductory Study by William L. MacDonald (review). Technology and Culture, 7(4):520–521, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894541/pdf.

Meighan:1968:AGG

[Mei68] Clement W. Meighan. Ancient Greek Gadgets and Machines by Robert S. Brumbaugh (review). Technology and Culture, 9(2):218– 219, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894302/pdf.

Mendelsohn:1966:SNC

[Men66] Everett Mendelsohn. Science in the Nineteenth Century America: a Documentary History ed. by Nathan Reingold (review). Technology and Culture, 7(4):543–544, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894555/pdf.

Merdinger:1965:IHC

[Mer65] Charles J. Merdinger. An Illustrated History of Civil Engineering by J. P. M. Pannell (review). Technology and Culture, 6(3):447–448, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894682/
pdf.

Mesthene:1965:UCH

[Mes65] Emmanuel G. Mesthene. On understanding change: The Harvard University Program on Technology and Society. Technology and Culture, 6(2):222–235, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894734/pdf.

Mesthene:1966:EUH

[Mes66] Emmanuel G. Mesthene. An experiment in understanding: The Harvard program two years after. Technology and Culture, 7(4):475–492, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894531/
pdf.

Mesthene:1969:CC

[Mes69a] Emmanuel G. Mesthene. A comment on the comments. Technology and Culture, 10(4):535–536, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892399/pdf.

Mesthene:1969:SGI

[Mes69b] Emmanuel G. Mesthene. Some general implications of the research of the Harvard University Program on Technology and Society. Technology and Culture, 10(4):489–513, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892423/pdf.

Meursinge:1964:HRU

[Meu64] John H. Meursinge. Human Resources in the Urban Economy ed. by Mark Perlman (review). Technology and Culture, 5(4):625, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894838/pdf.

Meursinge:1966:OHT

[Meu66] John H. Meursinge. Overlapping histories of technology. Technology and Culture, 7(4):517–518, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894539/pdf.

Meyerhof:1960:CBR

[Mey60] G. G. Meyerhof. The Canal Builders by Robert Payne (review). Technology and Culture, 1(2):185–186, Spring 1960. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895256/pdf.

Michael:1964:WNP

[Mic64] Donald N. Michael. Who Needs People? Automation and Your Future by Robert E. Cubbedge (review). Technology and Culture, 5 (1):104–106, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894937/pdf.

Michael:1966:AAS

[Mic66] Donald N. Michael. The Age of Automation by Sir Leon Bagrit (review). Technology and Culture, 7(2):257–261, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895194/pdf.

Middleton:1963:STA

[Mid63] W. E. K. Middleton. Science and Technology by Andrew G. van Melsen (review). Technology and Culture, 4(1):120–122, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895062/pdf.

Middleton:1968:EMB

[Mid68] W. E. K. Middleton. The Evolution of the Microscope by S. Bradbury (review). Technology and Culture, 9(3):488–489, July 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894270/pdf.

Miller:1961:MSA

[Mil61] Genevieve Miller. Medicine and Society in America: 1660–1860 by Richard Harrison Shryock (review). Technology and Culture, 2 (3):277–278, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895475/pdf.

Miller:1962:WFI

[Mil62] Joseph A. Miller. Western Forest Industry. An Economic Outlook by John A. Guthrie and George A. Armstrong (review). Technology and Culture, 3(2):192–194, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895390/pdf.

Miles:1963:P

[Mil63] Wyndham D. Miles. The Polaris. Technology and Culture, 4(4):478– 489, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894965/
pdf.

Miller:1964:PCI

[Mil64] Stanley S. Miller. Promotion and Control of Industry in Postwar France by John Sheahan (review). Technology and Culture, 5(2):270– 271, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894910/pdf.

Mingay:1965:AED

[Min65a] G. E. Mingay. Agriculture in Economic Development ed. by Carl K. Eicher and Lawrence Witt (review). Technology and Culture, 6(1):137–138, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894798/pdf.

Mingay:1965:CSC

[Min65b] G. E. Mingay. Capital, Saving and Credit in Peasant Societies ed. by Raymond Firth and B. S. Yamey (review). Technology and Culture, 6(2):293–294, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894755/pdf.

Moeller:1966:LOL

[Moe66] Walter O. Moeller. The Lanifricarius and the Officinae Lanifricariae at Pompeii. Technology and Culture, 7(4):493–496, Fall 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894532/pdf.

Moeller:1968:AP

[Moe68] Walter O. Moeller. The “Atramentarius” at Pompeii. Technology and Culture, 9(1):51–54, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894343/pdf.

Moeller:1969:MWP

[Moe69] Walter O. Moeller. The male weavers at Pompeii. Technology and Culture, 10(4):561–566, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892402/pdf.

Montgomery:1967:AIV

[Mon67] David Montgomery. Age of Industrial Violence, 1910–15: The Activities and Findings of the United States Commission on Industrial Relations by Graham Adams, Jr. (review). Technology and Culture, 8(2):234–237, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894471/pdf.

Morse:1961:PEP

[Mor61] Chandler Morse. Possibilities of Economic Progress by A. J. Youngson (review). Technology and Culture, 2(2):196–200, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895563/pdf.

Morris:1963:SCJ

[Mor63a] Bertram Morris. A Search for Civilization by John U. Nef (review). Technology and Culture, 4(3):369–370, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894994/pdf.

Morris:1963:TRS

[Mor63b] Bertram Morris. Toward a Reasonable Society by Clarence E. Ayres (review). Technology and Culture, 4(1):118–120, Winter 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895061/pdf.

Morehouse:1967:CFD

[Mor67] Ward Morehouse. Confronting a four-dimensional problem: Science, technology, society, and tradition in India and Pakistan. Technology and Culture, 8(3):363–375, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894433/pdf.

Morehouse:1969:MPI

[Mor69a] Ward Morehouse. Manufacturing Problems in India: The Cummins Diesel Experience by Jack Baranson (review). Technology and Culture, 10(2):331–335, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892383/pdf.

Morrison:1969:LMR

[Mor69b] Jean Morrison. L’Évolution de la maison rurale laurentienne by Georges Gauthier-Larouche (review). Technology and Culture, 10 (3):446–447, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901628/pdf.

Moszynski:1967:PWP

[Mos67] J. R. Moszynski. Przemysl Wiejski na Podhalu (“Rural Industry in the Polish Tatra Highlands”) by Jan Reychman, Stefan Reychman, and: Technika Motoryzacyjna w Polsce w Okresie Miedzywojennym (“Motorization in Poland between the Wars”) by Kazimier Groniowski (review). Technology and Culture, 8(3):410–411, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894444/pdf.

Multhauf:1962:HUS

[Mul62] Robert P. Multhauf. A History of the United States Weather Bureau by Donald R. Whitnah (review). Technology and Culture, 3 (3):332–333, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895142/pdf.

Multhauf:1965:MCH

[Mul65a] Robert P. Multhauf. A museum case history: The Department of Science and Technology of the United States Museum of History and Technology. Technology and Culture, 6(1):47–58, Winter 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894778/pdf.

Multhauf:1965:SAC

[Mul65b] Robert P. Multhauf. Sal ammoniac: a case history in industrialization. Technology and Culture, 6(4):569–586, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894639/pdf.

Mullen:1968:CCU

[Mul68] Pierce C. Mullen. The Convergent Century: The Unification of Science in the Nineteenth Century by Harold Sharlin (review). Technology and Culture, 9(1):111–113, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894358/pdf.

Mullen:1969:SSN

[Mul69] Jerry Mac Mullen. Ships That Sail No More; Marine Transportation from San Diego to Puget Sound, 1910–1940 by Giles T. Brown (review). Technology and Culture, 10(2):321–322, April 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892376/pdf.

Mumford:1960:TM

[Mum60] Lewis Mumford. Tools and the man. Technology and Culture, 1(4): 320–334, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895608/pdf.

Mumford:1961:HNC

[Mum61] Lewis Mumford. History: Neglected clue to technological change. Technology and Culture, 2(3):230–236, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895459/pdf.

Mumford:1964:ADT

[Mum64] Lewis Mumford. Authoritarian and democratic technics. Technology and Culture, 5(1):1–8, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894918/pdf.

Mumford:1965:MF

[Mum65] Lewis Mumford. Man the finder. Technology and Culture, 6(3):375– 381, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894672/pdf.

Mumford:1966:TNM

[Mum66] Lewis Mumford. Technics and the nature of man. Technology and Culture, 7(3):303–317, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894577/pdf.

Murphy:1969:DPC

[Mur69] Larry J. Murphy. The Development of Punch Card Tabulation in the Bureau of the Census, 1890–1940 by Leon D. Truesdell (review). Technology and Culture, 10(3):455–457, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901634/pdf.

Mayo-Wells:1963:OST

[MW63] Wilfrid J. Mayo-Wells. The origins of space telemetry. Technology and Culture, 4(4):499–514, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894967/pdf.

Neale:1964:GSI

[Nea64] Walter C. Neale. Gopalpur: a South Indian Village by Alan R. Beals (review). Technology and Culture, 5(1):131–133, Winter 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894952/pdf.

Needham:1964:CPC

[Nee64] Joseph Needham. Chinese priorities in cast iron metallurgy. Technology and Culture, 5(3):398–404, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894858/pdf.

Needham:1969:AMR

[Nee69] Joseph Needham. Alchemy, Medicine, and Religion in the China of A.D. 320; the Nei P’ien of Ko Hung ed. by J. R. Ware (review). Technology and Culture, 10(1):90–92, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/892322/pdf.

Nelson:1966:RSP

[Nel66] R. J. Nelson. Recognition: a Study in the Philosophy of Artificial Intelligence by Kenneth M. Sayre (review). Technology and Culture, 7(1):131–132, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893845/pdf.

Netschert:1965:DUS

[Net65] Bruce C. Netschert. Direct Use of the Sim’s Energy by Farrington Daniels (review). Technology and Culture, 6(3):501–503, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894709/pdf.

Newhall:1966:SBP

[New66a] Beaumont Newhall. A Source Book of Photographic History (“Quellendarstellungen zur Geschichte der Fotografie”) by Wolfgang Baier (review). Technology and Culture, 7(2):237–240, Spring 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895184/pdf.

Newhall:1966:SVH

[New66b] Beaumont Newhall. Stereo Views: a History of Stereographs in America and Their Collection by William Culp Darrah (review). Technology and Culture, 7(2):240–241, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895185/pdf.

Nieburg:1962:AFN

[Nie62] H. L. Nieburg. America Faces the Nuclear Age ed. by Johnson E. Fairchild and David Landman (review). Technology and Culture, 3 (3):358–359, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895154/pdf.

Nieburg:1964:PPR

[Nie64] H. L. Nieburg. Project Plowshare by Ralph Sanders (review). Technology and Culture, 5(1):100–101, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894934/pdf.

Nieburg:1965:PAF

[Nie65] Harold L. Nieburg. The Peaceful Atom in Foreign Policy by Arnold Kramish (review). Technology and Culture, 6(1):156–157, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894810/pdf.

Nieburg:1967:CAH

[Nie67a] H. L. Nieburg. Contracting for Atoms by Harold Orlans (review). Technology and Culture, 8(4):512–514, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894415/pdf.

Nieburg:1967:IIP

[Nie67b] H. L. Nieburg. The Intruders: The Invasion of Privacy by Government and Industry by Edward V. Long (review). Technology and Culture, 8(4):514–515, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894416/pdf.

Nieburg:1967:RDE

[Nie67c] H. L. Nieburg. The Research and Development Effort in Western Europe, North America and the Soviet Union by C. Freeman, A. Young (review). Technology and Culture, 8(2):251–252, April 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894479/pdf.

Nieburg:1969:SWR

[Nie69] H. L. Nieburg. The Silent Weapons: The Realities of Chemical and Biological Warfare by Robin Clarke (review). Technology and Culture, 10(2):341–342, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892387/pdf.

Niles:1961:TCS

[Nil61] Alfred S. Niles. The Tale of the Comet: The Story of the British Jet Airliner by Derek D. Dempster (review). Technology and Culture, 2(3):291–293, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895483/pdf.

Nixon:1964:TNP

[Nix64] Frank Nixon. Thomas Newcomen: The Prehistory of the Steam Engine by L. T. C. Rolt (review). Technology and Culture, 5(1):76–80, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894926/
pdf.

Nixon:1965:BIA

[Nix65] Frank Nixon. Britain’s Imperial Air Routes, 1918 to 1939: The Story of Britain’s Overseas Airlines by Robin Higham (review). Technology and Culture, 6(2):317–318, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894768/pdf.

Nunis:1963:OHH

[Nun63] Doyce B. Nunis, Jr. Oral history and the history of technology. Technology and Culture, 4(2):149–154, Spring 1963. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894998/pdf.

ODea:1969:HLL

[O’D69] W. T. O’Dea. A Heritage of Light: Lamp and Lighting in the Early Canadian Home by Loris S. Russell (review). Technology and Culture, 10(2):330–331, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892382/pdf.

Ogden:1966:KNO

[Ogd66] Warren G. Ogden, Jr. A. K. Nartov, an Outstanding Machine Builder of the 18th Century by A. S. Britkin and S. S. Vidonov (review). Technology and Culture, 7(3):426–428, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894597/pdf.

Oliver:1960:HWT

[Oli60] John W. Oliver. A History of Western Technology by Friedrich Klemm (review). Technology and Culture, 1(2):178–181, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895254/pdf.

Olson:1969:IIA

[Ols69] Frederick I. Olson. Industrialization, Immigrants, and Americanizers: The View from Milwaukee, 1866–1921 by Gerd Korman (review). Technology and Culture, 10(1):109–111, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892332/pdf.

Ong:1962:C

[Ong62] Walter J. Ong, S.J. Commentary. Technology and Culture, 3(4):459– 462, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895102/
pdf.

Outwater:1962:PET

[Out62] J. Ogden Outwater, Jr. Possible evidences of a textile manufacturing center at Palenque. Technology and Culture, 3(2):161–166, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895378/pdf.

Outwater:1965:AUS

[Out65] John O. Outwater. The Aztecs under Spanish Rule by Charles Gibson (review). Technology and Culture, 6(2):274–276, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894744/pdf.

Owen:1965:TPR

[Owe65] Wilfred Owen. Transportation and Politics by Roy I. Wolfe (review). Technology and Culture, 6(2):303–305, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894761/pdf.

Packer:1968:SDA

[Pac68] James E. Packer. Structure and design in Ancient Ostia: a contribution to the study of Roman Imperial architecture. Technology and Culture, 9(3):357–388, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894255/pdf.

Panseri:1960:M

[Pan60] Carlo Panseri. Metallurgy. Technology and Culture, 1(4):389–394, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895618/
pdf.

Parsons:1966:IFE

[Par66] Kenneth H. Parsons. Individual Freedom and the Economic Organization of Agriculture by Harold F. Breimeyer (review). Technology and Culture, 7(4):568–570, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894568/pdf.

Patterson:1962:VRC

[Pat62] William H. Patterson. Virginia Railroads and the Civil War by Angus James Johnston II (review). Technology and Culture, 3(3):337–339, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895145/
pdf.

Paul:1965:CWF

[Pau65] Rodman W. Paul. Conestoga Wagon 1750–1850: Freight Carrier for 100 Years of America’s Westward Expansion by George Shumway, et al. (review). Technology and Culture, 6(3):469–471, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894694/pdf.

Paul:1967:CWF

[Pau67a] Rodman W. Paul. Conestoga Wagon 1750–1850: Freight Carrier for 100 Years of America’s Westward Expansion by George Shumway, et al. (review). Technology and Culture, 8(3):405–406, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894440/pdf.

Paul:1967:CTA

[Pau67b] Rodman W. Paul. The Company Town in the American West by James B. Allen (review). Technology and Culture, 8(1):108–109, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894507/
pdf.

Payen:1963:SHT

[Pay63] Jacques Payen. A Short History of Technology by T. K. Derry and Trevor I. Williams (review). Technology and Culture, 4(1):53–54, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895034/
pdf.

Payen:1964:GHP

[Pay64] Jacques Payen. La Genèse de l’helice propulsive by Paul Augustin-Normand (review). Technology and Culture, 5(3):450–452, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894872/pdf.

Pearsall:1965:RPE

[Pea65] Marion Pearsall. Resources and People in East Kentucky: Problems and Potentials of a Lagging Economy by Mary Jean Bowman and W. Warren Haynes (review). Technology and Culture, 6(1):142–144, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894801/
pdf.

Pearce:1966:TSE

[Pea66a] Lorraine W. Pearce. Tin in Social and Economic History by E. S. Hedges (review). Technology and Culture, 7(1):76–77, Winter 1966.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893861/pdf.

Pearson:1966:PPS

[Pea66b] Lee M. Pearson. The “Princeton” and the “Peacemaker”: a study in Nineteenth-Century Naval research and development procedures. Technology and Culture, 7(2):163–183, Spring 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895171/pdf.

Pendray:1963:PRD

[Pen63] G. Edward Pendray. Pioneer rocket development in the United States. Technology and Culture, 4(4):384–392, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894959/pdf.

Penick:1966:RRM

[Pen66] James L. Penick, Jr. Readings in Resource Management and Conservation ed. by Ian Burton and Robert W. Kates (review). Technology and Culture, 7(2):244–245, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895188/pdf.

Perry:1963:ATT

[Per63] Robert L. Perry. The Atlas, Thor, and Titan. Technology and Culture, 4(4):466–477, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894964/pdf.

Perry:1968:AAC

[Per68] Robert L. Perry. Aeronautics and Astronautics, 1966: Chronology in Science, Technology, and Policy by the Science and Technology Division, Library of Congress (review). Technology and Culture, 9(3):510–512, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894284/pdf.

Perrot:1969:TSG

[Per69] Paul N. Perrot. Throwing stones at glass museums. Technology and Culture, 10(2):174–177, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892354/pdf.

Pettit:1965:LLE

[Pet65] Joseph M. Pettit. Listen to Leaders in Engineering ed. by Albert Love and James Saxon Childers (review). Technology and Culture, 6(4):679–681, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894664/pdf.

Phillips:1963:LSM

[Phi63] Albert J. Phillips. The Lead Smelting Mills of the Yorkshire Dales by Robert Clough (review). Technology and Culture, 4(2):219, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895009/pdf.

Piaskowski:1961:PRE

[Pia61] J. Piaskowski. Preliminary report on early iron technology in Poland. Technology and Culture, 2(3):245–248, Summer 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895462/pdf.

Piel:1962:C

[Pie62] Gerard Piel. Commentary. Technology and Culture, 3(4):463–465, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895103/
pdf.

Piel:1963:R

[Pie63] Gerard Piel. Response. Technology and Culture, 4(2):256, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895024/pdf.

Pierce:1964:HFT

[Pie64] Bernard F. Pierce. Human Factors In Technology ed. by Edward Bennett, James Degan, and Joseph Spiegel (review). Technology and Culture, 5(4):632–634, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894842/pdf.

Pierce:1966:HEG

[Pie66] Bernard F. Pierce. Human Engineering Guide for Equipment Designers by Wesley E. Woodson and Donald W. Conover (review). Technology and Culture, 7(1):124–126, Winter 1966. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893842/pdf.

Pilling:1966:WFA

[Pil66] Arnold R. Pilling. The World of the First Australians by Ronald M. Berndt and Catherine H. Berndt (review). Technology and Culture, 7 (4):576, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894572/pdf.

Pizor:1968:GSD

[Piz68] Faith K. Pizor. The great steam duck. Technology and Culture, 9(1):86–89, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894348/pdf.

Poole:1962:PLP

[PR62] J. B. Poole and R. Reed. The preparation of leather and parchment by the Dead Sea Scrolls community. Technology and Culture, 3(1):1–26, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895411/
pdf.

Preece:1962:RDa

[Pre62a] Warren E. Preece. A report of the discussion. Technology and Culture, 3(4):466–485, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895104/pdf.

Preece:1962:RDb

[Pre62b] Warren E. Preece. A report of the discussion. Technology and Culture, 3(4):524–534, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895109/pdf.

Preece:1962:RDc

[Pre62c] Warren E. Preece. A report of the discussion. Technology and Culture, 3(4):558–562, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895113/pdf.

Preece:1962:RDd

[Pre62d] Warren E. Preece. A report of the discussion. Technology and Culture, 3(4):624–635, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895119/pdf.

Preece:1962:RDe

[Pre62e] Warren E. Preece. A report of the discussion. Technology and Culture, 3(4):647–650, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895122/pdf.

Prufer:1962:RMT

[Pru62] Olaf H. Prufer. Races of Mankind: Their Origin and Mankind by Calvin Kephart (review). Technology and Culture, 3(1):105–108, Winter 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895426/
pdf.

Prufer:1965:BCS

[Pru65] Olaf H. Prufer. The Bark Canoes and Skin Boats of North America by Edwin Tappan Adney and Howard I. Chapelle (review). Technology and Culture, 6(3):460–462, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894689/pdf.

Pi-Sunyer:1964:PCG

[PS64] Oriol Pi-Sunyer. Penny Capitalism: a Guatemalan Indian Economy by Sol Tax (review). Technology and Culture, 5(2):274–276, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894912/pdf.

Pi-Sunyer:1965:NNL

[PS65] Oriol Pi-Sunyer. New Nations by Lucy Mair (review). Technology and Culture, 6(1):145–147, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894803/pdf.

Pi-Sunyer:1966:ICG

[PS66] Oriol Pi-Sunyer. Indian Crafts of Guatemala and El Salvador by Lilly de Jongh Osborne (review). Technology and Culture, 7(4):577–579, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894573/
pdf.

Pi-Sunyer:1967:MIB

[PS67] Oriol Pi-Sunyer. The Military Intellectuals in Britain by Robin Higham (review). Technology and Culture, 8(3):420–422, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894450/pdf.

Pi-Sunyer:1968:LAT

[PS68] Oriol Pi-Sunyer. The Latin American Tradition by Charles Wagley (review). Technology and Culture, 9(4):633–635, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894242/pdf.

Pi-Sunyer:1964:CRT

[PSD64] Oriol Pi-Sunyer and Thomas De Gregori. Cultural resistance to technological change. Technology and Culture, 5(2):247–253, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894897/pdf.

Pi-Sunyer:1967:LHM

[PSD67] Oriol Pi-Sunyer and Thomas R. De Gregori. The Liddell Hart Memoirs: The Later Years by B. H. Liddell Hart (review). Technology and Culture, 8(3):423–425, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894451/pdf.

Puhvel:1964:IEI

[Puh64] Jaan Puhvel. The Indo-European and Indo-Aryan plough: a linguistic study of technological diffusion. Technology and Culture, 5 (2):176–190, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894891/pdf.

Pursell:1962:TTF

[Pur62] Carroll W. Pursell, Jr. Tariff and technology: The foundation and development of the American tin-plate industry, 1872–1900. Technology and Culture, 3(3):267–284, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895128/pdf.

Pursell:1964:HSU

[Pur64] Carroll W. Pursell, Jr. A History of Steamboating on the Upper Missouri River by William E. Lass (review). Technology and Culture, 5(4):613–614, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894833/pdf.

Pursell:1965:CME

[Pur65a] Carroll W. Pursell, Jr. Cassier’s Magazine, Engineering Monthly, 1891–1913. Subject Index comp. by Eugene S. Ferguson (review). Technology and Culture, 6(3):538, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894728/pdf.

Pursell:1965:TUS

[Pur65b] Carroll W. Pursell, Jr. Tanning in the United States to 1850: a Brief History by Peter C. Welsh (review). Technology and Culture, 6(3):462–463, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894690/pdf.

Pursell:1968:SSF

[Pur68a] Carroll W. Pursell, Jr. Safer Skyways: Federal Control of Aviation, 1926–1966 by Donald R. Whitnah (review). Technology and Culture, 9(3):506–507, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894282/pdf.

Pursell:1968:MEA

[Pur68b] Carroll W. Pursell, Jr. The Mechanical Engineer in America, 1830– 1910: Professional Cultures in Conflict by Monte A. Calvert (review). Technology and Culture, 9(3):496–497, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894275/pdf.

Pursell:1968:PGS

[Pur68c] Carroll W. Pursell, Jr. A preface to government support of research and development: Research legislation and the National Bureau of Standards, 1935–41. Technology and Culture, 9(2):145–164, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894294/pdf.

Pursell:1968:SRG

[Pur68d] Carroll W. Pursell, Jr. Some recent government publications concerning technology. Technology and Culture, 9(1):90–93, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894349/pdf.

Pursell:1969:CCS

[Pur69a] Carroll W. Pursell, Jr. Christopher Colles’s steam engine for the New York Water Works, 1775. Technology and Culture, 10(4):567–569, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892403/
pdf.

Pursell:1969:ELP

[Pur69b] Carroll W. Pursell, Jr. Ebenezer A. Lester’s pendulum steam engine. Technology and Culture, 10(1):65–67, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892317/pdf.

Putnam:1967:NSH

[Put67] William D. Putnam. In the Name of Science by H. L. Nieburg (review). Technology and Culture, 8(2):249–251, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894478/pdf.

Quimby:1965:DBG

[Qui65] Ian M. G. Quimby. Drawing of bobbin girl, by Winslow Homer. Technology and Culture, 6(2):261–262, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894738/pdf.

Rae:1960:AAT

[Rae60a] John B. Rae. Automotive and aeronautical technology. Technology and Culture, 1(4):411–414, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895622/pdf.

Rae:1960:KHT

[Rae60b] John B. Rae. The “Know-How” tradition: Technology in American history. Technology and Culture, 1(2):139–150, Spring 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895249/pdf.

Rae:1961:EAF

[Rae61a] John B. Rae. From Engines to Autos: Five Pioneers in Engine Development and Their Contributions to the Automotive Industry by Eugen Diesel, et al. (review). Technology and Culture, 2(2):189–191, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895560/
pdf.

Rae:1961:MWH

[Rae61b] John B. Rae. Monopoly on Wheels: Henry Ford and the Selden Automobile Patent by William Greenleaf (review). Technology and Culture, 2(3):289–291, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895482/pdf.

Rae:1961:SEH

[Rae61c] John B. Rae. Science and engineering in the history of aviation. Technology and Culture, 2(4):391–399, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895437/pdf.

Rae:1962:HAP

[Rae62] John B. Rae. History of the Aircraft Piston Engine. A Brief Outline by S. D. Heron (review). Technology and Culture, 3(2):209–210, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895397/
pdf.

Rae:1963:C

[Rae63] John B. Rae. Commentary. Technology and Culture, 4(2):174–176, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895001/
pdf.

Rae:1965:HCH

[Rae65] John B. Rae. Herbert Clark Hoover (1874–1964). Technology and Culture, 6(1):108–109, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894783/pdf.

Rae:1966:DUS

[Rae66] John B. Rae. Dynamics of the United States Automobile Industry by Charles E. Edwards (review). Technology and Culture, 7(4):549–551, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894559/
pdf.

Rae:1968:GIT

[Rae68] John B. Rae. Government, industry, and technology: The SHOT-OAH program April 1968. Technology and Culture, 9(4):576–579, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894216/
pdf.

Rae:1969:ATB

[Rae69] John B. Rae. Air Transportation, 1915 and Beyond: a Systems Approach ed. by B. A. Shriever, William W. Seifert (review). Technology and Culture, 10(2):328–329, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892380/pdf.

Ramo:1969:CAC

[Ram69] Simon Ramo. Comment: The anticipation of change. Technology and Culture, 10(4):514–521, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892396/pdf.

Rand:1963:TSA

[Ran63] Walter Rand. Technology, Science and Art: Common Ground (review). Technology and Culture, 4(2):243–247, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895021/pdf.

Rasmussen:1960:FT

[Ras60] Wayne D. Rasmussen. Food technology. Technology and Culture, 1(4): 376–381, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895616/pdf.

Rasmussen:1961:FAA

[Ras61] Wayne D. Rasmussen. The Farmer’s Age: Agriculture, 1815–1860 by Paul W. Gates (review). Technology and Culture, 2(2):185–187, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895558/
pdf.

Rasmussen:1964:AIA

[Ras64a] Wayne D. Rasmussen. Agrindus: Integration of Agriculture and Industries by Haim Halperin (review). Technology and Culture, 5(4): 631–632, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894841/pdf.

Rasmussen:1964:FHH

[Ras64b] Wayne D. Rasmussen. The F. Hal Higgins Library of Agricultural Technology. Technology and Culture, 5(4):575–577, Fall 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894820/pdf.

Rasmussen:1966:LAD

[Ras66] Wayne D. Rasmussen. The Lifeline of America: Development of the Food Industry by Edward C. Hampe, Jr. and Merle Wittenberg (review). Technology and Culture, 7(1):80–81, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893863/pdf.

Rasmussen:1967:CAS

[Ras67] Wayne D. Rasmussen. The College of Agriculture: Science in the Public Service by Charles E. Kellogg, David C. Knapp (review). Technology and Culture, 8(1):121–122, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894514/pdf.

Rasmussen:1968:AAA

[Ras68a] Wayne D. Rasmussen. Advances in American agriculture: The mechanical tomato harvester as a case study. Technology and Culture, 9(4):531–543, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894212/pdf.

Rasmussen:1968:FCS

[Ras68b] Wayne D. Rasmussen. Food and Civilization: a Symposium ed. by Seymour M. Farber, et al. (review). Technology and Culture, 9 (4):627–629, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894239/pdf.

Ravetz:1965:MTA

[Rav65] Alison Ravetz. Modern technology and an ancient occupation: Housework in present-day society. Technology and Culture, 6(2):256–260, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894737/
pdf.

Ravetz:1966:IST

[Rav66] J. R. Ravetz. The Impact of Science on Technology ed. by Aaron W. Warner, et al. (review). Technology and Culture, 7(2):255–257, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895193/
pdf.

Razak:1964:SBL

[Raz64] Kenneth Razak. The Story of the Britinsh Light Aeroplane by Terence Boughton (review). Technology and Culture, 5(3):457–459, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894877/pdf.

Reingold:1960:UPO

[Rei60] Nathan Reingold. U.S. Patent Office records as sources for the history of invention and technological property. Technology and Culture, 1(2):156–167, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895251/pdf.

Reingold:1962:MFM

[Rei62a] Nathan Reingold. Matthew Fontaine Maury & Joseph Henry: Scientists of the Civil War by Patricia Johns (review). Technology and Culture, 3(2):189–190, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895388/pdf.

Reiss:1962:MMF

[Rei62b] Richard F. Reiss. Mathematical Machines by Francis J. Murray (review). Technology and Culture, 3(2):201–202, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895394/pdf.

Reiss:1963:AAG

[Rei63] Richard F. Reiss. Anatomy of Automation by George H. Amber and Paul S. Amber (review). Technology and Culture, 4(3):365–366, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894991/
pdf.

Reiss:1964:MRJ

[Rei64] Richard F. Reiss. The Minds of Robots by James T. Culbertson (review). Technology and Culture, 5(4):635–638, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894843/pdf.

Reinartz:1965:ISI

[Rei65a] Leo F. Reinartz. Iron and Steel in the Industrial Revolution by Thomas S. Ashton (review). Technology and Culture, 6(1):118–121, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894787/
pdf.

Reingold:1965:FRL

[Rei65b] Nathan Reingold. The Future of the Research Library by Verner W. Clapp (review). Technology and Culture, 6(1):161–162, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894813/pdf.

Reingold:1967:ECU

[Rei67] Nathan Reingold. Engineering at Cambridge University, 1783–1965 by T. J. N. Hilken (review). Technology and Culture, 8(4):499–500, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894406/
pdf.

Reti:1961:WTA

[Ret61] Ladislao Reti. Der Weg aus der Tretmuhle. Ein Abriss der Geschichte der Technik der Neueren Zeit by Franz Hendrichs (review). Technology and Culture, 2(1):41–42, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895578/pdf.

Reti:1963:FGM

[Ret63] Ladislao Reti. Francesco di Giorgio Martini’s treatise on engineering and its plagiarists. Technology and Culture, 4(3):287–298, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894973/pdf.

Reti:1965:PFD

[Ret65] Ladislao Reti. A postscript to the Filarete discussion: On horizontal waterwheels and smelter blowers in the writings of Leonardo da Vinci and Juanelo Turriano. Technology and Culture, 6(3):428–441, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894678/

pdf. See [Spe63, Smi64f, Eic64, Spe64].

Reti:1967:CJT

[Ret67a] Ladislao Reti. The Codex of Juanelo Turriano (1500–1585). Technology and Culture, 8(1):53–66, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894489/pdf.

Reti:1967:LVC

[Ret67b] Ladislao Reti. The Leonardo da Vinci Codices in the Biblioteca Nacional of Madrid. Technology and Culture, 8(4):437–445, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894396/pdf.

Reti:1967:EEH

[Ret67c] Ladislao Reti. On the efficiency of early horizontal waterwheels. Technology and Culture, 8(3):388–394, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894436/pdf.

Richman:1968:HVN

[Ric68] Irwin Richman. Hopewell Village National Historic Site by John C. W. Riddle (review). Technology and Culture, 9(2):213–217, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894301/pdf.

Riepe:1965:SAI

[Rie65] Dale Riepe. The Scientific Age: The Impact of Science on Society by L. V. Berkner (review). Technology and Culture, 6(4):681–683, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894665/
pdf.

Riepe:1967:MMM

[Rie67] Dale Riepe. Men, Machines, and Modern Times by Elting E. Morison, and: The Evolving Society ed. by Alice Mary Hilton, and: Science, Technology, and Human Values by Cornelius A. Benjamin (review). Technology and Culture, 8(4):524–532, October 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894422/pdf.

Riepe:1968:PCE

[Rie68a] Dale Riepe. Philosophy and Cybernetics ed. by Frederick J. Crosson, Kenneth M. Sayre (review). Technology and Culture, 9(4):625–627, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894238.

Riepe:1968:NUS

[Rie68b] Dale Riepe. The New Utopians: a Study of System Design and Social Change by Robert Boguslaw (review). Technology and Culture, 9(3):521–523, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894290/pdf.

Ringo:1964:BSJ

[Rin64] Miriam K. Ringo. Business and Society by Joseph W. McGuire (review). Technology and Culture, 5(3):478–480, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894887/pdf.

Ringo:1965:GMN

[Rin65] Miriam K. Ringo. Government and Manpower by National Manpower Council, and: Public Policies and Manpower Resources by National Manpower Council (review). Technology and Culture, 6(3):515–518, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894717/
pdf.

Ringo:1966:MPD

[Rin66] Miriam K. Ringo. Manpower Policies for a Democratic Society by National Manpower Council (review). Technology and Culture, 7 (2):261–262, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895195/pdf.

Rinker:1968:MTW

[Rin68] Harry L. Rinker. Moravians in Two Worlds by Gillian Lindt Gollin (review). Technology and Culture, 9(4):608–609, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894227/pdf.

Roach:1963:FRB

[Roa63] Robert D. Roach, Jr. The first rocket-belt. Technology and Culture, 4(4):490–498, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894966/pdf.

Robinson:1965:TCS

[Rob65] Ormsbee W. Robinson. The Two Cultures: And a Second Look by C. P. Snow (review). Technology and Culture, 6(1):162–164, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894814/pdf.

Robinson:1966:SMP

[Rob66] Douglas H. Robinson. Space Medicine in Project Mercury by Mae Mills Link (review). Technology and Culture, 7(3):446–447, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894609/pdf.

Roberts:1969:BPM

[Rob69] Kenneth D. Roberts. British Plane Makers from 1700 by William L. Goodman (review). Technology and Culture, 10(3):447–448, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/901629/pdf.

Rock:1966:SIM

[Roc66] Vincent P. Rock. Space: Its Impact on Man’s Society ed. by Lillian Levy (review). Technology and Culture, 7(1):118–120, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893885/pdf.

Rodnick:1961:MIA

[Rod61] David Rodnick. The Modernization of Iran, 1921–41 by Amin Banani (review). Technology and Culture, 2(4):420–422, Fall 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895449/pdf.

Rodnick:1962:IIK

[Rod62a] David Rodnick. The Industrialization of Iraq by Kathleen M. Langley (review). Technology and Culture, 3(3):369–370, Summer 1962.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895159/pdf.

Rodnick:1962:PPP

[Rod62b] David Rodnick. The Philippines: Public Policy and National Economic Development by Frank Golay (review). Technology and Culture, 3(2):223–226, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895404/pdf.

Rodnick:1965:IDG

[Rod65] David Rodnick. International Development: Growth and Change by H. W. Singer (review). Technology and Culture, 6(2):294–295, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894756/pdf.

Rodnick:1967:TFD

[Rod67] David Rodnick. Trade, Finance and Development in Pakistan by J. Russell Andrus, Azizali F. Mohammed (review). Technology and Culture, 8(1):124, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894516/pdf.

Rogers:1960:TTM

[Rog60] Grace L. Rogers. Textiles and textile machinery. Technology and Culture, 1(4):394–401, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895619/pdf.

Rogers:1969:IDS

[Rog69] Everett M. Rogers. Innovation Diffusion as a Spatial Process by Torsten Hägerstrand (review). Technology and Culture, 10(3):480– 482, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901644/pdf.

Roller:1964:ESJ

[Rol64] Duane Roller. The Excitement of Science by John Rader Platt (review). Technology and Culture, 5(1):117–118, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894944/pdf.

Rolt:1965:SHS

[Rol65] L. T. C. Rolt. A Short History of the Steam Engine by H. W. Dickinson (review). Technology and Culture, 6(1):115–118, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894786/pdf.

Rolt:1967:LES

[Rol67] L. T. C. Rolt. Lives of the Engineers: Selections from Samuel Smiles ed. by Thomas Parke Hughes (review). Technology and Culture, 8(1):97–99, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894500/pdf.

Rolt:1968:CBE

[Rol68] L. T. C. Rolt. The Cornish Beam Engine by D. B. Barton (review). Technology and Culture, 9(3):492–494, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894273/pdf.

Rossman:1961:SSS

[Ros61] Joseph Rossman. Salt and Salt Shakers by Arthur G. Peterson (review). Technology and Culture, 2(1):71, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895591/pdf.

Rosenbloom:1963:OAS

[Ros63] Richard S. Rosenbloom. Organization, Automation, and Society: The Scientific Revolution in Industry by Robert A. Brady (review). Technology and Culture, 4(1):108–110, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895056/pdf.

Rosenbloom:1964:PWI

[Ros64a] Richard S. Rosenbloom. Post-War Investment, Location and Size of Plant by P. Sargant Florence (review). Technology and Culture, 5(1):110–111, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894940/pdf.

Rosenbloom:1964:MMS

[Ros64b] Richard S. Rosenbloom. Men and machines: Some 19th-Century analyses of mechanization. Technology and Culture, 5(4):489–511, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894815/pdf.

Rosegger:1965:IPP

[Ros65a] Gerhard Rosegger. Industrial Progress in Poland, Czechoslovakia, and East Germany, 1937–1962 by Alfred Zauberman (review). Technology and Culture, 6(4):675–677, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894662/pdf.

Rosegger:1965:IBH

[Ros65b] Gerhard Rosegger. Industrialization of Bosnia-Hercegovina, 1878– 1918 by Peter F. Sugar (review). Technology and Culture, 6(2):289– 291, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894753/pdf.

Rosenbloom:1966:SIE

[Ros66] Richard S. Rosenbloom. The Sources of Increased Efficiency: a Study of DuPont Rayon Plants by Samuel Hollander (review). Technology and Culture, 7(1):121–123, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893886/pdf.

Rosegger:1967:IDI

[Ros67] Gerhard Rosegger. International Differences in Industrial Structure by Joe S. Bain (review). Technology and Culture, 8(3):416–418, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894448/pdf.

Rosegger:1968:VED

[Ros68] Gerhard Rosegger. Vorgeschichte, Entstehung und Demontage der Reichswerke im Salzgittergebiet by Matthias Riedel (review). Technology and Culture, 9(3):500–502, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894278/pdf.

Rosegger:1969:TGB

[Ros69] Gerhard Rosegger. Technik als geistige Bewegung in den Anfängen des deutschen Industriestaates by Gustav Goldbeck (review). Technology and Culture, 10(4):615–616, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892415/pdf.

Rotsch:1960:FTE

[Rot60] Melvin M. Rotsch. The Functional Tradition in Early Industrial Buildings by J. M. Richards (review). Technology and Culture, 1 (3):278–280, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895643/pdf.

Rotsch:1965:ESK

[Rot65] Melvin M. Rotsch. Die Entwicklung des Stahlhochbaus by Karl H. Wittek (review). Technology and Culture, 6(2):279, Spring 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894747/pdf.

Rotsch:1969:ABM

[Rot69] Melvin M. Rotsch. American Building. Materials and Techniques from the First Colonial Settlements to the Present by Carl W. Condit (review). Technology and Culture, 10(3):442–445, July 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901627/pdf.

Rubin:1967:RAE

[Rub67] Julius Rubin. Railroads and American Economic Growth: Essays in Econometric History by Robert William Fogel (review). Technology and Culture, 8(2):228–234, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894470/pdf.

Ryan:1967:TAR

[Rya67] James E. Ryan. Typefounding in America, 1787–1825 by Rollo G. Silver (review). Technology and Culture, 8(1):102–103, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894502/pdf.

Saloutos:1964:PRI

[Sal64] Theodore Saloutos. The Populist Response to Industrial America by Norman Pollack (review). Technology and Culture, 5(3):455–457, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894876/
pdf.

Saloutos:1965:AME

[Sal65] Theodore Saloutos. Aridity and Man ed. by Carle Hodge (review). Technology and Culture, 6(1):138–140, Winter 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894799/pdf.

Samolin:1965:TSC

[Sam65] William Samolin. Technical studies of Chinese and Eurasian archeological objects. Technology and Culture, 6(2):249–255, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894736/pdf.

Sanders:1960:STC

[San60] Ralph Sanders. Science and Technology in Contemporary War by Maj. Gen. G. I. Pokrovsky (review). Technology and Culture, 1 (3):286–288, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895646/pdf.

Sanders:1961:IAP

[San61a] R. Sanders. The Impact of Air Power: National Security and World Politics ed. by Eugene M. Emme (review). Technology and Culture, 2(1):75–77, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895594/pdf.

Sanders:1961:NNE

[San61b] Ralph Sanders. Nasser’s New Egypt by Keith Wheelock (review). Technology and Culture, 2(2):201–202, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895564/pdf.

Sanders:1963:DPP

[San63] Ralph Sanders. Defense of Project Plowshare. Technology and Culture, 4(2):252–255, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895023/pdf.

Sanders:1964:RWP

[San64] Ralph Sanders. The Revolution in World Politics ed. by Morton A. Kaplan (review). Technology and Culture, 5(1):122–123, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894947/pdf.

Sandler:1968:ANP

[San68] Stanley Sandler. Die Architektur der Nordatlantik-Passagierdampfer 1837–1914 by Peter Kleine (review). Technology and Culture, 9 (4):617–618, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894234/pdf.

Scamehorn:1961:BEB

[Sca61] Howard L. Scamehorn. Big Eight: a Biography of an Airplane by Richard G. Hubler (review). Technology and Culture, 2(3):293–295, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895484/
pdf.

Scamehorn:1965:RSF

[Sca65a] Lee Scamehorn. Revolution in the Sky: Those Fabulous Lockheeds, the Pilots Who Flew Them by Richard Sanders Allen (review). Technology and Culture, 6(1):148–150, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894805/pdf.

Scamehorn:1965:FNC

[Sca65b] Lee Scamehorn. The First Nonstop Coast-to-Coast Flight and the Historic T-2 Airplane by Louis S. Casey (review). Technology and Culture, 6(4):666–668, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894657/pdf.

Scamehorn:1966:SFC

[Sca66] Lee Scamehorn. Space Flight: Countdown for the Future by Eugen Sänger (review). Technology and Culture, 7(2):248–250, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895190/pdf.

Scamehorn:1967:HSF

[Sca67] Lee Scamehorn. A History of Space Flight by Eugene M. Emme (review). Technology and Culture, 8(2):248–249, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894477/pdf.

Scamehorn:1968:AHB

[Sca68a] Lee Scamehorn. The Aeronauts: a History of Ballooning, 1783–1903 by L. T. C. Rolt (review). Technology and Culture, 9(2):245–247, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894318/
pdf.

Scamehorn:1968:BGI

[Sca68b] Lee Scamehorn. With Brass and Gas: an Illustrated and Embellished Chronicle of Ballooning in Mid-nineteenth Century America by Munson Baldwin (review). Technology and Culture, 9(3):504–505, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894280/pdf.

Schmookler:1960:ETI

[Sch60] Jacob Schmookler. An economist takes issue. Technology and Culture, 1(3):214–220, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895629/pdf.

Schofield:1962:PBO

[Sch62a] Robert E. Schofield. Pierre Bouguer’s Optical Treatise on the Gradation of Light by W. E. Knowles Middleton (review). Technology and Culture, 3(2):183–184, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895384/pdf.

Schofield:1962:JWT

[Sch62b] Robert E. Schofield. Josiah wedgwood and the technology of glass manufacturing. Technology and Culture, 3(3):285–297, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895129/pdf.

Schofield:1963:GFA

[Sch63a] Robert E. Schofield. Guide to Federal Archives Relating to the Civil War by Kenneth W. Munden and Henry Putney Beers (review). Technology and Culture, 4(3):356–357, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894986/pdf.

Schofield:1963:LCY

[Sch63b] Robert E. Schofield. Lavoisier — The Crucial Year: The Background and Origin of His First Experiments on Combustion in 1772 by Henry Guerlac (review). Technology and Culture, 4(1):73–74, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895043/pdf.

Scherer:1964:FMF

[Sch64a] F. M. Scherer. Free Men and Free Markets by Robert Theobald (review). Technology and Culture, 5(3):481–483, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894888/pdf.

Schrier:1964:TTT

[Sch64b] Elliot Schrier. Toward technology transfer: The Engineering Foundation Research Conference on “Technology and the Civilian Economy”. Technology and Culture, 5(3):344–358, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894851/pdf.

Scherer:1965:IIW

[Sch65a] F. M. Scherer. Invention and innovation in the Watt–Boulton steamengine venture. Technology and Culture, 6(2):165–187, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894731/pdf.

Schofield:1965:CEH

[Sch65b] Robert E. Schofield. Comment: On the equilibrium of a heterogeneous social system. Technology and Culture, 6(4):591–595, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894641/pdf.

Schofield:1966:EER

[Sch66] Robert E. Schofield. Early Engineering Reminiscences (1815–40) of George Escol Sellers ed. by Eugene S. Ferguson (review). Technology and Culture, 7(2):233–234, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895182/pdf.

Scoville:1961:FED

[Sco61] Warren C. Scoville. France and the Economic Development of Europe 1800–1914 by Rondo E. Cameron (review). Technology and Culture, 2(4):403–405, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895440/pdf.

Scott:1966:BSR

[Sco66] Robert F. Scott. Brough Superior — the Rolls Royce of Motor Cycles by Ronald H. Clark (review). Technology and Culture, 7(1):94–95, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893871/
pdf.

Seligman:1969:LCE

[Sel69] Ben B. Seligman. Labor in a Changing Economy ed. by William Haber (review). Technology and Culture, 10(4):623–624, October 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892420/pdf.

Severinghaus:1965:PFM

[Sev65] Elmer L. Severinghaus. Population and Food by Michele Cepede, et al. (review). Technology and Culture, 6(3):492–494, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894705/pdf.

Sharlin:1963:HEE

[Sha63a] Harold I. Sharlin. A History of Electrical Engineering by Percy Dunsheath, and: The Development of Electrical Technology in the 19th Century by W. James King (review). Technology and Culture, 4 (3):340–343, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894979/pdf.

Sharlin:1963:MFM

[Sha63b] Harold I. Sharlin. My Father Marconi by Degna Marconi (review). Technology and Culture, 4(1):100–102, Winter 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895052/pdf.

Shapiro:1964:OSP

[Sha64] Sheldon Shapiro. The origin of the suction pump. Technology and Culture, 5(4):566–574, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894819/pdf.

Sharlin:1965:LHH

[Sha65] Harold I. Sharlin. Lightning in His Hand: The Life Story of Nikola Tesla by Inez Hunt, Wanetta W. Draper (review). Technology and Culture, 6(4):664–665, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894655/pdf.

Shankland:1967:CPE

[Sha67a] R. S. Shankland. The Collected Papers of Enrico Fermi, Vol. II: U.S. 1939–1954 (review). Technology and Culture, 8(1):119–120, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://www.jstor.org/stable/3101549;   https:
//muse.jhu.edu/pub/1/article/894513/pdf.

Shapiro:1967:OEP

[Sha67b] Theresa R. Shapiro. Output, Employment, and Productivity in the United States after 1800; Studies in Income and Wealth by the Conference on Research in Income and Wealth (review). Technology and Culture, 8(2):237–239, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894472/pdf.

Shelby:1961:CGC

[She61a] L. R. Shelby. The Construction of Gothic Cathedrals: a Study of Medieval Vault Erection by John Fitchen (review). Technology and Culture, 2(4):400–402, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/895438/pdf.

Shelby:1961:MMT

[She61b] Lonnie R. Shelby. Medieval masons’ tools: The level and the plumb rule. Technology and Culture, 2(2):127–130, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895547/pdf.

Shelby:1963:KTZ

[She63] L. R. Shelby. Kulturgeschichte des technischen Zeichnens by Alois Nedoluha (review). Technology and Culture, 4(2):217–218, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895008/pdf.

Shelby:1965:HKW

[She65a] L. R. Shelby. The History of the King’s Works ed. by H. M. Colvin (review). Technology and Culture, 6(2):269–271, Spring 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894742/pdf.

Shelby:1965:MMT

[She65b] L. R. Shelby. Medieval masons’ tools. II. Compass and square. Technology and Culture, 6(2):236–248, Spring 1965. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894735/pdf.

Shepard:1965:MBL

[She65c] Herbert A. Shepard. Managerial Behavior by Leonard R. Sayles (review). Technology and Culture, 6(2):309–310, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894764/pdf.

Sherwani:1965:LWU

[She65d] J. K. Sherwani. Land and Water Use ed. by Wynn Thorne (review). Technology and Culture, 6(3):496–499, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894707/pdf.

Shepherd:1968:TCP

[She68] Dorothy G. Shepherd. The Traditional Crafts of Persia; Their Development, Technology, and Influence on Eastern and Western Civilizations by Hans E. Wulff (review). Technology and Culture, 9(3):476– 480, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894266/pdf.

Shelby:1969:SKP

[She69] Lon R. Shelby. Setting out the keystones of pointed arches: a note on medieval “Baugeometrie”. Technology and Culture, 10(4):537–548, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892400/
pdf.

Shryock:1963:MMW

[Shr63] Richard H. Shryock. From Miasmas to Molecules by W. Barry Wood, Jr., M.D. (review). Technology and Culture, 4(3):368–369, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894993/pdf.

Shrader:1968:DHE

[Shr68] Erwin F. Shrader. The Development of High-Energy Accelerators ed. by M. Stanley Livingston (review). Technology and Culture, 9(2):259– 260, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894327/pdf.

Shukhardin:1967:CAH

[Shu67] S. Shukhardin. “Class approach” to the history of technology. Technology and Culture, 8(4):490, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894401/pdf.

Silliman:1969:WSF

[Sil69] Robert H. Silliman. William Shipley, Founder of the Royal Society of Arts by D. G. C. Allan (review). Technology and Culture, 10(2):309–310, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892367/pdf.

Singer:1960:HHT

[Sin60a] Charles Singer. How “A history of technology” came into being. Technology and Culture, 1(4):302–311, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895606/pdf.

Singer:1960:ASF

[Sin60b] Dorothea Waley Singer. Alchemy: the story of the fascination of gold, and the attempts of chemists, mystics, and charlatans to find the Philosophers’ Stone by E. J. Holmyard (review). Technology and Culture, 1(2):182–185, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895255/pdf.

Sinclair:1966:SSU

[Sin66] Bruce Sinclair. Scientific Societies in the United States by Ralph S. Bates (review). Technology and Culture, 7(4):544–546, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894556/pdf.

Sinclair:1969:TSW

[Sin69a] Bruce Sinclair. At the turn of a screw: William Sellers, the Franklin Institute, and a Standard American thread. Technology and Culture, 10(1):20–34, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892312/pdf.

Sinclair:1969:ASA

[Sin69b] Bruce Sinclair. American Science in the Age of Jackson by George H. Daniels (review). Technology and Culture, 10(1):101–102, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892328/pdf.

Skolnik:1961:UAW

[Sko61] Herman Skolnik. U. S. Army in World War II. The Technical Services. The Chemical Warfare Service: Organizing for War by Leo P. Brophy and George J. B. Fisher, and: From Laboratory to Field by Leo P. Brophy, et al. (review). Technology and Culture, 2(4):417–419, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895447/
pdf.

Skolimowski:1966:STT

[Sko66] Henryk Skolimowski. The structure of thinking in technology. Technology and Culture, 7(3):371–383, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894582/pdf.

Skolnikoff:1967:GAR

[Sko67] Eugene B. Skolnikoff. Government and Allocation of Resources to Science, Fundamental Research and the Policies of Government, Government and Technical Innovation, and The Social Sciences and the Policies of Government by the Organization for Economic Cooperation and Development (review). Technology and Culture, 8(1):126– 132, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894518/pdf.

Skramstad:1969:GCI

[Skr69] Harold Skramstad. The Georgetown Canal incline. Technology and Culture, 10(4):549–560, October 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892401/pdf.

Smith:1960:RMM

[Smi60] Cyril Stanley Smith. Reply to Mr. Martin Burgess on nail-making methods. Technology and Culture, 1(3):289–291, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895647/pdf.

Smith:1961:HML

[Smi61a] Cyril Stanley Smith. A History of Metals by Leslie Aitchison (review). Technology and Culture, 2(2):166–170, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895550/pdf.

Smith:1961:ISP

[Smi61b] Cyril Stanley Smith. The interaction of science and practice in the history of metallurgy. Technology and Culture, 2(4):357–367, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895434/pdf.

Smith:1961:HMI

[Smi61c] Thomas M. Smith. A History of Mechanical Inventions by Abbott Payson Usher (review). Technology and Culture, 2(1):34–37, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895576/pdf.

Smith:1962:HPD

[Smi62] Cyril Stanley Smith. A History of Platinum by Donald McDonald (review). Technology and Culture, 3(1):88–89, Winter 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895416/pdf.

Smith:1963:LEM

[Smi63] C. O. Smith, Jr. L’orange d’Espagne sur les marchés européens: Le problème oranger espagnol by Max Liniger-Goumaz (review). Technology and Culture, 4(2):237–238, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895018/pdf.

Smiley:1964:MFF

[Smi64a] Francis E. Smiley, Jr. Mining Frontiers of the Far West by Rodman W. Paul (review). Technology and Culture, 5(2):261–262, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894903/pdf.

Smith:1964:PEE

[Smi64b] Alfred G. Smith. Profiles in Ethnology by Elman R. Service (review). Technology and Culture, 5(1):127–131, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894951/pdf.

Smith:1964:SCH

[Smi64c] Cyril Stanley Smith. Scientific Change: Historical studies in the intellectual, social and technical conditions for scientific discovery and technical invention, from antiquity to the present ed. by Alistair Crombie (review). Technology and Culture, 5(4):596–598, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894823/pdf.

Smith:1964:CAS

[Smi64d] Cyril Stanley Smith. The Coming of the Age of Steel by Theodore A. Wertime (review). Technology and Culture, 5(1):87–88, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894929/pdf.

Smith:1964:DCS

[Smi64e] Cyril Stanley Smith. The discovery of carbon in steel. Technology and Culture, 5(2):149–175, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894890/pdf.

Smith:1964:GIF

[Smi64f] Cyril Stanley Smith. Granulating iron in Filarete’s smelter. Technology and Culture, 5(3):386–390, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894856/pdf. See [Spe63, Spe64].

Smith:1965:AAM

[Smi65a] Cyril Stanley Smith. Analysis of Ancient Metals by Earle R. Caley (review). Technology and Culture, 6(4):650–651, Fall 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894649/pdf.

Smith:1965:GJA

[Smi65b] Cyril Stanley Smith. “The Glazier” by Jost Amman. Technology and Culture, 6(1):110–111, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894784/pdf.

Smith:1965:APU

[Smi65c] Thomas M. Smith. Abbott Payson Usher (1883–1965). Technology and Culture, 6(4):630–632, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894645/pdf.

Smith:1966:GGJ

[Smi66a] Cyril Stanley Smith. Greek Gold: Jewelry from the Age of Alexander by Herbert Hoffmann and Patricia F. Davidson (review). Technology and Culture, 7(3):413–415, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894591/pdf.

Smith:1966:AAC

[Smi66b] Goldwin Smith. Astronautics and Aeronautics, 1964: Chronology on Science, Technology, and Policy by National Aeronautics and Space Administration Historical Staff (review). Technology and Culture, 7 (3):445–446, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894608/pdf.

Smith:1967:GEA

[Smi67a] Cyril Stanley Smith. Gladius. Études sur les Armes Anciennes, l’Armement, l’Art Militaire et la Vie Culturelle en Orient et Occident (review). Technology and Culture, 8(1):86–88, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894494/pdf.

Smith:1967:TTA

[Smi67b] Cyril Stanley Smith. Tatnall on Testing. An Autobiographical Account of Adventures under 13 Vice Presidents by F. G. Tatnall (review). Technology and Culture, 8(1):107, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894505/pdf.

Smith:1967:JS

[Smi67c] Cyril Stanley Smith. A jeweler’s shop (1533). Technology and Culture, 8(2):207–209, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894461/pdf.

Smith:1967:CCU

[Smi67d] Thomas M. Smith. The Challenge of the Computer Utility by Douglas F. Parkhill (review). Technology and Culture, 8(4):515–516, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894417/pdf.

Smith:1967:PCS

[Smi67e] Cecil O. Smith, Jr. Pioneer Class: Sketches of the Lives of All Members of the Class of 1871 of Worcester Polytechnic Institute by Donald E. Johnson, James E. Mooney (review). Technology and Culture, 8(2):239–240, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894473/pdf.

Smith:1968:HTM

[Smi68a] Cyril Stanley Smith. A History of Tin Mining and Smelting in Cornwall by D. B. Barton (review). Technology and Culture, 9(3):491–492, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894272/
pdf.

Smith:1968:GLI

[Smi68b] Goldwin Smith. Greater London: an Industrial Geography by J. E. Martin (review). Technology and Culture, 9(1):120–122, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894363/pdf.

Smith:1968:PMA

[Smi68c] Richard K. Smith. The Politics of Military Aviation Procurement, 1926–1934: a Study in the Political Assertation of Consensual Values by Edwin H. Rutkowski (review). Technology and Culture, 9(2):248– 249, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894320/pdf.

Smith:1969:TCS

[Smi69a] Cyril Stanley Smith. Traditional Crafts and Skills. Life and Work in Mediaeval and Renaissance Times by Václav Husa, et al. (review). Technology and Culture, 10(1):92–94, January 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892323/pdf.

Smith:1969:HTI

[Smi69b] N. A. F. Smith. History of technology at Imperial College, London. Technology and Culture, 10(2):172–173, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892353/pdf.

Smith:1969:TDM

[Smi69c] Richard K. Smith. The TFX Decision: McNamara and the Military by Robert J. Art, and: The SST: Here It Comes Ready or Not: The Story of the Controversial Supersonic Transport by Don Dwiggins (review). Technology and Culture, 10(1):120–124, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892339/pdf.

Smith:1969:TIS

[Smi69d] Cecil O. Smith, Jr. Technological Innovation and Society ed. by Dean Morse and Aaron W. Warner (review). Technology and Culture, 10(3):470–473, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901641/pdf.

Snyder:1962:AAA

[Sny62] Melvin H. Snyder, Jr. Aeronautics and Astronautics: an American Chronology of Science and Technology in the Exploration of Space, 1915–1960 by Eugene M. Emme (review). Technology and Culture, 3(2):210–212, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895398/pdf.

Snyder:1966:BDB

[Sny66] Melvin H. Snyder, Jr. Billion Dollar Battle by Harold Mansfield (review). Technology and Culture, 7(1):113–115, Winter 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893883/pdf.

Solo:1968:MME

[Sol68] Robert A. Solo. The meaning and measure of economic progress. Technology and Culture, 9(3):389–414, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894256/pdf.

Sonnenfeld:1963:RCG

[Son63] Joseph Sonnenfeld. Readings in Cultural Geography ed. by Phillip L. Wagner and Marvin W. Mikesell (review). Technology and Culture, 4(2):247–251, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895022/pdf.

Spears:1962:USA

[Spe62a] James R. H. Spears. United States Army in World War II. The Technical Services. The Ordnance Department: Procurement and Supply by Harry C. Thomson and Lido Mayo (review). Technology and Culture, 3(3):340–341, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895146/pdf.

Spence:1962:EUE

[Spe62b] Clark C. Spence. Early uses of electricity in American agriculture. Technology and Culture, 3(2):142–160, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895377/pdf.

Spencer:1963:FDF

[Spe63] John R. Spencer. Filarete’s description of a Fifteenth Century Italian iron smelter at Ferriere. Technology and Culture, 4(2):201–206, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895003/pdf.

See comments [Smi64f, Eic64, Spe64, Ret65].

Spencer:1964:RDF

[Spe64] John R. Spencer. Response: In defense of Filarete. Technology and Culture, 5(3):406–407, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894860/pdf. See [Spe63, Ret65].

Spencer:1969:MDG

[Spe69] Daniel L. Spencer. Modernization: The Dynamics of Growth ed. by Myron Weiner (review). Technology and Culture, 10(1):128–130, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892343/
pdf.

Spier:1960:AML

[Spi60] R. F. G. Spier. Adventures with the Missing Link by Raymond A. Dart and Dennis Craig (review). Technology and Culture, 1(3):254–256, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895633/
pdf.

Spier:1964:HNS

[Spi64a] Robert F. G. Spier. Human Nature and The Study of Society: The Papers of Robert Redfield ed. by Margaret Park Redfield (review). Technology and Culture, 5(1):126–127, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894950/pdf.

Spier:1964:PBC

[Spi64b] Robert F. G. Spier. Prehistory and the Beginnings of Civilization by Jacquetta Hawkes and Sir Leonard Woolley (review). Technology and Culture, 5(3):443–445, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894868/pdf.

Spier:1968:EER

[Spi68a] Robert F. G. Spier. Ethnologia Europaea: Revue Internationale d’Ethnologie Européenne (review). Technology and Culture, 9(4):635– 636, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894243/pdf.

Spier:1968:SAH

[Spi68b] Robert F. G. Spier. The Stone Age Hunters by Grahame Clark (review). Technology and Culture, 9(4):601–602, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894222/pdf.

Spier:1969:MCP

[Spi69] Robert F. G. Spier. The Material Culture of the Peoples of the Gwembe Valley by Barrie Reynolds (review). Technology and Culture, 10(2):343–344, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892389/pdf.

Sporn:1963:ERB

[Spo63] Philip Sporn. Economic Redevelopment in Bituminous Coal by C. L. Christenson (review). Technology and Culture, 4(2):224–227, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895013/pdf.

Smith:1961:ISF

[SS61] Cyril Stanley Smith and Anneliese Sisco. Iron and steel in 1621: The “Fidelle ouverture de l’art de serrurier” of Mathurin Jousse. Technology and Culture, 2(2):131–145, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895548/pdf.

Stern:1962:SME

[Ste62] N. S. Stern. Siegfried Marcus, Ein Erfinderleben by Gustav Goldbeck (review). Technology and Culture, 3(1):93–95, Winter 1962. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895420/pdf.

Steinitz:1963:WLV

[Ste63] Kate T. Steinitz. The World of Leonardo da Vinci, Man of Science, Engineer and Dreamer of Flight by Ivor B. Hart (review). Technology and Culture, 4(1):84–88, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895048/pdf.

Stephenson:1965:MGF

[Ste65] C. B. Stephenson. Of Men and Galaxies by Fred Hoyle (review). Technology and Culture, 6(3):503–504, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894710/pdf.

Stearns:1967:HTH

[Ste67] Peter N. Stearns. Hard Times: Human Documents of the Industrial Revolution by E. Royston Pike (review). Technology and Culture, 8(4):497–499, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894405.

Stearns:1968:ESB

[Ste68a] Peter N. Stearns. An Expanding Society: Britain, 1830–1900 by G. S. R. Kitson Clark (review). Technology and Culture, 9(4):618–620, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894235/
pdf.

Stearns:1968:GTH

[Ste68b] Peter N. Stearns. “Golden Times”: Human Documents of the Victorian Age by E. Royston Pike (review). Technology and Culture, 9(3):494–495, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894274/pdf.

Stearns:1968:ATB

[Ste68c] Peter N. Stearns. The Age of Transition: Britain in the Nineteenth and Twentieth Centuries by D. F. MacDonald (review). Technology and Culture, 9(1):123–125, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894365/pdf.

Stover:1962:C

[Sto62a] Carl F. Stover. Commentary. Technology and Culture, 3(4):617–619, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895117/
pdf.

Stover:1962:I

[Sto62b] Carl F. Stover. Introduction. Technology and Culture, 3(4):383–384, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895097/
pdf.

Stockwell:1963:WAP

[Sto63] Richard E. Stockwell. The Weapons Acquisition Process: an Economic Analysis by Merton J. Peck and Frederic M. Scherer (review). Technology and Culture, 4(2):233–234, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895016/pdf.

Stockwell:1964:AAE

[Sto64a] R. E. Stockwell. Astronautical and Aeronautical Events of 1962, and: Project Mercury: a Chronology by James M. Grimwood of the National Aeronautics and Space Administration, and: Mercury Project Summary, Including Results of the Fourth Manned Orbital Flight, May 15 and 16, 1963 by various administrators and scientists of the National Aeronautics and Space Administration (review). Technology and Culture, 5(3):459–461, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894878/pdf.

Stockwell:1964:MMV

[Sto64b] R. E. Stockwell. Mariner: Mission to Venus by the Staff, Jet Propulsion Laboratory, California Institute of Technology (review). Technology and Culture, 5(2):267–268, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894908/pdf.

Stockwell:1965:EYG

[Sto65a] Richard E. Stockwell. The Early Years, Goddard Space Flight Center: Historical Origins and Activities through December 1962 by the Staff, Goddard Space Flight Center. National Aeronautics and Space Administration (review). Technology and Culture, 6(1):150, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894806/pdf.

Stockwell:1965:HRT

[Sto65b] Richard E. Stockwell. The History of Rocket Technology ed. by Eugene M. Emme (review). Technology and Culture, 6(4):668–669, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894658/pdf.

Stockwell:1968:SFW

[Sto68] Richard E. Stockwell. Space Frontier by Wernher von Braun (review). Technology and Culture, 9(2):251–252, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894323/pdf.

Stockwell:1969:BSC

[Sto69a] Richard E. Stockwell. The Beginnings of Satellite Communications by J. R. Pierce (review). Technology and Culture, 10(2):329–330, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/892381/
pdf.

Stover:1969:ALE

[Sto69b] John F. Stover. American Locomotives: an Engineering History, 1830–1880 by John H. White, Jr. (review). Technology and Culture, 10(2):316–317, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892372/pdf.

Strayer:1963:MTS

[Str63] Joseph R. Strayer. Medieval Technology and Social Change by Lynn White, jr (review). Technology and Culture, 4(1):62–65, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895038/pdf.

Strassmann:1964:RIC

[Str64] W. Paul Strassmann. The risks of innovation in 20th century manufacturing methods. Technology and Culture, 5(2):215–223, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894893/pdf.

Strassmann:1967:BBI

[Str67] W. Paul Strassmann. The British Building Industry: Four Studies in Response and Resistance to Change by Marian Bowley (review). Technology and Culture, 8(1):113–115, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894510/pdf.

Susskind:1965:VPZ

[Süs65] Charles Süsskind. Vědecké Práce Zemědělského Muzea (“Scientific Studies of the Agricultural Museum”) (review). Technology and Culture, 6(1):136–137, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894797/pdf.

Susskind:1968:WBT

[Süs68a] Charles Süsskind. The Wind and Beyond: Theodore von Kármán, Pioneer in Aviation and Pathfinder in Space by Theodore von Kármán and Lee Edson (review). Technology and Culture, 9(3):507–509, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894283/pdf.

Susskind:1968:WWD

[Süs68b] Charles Süsskind. Was ist und wozu dient die Elektrodynamik? by Walter Gerlach (review). Technology and Culture, 9(2):236–237, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894312/pdf.

Sutton:1966:PFJ

[Sut66a] George P. Sutton. Problems of Flight by Jet Propulsion — Interplanetary Flights by F. A. Tsander (review). Technology and Culture, 7(1):115–118, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893884/pdf.

Sutton:1966:RFE

[Sut66b] George P. Sutton. Rocket Flight Engineering by Eugen Sänger (review). Technology and Culture, 7(3):444–445, Summer 1966. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894607/pdf.

Svoboda:1967:GAM

[Svo67] Joseph Svoboda. Guide to the Archives and Manuscript Collections of the American Philosophical Society comp. by Whitfield J. Bell, Jr., Murphy D. Smith (review). Technology and Culture, 8(4):540–542, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894428/
pdf.

Swain:1967:HDS

[Swa67a] Donald C. Swain. High Dams and Slack Waters: TVA Rebuilds a River by Wilmon Henry Droze (review). Technology and Culture, 8(1):109–111, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894508/pdf.

Swain:1967:SSU

[Swa67b] Donald C. Swain. Science and Society in the United States ed. by David D. Van Tassel, Michael G. Hall (review). Technology and Culture, 8(3):432–433, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894455/pdf.

Swenson:1965:PPR

[Swe65] Loyd S. Swenson, Jr. Pride and Power: The Rationale of the Space Program by Vernon Van Dyke (review). Technology and Culture, 6(2):320–322, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894770/pdf.

Swenson:1967:SPA

[Swe67] Loyd S. Swenson, Jr. The Structure and Performance of the Aerospace Industry by Herman O. Stekler (review). Technology and Culture, 8(1):117–118, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894512/pdf.

Swenson:1968:MDD

[Swe68] Loyd S. Swenson, Jr. The Moon-Doggle: Domestic and International Implications of the Space Race by Amitai Etzioni, and: The Images of Space by Harold Leland Goodwin (review). Technology and Culture, 9(2):252–254, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894324/pdf.

Tager:1969:PPC

[Tag69] Jack Tager. Progressive Politics and Conservation: The Ballinger-Pinchot Affair by James Penick, Jr. (review). Technology and Culture, 10(3):453–455, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901633/pdf.

Tangri:1965:NRI

[Tan65] Shanti S. Tangri. Natural Resources and International Development ed. by Marion Clawson (review). Technology and Culture, 6(4):672– 674, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894661/
pdf.

Tascher:1966:URS

[Tas66] John Tascher. U.S. Rocket Society Number Two: The story of the Cleveland Rocket Society. Technology and Culture, 7(1):48–63, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893852/
pdf.

Taviss:1969:CNT

[Tav69] Irene Taviss. The Conquest of Nature: Technology and Its Consequences by R. J. Forbes (review). Technology and Culture, 10(2):304– 307, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892365/pdf.

Taylor:1967:ACI

[Tay67] C. F. Taylor. Aviation, the Creative Ideas by Oliver Stewart (review). Technology and Culture, 8(3):414–415, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894446/pdf.

Teichroew:1967:COW

[Tei67] Daniel Teichroew. Changing Organizations by Warren G. Bennis (review). Technology and Culture, 8(1):143–146, January 1967. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894523/pdf.

Teichroew:1968:ICM

[Tei68] Daniel A. Teichroew. The Impact of Computers on Management ed. by Charles A. Myers (review). Technology and Culture, 9(3):512–513, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894285/
pdf.

Thai:1962:C

[Tha62] Vu Van Thai. Commentary. Technology and Culture, 3(4):620–623, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895118/
pdf.

Thackray:1969:LSB

[Tha69] Arnold Thackray. The Lunar Society of Birmingham. University of Birmingham Historical Journal (review). Technology and Culture, 10(2):310–312, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892368/pdf.

Theobald:1962:LTP

[The62] Robert Theobald. Long-term prospects and problems. Technology and Culture, 3(4):601–616, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895116/pdf.

CMSTICM:1963:MST

[The63] The Committee on Museums of Science and Technology (Chairman, Torsten Althin) of The International Council of Museums (UN-ESCO). Museums of science and technology. Technology and Culture, 4(1):130–147, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895067/pdf.

Thompson:1960:PRA

[Tho60] George Raynor Thompson. The Pulse of Radar. The Autobiography of Sir Robert Watson-Watt by Sir Robert Watson-Watt (review). Technology and Culture, 1(2):191–194, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895260/pdf.

Thompson:1961:EAE

[Tho61a] B. C. Thompson. Energy in the American Economy, 1850–1975: an Economic Study of its History and Prospects by Sam H. Schurr and Bruce C. Netschert (review). Technology and Culture, 2(3):285–287, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895480/
pdf.

Thompson:1961:AER

[Tho61b] B. C. Thompson. The Atom and the Energy Revolution by Norman Lansdell (review). Technology and Culture, 2(1):83–85, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895598/pdf.

Thomas:1965:GEF

[Tho65a] Wayne Thomas. Giant Enterprise: Ford, G.M., and the Automobile Industry by Alfred D. Chandler, Jr. (review). Technology and Culture, 6(2):284–285, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894750/pdf.

Thompson:1965:NRU

[Tho65b] Burlock C. Thompson. Natural Resources for U.S. Growth: a Look Ahead to the Year 2000 by Hans H. Landsberg (review). Technology and Culture, 6(3):499–501, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894708/pdf.

Thorndike:1965:WT

[Tho65c] Lynn Thorndike. Words in theophilus. Technology and Culture, 6 (3):442–443, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894679/pdf.

Thompson:1966:EPC

[Tho66] B. C. Thompson. Energy: Its Production, Conversion and Use in the Service of Man by Philip Sporn (review). Technology and Culture, 7(1):77–80, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893862/pdf.

Tibensky:1967:HST

[Tib67] Jan Tibensky. The history of science and technology in Slovakia. Technology and Culture, 8(4):480–484, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894399/pdf.

Toffler:1966:ACM

[Tof66] Alvin Toffler. The Accidental Century by Michael Harrington (review). Technology and Culture, 7(2):252–255, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895192/pdf.

Toffler:1967:FPE

[Tof67] Alvin Toffler. Future Perfect ed. by H. Bruce Franklin (review). Technology and Culture, 8(1):151–153, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894527/pdf.

Tokaty:1963:SRT

[Tok63] G. A. Tokaty. Soviet rocket technology. Technology and Culture, 4(4): 515–528, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894968/pdf.

Tokaty:1968:RSF

[Tok68] G. A. Tokaty. Russian Solid-Fuel Rockets by V. N. Sokol’skii (review). Technology and Culture, 9(4):621–625, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894237/pdf.

Topping:1965:BRA

[Top65a] A. D. Topping. The British Rigid Airship, 1908–1931: a Study in Weapons Policy by Robin Higham (review). Technology and Culture, 6(2):312–315, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894766/pdf.

Topping:1965:HAB

[Top65b] A. D. Topping. The History of Airships by Basil Clarke (review). Technology and Culture, 6(2):315–316, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894767/pdf.

Topping:1966:AAM

[Top66a] A. D. Topping. The Airships Akron and Macon: Flying Aircraft Carriers of the United States Navy by Richard K. Smith (review). Technology and Culture, 7(4):554–558, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894562/pdf.

Topping:1966:ISR

[Top66b] A. D. Topping. The inaccurate “Slide rule”. Technology and Culture, 7(3):408–409, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894588/pdf.

Towle:1963:GLC

[Tow63] Edward L. Towle. The Great Lakes Car Ferries by George W. Hilton (review). Technology and Culture, 4(3):357–359, Summer 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894987/pdf.

Towle:1964:GLB

[Tow64] Edward L. Towle. Great Lakes Basin ed. by Howard J. Pincus (review). Technology and Culture, 5(3):467–469, Summer 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894882/pdf.

Towle:1965:PHG

[Tow65] Edward L. Towle. A Pictorial History of the Great Lakes by Harlan Hatcher and Erich Walter (review). Technology and Culture, 6(1):134–136, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894796/pdf.

Trachtenberg:1964:THV

[Tra64] Alan Trachtenberg. Technology and human values: The penn state CCLE conference. Technology and Culture, 5(3):359–376, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894853/pdf.

Trachtenberg:1965:MTF

[Tra65] Alan Trachtenberg. Mark Twain’s Fable of Progress: Political and Economic Ideas in “A Connecticut Yankee” by Henry Nash Smith (review). Technology and Culture, 6(3):527–530, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894723/pdf.

Tyler:1962:WWC

[Tyl62] Ralph W. Tyler. Where we came out. Technology and Culture, 3(4): 651–658, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895123/pdf.

Sun:1964:SCC

[tZS64] E tu Zen Sun. Science and Civilization in China. Vol. IV, Physics and Physical Technology; Part I: Physics by Joseph Needham (review). Technology and Culture, 5(1):89–92, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894930/pdf.

Sun:1965:WCC

[tZS65] E tu Zen Sun. Wu Ch’i-chün: Profile of a Chinese scholartechnologist. Technology and Culture, 6(3):394–406, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894674/pdf.

Usher:1960:IMB

[Ush60] Abbott Payson Usher. The industrialization of Modern Britain. Technology and Culture, 1(2):109–127, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895247/pdf.

Usher:1961:HMM

[Ush61] Abbott Payson Usher. History of the Milling Machine: a Study in Technical Development by Robert S. Woodbury (review). Technology and Culture, 2(1):43–44, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895579/pdf.

Usher:1964:WWM

[Ush64] Abbott Payson Usher. The Waldstein Woolen Mill. Noble Entrepreneurship in Eighteenth Century Bohemia by Herman Freudenberger (review). Technology and Culture, 5(3):449–450, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894871/pdf.

Uyeki:1964:CSS

[Uye64] Eugene S. Uyeki. Conference on Space, Science, and Urban Life (review). Technology and Culture, 5(4):626–628, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894839/pdf.

Uyeki:1968:MTP

[Uye68a] Eugene S. Uyeki. Metropolitan Transportation Politics and the New York Region by Jameson W. Doig (review). Technology and Culture, 9(3):514–516, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894286/pdf.

Uyeki:1968:UTP

[Uye68b] Eugene S. Uyeki. The Urban Transportation Problem by J. R. Meyer, et al. (review). Technology and Culture, 9(2):240–242, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894315/pdf.

Vatter:1968:DGN

[Vat68] Harold G. Vatter. The Dynamics of Growth in New England’s Economy, 1870–1964 by Robert W. Eisenmenger (review). Technology and Culture, 9(4):606–608, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894226/pdf.

vonBraun:1963:RJJ

[vB63] Wernher von Braun. The Redstone, Jupiter, and Juno. Technology and Culture, 4(4):452–465, Fall 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894963/pdf.

Vernon:1964:IGW

[Ver64] Raymond Vernon. Industrial Growth and World Trade by A. Maizels (review). Technology and Culture, 5(2):268–270, Spring 1964. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894909/pdf.

Virginsky:1968:BSN

[Vir68] V. S. Virginsky. The birth of steam navigation in Russia and Robert Fulton. Technology and Culture, 9(4):562–569, October 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894214/pdf.

Vivian:1968:ATK

[Viv68] C. H. Vivian. The Art of Tunnelling by Karoly Szechy (review). Technology and Culture, 9(1):130–132, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894369/pdf.

Vogel:1965:ANH

[Vog65] Robert M. Vogel. Assembling a new Hall of Civil Engineering. Technology and Culture, 6(1):59–73, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894779/pdf.

Vogel:1967:IAS

[Vog67] Robert M. Vogel. Industrial archeology at the Smithsonian institution: an interim report. Technology and Culture, 8(3):379–387, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894435/pdf.

Vogel:1969:SSE

[Vog69] Robert M. Vogel. The Stationary Steam Engine by George Watkins (review). Technology and Culture, 10(4):610–613, October 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892413/pdf.

vonWeiher:1967:VTT

[vW67] Sigfrid von Weiher. Versuch über Telegraphie und Telegraphen by J. L. Boeckmann (review). Technology and Culture, 8(2):222–223, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894467/
pdf.

Wailes:1967:IMS

[Wai67] Rex Wailes. The industrial monuments survey. Technology and Culture, 8(2):199–203, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894459/pdf.

Wailes:1968:WSE

[Wai68] Rex Wailes. Windmills in Southern Europe. Technology and Culture, 9(3):456–458, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894262/pdf.

Wallace:1969:CEU

[Wal69] James C. Wallace. Comment: The engineering use of human beings. Technology and Culture, 10(1):17–19, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892311/pdf.

Warner:1968:GSC

[War68] Sam Bass Warner, Jr. The Growth of Seaport Cities 1790–1825 ed. by D. T. Gilchrist (review). Technology and Culture, 9(2):229–230, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894308/
pdf.

Watson:1961:ADB

[Wat61a] Sara Ruth Watson. An appreciation of David Bernard Steinman (1886–1960). Technology and Culture, 2(1):23–27, Winter 1961. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895573/pdf.

Watson:1961:ACE

[Wat61b] Sara Ruth Watson. The American Civil Engineer by Daniel H. Calhoun (review). Technology and Culture, 2(1):61–63, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895587/pdf.

Watson:1963:EET

[Wat63] W. B. Watson. Europe Emerges: Transition Toward an Industrial World-Wide Society, 600–1750 by Robert L. Reynolds (review). Technology and Culture, 4(1):66–67, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895039/pdf.

Watson:1965:JRL

[Wat65] S. R. Watson. John Rennie, 1761–1821: The Life and Work of a Great Engineer by Cyril T. G. Boucher (review). Technology and Culture, 6(1):121–123, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894788/pdf.

Watt:1968:GNE

[Wat68] Kenneth E. F. Watt. Governing Nature by Earl Finbar Murphy (review). Technology and Culture, 9(3):520–521, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894289/pdf.

Weinberg:1962:DAE

[Wei62] Meyer Weinberg. The Development of the American Economy by August C. Bolino (review). Technology and Culture, 3(2):185–188, Spring 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895386/
pdf.

Weinberg:1963:TAI

[Wei63] Meyer Weinberg. Toys in America by Inez McClintock and Marshall McClintock (review). Technology and Culture, 4(2):222–223, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895011/pdf.

Weinberg:1964:IFM

[Wei64] Meyer Weinberg. Insull by Forrest McDonald (review). Technology and Culture, 5(1):108–110, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894939/pdf.

Weinberg:1966:ECC

[Wei66a] Meyer Weinberg. Economic Change in the Civil War Era ed. by David T. Gilchrist and W. David Lewis (review). Technology and Culture, 7(1):106–108, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893878/pdf.

Weinberg:1966:TSI

[Wei66b] Meyer Weinberg. Television and Society: an Inquest and Agenda for Improvement by Henry J. Skornia, and: Television: a World View by Wilson P. Dizard (review). Technology and Culture, 7(4):574–575, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894571/
pdf.

Welsh:1963:CRC

[Wel63] Peter C. Welsh. A craft that resisted change: American tanning practices to 1850. Technology and Culture, 4(3):299–317, Summer 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894974/pdf.

Welsh:1965:MEA

[Wel65] Peter C. Welsh. A Museum of Early American Tools by Eric Sloane (review). Technology and Culture, 6(1):125–126, Winter 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894790/pdf.

Welsh:1966:MWP

[Wel66] Peter C. Welsh. The metallic woodworking plane: an American contribution to hand-tool design. Technology and Culture, 7(1):38–47, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/893851/
pdf.

Welsh:1967:TCC

[Wel67] Peter C. Welsh. Traditional Country Craftsmen by J. Geraint Jenkins (review). Technology and Culture, 8(1):104–105, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894503/pdf.

Wertime:1964:AIE

[Wer64a] Theodore A. Wertime. Asian influences on European metallurgy. Technology and Culture, 5(3):391–397, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894857/pdf.

Wertime:1964:MIC

[Wer64b] Theodore A. Wertime. Men of Iron: The Crowleys in the Early Iron Industry by M. W. Flinn (review). Technology and Culture, 5(2):257–258, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894900/pdf.

Wertime:1965:CKJ

[Wer65] Theodore A. Wertime. The Copper King by J. R. Harris (review). Technology and Culture, 6(2):278–279, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894746/pdf.

Wertime:1966:HRM

[Wer66a] Theodore A. Wertime. Human resources and manpower needs. Technology and Culture, 7(1):67–68, Winter 1966. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893854/pdf.

Wertime:1966:MME

[Wer66b] Theodore A. Wertime. Marshall Mcluhan and the extensions of Man. Technology and Culture, 7(4):504–510, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894534/pdf.

Wertime:1967:NLA

[Wer67] Theodore A. Wertime. A new look at ancient technology: The Smithsonian program. Technology and Culture, 8(4):485–489, October 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894400/pdf.

Wertime:1968:CCC

[Wer68a] Theodore A. Wertime. Culture and continuity: a commentary on Mazlish and Mumford. Technology and Culture, 9(2):203–212, April 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/894300/pdf.

Wertime:1968:HEW

[Wer68b] Theodore A. Wertime. Hans Eberhard Wulff (1907–67). Technology and Culture, 9(3):459–461, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894263/pdf.

White:1960:TMA

[Whi60] Lynn White, Jr. Technology in the Middle Ages. Technology and Culture, 1(4):339–344, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895610/pdf.

White:1961:EME

[Whi61] Lynn White, Jr. Eilmer of malmesbury, an eleventh century aviator: a case study of technological innovation, its context and tradition. Technology and Culture, 2(2):97–111, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895545/pdf.

White:1962:AIC

[Whi62] Lynn White, Jr. The act of invention: Causes, contexts, continuities and consequences. Technology and Culture, 3(4):486–500, Fall 1962.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895105/pdf.

White:1963:DPL

[Whi63] Lynn White, Jr. Documents pour l’Histoire des Techniques, Cahier No. 1 (review). Technology and Culture, 4(1):51, Winter 1963. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895032/pdf.

White:1964:TR

[Whi64] Lynn White, Jr. Theophilus redivivus. Technology and Culture, 5(2):224–233, Spring 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894894/pdf.

White:1967:FSD

[Whi67] John H. White, Jr. Farewell to Steam by David Plowden (review). Technology and Culture, 8(3):406, July 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894441/pdf.

White:1968:DSG

[Whi68a] John H. White, Jr. Diesel Spotter’s Guide by Jerry A. Pinkepank, and: Loco 1 — the Diesel ed. by Harold H. Carstens (review). Technology and Culture, 9(4):616–617, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894233/pdf.

White:1968:NGL

[Whi68b] John H. White, Jr. The Narrow-Gauge Locomotives: The Baldwin Catalog of 1877 (review). Technology and Culture, 9(4):614–615, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894232/
pdf.

White:1968:RCY

[Whi68c] John H. White, Jr. The Railroad Caboose: Its 100 Year History, Legend and Lore by William F. Knapke (review). Technology and Culture, 9(4):614, October 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894231/pdf.

White:1968:SNO

[Whi68d] John H. White, Jr. Septimus Norris and the origins of the ten-wheel locomotive. Technology and Culture, 9(1):55–62, January 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894344/pdf.

White:1968:HMZ

[Whi68e] Lynn White, Jr. Das Hausbuch der Mendelschen Zwölfbrüderstiftung zu Nürnberg: Deutsche Handwerkerbilder des 15. und 16. Jahrhunderts ed. by Wilhelm Treue, et al. (review). Technology and Culture, 9(3):480–481, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894267/pdf.

White:1968:IP

[Whi68f] Lynn White, Jr. The invention of the parachute. Technology and Culture, 9(3):462–467, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894264/pdf.

White:1969:KBF

[Whi69] Lynn White, Jr. Kyeser’s “Bellifortis”: The first technological treatise of the Fifteenth Century. Technology and Culture, 10(3):436–441, July 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/901626/
pdf.

Wiebe:1964:OTD

[Wie64] Robert Wiebe. The Origins of Teapot Dome: Progressives, Parties, and Petroleum, 1909–1921 by J. Leonard Bates (review). Technology and Culture, 5(4):614–616, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894834/pdf.

Wik:1960:AM

[Wik60] Reynold Millard Wik. Agricultural machinery. Technology and Culture, 1(4):381–388, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895617/pdf.

Wik:1961:GSP

[Wik61] Reynold M. Wik. God Speed the Plow: The Coming of Steam Cultivation to England by Clark C. Spence (review). Technology and Culture, 2(2):187–189, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895559/pdf.

Wik:1962:HFS

[Wik62] Reynold Millard Wik. Henry Ford’s science and technology for rural America. Technology and Culture, 3(3):247–258, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895126/pdf.

Wik:1964:FDR

[Wik64] Reynold M. Wik. Ford: Decline and Rebirth, 1933–1962 by Frank E. Hill and Allan Nevins (review). Technology and Culture, 5(4):616– 618, Fall 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894835/
pdf.

Williams:1960:BIE

[Wil60a] Trevor I. Williams. The Brewing Industry in England, 1700–1830 by Peter Mathias (review). Technology and Culture, 1(2):188–189, Spring 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895258/
pdf.

Williams:1960:EAC

[Wil60b] Hermann Warner Williams, Jr. European Armour, circa 1066 to circa 1700 by Claude Blair (review). Technology and Culture, 1(3):262–263, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895636/
pdf.

Williams:1961:GBI

[Wil61a] Trevor I. Williams. Guinness’s Brewery in the Irish Economy, 1759– 1876 by Patrick Lynch and John Vaizey (review). Technology and Culture, 2(3):273–275, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895473/pdf.

Williams:1961:WRR

[Wil61b] Trevor I. Williams. Williams reviews the review issue. Technology and Culture, 2(3):254–260, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895465/pdf.

Williamson:1961:EIR

[Wil61c] Harold F. Williamson. The Electric Interurban Railways in America by George W. Hilton and John F. Due (review). Technology and Culture, 2(3):284–285, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895479/pdf.

Wilkinson:1963:BBE

[Wil63] Norman B. Wilkinson. Brandywine borrowings from European technology. Technology and Culture, 4(1):1–13, Winter 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895028/pdf.

Williams:1964:RBH

[Wil64a] Trevor I. Williams. The Red Barrel: a History of Watney Mann by Hurford Janes (review). Technology and Culture, 5(3):452–453, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894873/
pdf.

Williamson:1964:SSC

[Wil64b] Harold F. Williamson. Strategy and Structure: Chapters in the History of the Industrial Enterprise by Alfred D. Chandler, Jr. (review). Technology and Culture, 5(1):107, Winter 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894938/pdf.

Wilkinson:1965:DRB

[Wil65a] Norman B. Wilkinson. David Rittenhouse by Brooke Hindle (review). Technology and Culture, 6(3):465–468, Summer 1965. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894692/pdf.

Wilkinson:1965:MPJ

[Wil65b] Norman B. Wilkinson. Making Powder, by Jean Appier Hanzelet. Technology and Culture, 6(4):633–635, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894646/pdf.

Wilkinson:1965:AFW

[Wil65c] Norman B. Wilkinson. In anticipation of Frederick W. Taylor: a study of work by Lammot du Pont, 1872. Technology and Culture, 6(2):208–221, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894733/pdf.

Williamson:1967:WFW

[Wil67] Harold F. Williamson. The Wire That Fenced the West by Henry D. McCallum, Frances T. McCallum (review). Technology and Culture, 8(1):107–108, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894506/pdf.

Wilson:1968:CAA

[Wil68] Eugene E. Wilson. Carrier Admiral by Admiral J. J. “Jocko” Clark, U.S.N. (Ret.), and Clark G. Reynolds (review). Technology and Culture, 9(3):505–506, July 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894281/pdf.

Wiles:1969:SMF

[Wil69a] Peter Wiles. Soviet Metal-fabricating and Economic Development: Practice versus Policy by David Granick (review). Technology and Culture, 10(1):117–118, January 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892337/pdf.

Wilson:1969:LEP

[Wil69b] Eugene E. Wilson. The Liberty Engine, 1918–1942 by Philip S. Dickey III (review). Technology and Culture, 10(3):464–468, July 1969. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901639/pdf.

Wisnewsky:1961:NIP

[Wis61] Edward Wisnewsky. Nationalized Industry and Public Ownership by William A. Robson (review). Technology and Culture, 2(4):422–426, Fall 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895450/
pdf.

Wish:1963:IWH

[Wis63] Harvey Wish. The Image, or What Happened to The American Dream by Daniel Boorstin (review). Technology and Culture, 4(2):215–216, Spring 1963. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895007/
pdf.

Wisdom:1966:NCC

[Wis66] J. O. Wisdom. The need for corroboration: Comments on Agassi’s paper. Technology and Culture, 7(3):367–370, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894581/pdf. See [Aga66,

Aga67].

Wisdom:1967:RMD

[Wis67] J. O. Wisdom. Rules for making discoveries: Reply to Dr. Agassi. Technology and Culture, 8(2):210–212, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894462/pdf. See [Aga67].

Wolfe:1965:SMW

[Wol65] Roy I. Wolfe. Strategy for Mobility by Wilfred Owen (review). Technology and Culture, 6(2):305–307, Spring 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/894762/pdf.

Wolff:1966:DST

[Wol66] Klaus H. Wolff. The Development of the Spanish Textile Industry, 1750–1800 by James Clayburn LaForce, Jr. (review). Technology and Culture, 7(4):528–529, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894547/pdf.

Wolfe:1967:IHI

[Wol67] Roy I. Wolfe. The Impact of Highway Investment on Development by George W. Wilson, et al. (review). Technology and Culture, 8(2):253– 255, April 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894480/pdf.

Woodbury:1960:CMH

[Woo60a] Robert S. Woodbury. Contributions from the Museum of History and Technology, Papers 1–11 (review). Technology and Culture, 1 (3):269–271, Summer 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895639/pdf.

Woodbury:1960:LEW

[Woo60b] Robert S. Woodbury. The legend of Eli Whitney and interchangeable parts. Technology and Culture, 1(3):235–253, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895632/pdf.

Woodbury:1960:SFH

[Woo60c] Robert S. Woodbury. The scholarly future of the history of technology. Technology and Culture, 1(4):345–348, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895611/pdf.

Woodbury:1961:EDB

[Woo61a] Robert S. Woodbury. Die Entwicklung der Drehbank bis zum Jahre 1939 by Karl Wittmann (review). Technology and Culture, 2(2):176– 178, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895554/pdf.

Woodbury:1961:PVM

[Woo61b] Robert S. Woodbury. Precision Valley: The Machine Tool Companies of Springfield, Vermont by Wayne G. Broehl, Jr. (review). Technology and Culture, 2(2):180–182, Spring 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895556/pdf.

Woodman:1965:MGT

[Woo65] Harold D. Woodman. The Machine in the Garden: Technology and the Pastoral Ideal in America by Leo Marx (review). Technology and Culture, 6(4):661–664, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894654/pdf.

Woodbury:1966:SHM

[Woo66a] Robert S. Woodbury. A Short History of Machine Tools by L. T. C. Rolt (review). Technology and Culture, 7(2):231–233, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/895181/pdf.

Woodbury:1966:PBM

[Woo66b] Robert S. Woodbury. The Portsmouth Blockmaking Machinery by K. R. Gilbert (review). Technology and Culture, 7(2):230–231, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895180/pdf.

Woodruff:1968:IST

[Woo68] William Woodruff. The Industrial Society: Three Essays on Ideology and Development by Raymond Aron (review). Technology and Culture, 9(1):101–102, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894352/pdf.

Woodbury:1969:CMH

[Woo69a] Robert S. Woodbury. Contributions from the Museum of History and Technology by Smithsonian Institution (review). Technology and Culture, 10(2):350–352, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892393/pdf.

Woodbury:1969:MTC

[Woo69b] Robert S. Woodbury. Machine Tools (Catalogue of Exhibits [Science Museum, London] with Historical Introduction) by K. R. Gilbert (review). Technology and Culture, 10(2):352, April 1969. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892394/pdf.

Wright:1968:LIE

[Wri68] J. E. Wright. The Library of the Institution of Electrical Engineers in London. Technology and Culture, 9(2):191–196, April 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894298/pdf.

Wulff:1965:NDS

[Wul65] Hans E. Wulff. Notes on Damascene steel and Pamor. Technology and Culture, 6(4):627–629, Fall 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/

article/894644/pdf.

Wulff:1966:DLP

[Wul66a] Hans E. Wulff. Door locks in Persia. Technology and Culture, 7(4): 497–503, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894533/pdf.

Wulff:1966:PRN

[Wul66b] Hans E. Wulff. A postscript to Reti’s notes on Juanelo Turriano’s water mills. Technology and Culture, 7(3):398–401, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894586/pdf.

Watson-Watt:1962:TMW

[WW62] Sir Robert Watson-Watt. Technology in the modern world. Technology and Culture, 3(4):385–393, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895098/pdf.

Woodruff:1966:EGM

[WW66] William Woodruff and Helga Woodruff. Economic growth: Myth or reality: The interrelatedness of continents and the diffusion of technology, 1860–1960. Technology and Culture, 7(4):453–474, Fall 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894530/pdf.

Wynar:1960:PPS

[Wyn60] Lubomyr Wynar. Pioneers in Printing by Sean Jennett, and: U Istokov Russkogo Knigopechatanija (Sources of Russian Printing) (review). Technology and Culture, 1(3):264–266, Summer 1960. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895637/pdf.

Yabuuti:1967:TKK

[Yab67] K. Yabuuti. T’ien-kung K’ai-wu: Chinese Technology in the Seventeenth Century by Sung Ying-hsing (review). Technology and Culture, 8(1):92–94, January 1967. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894497/pdf.

Yamamura:1968:GGR

[Yam68] Kozo Yamamura. Gendai Gizitsu-ron by Keiichi Yamada (review). Technology and Culture, 9(1):105–106, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894355/pdf.

Zapffe:1962:HIS

[Zap62] Carl A. Zapffe. A History of Iron and Steelmaking in the United States by the Metallurgical Society (review). Technology and Culture, 3(3):320–321, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895137/pdf.

Zimmerman:1964:FCP

[Zim64] Joseph F. Zimmerman. Federal Conservation Policy, 1921–1933 by Donald C. Swain (review). Technology and Culture, 5(3):465–467, Summer 1964. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/894881/
pdf.

Zimmerman:1965:LUP

[Zim65] Joseph F. Zimmerman. Land Use Policy and Problems in the United States ed. by Howard Ottoson (review). Technology and Culture, 6(1):140–142, Winter 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894800/pdf.

Zuckerman:1968:IPS

[Zuc68] Harriet Zuckerman. Invention and the Patent System by S. C. Gilfillan (review). Technology and Culture, 9(1):107–109, January 1968. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894356/pdf.

Zvorikine:1960:SHT

[Zvo60] A. Zvorikine. The Soviet history of technology. Technology and Culture, 1(4):421–425, Fall 1960. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895625/pdf.

Zvorikine:1961:CUC

[Zvo61a] A. Zvorikine. Concerning a unified concept of technical progress: Zvorkine on Drucker’s “Work and Tools”. Technology and Culture, 2(3):249–253, Summer 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/ 1/article/895463/pdf. See comments [Dru61a].

Zvorikine:1961:HTS

[Zvo61b] A. Zvorikine. The history of technology as a science and as a branch of learning: a Soviet view. Technology and Culture, 2(1):1–4, Winter 1961. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/895569/pdf.

Zvorikine:1962:TLD

[Zvo62] A. Zvorikine. Technology and the laws of its development. Technology and Culture, 3(4):443–458, Fall 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895101/pdf.